A Complete Bibliography of TEST: An Official Journal of the Spanish Society of Statistics

and Operations Research

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 August 2025 Version 1.16

Title word cross-reference

0 [IPR11]. 1 [IPR11]. α [PJPW24]. AR(1) [HL09]. ARCH(p) [Stu01]. arcsinh [RJP11]. B [CRV21a, CRV21b]. D [FGS21]. δ [GLS12, GHF22, LBSM13, LBSM15]. δ ≥ 0 [LBSM15]. δ ≤= 0 [LBSM13]. ℓ_{1} [Flo15, BV23]. ϵ [GVGP08]. F [GP08, Pau11]. g [MS17, Pap18]. DD^{G} [CAFBdlF17]. k [DJ21, HC18, NLP17, SAE12]. L^{2} [EH20b, SR20, EH20a, JG20, Mei20, Ric20]. L_{1} [Lou05, Ant10, FL10, Lug10, SBvdG10a, SBvdG10b, SZ10, dB10]. L_{2} [dW02]. l_{∞} [IPT98]. L_{p} [LSW24]. M [Arc05, BM17, CJMW24, Kla05]. G^{0} [FG20]. μ [Aut08]. n [DJ21, HC18, NLP17, SAE12]. p [AABL18, DRB99, GP14, WG07, dlHRB01]. p > n [FB22]. ψ [HCS17]. R [MPS00]. sinh [RJP11]. \sqrt{ΔP} (\sqrt{Δ}) [Kok94]. T [Yan99, AVBI94, BA24, LWML15, RJP11, WTZL17, Wan19, WL22]. T^{2} [DLZ24]. × [AQGSM05, LC03, Pér94]. U [BHJ24, NSS13, SDZ20]. φ [MP01]. L^{2} [GZCZ19]. M [GP14]. W [GLSU15].

-classifier [CAFBdlF17]. -Contaminated [GVGP08]. -distance [Lou05]. -divergence [MP01]. -estimation [CJMW24]. -estimator [Arc05, GP14]. -estimators [Arc05, BM17]. -fold [AABL18]. -mixing [HCS17]. -norm [BV23, IPT98]. -norm-based [GZCZ19]. -optimal [FGS21]. -out-of- [DJ21, HC18, NLP17, SAE12]. -penalization [Ant10, FL10, Lug10, SBvdG10b, SZ10, dB10, SBvdG10a]. -priors [MS17, Pap18]. -quantile [LSW24]. -ratio [Pau11, Yan99]. -record [GLS12, LBSM13, LBSM15]. -recurrence [MPS00]. -shock [GHF22]. -stable [PJPW24]. -statistical [CRV21a, CRV21b]. -statistics [EH20a, JG20, Mei20, Ric20, BHJ24, EH20b, SR20]. -structure [SDZ20]. -tests [GP08, NSS13]. -thresholding [Aut08]. -time [Stu01]. -value [DRB99, WG07]. -values [dlHRB01].

19 [SMS23]. 1985-1997 [GPR00].

25-year-old [DT19b]. 25-years-old [BR19, DT19a, Kot19, Mor19, WZ19, del19a].

A. [GM98a]. ability [TA06]. absolute [LLZZ14, WY23]. absorption [AG15]. abstract [CFM23]. accelerated [YZ19, ZNSW19]. accidents [MD03]. Accounting [MG21]. Acknowledgement [Ano07, Ano09, Ano10, Ano11]. Acknowledgment [Ano08]. across [LLVR21]. actions [AD99, FMR24]. Active [GYP20, Gho20, LDR20a, LDR20b]. activities [LLVR21]. Adaptive [CSL22, CI10, ANZ22, CL15, LZYZ23, VPP22]. adaptive-to-model [LZYZ23]. adaptively [HQ23]. Additive [Eil20, GS20, Kne20, Woo20b, BB03, BM17, CDZ24, FBGM13, FGFBGM23a, FGFBGM23b, Goi19, HS22, IPPC13, KKLU19a, KKLU19b, LZ21, LNZ23, Rei19, Sch19, SLH99, SRHD19, WL17, WW12, Woo20a, ZLYH16]. adequacy [AHM18]. adjustability [MC09]. adjusted [CWZ21]. admissibility [VM99]. advances [CSN21, Huc21, Mar21, PGP21a, PGP21b, Sce21]. advantages [Are07, Bal07c, Hsi07a, Hsi07b, Mai07, Ner07, PS07, Shi07, Sic07, WM07]. Affine [DF19]. after [FRS24]. against [Kla05, MS10]. age [AF24, FC22]. aggregated [HPO04]. aggregation [HL09]. algorithm [BPD16, BP17, BBK97, EM15, LQR97, dRF92]. algorithms [CFP^{+}96, KPB^{+}00]. allocation [LD93, SRDMLF08, ZZ23]. allowing [CZ18]. Almost [SB92, CRV12, MNOP02]. alternating [GM24]. Alternative [De 07, BSNSS23, Cíz13, Rue92, Yek15]. alternatives [HPF12, MS10]. always [CF20]. Amari [ZNSW19]. among [HN18]. amplitude [Van12]. analyse [CDM11, GCY21]. analyses [GLGLM01]. Analysis [KWWZ23, LKM23, MCA11, QDG25b, sS12, TPB20, AAMDR20, AD23, AMO23, Are07, Ast14, ACR17, Bal07c, BCS20, Bar09, BFP14a, BFP14b, BB03, BMP^{+}94, Ber14, BT11, BS94, Bia14, BM14, BM24, CGB17, CR94, CFMPL24, DNR07, Deh14, Del24, Dry24, EY00, EdOS20, FMP18, FS12, GS24, GDS03, GMNH24, Gol25, GF16, Hog09b, HR14a, HR14b, Hor24, Hsi07a, Hsi07b, HP14, Hyn25, ID02, Kir14, Kok14, Lil25, LWML15, LA05, LMS^{+}99, LP25, Mai07, MP14, MGP00, MRMPEG07, MP93, Ner07, Nik24, Osi99, OMRM24a, OMRM24b, Paa14, PS25, PS07, PORCGP00, Pér94, PRSW16, PMPS11, PJH^{+}24a, PJH^{+}24b, QDG25a, Ras95, RPL01, Sch96, Shi07, SZ24, Sic07, SWZ25, SSG24, Stu01, SZL24, Tra14, Van12, VS14, Vov93, WL20, WYH24, WM07, WW06, WHS24a]. analysis [WHS24b, ZS24, vEZ94a, vEZ94b, MCA24]. Analytic [AD99]. analyze [BKK08, GMC08, KC08, Li08a, Li08b, MWN08, Sch08, Str08, TC08a, TC08b]. analyzers [WL22]. Analyzing [GG08]. ANCOVA [MK25]. and/or [ABA^{+}02]. ANOVA [Bia95, CAFB10, San97]. applicability [GSS11]. Application [FMR24, ARM08, AF07, BPY18, Ban18, BK22, Ber11, BCF25, Bic08, Bra11, Bri08, CGB17, CSS18, CGPV08, CBB^{+}95, Daw08, DPR11a, DPR11b, Fue08, GSG^{+}08a, GSG^{+}08b, HZY22, JSV16, Jol08, LM11, LM18, LPL15, LKM23, LB18, LLVR21, Pap11, Sch18, SMS23, Spe19, SRDMLF08, Vel11, WSCH15, WBG18a, WBG18b, WJ08]. applications [AABL18, BA24, DGSM11, DNR07, DGGJ05, DG22, Duc05, Gam14, HCS17, HMZ09, JGMRMG18, KWWZ23, MCL16, Men94b, NA23, PRSW16, WXH^{+}14, WL17, WWHY18, YZHG19, ZLWH17, ZF18]. applied [GS13, GMM19]. appraisal [Arn07, Bal07a, Bal07b, BL07, CS07, Dem07, Gui07, Jos07, KC07, Kun07, Nag07, NC07]. approach [AOV99, ANZ22, dZBT03, Ber05, BP21, Che07, CDM11, CBV24, DGSV98, FGV02, FMP18, GB16a, GG11, GF16, HX13, HG08, KN13, Kon13, LP18, LZYZ23, LDLDMF18, MLLC19, MT08, MMS21, NG93, NA07, NAV25, NARPV99, PSK25, PL03, RCSN22a, RCSN22b, Rom94, SAF21b, VD96, VP15, Vél01, YWLZ15, ZG07, dBIM20, SAF21a]. approaches [AMAEV13, SJD21, WG07]. Approximate [LAGR20, CM12, RH17, dBIM20]. approximation [AD99, CPW21, FFR16, GP03, Kon13]. Approximations [GP08, MS10, AABL18, Are14, Gef09, MP01, San97, WVT23]. AR-error [FF02]. Aranda [PRB20]. arbitrary [LGL25]. ARCH [MO99, ZZY25]. Archimedean [EH11, GNZ11b, JD11, Lam11, Seg11, Tsu11, Val11, WE11, GNZ11a]. arcsinh [Pew18]. area [Bel14, BLM16, Dag01, DGSM11, DKMR11, ELP20, ELP21, ELLV^{+}23, GS13, GMM19, HL14, HMS18, JL06, LMJ25, Mor14, PST14a, PST14b, RMG10, STPC12, SU14, SKR18, TJT16]. area-level [BLM16, ELP20, ELP21]. areas [MCN22, UMG09]. ARH [RMME19]. ARIMA [GM95]. ARIMA-based [GM95]. arising [Jon04]. ARMA [GJL96]. ARMAX [FF13]. armed [IV05]. arrays [dBJM09]. arrival [GHF23]. arrivals [SW16]. ART [HQ23]. artificial [MHSB20]. aspect [Ban18, LM18, Sch18, WBG18a, WBG18b]. Aspects [AD23]. assertion [GM98a]. Assessing [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, MMI00, WJ08, Zha14, LMJ25, PPB23]. Assessment [RAP12, De 07]. assessments [MP93]. assignment [VPR15]. associated [Mal98]. association [AD23, BdCPG14, BW18, CWZ21, vdL04]. assumption [NG93]. assurance [IP94]. asymmetric [IKvdH23]. Asymptotic [AQ01, BBGMPG11, FRS24, GS96, GS07, LLT18, LdUÁdCIP12, Men94b, Rah09, SS97, AH17, AHKS08, BR24, Duc05, Flo15, GLS12, Mur16, WVT23, dBCAM^{+}00]. Asymptotically [Kar00, CG02, DGG14, EMJ20]. Asymptotics [Kal22, WWY^{+}19, KB17]. atomic [BC07]. auctions [AJS04]. Augmenting [BKK08, GMC08, KC08, Li08a, Li08b, MWN08, Sch08, Str08, TC08a, TC08b]. Aumann [Ter08]. Auto [BdCPG14]. Auto-association [BdCPG14]. Autocorrelation [Raï10]. Autocorrelation-based [Raï10]. autocorrelations [Cha95]. autocovariances [Duc05]. Automated [WL20]. Automatic [VS09, Pér94, Ten22]. autoregression [Are14, CMGA^{+}24]. autoregressive [Ane12, Dol12, Dou12, FFZdC15, Fok12, Gal12, Gao12, Hei12, HHKM12, HS22, Ked12, LW24, MU23a, MU23b, PAT04, PPT24a, PPT24b, RCN09, RCN17, SL02, SYL25, Tjø12a, Tjø12b, WS24]. auxiliary [Bou94, BEMP20]. average [ELLV^{+}23, HSC10, PPT24a, PPT24b, RCN09, RCN17, Tse02]. Averages [MGN04]. averaging [ZPH18].

backfitting [SLH99]. backs [AV00]. balanced [Sha01, TS05, VH07, Ye95]. band [CY15, CJMW24, PPST96, ZY18]. Bandit [IV05]. bands [FPI96, MP93, WWY16]. Bandwidth [BCS23, RFFC17, AACRC19, BGN24, CD10, CASS19, MS11, MS15, YW08]. Barron [BBBV02]. Barron-type [BBBV02]. Bartlett [MC09, SK03]. based [AHMJ92, AH10, AN19, BBC10, BCN15, BDMPF22, BMRR21, BHGR17, BK22, BC07, BSFB20, BD00, BCD^{+}16, BM24, BP08, BA24, BP21, CM12, CG19, Cru10, CM22, DKMR11, DC11, DF19, DV25, Del24, DJ21, DFP21, DP18, Dry24, EMJ20, EM15, FP20, FPRMA04, Fer04, FC21, FGFBGM23a, FGFBGM23b, GP14, GDS03, GR04, GVM25, GB16b, Gre11, GZCZ19, GM94, He11, HX13, HJG19, Hig10, HHKM12, Hog09b, Hor24, HGV13, ICJ02, Ish11, KN13, LZYZ23, LQR97, MHSB20, Mar15b, Mat10, MG21, MC09, NBGP22, NAV25, Ozt19, Par11, PGB12, Pew18, Raï10, RSMJ19, Riv04, RMG10, Sah07, SDZ20, San10, SA21, SJD21, SGL14, She13, SM23, SSG24, TWHZ12, Ten22, Ton11, TPB20, VMS08, WLL15, WWT22, WH10a, WH10b, WHS24a, WHS24b, ZCL16, ZBS11a, ZBS11b]. based [ZLWH17, dLNA21, dUÁ11, dW02, dBIM20, GM95]. bases [KPB^{+}00]. Bayes [WT95, WT96, vEZ94a, AD99, De 07, LW22, MP93, O’H97, RC96, SJD21, TJT16, TW96, vEZ94b]. Bayesian [AHA03, AJS04, ALB22a, ALB22b, ACR17, BB03, BMP^{+}94, dZBT03, Ber05, BT11, BS94, Bia95, Bia97, BCD24, CS16a, CSR08, CIS18, CR94, CMGA^{+}24, dCCIS21, CGPV08, CD16, Dag01, DGSM11, DDP06, EY00, Fan01, FT10, GSCB92, GP94, GR07, GMRV19, GDVPP06, GVS98, GPSB^{+}97, HV14, HG08, Jah03, LD96, LAGR20, LIAV02, Mac23, MWAV19, MIR03, Men94a, Men94b, MGP00, Men99, MMR08, Mor22, Mor05, MG08a, MG08b, MMVP21, MM22, NG93, Osi99, PM25, Pap18, PC20, PAT04, Pér94, PPST96, RC96, Rod94, Rue92, RSSSL25, RSM06, Sal22, Sch96, Scu22, SDM20, Spe22, SMT22, Van95, VD96, VH07, Vil17, WGP07, WYH24, WW06, Yek15, ZB23]. Bayesianity [MEW01]. Bayesians [KGP^{+}93]. be [FS12]. behavior [BBGMPG11, RGEGMI13]. behaviour [MMR04, Tem00]. Behrens [NG93]. belonging [FJL24]. Benchmarked [UMG09]. benchmarking [Bel14, DGSM11, GS13, HL14, Mor14, PST14a, PST14b, SU14]. Benford [BCD23, CL23]. Berkson [GK19, XZ18]. Bernoulli [CTC12, LLTS24]. Bernstein [JSV16]. best [BLM16]. Beta [RCN09, RCN17, BCN15, FFZdC15, LLTS24, RS11, WW06, Sim22]. between [dCCIS21, CBB^{+}95, DDP06, DOT19, GLGLM01, NA23, XTZ20, vdL04]. beyond [HQ23, SP15]. bi [CSL22]. bi-level [CSL22]. Bias [DGGG08, EBGGY17, GF06, BPZ20, BCG09, CG09, CB07, JGMRMPMR05]. Bias-corrected [EBGGY17]. Bias-reduced [DGGG08]. biased [BC21, CA01, LL23, XTZ20, dUÁ02]. Bickel [LLG14]. Bienaymé [IKvdH23]. big [BC21, Büh19, Cao19, Del19b, GP19a, GP19b, GS19, Mar19, NS19, RACC19, SH19, Tsa19, VZ19]. big-but-biased [BC21]. Bin [DL04]. binary [DOT19, HL21, LB18, LLTS24, YH19]. Bingham [BR24]. BINMA [RSMJ19]. binomial [CP98, DDP06, HMS18, Kou98, LJW^{+}19, Men12, MMVP21]. binomial-logit [HMS18]. bioassays [SRDMLF08]. bioequivalence [Tse02]. biplot [CVO02]. Birnbaum [TPB20]. Bisexual [MJR08, GMdP11, MMR04, MMR08]. bivariate [BGLV23, BK22, Bia97, CZ18, DOT19, EG12, EBGGY17, Kou98, LLT18, LLTS24, SJD21]. Block [LWZZ23, CR17, CR22, LQR97, ZKR18]. Block-diagonal [LWZZ23]. blocking [GM94]. blocks [Rad09]. bonus [GDVPP06]. bonus-malus [GDVPP06]. Bootstrap [ARVV18, BCN15, CL10, DP18, VFVFGM07, WTZL17, ANZ22, BC07, BKK08, BZ17, Cao99, Cha17, DBZ17a, DBZ17b, FvdW08, GMC08, Guo08, KC08, Li08a, Li08b, LY17, LS17, LN17, MMI00, MWN08, Pew18, Rad98, Rad04, Rad09, RSW08a, RSW08b, SH08, Sch08, Str08, Tro08, TC08a, TC08b, VS09, Yek08, dBJM09]. Bootstrap- [DP18, WTZL17]. Bootstrap-based [BCN15]. Bootstrapping [AE06, FMM21, MO99, Nig06]. both [Ras95]. bounded [KWWZ23]. bounds [BGLM19, Flo15, GNDR09, Ryc19]. Box [Yan99]. branching [GMdP04, GMdP05, GMdP11, MJR08, MMR08, Rah09]. breakdown [RGEGMI13]. breaks [AE06]. bring [LMJ25]. building [Goi19, KKLU19a, KKLU19b, Rei19, Sch19, SRHD19]. Bump [CS23]. Burr [AHMJ92, AZ04]. business [BS94].

C [Rom94]. calculation [LC03]. Calibration [BR19, Kot19, Mor19, WH14, WZ19, del19a, DT19a, DT19b, EY00, GCS95, PR98]. call [AV00]. campaigns [PPB23]. can [FS12]. Canonical [Lop10, AD23, AN19, DNR07, KLYZ17]. capability [SBC^{+}98]. capture [LLLZ24]. capture-recapture [LLLZ24]. Carlo [CFP^{+}96, SJD21]. Case [Bül97, BGLV23, LBSM15, MMVP21, RI92, SRDMLF08, VD96, LBSM13]. cases [CIS18, Gho97]. casewise [ALYZ15b, ALYZ15a, CÖ15, Far15, Mar15a, RV15, Van15, Wel15]. categorical [BdCPG14, CG19, Di 12, LA05, Qin25]. categorized [MP00]. causal [Pea03, PLZ25, PPB23, SP15]. causal-noncausal [PLZ25]. cellwise [ALYZ15b, ALYZ15a, CÖ15, Far15, Mar15a, RV15, Van15, Wel15]. Censored [MCL16, MvdG16, AHMJ92, AH10, BBC10, BB03, BV23, Bha20, CJV05, CdU07, CM22, DP18, Fer04, HJV18, HC18, HPF12, LdUÁdCIP12, MHSB20, MFBG15, MLLC19, PK09, RAVL15, sS12, SG04, TWHZ12, TV23, VP17, WL22, WLL25, WLVF25, XTZ20, dUÁ02]. censoring [Arn07, Bal07a, Bal07b, BL07, BPD16, BP17, BSFB20, BCG09, BCGC16, CS07, CM10, CSS18, CI10, CBV24, DV25, Dem07, Gef09, GGQ24, Gui07, Jos07, KC07, Kun07, Nag07, NC07, WY09]. censorship [Yua05]. Central [PR05, BPY18, Bar97, Rom94]. chains [BC07, BCF25, Gho22, Rad04]. challenges [Are07, Bal07c, Hsi07a, Hsi07b, Mai07, Ner07, PS07, Shi07, Sic07, WM07]. Change [BHJ24, GN24, WG07, Ast14, Ber14, Deh14, DK23, GW17, HR14a, HR14b, HP14, Kir14, Kok14, LLT18, MP14, PLQ23, Tra14, WLL15, ZZT24]. Change-point [GN24, WG07, DK23, WLL15]. changepoint [PW20]. changes [HHKM12, Jar15]. characteristic [Gam14, JGMRMG18, MT08]. characteristics [EM21, EM22]. Characterization [CL94, AZ04, BMRR14, BE20, CM22]. characterization-based [CM22]. Characterizations [IPT98, BFR00]. characterizes [AF07]. charts [LC06]. Chebyshev [IKvdH23]. checking [DGSV98, GK19, LLHL24]. checks [Lie12]. Chentsov [ZNSW19]. chi [AQGSM05, CMLB05, TW14]. chi-processes [TW14]. chi-square [CMLB05]. chi-squared [AQGSM05]. choice [DDP06, MMVP21]. Cholesky [LGW19]. choose [GP14]. Choosing [JV98]. circuits [MPU13]. circula [KPJ22]. Circular [DFPT18, HLM22, MLG16, MVFFP21, RP09]. claim [BK22]. class [AD99, AAMDR20, Are14, BCCAVG22, CG02, CG19, EMJ20, ELORM15, EdOS20, FRV21, FS12, GHF23, HJG19, Kok94, MWAV19, MMR08, QQP21, RS11, She13, WWHY18]. classes [PPST96, Rom94, SP92, Kla05]. classical [Ast14, Ber14, Deh14, HR14a, HR14b, HP14, Kir14, Kok14, MP14, Tra14]. classification [BCS20, BP08, GLW25, SGL14]. classifier [CAFBdlF17]. classifying [Agu16, AV16, BL16a, FB16, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. climate [KWWZ23]. clinical [ANZ22]. close [GP08, Jon96]. cloud [PLRC13]. cluster [May02]. clustered [Li17, YWLZ15]. clustering [BA24, CP12, Cey14, FP20, LQR97, RGEGMI13, WL22, WLL25, WG21]. co [AE06]. co-breaks [AE06]. coca [LB18]. coefficient [AGV17, AGV14, Bra20, CDZ24, DGGG08, GWHY14, HJV18, LW24, LGW19, Pru20, TZ15, YWLZ15, YLL19, ZLWH17]. coefficients [AH17, FFO20, LGL25, TS05, WC98]. coherence [VM99]. Coherent [DDM^{+}95, NR10, Nav16, NLP17, NM20, ZZ23]. cohort [WZYW18]. cointegration [AE06]. collected [AVR00]. columns [LNZ23]. combination [DDM^{+}95, Mur16]. combinations [LHB^{+}17]. combinatorial [DL04]. combined [KL14]. combining [FB22]. Comment [Bra11, Lug10, Raj12]. Comments [Agu16, Ane12, AV16, ARM08, Ant10, Are07, AG16, Arn07, Ast14, BL07, Bal07c, Ban18, BR19, Bel14, Ber14, Ber11, Bia14, BL16a, Bic07, Bic08, BM14, BM24, BKK08, BZ17, Bri08, BL16b, Büh19, Cab09, Cao07, Cao19, Cao23, Car10, CM07, Cas12, CS07, CG15, Cha17, CJ23, CGR10, CÖ15, Cru10, CSN21, CA13, DW09, Daw08, DC11, Deh14, Del19b, Del24, Dem07, Det13, Dol12, Dou12, DFK12, Dry24, Efr07, Eil20, EH11, FZZ10, FL10, Far15, FB13, FB16, FvdW08, FH19, Fok12, Fue08, GYP20, Gal12, Gam13, Gao12, GS24, Gel15, GS19, GW16, Gho20, GMNH24, Goi19, Gol25, GMC08, GMMC10, GT18, Gre19, GS20, Gui07, Guo08, HL14, Hal07, HW23, He11, Hec10, Hei12, Her14, Hig10, Hog09a]. Comments [Hoo10, HM16, Hor07, Hor24, Huc21, HP14, Hyn25, Ish11, JD11, JN09, JG20, Jol08, Jos07, KC07, Ked12, KC08, KC09, Kir14, Kne20, Kok14, Kot19, Kun07, LW07, LM11, Lam11, Lam23, LM18, Li08a, Li08b, LL09, Lil25, Lin15, Lit09, LY17, LS17, LN17, LP25, LR12, Mai07, Mam07, Mar21, Mar15a, Mar19, MF19, MP14, MB18, Mat10, McK09, MWN08, Mei13, Mei20, Mol09, Mor22, Mor14, Mor19, Mül07, MK09, Muñ14, NS19, Nag07, Ner07, NC07, Paa14, PS25, Pap11, Par11, PS07, PZ09, Pin14, Rei19, RACC19, Ric20, Rit23, RV15, RM15, SF18, Sal22, SS09, SSGM23, SNC16, San10, SH08, Sce21, Sch08, Sch18, Sch19, Scu22, Seg11, Sen10, SH19, Shi07, SZ24, Sic07, Spe09, Spe13b, Spe13a]. Comments [Spe22, Sta12, SRHD19, SU14, SSG24, Str08, SZ10, SR20, Tha09, TL14, TF23, TS15, Ton11, Tra14, Tro08, Tsa19, Tsu11, Uga09, Val11, Van15, VZ19, Van13b, Van13a, Vel09, Vel11, Vel13, VS14, Wag16, Wan10, WE11, WM07, Wel15, WJ08, WZ19, Yek08, ZL16, Zha16, ZS24, dUÁ11, dUÁ13, dUÁ23, dB10, del19a]. common [BLBB03, Hay14, MCA11, MCA24]. compact [CPW21]. comparative [AHMJ92]. Comparing [BW18, BCS15a, BCS15b, CG15, FG20, Gel15, Lin15, RM15, TS15, Pau11]. Comparison [Cey14, GS06, MG08a, MG08b, NLP17, TV23, CGPV08, FZWZ15, Oht98, PC20, RH17, SLH99, VFVFGM07]. Comparisons [BMRR21, NR10, NM20, HF18, Nav16, SG04, ZZ23]. Compatible [CGPV08, AG98]. competing [Gef09, SGR07, WG07]. competitiveness [AFO22a, AFO22b]. Complete [BR24, WSCH15, WXH^{+}14]. completeness [MNOP02]. complex [CIS18, CAM15, DP23, HLM22, JGMRMPMR05]. complex-valued [HLM22]. complexity [BGLV23]. component [AOV99, ARS22, Bar09, DJ21, GF16, Jah03, LMS^{+}99, MK14]. components [AD23, CGB17, FBKV14, GM03, NRS06, NLP17, NSS13, RAP12]. Composite [WLVF25, HMW23, JS22, MP00, Rya12, Yek15]. Compositional [EPG19a, ELP20, ELP21, ID02, EPG19b, FH19, Gre19, MF19]. compositions [ELLV^{+}23]. compound [Bha20, CR17, CR22, RSSSL25]. Computation [BP17, Eil20, GS20, Kne20, SS92, Woo20a, Woo20b]. computationally [BA24]. computed [GS96]. computing [GDVPP06]. concave [SP92]. concentrated [KPB^{+}00]. concept [CRV21a, CRV21b, Thà23]. concepts [VM99]. conceptual [BBLC23]. concurrent [FGFBGM23a, FGFBGM23b, GM95]. condition [AQGSM05]. Conditional [Di 12, GGQ24, SYL25, AG98, AFO22a, AFO22b, BGN24, CRV12, CY15, DGG14, GR18, GMdP11, JSV16, JGMRMPMR05, LKM23, LP18, LdUÁdCIP11, LdUÁdCIP12, MVYA19, Mar15b, NRS06, PJPW24, VCS00, ZFX15]. conditionally [JGLM20, MNP24]. Conditioning [LIAV02]. conditions [BMRR14, CRdUÁH19, Li17, VFVF00]. condor [LDLDMF18]. Confidence [LX16, MC09, BBC10, CY15, CJMW24, GWHY14, JvdG17, LPQ11, PR98, RC94, Tse02, WWY16, ZY18, ZLYH16]. Conformal [Mar24]. conjugate [BLBB03, GPSB^{+}97, MGP00, San97]. connection [MC09, WTZL17]. connections [NA23]. Connor [EG13]. consecutive [SAE12, SKS13]. conservative [FPI96]. Considerations [Hog09b, Rod94]. Consistency [GR07, MS11, MS15, RCT14]. consistent [BCDG08]. constant [GP14]. Constrained [TJT16, PC20, PPST96, ZB23]. constraint [Kon13]. constraints [BRV20, CALF15]. constructed [FS12]. Construction [LPL15, BDMPF22, Mac23]. Contaminated [GVGP08]. contamination [ALYZ15b, ALYZ15a, ÁEdBCAM16, CÖ15, Far15, Mar15a, RV15, Van15, Wel15]. contemporaneous [GN99]. context [LD96]. contingency [GVGP08, PC20, Pér94]. continuation [Bar97]. continuous [BQ04, GNDR09, HMV05, LLG14, PMPS11, Vov93]. contoured [CR17, CR22]. contralateral [Lop08]. contrast [ELORM15]. contrats [Cle02]. Contributions [dBCAM^{+}00, CA01, FF12]. Control [FvdW08, Guo08, RSW08a, RSW08b, SH08, Tro08, Yek08, Mar24, RH17, SBC^{+}98, Xia17]. controlled [GMdP04, GMdP05]. Controlling [ZF18]. convenience [CBBRMB19]. converge [Sch98]. Convergence [Arc05, AM16, CRV12, CRV21a, CRV21b, HMV05, WXH^{+}14, WSCH15, dS18]. convex [FPRG17]. convex-type [FPRG17]. convexity [CFMPL24]. Convolution [WMS25]. Copula [BGLV23, BK22, DV25, SDM20, DOT19, EH11, EM15, FFO20, GT11, GR04, GNZ11a, GNZ11b, JD11, JSV16, KN13, Lam11, RF24, SJD21, Seg11, Tsu11, Val11, WE11, dUÁV13]. Copula-based [BK22, EM15, KN13, SJD21]. copula-graphic [dUÁV13]. copulas [PQ10, PQV12, SKS13]. coregionalization [GSBS04]. coregionalized [Mac23]. Correct [PKB23]. corrected [CB07, EBGGY17, LJC10]. Correction [AFO22a, Bül97, CRV21a, CR22, EM22, ELP21, FGFBGM23a, MU23a, MCA24, OMRM24a, PPT24a, RCSN22a, SAF21a, Raï10]. Corrections [WT96, vEZ94a, SK03]. correlated [BHGR17, GALT23, MVOC20, Men99, TPB20]. correlation [AD23, CDZ24, Csö02, CZ18, GW17, GS06, GLGLM01, HF19, HMZ09, Kra09, LLTS24, SP24, ZY18]. correlations [AFO22a, AFO22b, KLYZ17]. correspondence [Pap18]. corridor [GWHY14]. corridors [ZLYH16]. cost [CALF15]. count [Ane12, BK22, BW17, BdM22, DC11, Dol12, Dou12, Fok12, Gal12, Gao12, He11, Hei12, HHM24, Ish11, KWWZ23, Ked12, Nik24, Par11, PLZ25, SWMG18, Tjø12a, Tjø12b, Ton11, WH14, ZBS11a, ZBS11b, dUÁ11]. counting [CTC12, GLS12]. counts [SMS23]. couples [MJR08]. covariance [ASLFP13, CVO02, CR17, CR22, FJL24, GZCZ19, HF19, HC21, HN18, LWZZ23, MLG16, NG93, Spe19, SRDMLF08, Xia17, ZKR18]. Covariate [CWZ21, DNR07]. Covariate-adjusted [CWZ21]. covariates [BGN24, BFP14a, BFP14b, BBGMPG19, Bia14, BM14, CSL22, Cao23, HW23, HHM24, ID02, Lam23, LLLZ24, LCPJ23a, LCPJ23b, Paa14, SSGM23, TZ15, VS14, Yua05]. covariation [CGB17]. COVID [SMS23]. COVID-19 [SMS23]. Cox [FTSM22, Yan99]. CQR [TZ15]. Cramer [Rao01, FMR24]. Cramer-Rao [Rao01]. credible [Ber05]. credit [SAF21a, SAF21b]. crime [KWWZ23]. criteria [BCN15, BSFB20, BCS15a, BCS15b, CG15, Gel15, Lin15, RM15, TS15]. Criterion [ZB23, OR98]. critical [AACRC19, EH20a, EH20b, GP03, JG20, Mei20, Ric20, SR20]. cross [Bel14, CZ18, De 08, HL14, Mor14, PST14a, PST14b, SU14, Yek15, dRF92]. cross-correlation [CZ18]. cross-products [De 08]. cross-sectional [Bel14, HL14, Mor14, PST14a, PST14b, SU14]. cross-tabulated [Yek15]. cross-validation [dRF92]. crossed [MSK23]. cub [IMP16]. cultivation [LB18]. cum [SB92]. cure [BP21, Cao23, CHV20, DV25, GJV20, HW23, Lam23, LAGR20, LCJC17, LCPJ23a, LCPJ23b, LC12, PM25, SSGM23]. currents [BCS20]. curvature [CS23]. curve [Men99, ZY18, ZNAG^{+}01]. curves [DGGL11, FZWZ15, TV23, VFVFGM07]. CUSUM [LC06]. cylinder [JMS21].

damages [MS03]. Data [BBC22a, Cao19, GP19a, PJH^{+}24a, QDG25b, AOV99, Agu16, AHMJ92, AGMT23, AV16, AVR00, Are07, ACR17, BQ04, BBC10, Bal07c, Ban18, BCS20, BOQ17, BFP14a, BFP14b, BK22, BT11, BCCAVG22, Bia14, BL16a, BdCPG14, BM14, BHJ24, BM24, BC21, BKK08, BdM22, Büh19, BKM22, CLH^{+}20, CJV05, CdU07, CG19, CA01, CAFB10, CM22, DW09, DC11, Del19b, Del24, DP18, Dry24, EPG19a, EPG19b, FRV21, FBGM13, FB16, Fer04, FH19, FM01, FS12, GP19b, GS96, GDS03, GK19, GS19, GCY21, Gol25, GMC08, Gre19, GWHY14, GM95, GG08, GM94, HQ23, He11, HJV18, HJG19, HC18, Hid99, HHM21, Hog09a, Hor24, HL21, HMW23, Hsi07a, Hsi07b, HZY22, Hyn25, HN18, IM09a, IM09b, Ish11, ID02, JS22, JV98, KWWZ23, KC08]. data [KC09, Kon13, LdUÁPF25, LM18, LPL15, Li08a, Li08b, Li17, LHB^{+}17, LDR20a, LDR20b, LLTS24, LdUÁdCIP11, LdUÁdCIP12, Lil25, LCB19, Lit09, LA05, LNZ23, LLLZ24, LMS^{+}99, LP25, LGW19, MMI00, Mai07, MHSB20, Mar19, MF19, Mar15b, MLLC19, MWN08, MNP24, NS19, Ner07, NARPV99, OPV21, Paa14, PS25, PM25, Par11, PS07, PMPS11, Qin25, QDG25a, RFFC17, RACC19, SCZ23, SNC16, Sch08, Sch18, SGL14, SLP21, SMS23, sS12, She13, SH19, Shi07, Sic07, SG04, SWZ25, SSG24, Str08, SKS13, SVY20, SMT22, TA06, TV23, Ton11, Tsa19, TPB20, TC08a, TC08b, Uga09, VZ19, Van12, VS14, WBG18a, WBG18b, Wan19, WL22, WM07, WH14, WG21, WZ12, WHS24a, WHS24b, XZHW16, XTZ20, YRR15, YWZ16a, YWZ16b, Yek15, YWLZ15, YZHG19, Yua24, ZKR18, Zha14, ZL16, Zha16, ZBS11a, ZBS11b, ZZT24]. data [ZNAG^{+}01, dUÁ02, dUÁ11, dRF92, Büh19, Del19b, GYP20, GP19b, GS24, GS19, GMNH24, Mar19, NS19, PJH^{+}24b, RACC19, SH19, SZ24, Tsa19, VZ19, ZS24, Gho20]. Data-driven [BBC22a, SCZ23]. Dating [GW17]. David [Bül97]. DC [LL23]. debiased [HMW23, LL23]. Decision [BP08, Bül97, FI03, MIR03, RI92]. Decisions [Rab98, ZF18]. decomposition [LGW19]. deconvolution [YW08]. decorrelated [LSW24]. decreasing [MLDJ16]. default [SAF21a, SAF21b]. definition [Van12]. degenerate [Lef03]. degradation [FC21, FC22]. degradation-based [FC21]. degrees [GHF08]. deletion [SRGS00]. dense [CLH^{+}20]. densities [BCDG08, CRdUÁH19, GT11, KPJ22, SS92]. Density [FFL13, BBBV02, BPZ20, BCG09, BCGC16, Cey14, CS23, CL15, CS97a, DBC^{+}97, GB16a, Jon96, LS09, Lou05, PPST96, RFFC17, SYL25, Ten22, VS09, YW08]. dependence [AAMDR20, AQ01, ARVV18, EBGGY17, FvdW08, FF12, FMP18, FF21, FFO20, Guo08, MS03, MF05, NBGP22, RSW08a, RSW08b, SH08, Tro08, VFVF00, Yek08]. dependent [Ber11, BD00, Bra11, BKM22, BR18, CRV12, CBV24, DV25, DPR11a, DPR11b, EM15, Gef09, GCY21, GALT23, GHF23, HZY22, LM11, LTWY23, LdUÁdCIP11, LdUÁdCIP12, MMI00, MMR08, MG21, NLP17, NPM22, Pap11, PN22, Tem00, Vel11, VFVFGM07, WXH^{+}14, YZWH17, ZJ08, dRF92]. depending [MJR08]. Depth [SDZ20, ACZ06, CFM23, DF19, SA21, SGL14, SLP21]. Depth-based [SDZ20, DF19, SGL14]. depths [EN25]. derivative [JSV16]. derivatives [LDLDMF18]. derived [AD99]. Deriving [Rab98]. design [BHGR17, Bha20, DOT19, Mon11, PKB23, SDM20, WZYW18]. designs [ASLFP13, CALF15, CD16, DFP21, FGS21, GS06, JGMRMPMR05, OR98, Pru20, RH17, SP24]. Detecting [Jar15, BCD24]. detection [BHJ24, DK23, GN24, NECA21, PW20, ZLWH17, ZZT24]. determination [BCD24, De 07]. developments [LA05]. deviation [WY23, YZWH17]. deviations [Arc02, Lou05]. Deville [BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. diagnostic [AW20, CP98, SRGS00]. diagnostics [DMUOG15, JGMRMPMR05, MRMPEG07, RS11, TLdPDG19, WG07]. diagonal [LWZZ23]. Difference [dAM03, AGMT23, FGFBGM23a, FGFBGM23b, XTZ20]. differences [dCCIS21]. different [AVR00]. Differentiating [RMLDG03]. diffusion [FFR16, GRT01, Lef03]. digraph [Cey14]. Dimension [FRV21, Agu16, AV16, BNY21, BL16a, FB16, FGV02, FRS24, LP18, LTWY23, LZYZ23, LLR^{+}24, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. dimensional [BR24, BPY18, BHJ24, BKK08, BZ17, BP08, CGX23a, CGX23b, Cha17, CJ23, DBZ17a, DBZ17b, FTSM22, GB16b, GMC08, HMW23, HN18, JvdG17, JS22, KC08, LWZZ23, Lef03, Li08a, Li08b, LHB^{+}17, LGL25, LY17, LS17, LN17, MWN08, NECA21, NAV21, PP19, Rit23, Sch08, SWZ25, Str08, TF23, TC08a, TC08b, VA25, Xia17, ZZT24, dUÁ23]. dimensions [MvdG16]. direct [Kon13, PPB23]. direction [GM24, WL17]. directional [CSN21, Huc21, Mar21, PGP21a, PGP21b, Sce21, ZF18]. Dirichlet [CP98, EG13]. disaggregating [GM95]. disaggregation [GN99]. discovery [FvdW08, Guo08, Mar24, OPV21, PDBNR23, RSW08a, RSW08b, SH08, Tro08, Xia17, Yek08, ZF18]. Discrete [GW99, LBW01, BCF25, DJ21, GNDR09, IP94, RCSO03, Rya12]. Discretized [Sim22]. discriminant [AMO23, DNR07, MRMPEG07]. discriminate [DOT19]. discriminating [GB16b]. discrimination [CD16]. discriminatory [De 07]. Discussion [ACW^{+}98, Pau11, MCN22, QDG25b]. disease [CBBRMB19]. disjoint [Rad09]. disparities [HV14]. disparity [KB17]. dispatch [Her14, Muñ14, Pin14, TL14, ZGGX14a, ZGGX14b]. dispersion [GPC10, KV25, MVYA19]. Distance [RC94, BCD^{+}16, FG20, GK19, Gre11, Lou05, XZ18, dW02]. distance-based [BCD^{+}16]. distances [Mit92]. distinguishable [FFF17]. distorted [NA23]. distortion [CDZ24]. distributed [DLZ24, HMW23, Sch98, dBJM09]. Distribution [ACW17, CMLB05, GB16b, GLW25, LBSM13, LBSM15, AVBI94, Are14, BBS18, dZBT03, BBK97, Duc05, FG20, GP94, GG21, GD10, GMdP04, HX13, HHKM12, HM10, JSV16, LPQ11, LdUÁdCIP12, LWML15, LX16, MCA11, MCA24, MF05, May02, NK06, PJPW24, PK09, RCSN22a, RCSN22b, Ryc19, VMS08, VCS00, Vil17, WWY16, ZY06, ZFX15, ZZY25]. Distribution-free [ACW17, GB16b, GLW25]. distributional [ARS22, BE20, Goi19, KKLU19a, KKLU19b, Rei19, Sch19, SRHD19]. Distributions [AG98, AT08, AZ04, BR24, BQ04, Bar97, BMRSL15, BCF25, BCCAVG22, Bia97, BA24, BR18, CQ05, CSR08, CTC12, CS97b, CAM15, CR17, CR22, CA13, Duc01, EG13, FB13, Fer99, GLSU15, GMM19, Gre11, GM03, Hay14, HF18, HSK05, IPPC13, IPT98, IP94, Jon04, JMS21, Kla05, Kou98, LD96, LBW01, MWY21, MWAV19, MT08, MPS00, NRS06, NA23, PGB12, Pew18, PP19, Pol13a, Pol13b, Rah09, RP09, Riv04, RCSO03, RJP11, Ryc19, Spe13b, Van13b, Vel13, ZCL16]. disturbances [FMM21]. DIVAS [GS24, GMNH24, PJH^{+}24b, SZ24, ZS24, PJH^{+}24a]. Divergence [BBBV02, CS97a, FGFBGM23a, FGFBGM23b, GB16a, JS01, LS09, LLR^{+}24, MP01, PRSW16, SYL25]. Divergence-type [BBBV02]. diverging [CSL22]. diversity [PRSW16]. domains [LLVR21]. domination [Thà23]. dose [RH17]. dose-escalation [RH17]. Doubly [MFBG15, Fer04, FJL24, sS12, She13, SVY20]. doubly-truncated [She13]. driven [BBC22a, SCZ23]. dropouts [Mar15b]. Dynamic [Car10, CGR10, FZZ10, GMMC10, Hec10, Hoo10, MY10a, MY10b, Sen10, Wan10, FFF17, WZ12, dNGL16]. Dynamical [RMME19, Cru10, Hig10, Mat10, San10, WH10a, WH10b].

E. [GM98a]. ECM [AE06]. economic [GN99, LLVR21]. edf [Pew18]. edf-based [Pew18]. Editorial [MHJG18]. Effect [HL09, HPO04, AFO22a, AFO22b, BJ13, BBGMPG11, Cha95, DP18, FGS21, LGL25, Yan99]. effective [Agu16, AV16, BL16a, FB16, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. effects [CLH^{+}20, LX16, MCL16, MSK23, PPB23, PKB23, RAP12, SP15, TLdPDG19, VH07, WY23, WZ12, Ye95]. Efficiency [KL14, DS95, JS01, Mur16, RCT14, SS97]. Efficient [XZHW16, BA24, CLH^{+}20, CJMW24, EMJ20, KCV23, Kar00, WL17, WMY20, WZYW18]. EGARCH [FWZ18]. eigenvectors [Lop24]. elaboration [CP98]. electricity [LPL15]. elementary [AF07]. elements [GLS12]. elevation [Ban18, LM18, Sch18, WBG18a, WBG18b]. Elfving [OR98]. Eliciting [EG13]. Eliminating [BFR00]. elliptical [DMUOG15, HS15, Osi99, RAP12, SL17]. elliptically [CR17, CR22]. emphasis [EH20a, EH20b, JG20, Mei20, Ric20, SR20]. Empirical [ACZ06, BLM16, LPQ11, LLZZ14, LZ21, MT08, MP93, SKS13, AABL18, BOQ17, BBS18, CQ05, CM12, CV09b, CV09a, CDZ24, Gam14, GP94, GR04, HJG19, LL09, LLLZ24, LGW19, McK09, MC09, PZ09, PQ10, PQV12, PRSW16, Rom94, SS09, SDZ20, SJD21, Spe09, Vel09, WT96, YZ19, ZLWH17, dBCAM^{+}00, WT95]. endpoint [GGS12, LPQ11]. English [AJS04]. enhanced [Agu16, AV16, BL16a, FB16, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. ensemble [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, WJ08]. Entropy [dLNA21, AZ04, PRSW16]. Entropy-based [dLNA21]. environments [MMR04]. epidemic [FFR16]. equality [AH17, Bor01, CRdUÁH19, FB22, GZCZ19, JGCRJJ22]. equation [Ava22, CSR08]. equations [BS97, CALF15, SDZ20, TPB20]. equicorrelated [Osi99]. equidispersion [BW17]. Equilibrated [EE92]. equilibrium [CGPV08]. equivalence [Ter08]. ergodic [Rya12]. Erratum [GS15a, Hog09b, MS15, RCN17]. error [ABS01, BGLM19, BEMP20, BKM22, CSS18, DKMR11, ELLV^{+}23, GK19, HHKM12, HZY22, HGV13, HM10, LLZZ14, LMJ25, MWY21, MMI00, NAV25, Oht98, Raï10, Sha01, TJT16, TS05, VMS08, WLL15, WMY20, WT95, WT96, WVT23, ZFX15, FF02]. error-based [WLL15]. errors [AVR00, BS97, BBBV02, BHGR17, BD00, CDZ24, CFRG10, FFZdC15, GJL96, GALT23, HCS17, MVOC20, Men99, NARPV99, Osi99, Raï10, RAVL15, RMME19, VFVFGM07, WYH24, WW12, YZWH17]. errors-in-variables [HCS17, RAVL15, WW12]. escalation [RH17]. estimate [CB07, May02, Ten22]. estimate-based [Ten22]. estimates [BR12, BKM22, HMZ09, MMI00, UMG09, YSV96]. Estimating [GS15a, GS15b, GGS12, MU23b, RSM06, SMFP13, Yan99, ZFX15, AV00, BS97, BBGMPG11, Cao99, EM15, HMZ09, SCZ23, SJD21, TPB20, YWLZ15, MU23a]. Estimation [AHMJ92, DKMR11, GL07, GdW12, GMdP05, Men12, MCN22, PJPW24, RS93, TS05, VCS00, WZ12, WZYW18, ZZF19, ALYZ15b, ALYZ15a, AG19, AT05, AW20, AGMT23, AGV14, AVR00, Ava22, BGN24, BS97, Bel14, BPZ20, BNY21, dZBT03, Ber05, BCF25, BBGMPG19, BC21, Bra20, BCG09, BCGC16, BKM22, CG09, CLH^{+}20, CJMW24, CJV05, Cao23, CGB17, CL15, CPZ12, CS97a, CS16b, CL10, CSS18, CA01, CFRG10, CÖ15, D’E96, Dag01, DGSM11, DBC^{+}97, DGG14, EBGGY17, EE92, EPS06, ELP20, ELP21, ELLV^{+}23, Fan01, Far15, FFL13, FRS24, FMGE^{+}99, GSCB92, Gan13, GG21, GR07, GS13, GB16a, GPC10, GM98b, GG11, GGQ24, GF06, GMC93, GMdP04, GLS12, GMM19, HL14, HW23, HJV18, HMS18, HMW23, JvdG17, JSV16, JL06, JGMRMG18, Jon96, JMS21, KCV23, KV25, KB17]. estimation [Lam23, LW24, LdUÁdCIP11, LMJ25, LCJC17, LCPJ23a, LCPJ23b, Lou05, LJW^{+}19, Mac23, Mar15a, MVFFP21, MM02, MMR08, Mor14, MW04, NAV25, Pér94, PST14a, PST14b, Qin25, Rad98, RFFC17, Rod94, RV15, STPC12, SSGM23, SL17, SA21, SU14, SAF21a, SAF21b, SKR18, TA06, TJT16, Van15, VMS08, VS09, Vil95, Vil17, WMY20, Wan22, WMS25, Wel15, WLVF25, XZHW16, YW08, YLL19, ZLWH17, ZPH18, dUÁ02, vEZ94a, vEZ94b]. estimations [HL21, LL23]. Estimative [Mit92]. Estimator [XYY22, Alv01, Arc05, CL10, Dub99, GP14, HF19, HC21, LS09, LdUÁdCIP12, MVYA19, PR05, RSF97, SYL25, TWHZ12, TW96, WWY^{+}19, dUÁV13]. Estimators [XTZ20, AM16, Arc02, Arc05, BBBV02, BR09, Bia95, BBGMPG11, BBC22b, BM17, BRV20, CG02, CdU07, Cíz13, DGGL11, DKMR11, DJ21, DGGG08, DS95, EMJ20, ELORM15, Flo15, FRS24, GS07, GMdP11, HL09, HGV13, JS01, LP18, LL23, LJC10, MGN04, MS11, MS15, Oht98, Rah09, RCT14, RMG10, Sha01, SB92, SCJS07, SVY20, VA25, WY23, WT96, dAM03, WT95]. Europe [GPR00]. EV [WSCH15]. evaluation [RC94, WMC^{+}96]. event [HZY22, LIAV02, PM25]. event-related [HZY22]. events [BT11]. evidence [GVS98, MM22]. evolutionary [MMS21]. evolving [CPW21]. Exact [BBC10, HG18, CAM15, Di 12, MCA11, MCA24, TS05]. exceedances [Fer99]. exceptionally [WVT23]. excess [AACRC19, BNOR08, GG11]. exchangeability [GM98b]. exchangeable [NRS06, ZKR18]. exciting [SYL25]. Exhaustivity [MNOP02]. existence [MC09]. exogenous [CS16b, GRT01]. expansions [BR24, GS96]. expectations [Ryc19, SK03, Ter08]. expected [DGGJ01]. experiment [Bha20]. experimental [CD16]. experiments [GLÁM21, Sch98]. expert [DDM^{+}95]. Exploratory [QDG25a, QDG25b, AAMDR20, Gol25, Hyn25, Lil25, LP25, PS25]. Exploring [CWB^{+}93, GSS11]. Exponential [GPSB^{+}97, Kok94, Fer04, Mal98, NRS06, NM20, NPM22, NBGP22, PAT04, Pom96, PK09, vEZ94a, vEZ94b]. exponentiality [CM22, Kla05]. exponentials [RSM06]. exponentiated [AH10]. exposure [Kon13, LGL25]. exposure-mediator [LGL25]. Extended [DLZ24, AE06, CW23, GPC10]. Extensions [Ast14, Ber14, Deh14, HR14a, HR14b, HP14, Kir14, Kok14, LIAV02, MP14, Tra14, GPSB^{+}97, Woo20a, Woo20b, Eil20, GS20, Kne20]. extent [RCSN22a, RCSN22b]. extrapolation [FPRMA04]. extrapolations [And97]. Extremal [MF14, AAMDR20, FF12, MF05, Tem00]. extreme [Bar97, CQ05, DGGL11, GS15a, GS15b, GdW12, MF05, NA23, dNGL16]. Extremes [DHJL15, FF13, HSC10, Ber11, Bra11, DPR11a, DPR11b, LM11, NA07, Nig06, Pap11, Vel11].

F [GM98a]. factor [DBC^{+}97, Fan01, LWML15, RPL01, VH07, WL20, WL22, YZHG19]. factorial [CWZ21, DFP21, GLÁM21]. factors [AFO22a, AFO22b, De 07, GRT01, LW22, O’H97, Ras95, RCT14, RC96]. failure [AHMJ92, BR18, CSL22, YZ19]. failure-dependent [BR18]. False [OPV21, FvdW08, Guo08, Mar24, PDBNR23, RSW08a, RSW08b, SH08, Tro08, WVT23, Xia17, Yek08, ZF18]. Families [Jon04, AG98, Cey14, Csö02, GPSB^{+}97, HF18, Kok94, Mal98, MMVP21, NBGP22, Pom96, RP09, Ryc19, dW02]. Family [SCJS07, Arc05, MPU13, SLL24]. Fast [CGB17, CFRG10, LQR97, LC06]. Fay [BEMP20]. feature [GLW25, JT22]. features [AG15, CS23]. Feedback [RS93]. Field [BDMPF22]. fields [BCCAVG22, CP12, CPW21, ELORM15, FPRMA04, GT18, Mac18a, Mac18b, Mac23, MB18, SF18, VS09]. filter [CR97, MGRA98]. filtering [CGCK11, ICJ02]. filters [SA21, Zha14]. financial [LDLDMF18]. find [Zha22]. Finetti [IPR11, VPR15]. Finetti-type [IPR11]. finite [BK22, CG19, Gho22, HF18, JGMRMG18, May02, MG21, Ozt19, Rum03, SCZ23, Sim22, SB92, SCJS07, SLH99, WWY16, dAM03]. First [Lef03, Are14, GVM25, LBW01]. first-order [Are14, GVM25]. First-passage [Lef03]. Fisher [HMV05, NG93]. Fit [Det13, Gam13, GMC13a, GMC13b, Mei13, Spe13a, Van13a, dUÁ13, Cab09, CCR21, CP98, CH09a, CH09b, CV16, DRB99, Duc01, FWZ18, GT11, GJV20, HX13, JN09, JT22, JGLM20, MHSB20, MT08, MVOC20, Mol09, MP00, MK09, Pew18, SLL24, Ten22, Tha09, VMS08, dW02, dBCAM^{+}00]. Fitting [AAMDR20, BBS18, BdM22, CA13, FB13, Pol13a, Pol13b, Spe13b, Van13b, Vel13]. fixed [BE20, Pér94]. Flexible [WLL25, LLTS24, MWAV19]. flotation [GLÁM21]. flow [HPO04]. fold [AABL18]. follow [CD16]. follow-up [CD16]. following [GJL96]. force [LLVR21]. forecast [MO99]. Forecasting [AOV99, Vov93, CS16b, Her14, HPO04, LPL15, Muñ14, Pin14, TL14, ZGGX14a, ZGGX14b]. forecasts [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, WJ08]. forest [AG16, BS16a, BS16b, BL16b, GW16, HM16, Wag16]. form [HC21]. forma [LL23]. forms [BOQ17, MQ01]. formulae [SK03]. formulation [FGFBGM23a, FGFBGM23b, RF24]. forward [ACR17]. Foundations [Bül97, RI92]. Fourier [FFO20, HHKM12, KPJ22]. fractal [FGV02]. fraction [GJV20, LC12]. fractional [FPRMA04, O’H97]. fragmentation [AG15]. frailty [CNJ08, MLDJ16]. framework [BCD23, BFP14a, BFP14b, Bia14, BM14, Cru10, FTSM22, GDVPP06, Hig10, Mat10, Paa14, San10, VS14, WGP07, WH10a, WH10b]. frameworks [DLZ24, GCY21]. Fredholm [PL03]. free [ACW17, CA13, FB13, GB16b, GLW25, LJW^{+}19, PW20, Pol13a, Pol13b, Spe13b, Van13b, Vel13]. freedom [GHF08]. frequencies [RCSO03]. frequency [Kon13]. frequentist [CM12, GVS98, MM22, MC09]. full [BBS18]. Fully [SJD21]. function [ASLFP13, BGN24, BBS18, BCGC16, BBK97, CQ05, CY15, CM12, EBGGY17, Gam14, GLSU15, GJV20, GM24, GG11, HJG19, JGMRMG18, LPQ11, LdUÁdCIP11, LCB19, MS11, MS15, MT08, May02, Mur16, PRB20, Rod94, RMME19, Sha01, TS05, Vil95, WWY16, WMY20, YLL19, ZZY25]. function-on-scalar [GM24]. Functional [FGV02, FGSV13, GCY21, QDG25b, AMAEV13, AOV99, Agu16, AV16, ARVV18, BCS20, BOQ17, BHGR17, BJ13, BBLC23, BL16a, BM24, CLH^{+}20, CB22, CDM11, CA01, CAFB10, CAFBdlF17, Del24, Dry24, EN25, FBGM13, FB16, FLS25, FPRMA04, FM01, Gan13, Gol25, GWHY14, HHM21, Hor24, HZY22, Hyn25, Lil25, LCB19, LMS^{+}99, LP25, NAV21, NAV25, OPV21, OMRM24a, OMRM24b, PS25, QDG25a, SNC16, SGL14, SLP21, SSG24, TJT16, Van12, WG21, WHS24a, WHS24b, WLVF25, YWZ16a, YWZ16b, ZL16, Zha16]. functionals [FFO20, JSV16]. Functions [DGGJ05, AD99, BPY18, EM21, EM22, GPC10, GM98b, GZCZ19, GRT01, GM03, Kal22, Kok94, MK14, Mal98, MP01, NBGP22, Rao01, RMLDG03, Ryc19, SRDMLF08]. further [GZCZ19]. future [AHA03, VK20]. fuzzy [Krä06].

Galton [MGP00, MMR04]. gamma [HB99, Kou98, GLSU15]. GARCH [CPZ12, CS16b, CZ18, FWZ18, Gam14, HJG19, HG08, LS09]. Gauß [IKvdH23]. Gaussian [Car10, CGR10, FZZ10, GMMC10, GT18, Hec10, Hoo10, Mac18a, MB18, MY10b, SF18, Sen10, Wan10, BDMPF22, Che07, CPW21, Duc01, FPI96, KL14, Mac18b, Mac23, MY10a, Tem00, WLL25]. Gegenbauer [ELORM15]. Gender [AGMT23]. General [NBGP22, BFP14a, BFP14b, BBGMPG19, Bia14, BM14, Bra20, dCCIS21, Cru10, DK23, DFP21, EdOS20, GJL96, GMC93, GHF22, GHF23, Hig10, MCA11, MCA24, MS11, MS15, Mat10, NLP17, Paa14, PP00, Rao01, RS11, San10, SP24, VS14, WYH24, WH10a, WH10b]. generalised [Vil17]. generalization [GD10]. Generalized [BNOR08, Eil20, FBGM13, FBKV14, GS20, GG04, Kne20, Woo20b, dUÁV13, dZBT03, Bic07, BR12, BRV20, BCD^{+}16, Cao07, CM07, CL94, CB07, Efr07, FJ07a, FJ07b, FC21, Gan13, GB16a, GPC10, GMM19, Hal07, HS15, Hor07, KCV23, LW07, LLHL24, LZ21, MK14, Mam07, MVYA19, MPU13, Mül07, NK06, Nav16, Ozk23, PK09, SLL24, VA25, WMY20, Woo20a, XZHW16, YH19, ZLYH16, Ava22, KL14]. generating [CQ05, HJG19, Mur16]. genome [CWZ21]. genome-wide [CWZ21]. genomic [BKK08, GMC08, KC08, Li08a, Li08b, MWN08, Sch08, Str08, TC08a, TC08b]. geodesic [FG20]. geographically [STPC12]. geometric [BSNSS23, Cey14, GD10]. geometrical [BCS20]. ghost [DP23]. Gibbs [RS93]. Gini [GSCB92]. given [AG98, CR94, PN22]. Global [BCD^{+}16, LL23, BW18]. Gompertz [Jah03]. good [RSF97]. Goodness [Cab09, CCR21, CH09a, CH09b, CV16, Det13, Duc01, FWZ18, Gam13, GJV20, GMC13a, GMC13b, JN09, JT22, JGLM20, MHSB20, MT08, Mei13, Mol09, MK09, Spe13a, Tha09, VMS08, Van13a, dUÁ13, dW02, CP98, DRB99, GT11, HX13, MVOC20, MP00, Pew18, SLL24, Ten22, dBCAM^{+}00]. Goodness-of-Fit [Det13, Gam13, GMC13a, GMC13b, Mei13, Spe13a, Van13a, dUÁ13, CH09a, CH09b, CV16, Duc01, FWZ18, GJV20, JT22, JGLM20, MHSB20, MT08, VMS08, dW02, GT11, HX13, MVOC20, MP00, Pew18, SLL24, Ten22, dBCAM^{+}00]. Goodness-of-fit-tests [Cab09, JN09, Mol09, MK09, Tha09]. GQL [RSMJ19]. GQL-based [RSMJ19]. graphic [dUÁV13]. Graphical [XYY22, PKB23]. graphics [CWB^{+}93]. Green [VCS00]. Greenwood [NA07]. gross [FP20, Kon13]. gross-exposure [Kon13]. Group [vEZ94a, vEZ94b, CTC12, DG95, FMR24, Pru20]. Group-Bayes [vEZ94a, vEZ94b]. grouped [AVR00, RFFC17]. groups [GALT23]. groupwise [LZYZ23, WL17]. grows [MS17]. growth [AG15, Men99, ZNAG^{+}01]. growth-curve [Men99]. growth-fragmentation [AG15]. guided [AG16, BS16a, BS16b, BL16b, GW16, HM16, Wag16]. Gutiérrez [GM98a].

Hamiltonian [SJD21]. Handling [Mar15b]. Hardy [CGPV08]. harmonics [MQ01]. Hasofer [NA07]. haystack [Zha22]. Hazard [BB03, CSS18, BPZ20, BR18, CJV05, CdU07, YSV96]. Heavy [MLLC19, BCF25, Vil17]. Heavy-tailed [MLLC19, BCF25, Vil17]. Hellinger [Gre11]. help [LPL15]. Herer [Ter08]. Herriot [BEMP20]. heterogeneity [BCD24, FT10, Mar15b, MLDJ16, Wan19]. heterogeneous [ACW17]. heteroscedastic [HM10, JGLM20, MK25, OR98]. heteroscedasticity [CB22, CBBRMB19, MSK23]. heteroskedastic [MNP24]. hidden [ABA^{+}02, CWZ21, PPB23, PDBNR23, Spe19]. Hierarchical [CS97b, DGSV98, GDVPP06, HB99, ICJ02, LPL15, ZB23]. High [BZ17, Cha17, DBZ17a, DBZ17b, LY17, LS17, LN17, NECA21, WVT23, BR24, BPY18, BHJ24, BKK08, BP08, CG09, CGX23a, CGX23b, CJ23, GB16b, GGS12, GMC08, HMW23, HN18, JvdG17, JS22, KC08, Kon13, LWZZ23, Li08a, Li08b, LHB^{+}17, LGL25, MWN08, MvdG16, NAV21, PP19, Rit23, Sch08, SWZ25, Str08, TF23, TC08a, TC08b, VA25, Xia17, ZZT24, dUÁ23]. High-dimensional [BZ17, Cha17, DBZ17a, DBZ17b, LY17, LS17, LN17, NECA21, BR24, BPY18, BP08, CGX23a, CGX23b, CJ23, GB16b, HMW23, HN18, JvdG17, LWZZ23, LHB^{+}17, LGL25, NAV21, PP19, Rit23, SWZ25, TF23, VA25, Xia17, ZZT24, dUÁ23]. high-frequency [Kon13]. High-order [WVT23, GGS12]. Higher [Are14, LW24]. Higher-order [Are14, LW24]. highly [CBBRMB19]. Hilbert [DG22]. Hill [Alv01, MGN04]. histograms [DL04]. historical [LPL15]. history [PM25]. HIV [MCL16]. Hölder [CP12]. holiday [CLH^{+}20]. Homogeneity [GALT23, DG22]. homogeneous [CPW21]. homoscedasticity [BPY18]. Honest [JvdG17]. Hosmer [SLL24]. Hotelling [DLZ24, FB22]. Huang [Hor24]. Huber [Flo15, GP14]. hunting [CS23]. Hybrid [ALB22a, ALB22b, BP17, BSFB20, RSM06, Mor22, Sal22, Scu22, Spe22]. hyper [Mon11]. hyper-sphere [Mon11]. hyperplanes [GB16b]. hypersphere [dMGP23]. hypotheses [LW22, MP00, Rya12, SP24, WYH24, Yek15]. Hypothesis [HQ23, AT05, BCD23, CR17, CR22, GMC93, GHF08, MM22, Rue92, ZZF19].

i.i.d [HJG19]. identically [Sch98, dBJM09]. Identification [SP15, YLL19, BQ04, PLQ23, WL17]. Identifying [Hay14]. ignorable [Mar15b]. ignorance [SP92]. II [BBC10, BPD16, BP17, CI10, WY09]. iid} [HQ23]. imaging [Zha14]. immigration [GMdP11]. imperfect [FC21, FC22]. Implementation [Bha20]. implementations [BW18]. Improper [HB99]. improved [DS95, Her14, Muñ14, Pin14, TL14, WVT23, ZGGX14a, ZGGX14b]. inactivity [NLP17]. INAR [BSNSS23, BW17, SW16, SWMG18]. Incentive [Cle02]. inclusion [ASS07, Dub99]. incomplete [PMPS11, YRR15]. inconsistency [MBB21]. Increasing [VPP22]. Independence [AF07, GR04, GR18, HHM21, KN13, LKM23, SGR07]. independences [Cas12, LR12, Raj12, Sta12, WS12a, WS12b]. independencies [DFK12]. Independent [Bar09, Bia97, GP08, Gef09, MNP24, Sch98, Vél01, dBJM09]. index [BCF25, BR12, CG02, CJMW24, dCCIS21, CPZ12, CW23, GSCB92, GM24, HL21, JS22, LZYZ23, MF05, PJPW24, SMFP13, TW14, WL17, XZ18, XZHW16, Yua24, ZFX15, ZZF19]. indicators [BW18]. indices [SBC^{+}98]. induced [FC21]. inequalities [BR18, IKvdH23, Rao01]. inequality [GSS11, MCN22]. Inference [AH10, EH11, GNZ11a, GNZ11b, GRT01, HC18, HZY22, JD11, Lam11, Li17, Seg11, Sta14, Tsu11, Val11, WE11, Woo20a, Woo20b, ZZY25, AJS04, AG15, BCF25, Bia97, BS19, Bic07, BZ17, Bra20, CGX23a, CGX23b, Cao07, CM07, CFP^{+}96, Cha17, CJ23, DC11, DBZ17a, DBZ17b, DP18, Efr07, FJ07a, FJ07b, Gho22, GMRV19, Gre11, Hal07, He11, Hor07, HMW23, Ish11, LW07, LAGR20, LCB19, LY17, LZ21, LNZ23, LLLZ24, LS17, LN17, LIAV02, LGW19, Mac23, Mam07, MG21, Mül07, Ozt19, PP00, PM25, Par11, Pea03, PDBNR23, PGB12, RSMJ19, Rit23, Rum03, SL02, TF23, Ton11, WGP07, WS24, WW12, WVT23, YZ19, ZBS11a, ZBS11b, ZLYH16, dUÁ11, dUÁ23, Eil20, GS20, Kne20]. Inferences [CW23, ZJ08, LX16]. inferential [ANZ22]. inferiority [JFCZ14]. infinite [FTSM22]. infinite-dimensional [FTSM22]. inflated [KWWZ23]. inflation [MFBG15]. Influence [DMUOG15, JGMRMPMR05, MRMPEG07, RS11, TLdPDG19, EdOS20, SZL24]. influencing [AFO22a, AFO22b]. Information [WC95, Bou94, Fer04, LD93, MBB21, WL20]. informations [HMV05]. informative [LPL15, MG21, MMS21, TWHZ12, VP17]. INGARCH [LLT18, SZL24]. initial [LC06]. inliers [AT05, AT08]. INMA [AW20]. innovation [ZZY25]. innovations [GR18]. institutional [AFO22a, AFO22b]. instrumental [CBV24, FL24, SP15]. instruments [FL24]. Insua [Bül97]. insurance [BK22, GDVPP06]. integer [CZ18, HSC10, SYL25]. integer-valued [CZ18, HSC10, SYL25]. integrability [AM16, CRV21a, CRV21b]. Integral [CSR08, CIS18, Rao01, RMLDG03]. Integrals [Krä06]. integrated [AABL18, EN25]. integrating [BCS20]. Integration [SLH99, BM17, GS24, GMNH24, PJH^{+}24a, PJH^{+}24b, SZ24, ZS24]. Integrative [SWZ25]. intensity [EM21, EM22, dBJM09]. intensive [DGSV98]. inter [GHF23]. inter-arrival [GHF23]. interaction [FGS21]. interactions [KKLU19a, KKLU19b, Goi19, Rei19, Sch19, SRHD19]. intercept [ABS01, YH19]. interference [Rab14]. Intermittent [MW04]. internally [LJC10]. interpretation [AVBI94, SL17, WC98]. Interquantile [HS22]. Interval [CPZ12, AD99, AHA03, Ber05, LJW^{+}19, RFFC17]. interval-grouped [RFFC17]. intervals [BBC10, GLGLM01, IKvdH23, LPQ11, MO99, MC09, VH07]. intractable [CCR21]. Intrinsic [Ber05, CKM04, LC06, O’H97]. Invariance [Yan95, DNR07, FMR24, MNOP02]. invariant [BDMPF22, DG95, DF19, MS10]. invasive [RCSN22a, RCSN22b]. Inverse [KL14, CM10, Duc01, FL08, Men12]. inverse-linear [Men12]. inverses [HF19]. investments [LDLDMF18]. irregular [MS10]. irregularly [MCL16]. issues [BBLC23, IMP16]. Item [TA06]. items [FC22, HFC18]. iterative [AV00, AVR00, FL08].

Jackknife [CDZ24, PQV12, YZ19, PQ10, SDZ20]. Jeffreys [Sch98]. Joint [MIR03, CMGA^{+}24, VP15]. Jorgensen [CL94]. Journal [Ano25a, Ano25b, Ano25c]. Jump [ZLWH17, SCZ23]. Jump-detection-based [ZLWH17]. junction [MD03].

Kalman [CR97, CGCK11, MGRA98]. Kaplan [LP18]. Kernel [DGGL11, Vil95, BR09, BD00, GVM25, LLTS24, LJC10, MS11, MS15, RFFC17, RMME19, Ten22, WWY^{+}19, YW08]. kernel-based [GVM25]. kernel-type [MS11, MS15]. Kolmogorov [LdUÁPF25, dBIM20]. Kotz [AZ04]. Koziol [VCS00]. Kriged [MGRA98]. Kriging [SCC24]. Kuk [Sah07]. Kullback [CS97a]. Kumaraswamy [PRB20]. Kurtosis [YRR15].

L [FMGE^{+}99]. L-estimation [FMGE^{+}99]. labour [LLVR21]. lagged [De 08]. Lambert [GLSU15]. Lancaster [Kou98]. Laplacian [XYY22]. Large [PP00, CRV21a, CRV21b, CGX23a, CGX23b, CJ23, CRdUÁH19, GHF08, HF19, JGCRJJ22, JGdUÁ24, Lou05, Rit23, Thà23, TF23, VPR15, WWHY18, YZWH17, Zha14, dUÁ23]. Large-sample [PP00]. large-scale [CGX23a, CGX23b, CJ23, Rit23, TF23, Zha14, dUÁ23]. Lasso [XYY22, HMW23]. latency [LCJC17]. Latent [BFP14a, BFP14b, Bia14, BM14, Paa14, VS14, FRV21, Mar15b, RF24]. lattice [VS09]. law [CRV21a, CRV21b, CL23, VPR15, WWHY18]. laws [FHT12, HCS17, MNP24, MPU13, TW14, Thà23]. layout [GLÁM21]. Learning [GYP20, CFS20, Gho20, LDR20a, LDR20b, SCC24, WL20]. Least [Sha01, And97, BBS18, Cíz13, GMdP11, LLZZ14, SS97, WY23]. least-absolute-deviation [WY23]. least-squares [And97]. left [CJV05, LdUÁPF25, sS12, VP17]. left- [VP17]. left-truncated [LdUÁPF25, sS12]. Lego [Goi19, Rei19, Sch19, SRHD19, KKLU19a, KKLU19b]. Leibler [CS97a]. Lemeshow [SLL24]. length [CSS18, XTZ20, dUÁ02]. length-biased [XTZ20, dUÁ02]. Level [CFM23, BLM16, CSL22, DGGL11, ELP20, ELP21, GR07, GNDR09, HMS18, YWLZ15]. levels [AGMT23, Fer99, VPP22]. leverages [ZG07]. life [BPD16, Bha20, BSFB20, Kla05, ZNSW19]. life-testing [BPD16, Bha20, BSFB20]. lifetime [AF24, BPZ20, BR18, HC18, Jah03, SAE12]. lifetimes [ARS22, BNOR08, DJ21, NPM22]. like [GLSU15]. Likelihood [CG19, PGB12, AG19, BS97, Bic07, BP21, CCR21, Cao07, CM07, CV09b, CV09a, CDZ24, CAM15, DJ21, DFPT18, Efr07, FJ07a, FJ07b, Gan13, GLS12, Hal07, Hog09b, Hor07, HMZ09, HC21, Jon96, LW07, LL09, LPQ11, LLZZ14, LZ21, LLLZ24, LGW19, Mam07, MVYA19, MCA11, MCA24, McK09, Mül07, PZ09, PQ10, PQV12, Rit13, RAVL15, RSF97, SS09, SDZ20, SL17, Spe09, Vel09, VK20, YZ19, ZY06]. Likelihood-based [CG19, PGB12, BP21, Hog09b]. likelihood-ratio [ZY06]. likelihoods [CM12]. Limit [BFFS09, Fer99, FHT12, PRSW16, TW14, BPY18, Bar97, MPU13, PR05, Rom94, dUÁ02]. Limiting [MMR04, HSK05]. Lindley [Tsa06]. line [GP14]. linear [And97, AQ01, ARVV18, AVR00, AMO23, BSNSS23, BJ13, BBGMPG11, BGLM19, BR12, BRV20, BCD^{+}16, CB22, CL94, CW23, CB07, DMUOG15, FLS25, FL08, FFZdC15, FL24, GCS95, GB16a, GPC10, GLGLM01, GMC93, HL21, KCV23, LHB^{+}17, LLHL24, LZ21, LZYZ23, LBW01, MWAV19, MVYA19, Men12, MPU03, MG08a, MG08b, MBB21, MvdG16, Mur16, NSS13, NAV25, Pap18, PKB23, sS12, SP15, SLL24, TZ15, TS05, TW96, UMG09, VA25, VPP22, WWY^{+}19, WMY20, WYH24, WW12, WLVF25, XZHW16, YH19, ZG07, ZFX15]. linear-linear [Men12]. linearity [BD00, FLS25]. Linex [Rod94]. Link [YH19, Mar24, PRB20, YLL19]. linking [SKR18]. links [LB18]. LM [Kla05]. load [LPL15]. loads [MCL16]. Local [DGG14, FF02, LdUÁdCIP11, SZL24, SRGS00, BW18, BCD^{+}16, BCS15a, BCS15b, CG15, DFPT18, EM21, EM22, Gel15, Jon96, Lin15, RM15, TZ15, TS15, VFVF00, ZPH18, dRF92]. Locally [WMY20, FMM21]. location [ALYZ15b, ALYZ15a, AG19, Bia97, CS16a, CM10, CHV20, CÖ15, DF19, Far15, Mar15a, RV15, SP92, San97, Van15, Wel15, dW02]. location-scale [Bia97, CHV20]. location-scale-shape [CS16a]. locus [Rab14]. Log [VP17, Pap18, SP92, FWZ18, CS16b]. log-concave [SP92]. log-GARCH [CS16b]. log-linear [Pap18]. Log-symmetric [VP17]. logistic [BBLC23, BBC22b, Pap18, SL17, TLdPDG19]. logit [AMAEV13, HMS18]. lognormal [GRT01, PORCGP00]. long [FMP18]. long-memory [FMP18]. Longitudinal [NARPV99, BFP14a, BFP14b, Bia14, BM14, DW09, FRV21, Hog09a, HL21, IM09a, IM09b, KC09, Lit09, Lop08, LGW19, Mar15b, MLLC19, Paa14, TA06, Uga09, VS14, Wan19, XZHW16]. look [Rad09]. loss [AD99, LC03, Men12, Rao01, Rod94, Sha01, TS05]. losses [RC94]. lossless [GLW25]. low [WVT23, dBJM09]. LRD [OMRM24a, OMRM24b]. Lung [SRDMLF08].

M [Arc02, GM98a, Kal22, STPC12]. M-estimators [Arc02]. M-quantile [STPC12]. M-type [Kal22]. M4 [MF14]. machine [WMS25]. mail [PPB23]. main [CAM15]. majorizing [Rom94]. Making [Bül97, RI92]. malus [GDVPP06]. Manifold [GYP20, Gho20, LDR20a, LDR20b, ZNSW19]. manifolds [OMRM24a, OMRM24b]. Mann [DP18]. MANOVA [GHF08, SP24]. manufactured [HFC18]. mapping [CBBRMB19, Rab14]. maps [Ban18, LM18, Sch18, WBG18a, WBG18b]. Marginal [BM17, Nik24, BCCAVG22, BBGMPG11, BBGMPG19, CS97a, DG22, XZHW16]. margins [Kou98, Pér94]. marked [GCY21]. market [DG22]. Markov [BFP14b, BDMPF22, Bia14, BM14, GT18, Mac18a, MB18, Paa14, SF18, VS14, BFP14a, BC07, BCF25, CWZ21, FRV21, Gho22, Mac18b, Mac23, Mar15b, PPB23, PORCGP00, PDBNR23, Rad04, RF24, Spe19]. Markovianity [Di 12]. marks [CDM11]. martingale [FGFBGM23a, FGFBGM23b]. mass [AACRC19]. matched [JFCZ14]. matched-pair [JFCZ14]. matching [BCS23, GCS95, MC09]. mating [MJR08, MMR08]. matrices [CVO02, CD10, DGGJ05, HF19, LWZZ23, NG93, PKB23, SP24, Spe19, Xia17]. matrix [Ava22, CGB17, CR17, CR22, GW17, HC21, HN18, JvdG17, LNZ23, WT95, WT96]. matrix-valued [LNZ23]. max [AAMDR20, SWMG18]. max-INAR [SWMG18]. max-mixture [AAMDR20]. Maxima [GLSU15, FHT12, Lop08, TW14]. maximal [DG95]. maximize [HQ23]. Maximum [DJ21, RAVL15, SL17, VK20, AZ04, GLS12, HMZ09, HC21, MVYA19, RSF97]. Maxiset [BR09, Che07]. Maxisets [Aut08, KPB^{+}00]. Mean [WT95, WT96, ASLFP13, AF24, AHKS08, CRV12, CRV21a, CRV21b, DKMR11, GR18, GPC10, GG11, GMdP04, HGV13, HN18, IKvdH23, LHB^{+}17, LdUÁdCIP11, LCB19, Oht98, PL03, Rah09, Rod94, SAE12, SB92, SCJS07, Zha14, ZFX15, dlH92, vEZ94a, vEZ94b]. means [FM01, FB22, GM98b, GW99, Jar15, SS92]. measure [Rom94, dW02]. measured [BEMP20]. measurement [ABS01, BKM22, CDZ24, CSS18, GK19, NAV25, Rit13, Sha01, TJT16, WMY20]. measurements [BJ13]. measures [BdCPG14, CFM23, CFMPL24, LLJ21, MCL16, MP93, Ras95, Sta14, ZKR18]. Measuring [RSSSL25]. mechanism [SMT22]. median [CGB17, VP15, Zha14]. mediation [LGL25, WY23]. mediator [LGL25]. Meier [LP18]. membership [Qin25]. memory [FMP18]. method [CVO02, DL04, EPS06, GM24, LLR^{+}24, PQ10, PQV12, Sah07]. methodology [Arn07, Bal07a, Bal07b, BL07, CS07, Dem07, Gui07, Jos07, KC07, Kun07, Nag07, NC07]. Methods [GYP20, Ast14, Ber14, BD00, Cao99, CV09b, CV09a, DW09, Deh14, DFP21, FT10, FL08, Gho20, HHKM12, Hog09a, HR14a, HR14b, HP14, IM09a, IM09b, JFCZ14, KC09, Kir14, Kok14, LL09, LDR20a, LDR20b, Lit09, MP14, McK09, Mor05, PZ09, RAVL15, SS09, SLH99, Spe09, Tra14, Uga09, Vel09]. microarray [FS12, GDS03, YZHG19, ZF18]. mild [FP20]. mineral [GLÁM21]. Minimal [MPS00]. minimax [OR98]. Minimum [GK19, Gre11, LS09, ELORM15, JS01, KB17, Oht98, SYL25, XZ18]. Mises [GP03]. Missing [DW09, Hog09a, IM09a, IM09b, KC09, Lit09, Uga09, BBGMPG11, BBGMPG19, Bra20, CR97, LLLZ24, Nik24, SVY20, SMT22, TZ15, TS05, Wan19, WL20, WLL25]. misspecification [WT95, YH19, WT96]. misspecified [CG19, TW96]. Mixed [Dub99, JL06, PLZ25, Qin25, BLM16, Cab09, CH09a, CH09b, ELP20, ELP21, GMM19, HMS18, JN09, Li17, MWAV19, MCL16, Mol09, MK09, NSS13, Ozk23, PKB23, RMG10, RAP12, SL17, SJD21, SWZ25, TLdPDG19, Tha09, UMG09, Wan19, YH19, ZG07, ZF18]. mixed-effects [MCL16, RAP12]. mixed-type [SWZ25]. mixing [CRdUÁH19, HCS17]. Mixture [Tem00, Wan19, AAMDR20, Ant10, BK22, Cao23, CHV20, CG19, FL10, HW23, Jah03, Lam23, LAGR20, LCJC17, LCPJ23a, LCPJ23b, Lug10, MWY21, MS17, NAV21, QQP21, RS93, SCZ23, SSGM23, SBvdG10a, SBvdG10b, SZ10, WLL25, YZHG19, ZCL16, dB10]. mixtures [ACW17, CS16a, CS97b, FBKV14, GW99, HF18, Nav16, RSM06, Ryc19, Sim22, WL20, WL22, ZCL16]. MM [Wan22]. Modal [AVR00]. Mode [AACRC19]. Model [CA13, FB13, LLHL24, Lie12, Pol13a, Pol13b, Spe13b, Van13b, Vel13, AHMJ92, ÁEdBCAM16, AG15, BPY18, BS97, BCN15, BT11, Bia95, BBGMPG19, BCDG08, BRV20, BdM22, BA24, BCGC16, BEMP20, BKM22, BR18, CSL22, CSR08, CIS18, CPZ12, Che07, CHV20, CGPV08, CD16, CFRG10, CZ18, DKMR11, EM15, Fan01, FP20, FFR16, FGV02, FFZdC15, FJL24, FGFBGM23a, FGFBGM23b, GP08, GK19, GN24, GM24, GMC93, GHF22, GMM19, GG04, GG08, HB99, Hid99, HV14, ID02, Jah03, JL06, JMS21, LD96, LW24, LLTS24, LWML15, LX16, LLJ21, LNZ23, LZYZ23, LLVR21, LC12, Mar15b, MPU03, MG21, MSK23, MMVP21, MvdG16, MLDJ16, NM20, NPM22, NA23, Ozk23, PSK25, PP00, PLQ23, PPB23, PDBNR23, PR98, PRB20, RS93, RAVL15, RSF97, SP92, SWMG18, Sim22, SS97]. model [SZL24, TW96, VH07, VCS00, VSM02, WY09, WSCH15, WMY20, WS24, WLVF25, YZWH17, Ye95, YZ19, Yua05, YZHG19, ZG07]. model-based [BA24, DKMR11, FP20, MG21]. model-fitting [CA13, FB13, Pol13a, Pol13b, Spe13b, Van13b, Vel13]. Model-free [CA13, FB13, Pol13a, Pol13b, Spe13b, Van13b, Vel13]. Modeling [FFO20, LLTS24, Lop08, Ban18, BCCAVG22, BW17, FC22, GSBS04, LM18, PM25, Sch18, SMS23, SMT22, VP15, WBG18a, WBG18b, WG21, ZCL16]. Modelling [MMS21, Spe19, BGLV23, GF06, Pap18, ZNAG^{+}01]. Models [Eil20, GS20, Kne20, Woo20b, AGV17, AHM18, AGV14, AQ01, AV00, AVR00, Ant10, ABS01, ABA^{+}02, BLBB03, BCD23, BFP14a, BFP14b, BB03, BHGR17, BJ13, BK22, BBGMPG11, BS19, Bia14, BD00, BM14, BR12, BM17, BCD^{+}16, BLM16, Bra20, BP21, Cab09, CCR21, CGX23a, CGX23b, CJMW24, CB22, CKM04, Cao23, CL94, CALF15, CS16b, CW23, CS97b, CNJ08, CH09a, CH09b, CJ23, CV16, CG19, CB07, Cru10, CWZ21, Dag01, DG95, DDP06, DMUOG15, DOT19, DV25, DP23, Det13, DGSV98, DS95, Duc05, EG13, EH11, EY00, ELP20, ELLV^{+}23, FL10, FT10, FRV21, FBGM13, FLS25, FSF23, FMM21, FFL13, FPI96, FWZ18, Gam13, Gam14, Gan13, GNZ11a, GNZ11b, GK19, GB16a, GR18, GPC10, Goi19, GJL96, GMC13a, GMC13b, GALT23, GHF23, HW23, HJV18]. models [Hig10, HHKM12, HMV05, HMS18, Hog09b, HL21, HS22, HCS17, HHM24, HG08, HM10, IMP16, IPPC13, ICJ02, JD11, JN09, JGLM20, KCV23, KNV21, KKLU19a, KKLU19b, KL14, Lam11, Lam23, LAGR20, LS09, LLT18, Li17, LLHL24, Lie12, LLG14, LZ21, LZYZ23, LMJ25, LCJC17, LCPJ23a, LCPJ23b, Lug10, LGW19, MK14, MWAV19, MVYA19, MFBG15, MLG16, Mat10, MCL16, MLLC19, MVOC20, Mei13, Men94a, Men99, MO99, Mol09, MK25, MP01, MBB21, MK09, NBGP22, NSS13, Paa14, PC20, PAT04, PORCGP00, PKB23, Pru20, PPT24a, PPT24b, QQP21, Raï10, Ras95, Rei19, Rit23, Rit13, RCN09, RS11, RCN17, RMG10, RAP12, RF24, SCZ23, SGR07, SSGM23, San97, San10, SJD21, Sch19, Seg11, SDM20, Sha01, sS12, SLH99, Spe13a, Spe19, SBvdG10a, SBvdG10b, SP15, SRHD19]. models [SL02, SKR18, SZ10, SYL25, SLL24, TZ15, TLdPDG19, Tha09, TF23, TJT16, Tsu11, UMG09, Val11, VA25, Van13a, VP17, VPP22, VS14, WG07, WE11, WXH^{+}14, WL17, Wan19, WL20, WYH24, WLL25, WW12, WH10a, WH10b, Woo20a, WZ12, WZYW18, XZ18, XZHW16, YWLZ15, YH19, YLL19, ZNSW19, ZZF19, ZLWH17, ZPH18, ZLYH16, ZB23, dS18, dUÁ13, dUÁ23, dB10, dBIM20, dNGL16, ELP21]. models-finite [SLH99]. Moderate [Arc02]. modified [LGW19]. Modular [KKLU19a, KKLU19b, Goi19, Rei19, Sch19, SRHD19]. molecular [PRSW16]. moment [CQ05, GGQ24, HJG19, Li17, Mur16]. moments [AM16, GGS12, dS18]. Monitoring [AHKS08, HHKM12]. monotone [YRR15]. monotonicity [PN22]. monotony [BRV20]. Monte [SJD21, CFP^{+}96]. mortality [BS94]. Mosimann [EG13]. most [AG98, MS10, MCA11, MCA24]. Motivation [IP94]. moving [HSC10, PPT24a, PPT24b, RCN09, RCN17]. MSE [Oht98]. Multi [BSFB20, IV05, BCD24, CM10, SWZ25, YWLZ15, dLNA21]. Multi-armed [IV05]. Multi-criteria-based [BSFB20]. multi-level [YWLZ15]. multi-response [SWZ25]. multi-sample [CM10, dLNA21]. multi-site [BCD24]. Multidimensional [WWT22, ZF18]. multifractal [GL07]. multifractional [OMRM24a, OMRM24b]. multinomial [EG13, JV98, LC03, PP19]. multinormality [HJG19, MS10]. multiparameter [Gho97]. Multipartition [PLQ23]. Multiple [ZZT24, BGN24, CGX23a, CGX23b, CJ23, CWZ21, Di 12, GW17, GDS03, GN99, LC12, MVFFP21, MRMPEG07, PLQ23, PPB23, Pru20, Rit23, RMME19, SG04, TF23, WL17, Wan19, Xia17, Zha14, dUÁ23]. multiple-index [WL17]. Multiplicative [CGCK11, BCGC16, ZZF19]. multipliers [GM24]. multiply [WL22]. Multivariate [CD10, FMGE^{+}99, MCA11, WG21, AH17, ALYZ15b, ALYZ15a, AG19, ARM08, AVBI94, ABA^{+}02, AFO22a, AFO22b, AZ04, BMRSL15, Bic08, Bri08, BA24, CSL22, CVO02, CAM15, CBBRMB19, CFRG10, CÖ15, Daw08, DF19, DL04, DK23, EH20a, EH20b, EH11, ELLV^{+}23, Far15, FRV21, FF13, FJL24, FB22, Fue08, GSBS04, GNZ11a, GNZ11b, GSG^{+}08a, GSG^{+}08b, GT18, HG08, IPT98, JD11, JG20, Jol08, Lam11, Mac18a, Mac18b, Mac23, MQ01, Mar15a, MB18, MF05, Mei20, MNP24, MPU03, NG93, Ric20, RCSO03, RV15, RF24, SF18, SA21, Seg11, SR20, Tsu11, Val11, Van15, WE11, Wan19, Wel15, WJ08, YRR15, YW08, ZKR18, MCA24]. multiway [CAFB10].

Nadaraya [CL10, GS07]. naïve [And97]. Natural [Mal98, GCY21, Pom96]. navigation [Ano25a, Ano25b, Ano25c]. Near [SP92, CAM15, GLS15, MCA11, MCA24]. near-exact [CAM15, MCA11, MCA24]. near-record [GLS15]. nearly [AG98]. needles [Zha22]. negative [CZ18, DDP06, Kou98, MMVP21]. neighborhood [BPD16, BP17]. nested [ELLV^{+}23, LMJ25, MMVP21, RP09, VH07]. nested-error [LMJ25]. network [EM21, EM22, Yua24]. networks [ALB22a, ALB22b, FFF17, Mor22, RSM06, Sal22, Scu22, Spe22]. neutral [RC96]. NN [NAV25]. nodes [FFF17]. noise [Che07, De 08, GJL96, PAT04]. noisy [BJ13]. Non [BSNSS23, And97, AVR00, Bou94, CFMPL24, FS98, Flo15, GHF08, Jar15, JFCZ14, Kal22, Mar15b, MLG16, MMVP21, PW20, RSMJ19, SDZ20, San97, VP17, WWT22]. non-asymptotic [Flo15]. non-conjugate [San97]. non-convexity [CFMPL24]. non-grouped [AVR00]. non-ignorable [Mar15b]. non-inferiority [JFCZ14]. non-informative [VP17]. Non-linear [BSNSS23, And97]. non-nested [MMVP21]. non-Normal [FS98]. non-normality [GHF08]. non-response [Bou94]. non-separable [MLG16]. non-simultaneous [Jar15]. non-smooth [Kal22, SDZ20]. non-stationary [PW20, RSMJ19, WWT22]. noncausal [PLZ25]. noncentral [CMLB05]. Nonequivalence [WY23]. nonexchangeable [IPR11]. nonignorable [SMT22]. nonindependent [Raï10]. Noninformative [DG95, Fan01, GY96, RC96]. nonlinear [BS19, Cru10, HS15, ICJ02, Li17, Osi99, RAP12, SMS23, Wan19, WMY20, YZWH17]. nonnested [CKM04]. Nonparametric [BGN24, Bic07, BC21, BCG09, BCGC16, Cao07, CM07, DC11, Efr07, FJ07a, FJ07b, FSF23, FPI96, FFF17, GPC10, GM98b, GMdP04, Hal07, He11, Hid99, Hor07, Ish11, LW07, LW12, LCJC17, LCPJ23a, Mam07, MVFFP21, Mon11, Mül07, Par11, RPL01, Ton11, ZBS11a, ZBS11b, dUÁ11, BBC10, BPZ20, Bor01, CS16a, dCCIS21, CM10, CHV20, CASS19, CA01, FGV02, FRS24, GR07, GMC93, HPF12, HM10, KCV23, KV25, LJC10, Lou05, NARPV99, SAF21a, SAF21b, VFVFGM07, WGP07, WXH^{+}14, Cao23, HW23, Lam23, LCPJ23b, SSGM23]. nonprobability [FGBB^{+}22]. nonregular [Gho97]. nonresponse [AV00, SCJS07]. Nonstationary [GSBS04, Kar00, NARPV99, Tem00]. norm [BR09, BV23, GZCZ19, IPT98]. Normal [FS98, AN19, BA24, BBK97, CVO02, CS97b, CAM15, GP08, GY96, Gre11, GG04, GG08, Hay14, Lop10, MK14, MFBG15, MBB21, RS93, Rod94, RSF97, SS92, ZCL16, ZY06]. normality [AHKS08, BE20, EH20a, EH20b, GS07, GSS11, GHF08, HX13, JG20, Kra09, LLT18, MQ01, Mei20, Men94b, Ric20, SR20, YRR15]. normalization [Nig06]. normalized [Men12]. note [ASS07, BBS18, GM98a, GDVPP06, KN13, PR98, Tsa06, WC98]. Notes [D’E96, Bor01]. notion [SLP21]. Novel [FGFBGM23b, FGFBGM23a]. NOVELIST [HF19]. NSD [WSCH15]. Nuisance [PW20, BFR00, dlH92]. Nuisance-parameter-free [PW20]. Null [ABS01, FS12, GVS98, LW22, MPS00, MP00, ZY06]. number [CSL22, CRdUÁH19, JGCRJJ22, JGdUÁ24, MD03, MPU13, MJR08, MS17]. numbers [CRV21a, CRV21b, Thà23, VPR15, WWHY18]. numerical [Sah07].

Objective [BT11, CD16, GMRV19, Mor05, MMVP21, PC20, Ber05, Kal22, MG08a, MG08b]. objects [BCS20]. observable [vdL04]. observations [BB03, Bra20, CR97, GS06, GLS12, HMV05, Lop08, Nik24, TS05, VP17, Vil17]. observed [EN25, MCL16, VK20]. obtained [BKM22, GM94]. occurrence [RCSN22a, RCSN22b]. odds [ZJ08]. offspring [GMdP04, MJR08, Rah09]. old [BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. OLS [AM16]. omnibus [dMGP23]. One [CNJ08, Fer04, BMRR21, Bia95, Cíz13, DF19, GALT23, HPF12, KWWZ23, LX16, LMJ25, MK14, Mor05, MM22, Pér94, ZY06]. One- [Fer04]. one-bring-one [LMJ25]. one-sample [DF19]. One-sided [CNJ08, HPF12, Mor05, MM22]. one-step [Cíz13]. one-way [Bia95, GALT23, LX16]. online [CGB17, WMS25]. only [MK14]. operate [CIS18]. operating [IV05]. Operational [Men94a]. operator [BSNSS23]. opinions [DDM^{+}95]. Optimal [ASLFP13, CALF15, CL15, DOT19, FS12, HL21, LSW24, LD93, May02, OR98, Pru20, RH17, SRDMLF08, Tse02, YSV96, YW08, Arc05, Bha20, BSFB20, FGS21, IV05, Ozk23, PR05, Rum03, Sch98]. optimality [JvdG17]. optimization [BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. optimizing [FRV21]. Optimum [WY09, BPD16, BP17, Fan01]. Optional [Sah07]. Oracally [CLH^{+}20]. Oracle [CJMW24]. Oracle-efficient [CJMW24]. Ordaz [PRB20]. order [ÁEdBCAM16, Are14, BMRR14, BC07, CG09, CDM11, CB07, DHJL15, EM21, EM22, GVM25, GGS12, GdW12, Jon04, Kar00, KN13, LW24, PC20, WVT23]. order-constrained [PC20]. ordered [LA05, Ryc19, Sim22]. ordering [AN19, HFC18, LW12]. orderings [BNOR08]. orders [FPRG17, LDLDMF18]. ordinal [CG19]. ordinary [BBC10]. oriented [JMS21]. orthant [FPRG17, WXH^{+}14]. orthant-dependent [WXH^{+}14]. Orthogonal [Pom96, BA24]. other [GLSU15]. outcomes [SJD21]. outlier [BQ04, NECA21]. outliers [BCCAVG22, Cha95, CR97, FP20, HPO04, ZG07, ZY06]. outputs [MPU13]. overdispersion [BW17]. overlapped [AMO23]. overview [BMP^{+}94, Cao99, CA01, LA05, ZNAG^{+}01].

P [AGV14, FPI96, GM98a]. P-splines [AGV14]. pair [JFCZ14]. paired [LLTS24]. pairs [ZJ08]. Panel [Are07, Bal07c, Hsi07a, Hsi07b, Mai07, Ner07, PS07, Shi07, Sic07, WM07, DC11, He11, Ish11, Par11, Ton11, WZ12, ZBS11a, ZBS11b, dUÁ11]. paradox [Tsa06]. Parameter [AW20, dZBT03, DG95, GdW12, HG08, JV98, LLT18, MVYA19, PW20, Rod94, SL17, dlH92, vdL04]. parameterization [Mac23]. parameters [BFR00, GL07, Men94a, MPU03, TA06]. Parametric [Pew18, ZNAG^{+}01, Arc05, AG15, CG02, CG09, CM12, CS16b, GJV20, Gho22, GF06, GRT01, JGLM20, KNV21, Lie12, MWAV19, MLLC19, MP01, NA07, Rue92, SGR07, SM23, VMS08, WGP07, WS24, XZ18]. parametrized [SRDMLF08]. Paretian [MNP24]. Pareto [dZBT03, DGG14, Riv04, Vil17]. Pareto-type [DGG14]. parsimonious [WLL25]. Partial [De 08, AQ01, ARVV18, GM98b, LTWY23, LD93, NAV25, OR98, TO95, ZFX15]. Partially [FFZdC15, BBGMPG11, BR12, CW23, EN25, FLS25, HL21, LLHL24, LZ21, LZYZ23, WMY20, WW12, WLVF25, XZHW16]. particles [SRDMLF08]. partition [CR94]. Partly [FL24, BRV20]. passage [Lef03]. pattern [GLÁM21, dNGL16]. patterns [CDM11]. Peña [GM98a]. penalization [Ant10, FL10, Lug10, SBvdG10a, SBvdG10b, SZ10, dB10]. Penalized [AMAEV13, BBC22b, HJV18, Rit13, Ava22, BGLV23, Gan13, VA25, Wan22]. pereira [MEW01]. pereira-stern [MEW01]. perfect [CF20]. Performance [CS97a, Flo15]. peri [LW22]. peri-null [LW22]. periodic [MU23a, MU23b]. permutation [DFP21, DP18, LKM23, LQR97]. permutation-based [DP18, LQR97]. permutations [HG18]. perspective [LD96]. Peru [LB18]. phase [Van12]. phenomenon [Rab14]. Pick [Mal98]. PID [GSCB92]. piece [BCCAVG22]. pilot [CD10]. pivotal [dLNA21]. plan [WY09]. planar [dLNA21]. Planck [NK06]. plane [PLRC13]. planned [FI03]. plans [BPD16, BSFB20]. Plug [BBGMPG19, BGN24, CD10, CASS19, GF16]. Plug-in [BBGMPG19, BGN24, CD10, CASS19, GF16]. Poincaré [GSS11]. point [Ast14, Ber14, BE20, BCCAVG22, BHJ24, CPW21, CDM11, D’E96, Deh14, DK23, Fer99, GN24, GCY21, GVM25, GVS98, GLÁM21, GLS15, HR14a, HR14b, HP14, Kir14, Kok14, MP14, PLRC13, PLQ23, Tra14, Van95, Vov93, WG07, WLL15, ZZT24]. point-referenced [BCCAVG22]. points [GW17, MHSB20, MMS21]. Poisson [AF07, BT11, BLM16, CTC12, DDP06, FGS21, HB99, Kou98, LLT18, MMVP21, RSSSL25, SW16]. Poisson-gamma [HB99]. Poissonity [JGdUÁ24]. policies [BMRR21, ZZ23]. policy [Ozk23]. Pólya [MPU13]. polynomial [EY00, FF02, LdUÁdCIP11, OR98, VFVF00]. polynomials [DGGJ01, Pom96]. population [EE92, May02, MJR08, MMR08, MG21, Ozt19, SB92, SCJS07, TA06, WWY16]. population-size [MMR08]. populations [AH10, Bor01, CVO02, dCCIS21, HFC18, HN18, JGMRMG18, JGCRJJ22, JGdUÁ24, WTZL17, dAM03]. portfolios [Kon13]. portmanteau [BBC22a]. positive [WVT23]. possible [GS06]. possibly [HM10, IKvdH23]. post [Ozt19]. post-stratified [Ozt19]. Posterior [GM03, dlHRB01, CM12, CS97a, DRB99, GW99, HX13, HSK05, SS92, SK03, WG07]. posteriors [HB99]. Postgrouped [EPS06]. potentials [HZY22]. poverty [HMS18, MCN22]. Power [LC03, PAHW24, De 07, GB16a, GG08, HQ23, Her14, LS09, MK14, MFBG15, MP01, Muñ14, Nig06, PGB12, Pin14, SYL25, TL14, WTZL17, ZGGX14a, ZGGX14b]. power-generalized [MK14]. power-normal [MFBG15]. powerful [MS10]. PPS [May02]. practice [DBC^{+}97]. precedence [NR10]. Precision [Ava22, JvdG17]. Predicting [MD03, Cao99]. Prediction [BGLM19, GM94, AHA03, BBC10, BCS23, BS94, BLM16, Fer04, Jah03, JL06, Mar24, Men99, PAT04, VK20, ZZY25]. predictions [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, WJ08]. Predictive [DS95, CA13, DRB99, DP23, FB13, HX13, Mit92, Pol13a, Pol13b, Spe13b, Van13b, Vel13, WG07, dlHRB01]. predictivistic [AVBI94]. predictors [BCS15a, BCS15b, CG15, Gel15, Lin15, NAV21, RM15, TS15]. pregnancy [CSS18]. premium [GGQ24]. premiums [GDVPP06]. preposterior [vEZ94a, vEZ94b]. Prequential [Vov93]. presence [ALYZ15b, ALYZ15a, BT11, BCG09, CÖ15, Far15, GS06, ID02, Mar15a, Nik24, RV15, Van15, VP17, Wel15]. Preservation [ARS22]. preserving [WS24]. presmoothed [CdU07]. prevalence [BKM22, MBB21]. principal [AD23, AOV99, CGB17, GF16, LMS^{+}99]. Prior [LJW^{+}19, MP93, RS93, BLBB03, CSR08, EG13, EY00, Fer04, LPL15, MC09, Sch98, Yan95, dlH92]. Prior-free [LJW^{+}19]. priors [BGLV23, Bia97, CKM04, CIS18, CR94, CGPV08, DG95, FS98, GR07, Gho97, GCS95, GY96, GVGP08, HB99, MS17, Pap18, PAHW24, PR98, Riv04, SP92, San97, Sch98, Ye95]. Privacy [WS24]. Privacy-preserving [WS24]. pro [LL23]. probabilistic [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, LJW^{+}19, WJ08]. probabilities [May02, Rum03, WMC^{+}96]. Probability [Agu16, AV16, BL16a, FB16, GCS95, SNC16, SAF21b, YWZ16a, YWZ16b, ZL16, Zha16, ASS07, BBGMPG11, Kou98, Men12, MP93, MC09, Sah07, VSM02, SAF21a]. Probability-enhanced [Agu16, AV16, BL16a, FB16, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. probit [AV00]. probit-type [AV00]. problem [BOQ17, BDMPF22, BNY21, CM10, CA01, DRB99, DF19, DT19a, GY96, GVS98, LdUÁPF25, NG93, PR05, Rum03, WLL15, BR19, DT19b, Kot19, Mor19, WZ19, del19a]. Problems [XYY22, Bou94, FL08, FS98, Lef03, SS92, dLNA21]. procedure [AV00, BQ04, D’E96, DK23, GM95, Riv04, RGEGMI13]. procedures [Bel14, FI03, HL14, KL14, MNP24, Mor14, MG08a, MG08b, PST14a, PST14b, SU14, VD96]. Process [Ban18, LM18, Sch18, WBG18a, WBG18b, AABL18, And97, AF07, BT11, BW17, CP98, GL07, Gam14, GSBS04, GR04, GJL96, GMdP04, GMdP05, GLS12, GLS15, GRT01, MGP00, MMS21, PLZ25, Rah09, RSSSL25, SBC^{+}98]. processes [AAMDR20, AW20, ASLFP13, ACZ06, AF07, BOQ17, BSNSS23, BNOR08, CTC12, Car10, CGR10, DHJL15, FZZ10, Fer99, FF13, FC21, FTSM22, GCY21, GVM25, GMMC10, GMdP11, Hec10, Hoo10, IV05, Kar00, Lef03, MF14, MMR04, MJR08, MMR08, MY10a, Rom94, Rya12, SW16, SWMG18, Sen10, TW14, VPR15, Vov93, Wan10, dBCAM^{+}00, MY10b]. Product [CdU07, dUÁ02, Goi19, KKLU19a, KKLU19b, Rei19, Sch19, SB92, SCJS07, SRHD19]. Product-limit [dUÁ02]. Product-type [CdU07]. products [De 08]. Progressive [Arn07, Bal07a, Bal07b, BL07, CS07, CM10, Dem07, Gui07, Jos07, KC07, Kun07, Nag07, NC07, BPD16, CI10, WY09]. progressively [BBC10, BP17, HC18, PK09]. projected [FRV21, XZ18]. projected-distance [XZ18]. projection [FGSV13, Lop24]. projections [NECA21]. proof [AF07]. proper [Cle02, GMRV19, HB99, Pér94]. Properties [O’H97, AQ01, ARS22, BC07, EN25, FPRG17, LdUÁdCIP12, MF14, NRS06, PR98, SLH99, SS97]. property [BE20, RSF97]. proportion [LJW^{+}19, PDBNR23]. proportional [BR18, NG93]. proportions [ELP20, ELP21, HMS18]. proposal [BNY21]. protected [SVY20]. Pseudo [HMZ09]. Publisher [PPT24a]. pursuit [FGSV13, Lop24].

QANOVA [DFP21]. Quadratic [BOQ17, Fan01, FJL24, MQ01, NBGP22, TS05]. quality [AFO22a, AFO22b, IP94, Pau11]. quantifying [dCCIS21]. quantile [AGMT23, AGV14, Are14, CG09, CMGA^{+}24, CL10, CASS19, DFP21, GF06, HS15, HMW23, HGV13, JS22, Kra09, LSW24, LGW19, Ots09, PJPW24, STPC12, Sta14, TWHZ12, TV23, WWY^{+}19, WLVF25, XTZ20, YLL19, dBCAM^{+}00]. quantile-based [DFP21, HGV13]. quantiles [BR18, CM12, GS15a, GS15b, HS15, SVY20, dAM03]. quantitative [Rab14]. quantities [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, WJ08]. quantized [HMV05]. quasi [HC21, MPS00, SG04]. quasi-maximum [HC21]. quasi-stationary [MPS00].

radioactive [SRDMLF08]. Randić [Yua24]. Random [BDMPF22, VPR15, AG16, AF24, BFFS09, BHGR17, BCCAVG22, BS16a, BS16b, BL16b, CRV12, CRV21a, CRV21b, Cey14, CPW21, DGGJ05, ELORM15, FPRMA04, Fer99, FHT12, GS15a, GS15b, GLSU15, GW16, GLGLM01, GGQ24, GT18, GNDR09, GM94, HG18, HLM22, HM16, HCS17, IPR11, Krä06, LX16, LC06, LC03, Mac18a, Mac18b, Mac23, MB18, MPU13, NECA21, NPM22, Nig06, Pru20, RMLDG03, SF18, SCJS07, TW14, Ter08, VH07, Vél01, VS09, Vil95, Wag16, WXH^{+}14, WSCH15, WZ12, WWHY18, Ye95, YH19, dBJM09, vdL04]. randomisation [LKM23]. randomized [BLBB03, Sah07, TO95]. randomly [CRV12, CM22, WLVF25]. randomness [GR04]. ranges [SG04]. rank [AFO22a, AFO22b, LLR^{+}24, Mur16, Ozt19, Sch96, She13]. rank-based [Ozt19, She13]. ranked [ASS07, HMZ09]. ranking [KLYZ17, VD96]. Rao [BKM22, Rao01]. rate [CJV05, CdU07, FvdW08, Guo08, Mar24, OPV21, PM25, RSW08a, RSW08b, SH08, Tro08, Xia17, Yek08, ZF18]. rates [GLS12, PORCGP00, WVT23, dS18]. ratio [Bic07, Cao07, CM07, CAM15, Efr07, FJ07a, FJ07b, Hal07, Hor07, JFCZ14, LW07, Mam07, MCA11, MCA24, Mül07, Pau11, Rit13, SB92, SCJS07, Yan99, ZY06, ZJ08]. ratio-cum-product [SB92]. ratios [GP08, PJPW24]. recapture [FT10, LLLZ24]. Reconciling [GVS98]. reconstruction [FPRMA04]. record [GLS12, GLS15, LBW01, LBSM13, LBSM15, VK20]. records [AHA03]. Recovering [PLRC13]. recovery [Xia17]. recurrence [MPS00]. recurrent [BT11]. Recursive [VFVF00, D’E96, GM95]. Reduced [BPZ20, CG09, DGGG08, Sch96]. reduced-bias [CG09]. Reducing [FC22, HGV13]. reduction [Agu16, AV16, BNY21, BL16a, FRV21, FB16, FGV02, FRS24, GF06, LP18, LTWY23, LZYZ23, LLR^{+}24, SNC16, YWZ16a, YWZ16b, ZL16, Zha16]. reduction-based [LZYZ23]. refereeing [Ano07, Ano08, Ano09, Ano10, Ano11]. Reference [FT10, FS98, Gho97, Rab98, BT11, EY00, PR98, Yan95, Ye95]. referenced [BCCAVG22]. regeneration [BC07]. regeneration-based [BC07]. regional [BKM22]. regions [Ber05, JvdG17]. Regression [BHGR17, GLGLM01, XYY22, AMAEV13, AGV14, ARVV18, Ant10, ABS01, ACR17, BPY18, BK22, BBLC23, BV23, BBGMPG19, BS19, BBC22b, BD00, BGLM19, BRV20, BdM22, BP21, CGX23a, CGX23b, CV09b, CV09a, CJ23, CASS19, CWB^{+}93, Det13, DS95, Dub99, EdOS20, ELLV^{+}23, FL10, FZWZ15, FLS25, FF02, FSF23, FGSV13, FFZdC15, FMM21, Flo15, FRS24, FGS21, Gam13, GN24, Goi19, GMC93, GMC13a, GMC13b, GWHY14, GLW25, HPF12, HS15, HMW23, HCS17, HM10, JSV16, JS22, JGMRMPMR05, KV25, KKLU19a, KKLU19b, LB18, LL09, LLZZ14, LSW24, Lie12, LL23, LJC10, LMJ25, LBW01, Lug10, MWY21, MHSB20, MFBG15, MLLC19, McK09, MVOC20, MVFFP21, Mei13, Mon11, MG08a, MG08b, NAV21, NAV25, Oht98, OR98, Osi99, Ots09, Pap18, PZ09, Pru20, PRB20, Rei19, Rit23, RS11]. regression [RMME19, SCZ23, STPC12, SS09, SL17, Sch96, Sch19, Spe09, Spe13a, SBvdG10a, SBvdG10b, SRHD19, SZ10, TWHZ12, TLdPDG19, TV23, TF23, TS05, TW96, VMS08, Van13a, VP17, Vel09, VFVF00, VFVFGM07, Vil95, WXH^{+}14, WSCH15, WLL15, WC98, WLVF25, Yan00, YZWH17, ZCL16, ZFX15, ZPH18, dS18, dUÁ13, dUÁ23, dB10]. regressions [BCN15, Cas12, DFK12, FL24, LR12, Raj12, Sta12, WS12a, WS12b]. regressor [BJ13]. regressors [MS17, RMME19]. regular [BCF25]. Regularity [CPW21]. Regularization [BLT^{+}06, BV23]. reinsurance [GGQ24]. Rejoinder [ALYZ15a, ALB22b, Bal07b, BFP14b, BS16b, BCS15b, CGX23a, CV09a, CH09b, DT19b, DBZ17b, DPR11a, EH20a, EPG19b, FJ07b, GP19b, GNZ11b, GSG^{+}08b, GMC13a, HR14b, Hsi07b, IM09b, KKLU19b, LDR20b, LCPJ23b, Mac18a, MY10b, PGP21b, PST14a, Pol13b, PJH^{+}24b, QDG25b, RSW08b, SBvdG10b, Tjø12a, TC08b, WBG18b, WS12a, WH10b, Woo20b, WHS24a, YWZ16b, ZBS11b, ZGGX14a]. related [ABA^{+}02, CM12, Duc01, HZY22]. relation [CBB^{+}95, GLGLM01]. relations [Car10, CGR10, FZZ10, GMMC10, Hec10, Hoo10, Jon96, MY10a, MY10b, Sen10, Wan10]. Relative [CJV05, BMRSL15, Cey14, CL15, LLZZ14, WLL15]. relaxing [HG08]. relevation [BMRR21, ZZ23]. Reliability [MS03, NRS06, Ozk23, BMRR21]. remaining [AF24]. Renewal [Rad04, BNOR08, FC21]. Rényi [PRSW16]. repair [CF20, FC21, FC22]. reparametrization [Yan95]. repeated [AJS04, MCL16, Ras95, Rit13, ZKR18]. replacement [BMRR21, Ozk23]. replicability [VPP22]. replication [BCD24, PAHW24]. reporting [ABA^{+}02]. representation [PKB23]. repulsive [QQP21]. Resampling [GDS03, Pau11, SS92]. Resampling-based [GDS03]. research [GPR00]. RESET [Ots09]. Residual [EdOS20, Stu01, Duc05, RPL01, SAE12]. residuals [BPY18, De 08, Gam14, ZG07]. Resistant [KV25]. respect [OR98]. response [ANZ22, BLBB03, Bou94, DOT19, Fan01, GM94, LC06, MVFFP21, Sah07, SWZ25, SMT22, TA06, TO95]. response-adaptive [ANZ22]. responses [BBGMPG11, BBGMPG19, GLÁM21, Qin25, Sim22, WLVF25, YH19]. restricted [DS95, LWML15, Oht98, RSF97, RMG10, WW12]. restrictions [HG08, TS05, UMG09]. result [MS11, MS15, Men94b]. results [BDMPF22, CM12, FRS24, GR07, NLP17, Tem00, YZWH17]. retention [SRDMLF08]. Return [GNDR09]. returns [DG22]. reversibility [Di 12]. reversible [SCZ23]. Review [GPSB^{+}97, BFP14a, BFP14b, Bia14, BM14, CV09b, CV09a, DW09, Det13, EH20a, EH20b, Gam13, GMC13a, GMC13b, Hog09a, IM09a, IM09b, JG20, KGP^{+}93, KC09, LL09, Lit09, McK09, Mei13, Mei20, MCN22, Paa14, Pea03, PZ09, Ric20, SS09, Spe09, Spe13a, SR20, Uga09, Van13a, Vel09, VS14, dUÁ13]. Revisiting [BH16, BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. Reweighted [Cíz13]. right [BBC10, CJV05, DP18, GGQ24, sS12, VP17, XTZ20]. right-censored [DP18, sS12, VP17, XTZ20]. Ríos [Bül97]. risk [GF06, Kon13, SAF21a, SAF21b]. risks [Gef09, SGR07, WG07]. rival [DOT19]. road [MD03]. Robust [AH17, ALYZ15b, ALYZ15a, AT05, ACR17, BS97, BV23, Bia95, BS19, BR12, BRV20, Bra20, Bül97, CR94, CÖ15, Far15, FP20, FZWZ15, GB16a, Gho22, KCV23, LCB19, LMS^{+}99, Mar15a, MM02, MPU03, Ras95, RV15, SA21, SVY20, VA25, Van15, WL17, WL22, Wel15, ZCL16, AD23, AE06, BMP^{+}94, BNY21, BBGMPG11, BBC22b, CGB17, CFRG10, DGG14, EBGGY17, Flo15, GF16, LWML15, MVYA19, MLLC19, RCT14, RI92, RAVL15, WC98]. Robustness [IMP16, GP94, HV14, JT22, JS01, PPST96, Rod94]. Rosenblatt [LLG14]. route [LMJ25]. rows [LNZ23]. rule [Oht98]. rules [Aut08, Cle02, GMRV19, WMC^{+}96]. runs [BH16].

S [RCT14]. S-estimators [RCT14]. Sample [De 07, AGMT23, BBC10, Bar97, BOQ17, BH16, BCD24, Cha95, CM10, DF19, EPG19a, EPG19b, EMJ20, FS98, Fer04, FH19, FHT12, GY96, GB16b, Gre19, LdUÁPF25, LPL15, LC03, MF19, MS17, PP00, PP19, SG04, SM23, SLH99, WTZL17, ZY06, dLNA21]. sampled [Car10, CGR10, FZZ10, GMMC10, HQ23, Hec10, Hoo10, MY10a, MY10b, Sen10, Wan10]. samples [AH10, BP08, FG20, Ozt19, XTZ20]. Sampling [SS92, ASS07, CM10, EPS06, HQ23, HMZ09, LD93, LC03, May02, Men12, MG21, RS93, Sah07, SKR18, Vil95, WZYW18]. Sampling-resampling [SS92]. sandwich [HC21]. sandwich-form [HC21]. Särndal’s [BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. Saunders [TPB20]. scalar [GM24, Gre11]. scale [Bia97, CGX23a, CGX23b, CS16a, CHV20, CS97b, CJ23, Csö02, LC12, Rit23, TF23, ZCL16, Zha14, dUÁ23, dW02]. scatter [ALYZ15b, ALYZ15a, AG19, CÖ15, Far15, Mar15a, RV15, Van15, Wel15]. scheme [BPD16, BSFB20]. schemes [BP17, LKM23]. science [Büh19, Cao19, Cru10, Del19b, GP19a, GP19b, GS19, Hig10, Mar19, Mat10, NS19, RACC19, San10, SH19, Tsa19, VZ19, WH10a, WH10b]. science-based [Cru10, Hig10, Mat10, San10, WH10a, WH10b]. Score [SGR07, LSW24]. Scoring [WMC^{+}96, Cle02, GMRV19]. screening [KLYZ17]. search [ACR17, BPD16, BP17, WZR94]. Second [BC07, EM21, CG09, CDM11, LBW01, EM22]. Second-order [BC07, EM21, CG09, EM22]. sectional [Bel14, HL14, Mor14, PST14a, PST14b, SU14]. SEIHR [FFR16]. selecting [BCS15a, BCS15b, CG15, Gel15, Lin15, RM15, TS15]. Selection [LLVR21, Ye95, BGN24, BCS23, BCN15, BCDG08, CSL22, CSR08, CIS18, CD10, DBC^{+}97, DL04, GM24, GLW25, Hid99, IV05, Kar00, KN13, LD96, LLJ21, May02, MG08a, MG08b, MS17, RFFC17, VD96, WZYW18, YW08, ZLYH16]. Selective [PDBNR23, ABA^{+}02]. selector [CASS19]. self [SYL25]. self-exciting [SYL25]. Semi [AG15, CG09, CS16b, GF06, NA07, NAV25, ARVV18, CG02, CFS20, FLS25, KNV21]. Semi-functional [NAV25, ARVV18, FLS25]. Semi-parametric [AG15, CG09, CS16b, GF06, NA07, CG02, KNV21]. semi-supervised [CFS20]. semicontinuous [RMLDG03]. Semiparametric [IPPC13, MWY21, Yua05, AHM18, ALB22a, ALB22b, BB03, CV16, FSF23, HF18, ID02, LLLZ24, Mor22, NAV21, PSK25, Sal22, Scu22, Spe22, SMT22, VP15, WZYW18, ZPH18]. semivariogram [GS07]. sense [CRV21b, CRV21a]. Sensitivity [PMPS11, Hog09b, MIR03, RSSSL25]. separability [GVM25]. separable [MLG16]. separation [FRV21, Spe19]. Sequence [Sta12, FHT12, HL09]. Sequences [Cas12, DFK12, LR12, Raj12, WS12a, WS12b, BFFS09, Di 12, HSC10, Tem00]. sequential [SDM20]. sequentially [FI03]. Serial [GR18, HHM21]. series [Ane12, BBC22a, Bel14, Ber11, BdCPG14, BW17, Bra11, Cao99, Cha95, DK23, Dol12, DPR11a, DPR11b, Dou12, FGV02, FF21, Fok12, Gal12, Gao12, GM95, GN99, HL14, Hei12, Hid99, HHM24, KWWZ23, KPJ22, Ked12, LM11, MU23a, MU23b, MM02, Mor14, MW04, Nik24, OMRM24a, OMRM24b, Pap11, PW20, PST14a, PST14b, PL03, RSMJ19, RPL01, SMS23, SU14, Stu01, Tjø12a, Tjø12b, Vel11, Vél01, WWT22, ZZY25]. Seshadri [Kok94]. set [ASS07, CL94, HMZ09, RC94]. sets [BQ04, CR94, CFM23, GR07, Krä06, Ter08, Tse02]. setting [CAM15, CAFBdlF17, JFCZ14]. setup [HS15]. Several [KGP^{+}93, BLBB03, GZCZ19]. Severe [CL23]. Shannon [Sch98]. Shape [AGV17, WHS24a, WHS24b, CS16a, PLRC13, dLNA21, BM24, Del24, Dry24, Hor24, SSG24]. Shape-based [WHS24a, WHS24b, BM24, Del24, Dry24, Hor24, SSG24]. shared [CNJ08]. Sharp [BR18, Flo15, IKvdH23, Ryc19]. shift [PL03]. shifts [AHKS08]. shock [GHF22, GHF23, Ozk23]. shocks [MS03]. Short [AT08, AT05, FMP18]. short- [FMP18]. Short-tailed [AT08]. short-tailedness [AT05]. Shrinkage [XYY22, HL21, HS22]. sided [CNJ08, HPF12, LC06, Mor05, MM22]. sign [RMLDG03]. significance [FS12, LNZ23, VPP22]. Simes [FB22]. Simple [Bül97, LLJ21, San97, BdM22, CIS18, CAFB10, PSK25, RI92, SCC24]. simplex [EdOS20, ID02]. simplicity [KL14]. simplified [BBGMPG11]. simulation [DGSV98, Sah07]. simulation-intensive [DGSV98]. Simultaneous [WWY16, BZ17, CY15, CJMW24, CALF15, Cha17, DBZ17a, DBZ17b, GWHY14, HN18, Jar15, LCB19, LY17, LS17, LN17, ZY18, ZLYH16]. Single [Bel14, HL14, JS22, Mor14, PST14a, PST14b, SU14, BR12, CJMW24, CW23, Fan01, GM24, HL21, LZYZ23, XZ18, XZHW16, ZFX15, ZZF19]. Single- [Bel14, HL14, Mor14, PST14a, PST14b]. Single-index [JS22, BR12, CJMW24, CW23, GM24, HL21, LZYZ23, XZ18, XZHW16, ZFX15, ZZF19]. singular [DGGJ05]. sinh [Pew18]. sinh-arcsinh [Pew18]. site [BCD24]. size [BCD24, De 07, FHT12, JMS21, LC03, MMR08, MS17]. Skew [RJP11, AN19, BA24, Gre11, GG04, LWML15, Lop10, RP09, ZCL16]. skew- [BA24, LWML15]. skew-normal [AN19, BA24, Gre11, Lop10, ZCL16]. skew-symmetric [RP09]. skew-t [Gre11]. Skewed [ABA^{+}02, GMM19, GG08, NA23, WTZL17]. skewness [BMRSL15, Sta14, VP15]. slope [Ban18, LM18, Sch18, WBG18a, WBG18b]. Small [ELP20, ELLV^{+}23, HMS18, RMG10, STPC12, SKR18, Bel14, BP08, Dag01, DGSM11, DKMR11, GS13, GMM19, HL14, JL06, LPL15, LMJ25, MCN22, Mor14, PST14a, PST14b, SU14, TJT16, UMG09, ELP21]. Smirnov [LdUÁPF25, dBIM20]. Smirnov-based [dBIM20]. Smirnov-type [LdUÁPF25]. Smith [GM98a]. Smooth [JMS21, BBBV02, CY15, CJMW24, Kal22, MP00, SDZ20, YSV96, ZY18]. Smoothed [LGW19, PQ10]. smoothers [FF02]. smoothing [DBC^{+}97, FGBB^{+}22, FL24, HGV13, JV98, Kal22, WMS25]. smoothness [FF21]. softplus [SZL24]. solutions [CSR08]. solving [Bou94]. Some [Ane12, BDMPF22, Bor01, CRV12, Dol12, Dou12, Fok12, Gal12, Gao12, GT18, Hei12, Ked12, Mac18a, Mac18b, MB18, MP01, SF18, SG04, SM23, SK03, Tjø12a, Tjø12b, ASS07, ASLFP13, Ast14, Ber14, BBLC23, Deh14, Dub99, FF12, GRT01, HF18, HR14a, HR14b, HP14, Kir14, Kok14, MS10, MP14, NBGP22, TS05, Tra14]. sources [AVR00]. Space [Her14, Muñ14, Pin14, TL14, ZGGX14a, ZGGX14b, EPG19a, EPG19b, FMP18, FH19, Gre19, MF19]. spaced [AOV99]. spaces [CFM23, CPW21, DG22, RMME19]. spacings [EMJ20, SM23]. Sparse [AMO23, HMW23, NAV21, XYY22, Agu16, AV16, ACW17, BBC22b, BL16a, FB16, GF16, GWHY14, PP19, SNC16, SLP21, YWZ16a, YWZ16b, ZL16, Zha16]. sparsely [Car10, CGR10, FZZ10, GMMC10, Hec10, Hoo10, MY10a, MY10b, Sen10, Wan10]. Spatial [FTSM22, ICJ02, RCSN22a, RCSN22b, SGL14, AAMDR20, BCCAVG22, BW18, BCS15a, BCS15b, CG15, Cey14, CDM11, DMUOG15, FFL13, GS96, GR07, Gel15, GLÁM21, HS22, LW24, Lin15, MPU03, RM15, TS15, WS24]. Spatial-temporal [ICJ02, FFL13]. spatially [GSBS04, MVOC20]. Spatio [FMP18, MLG16, Cru10, GCY21, GVM25, Hig10, Mat10, San10, WH10a, WH10b]. Spatio-temporal [FMP18, MLG16, Cru10, GCY21, GVM25, Hig10, Mat10, San10, WH10a, WH10b]. species [RCSN22a, RCSN22b]. Specification [KNV21, LZYZ23, MNP24, FGFBGM23a, FGFBGM23b, LW24, PKB23, WWT22]. Specifications [HHM24, JGLM20]. spectral [OMRM24a, OMRM24b, VS09]. speed [Her14, Muñ14, Pin14, TL14, ZGGX14a, ZGGX14b]. sphere [Mon11]. Spherical [MQ01, EG12, IPT98]. spherically [WYH24]. spline [AMAEV13, HJV18]. Splines [Van95, AGV14, Kal22]. square [BBS18, CMLB05, GMdP11, WT95, WT96]. squared [AQGSM05, DKMR11, HGV13, Oht98]. squares [And97, Cíz13, GP08, Sha01, SS97]. Srivastava [Hor24]. stability [PJPW24]. stabilized [ANZ22]. stable [EBGGY17, MK14, MT08, MNP24, PJPW24]. stable-Paretian [MNP24]. stage [Bel14, GS13, HL14, Mor14, NARPV99, PST14a, PST14b, SU14]. standardization [BBS18]. state [FMP18]. state-space [FMP18]. stationarity [PL03]. stationary [BdCPG14, DHJL15, FMM21, FHT12, MPS00, MW04, PW20, RSMJ19, TW14, WWT22]. statistic [FS12, Yan00]. Statistical [BCD23, CGX23b, CFP^{+}96, CFMPL24, CJ23, EN25, GPR00, LNZ23, Ozt19, Rit23, Rum03, TF23, WW12, ZLYH16, dUÁ23, AABL18, AF07, BCS23, BCF25, CRV21a, CRV21b, CTC12, FL08, GSS11, ICJ02, SA21, Sch98, SCC24, WZR94, ZNSW19, CGX23a]. Statistics [Pea03, Are14, BLT^{+}06, BHJ24, Cao19, CAM15, CBB^{+}95, DHJL15, EH20b, FFF17, GP19a, GS96, GdW12, Jon04, MCA11, MCA24, MP00, MC09, NA07, PGP21a, SR20, dLNA21, Büh19, CSN21, Del19b, EH20a, GP19b, GS19, Huc21, JG20, Mar21, Mar19, Mei20, NS19, PGP21b, RACC19, Ric20, Sce21, SH19, Tsa19, VZ19]. Stein [BE20, Oht98]. Stein-rule [Oht98]. step [Cíz13, LLLZ24, PSK25, WY09]. step-stress [PSK25, WY09]. stern [MEW01]. Stochastic [HFC18, LDLDMF18, Nav16, PN22, ZZ23, ÁEdBCAM16, ASLFP13, AN19, FPRG17, FFR16, HF18, LW12, LLG14, MS03, NR10, SBC^{+}98, Thà23]. stock [DG22]. stopping [AHKS08, PR05, Rum03]. strategy [Bha20, EE92, Spe19, TO95]. stratified [LD93, Ozt19, Sah07]. stress [PSK25, WY09]. strictly [Cle02]. Strong [AABL18, Gef09, HCS17, VPR15, WWHY18]. Strongly [BCDG08, Tem00]. structural [AE06]. structure [BLBB03, CWB^{+}93, EPG19a, FF02, GCY21, MF05, Mar15b, MLG16, SDZ20, ZKR18, ZNSW19, EPG19b, FH19, Gre19, MF19]. structured [Goi19, KKLU19a, KKLU19b, Rei19, Sch19, SRHD19]. structures [FJL24, GM98b, NBGP22]. studies [BCD24, CWZ21, DW09, Hog09a, IM09a, IM09b, KC09, Lit09, PAHW24, RH17, Uga09, ZF18]. study [AHMJ92, CS97a, GZCZ19, Kra09, Sah07]. subgraphs [BDMPF22]. subgroup [SWZ25]. subject [CALF15, MS03, TS05]. subsample [Rad98]. Subsampling [Ber11, Bra11, DPR11a, DPR11b, LM11, Pap11, Vel11, FvdW08, Guo08, LSW24, RSW08a, RSW08b, SH08, Tro08, Yek08]. subset [TWHZ12]. subset-based [TWHZ12]. subspace [FJL24]. subspaces [GS24, GMNH24, PJH^{+}24a, PJH^{+}24b, SZ24, ZS24]. successful [BR19, DT19a, DT19b, Kot19, Mor19, WZ19, del19a]. Sufficiency [FI03]. sufficient [BMRR14, BNY21, FRS24, LLR^{+}24]. sum [CMLB05, PN22]. sums [BPY18, CRV12, CRV21a, CRV21b, De 08, GP08, HCS17, NPM22, WSCH15]. sup [BR09]. sup-norm [BR09]. superadditive [MMR04]. superfluous [Dub99]. superiority [WT95, WT96]. superpopulation [WWY16]. Supervised [BCS20, CFS20]. support [WMS25, Xia17]. Sure [KLYZ17, CRV12]. surface [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, GM94, Jol08, WJ08]. survey [AGMT23, BKM22, LA05, LLVR21]. surveys [AV00, FGBB^{+}22]. survival [BGN24, BCGC16, EM15, GJV20, GM98b, LC12, PORCGP00, SKS13]. Sylvester [Ava22]. symmetric [AZ04, CR17, CR22, IPPC13, MT08, PJPW24, RP09, RCSO03, VP17, WYH24]. symmetrized [LJC10]. symmetry [EG12, HLM22]. synchronous [FGFBGM23a, FGFBGM23b]. synergetic [FGS21]. system [ARS22, BR18, DJ21, Goi19, Her14, HC18, KKLU19a, KKLU19b, Muñ14, Pin14, Rei19, Sch19, SRHD19, TL14, ZGGX14a, ZGGX14b]. systems [MS03, NRS06, NR10, Nav16, NLP17, NM20, SAE12, ZZ23].

t [Gre11]. tables [AQGSM05, GVGP08, LC03, PC20, Pér94]. tabulated [Yek15]. Tail [BCF25, FF21, CG02, CPZ12, DGGG08, EBGGY17, GG21, GGQ24, PQ10]. tail-coefficient [DGGG08]. tailed [AT08, BCF25, MLLC19, Vil17]. tailedness [AT05]. tails [DGG14, GG11]. TCLUST [RGEGMI13]. techniques [SS92]. Temporal [GN99, Cru10, FMP18, FFL13, GCY21, GVM25, Hig10, HMS18, ICJ02, MLG16, Mat10, San10, WH10a, WH10b]. temporally [CPW21, HPO04]. Tensor [Lop24, GN24, Goi19, KKLU19a, KKLU19b, Rei19, Sch19, SRHD19]. terms [Bar97]. Test [GR04, LHB^{+}17, AQGSM05, AE06, BPY18, BH16, BMRR14, CB22, CAM15, De 08, DLZ24, GP03, GT11, GJV20, GVM25, GSS11, GJL96, GZCZ19, JT22, Kra09, LdUÁPF25, LWZZ23, LLT18, LW24, LLG14, MCA11, MCA24, MVOC20, MP00, PP19, SLL24, WTZL17, WWT22, XZ18, ZY06, ZZF19, GM98a]. Testing [AHM18, AFO22a, AFO22b, BE20, BD00, CR17, CRdUÁH19, DG22, EG12, FLS25, FJL24, FB22, GMC93, HPF12, HHM21, HC21, JGCRJJ22, JGdUÁ24, LGL25, MSK23, MK25, MLDJ16, PL03, Rya12, SP24, SW16, Xia17, ZKR18, AGV17, AT05, AACRC19, BPD16, Bha20, BSFB20, CQ05, CGX23a, CGX23b, CL23, Cey14, CJ23, CWZ21, FMR24, GDS03, GVS98, HQ23, HX13, HG18, HN18, JFCZ14, KN13, KNV21, KL14, LKM23, LZYZ23, MQ01, Mor05, MM22, Pew18, Rit23, RC94, RC96, Rue92, TF23, Tse02, WYH24, WW06, WZ12, Zha14, dUÁ23, CR22]. Tests [EH20b, HLM22, HM10, Kla05, SR20, AH17, AW20, AABL18, ACW17, AE06, BBC22a, Bic07, Bor01, Cab09, Cao07, CM07, CNJ08, CH09a, CH09b, CV16, CG19, Csö02, CM22, DF19, Det13, Di 12, Duc01, Efr07, FJ07a, FJ07b, dMGP23, FSF23, FWZ18, FGFBGM23a, FGFBGM23b, FB22, FS12, Gam13, GP08, GP14, GB16b, GR18, GMC13a, GMC13b, GALT23, GLW25, Hal07, HJG19, Hid99, Hor07, HHM24, JN09, JGLM20, LW07, LW12, MEW01, MS10, Mam07, MHSB20, MT08, Mei13, Mol09, MP01, Mül07, MK09, Mur16, NSS13, Pau11, PKB23, Raï10, Rit13, SGR07, She13, SM23, Spe13a, Ten22, Tha09, VMS08, Van13a, VFVFGM07, WTZL17, WH14, YRR15, Yek15, ZNSW19, dUÁ13, dW02, dBCAM^{+}00, EH20a, JG20, Mei20, Ric20]. their [AM16, Cas12, DFK12, Eil20, GS20, HF19, Kne20, LR12, PN22, Raj12, Sta12, WTZL17, WS12a, WS12b, Woo20a, Woo20b]. theorem [BPY18, FMR24, PR05, Rom94, VPR15]. theorems [BFFS09, CRV12, IPR11, PRSW16]. theoretical [WTZL17]. Theory [Bül97, Ane12, ACR17, BP08, CBB^{+}95, DBC^{+}97, Dol12, Dou12, Fok12, Gal12, Gao12, Hei12, Ked12, LBSM13, LBSM15, MCA11, MCA24, MIR03, RI92, TA06, Tjø12a, Tjø12b, WZR94, dBCAM^{+}00]. thinned [AF07]. thinning [BSNSS23]. third [CB07, GdW12]. third-order [CB07, GdW12]. three [NARPV99]. three-stage [NARPV99]. threshold [PAT04, SMS23, SL02, SYL25, Vil17]. Thresholding [KPB^{+}00, Riv04, Aut08]. tied [DP18]. Time [dNGL16, Ane12, BBC22a, Bel14, BMRR14, Ber11, BdCPG14, BW17, Bra11, BKM22, CSL22, Cao99, Cha95, DK23, Dol12, DPR11a, DPR11b, Dou12, FGV02, FF21, Fok12, Gal12, Gao12, GM95, GN99, HL14, Hei12, Her14, Hid99, HHM24, KWWZ23, Ked12, LM11, LLG14, LLVR21, LC12, MU23a, MU23b, MM02, MMS21, Mor14, MW04, MLDJ16, Muñ14, NM20, NPM22, Nik24, OMRM24a, OMRM24b, Pap11, PST14a, PST14b, Pin14, RSMJ19, RPL01, SMS23, SU14, Stu01, TL14, Tjø12a, Tjø12b, Vel11, WWT22, WG21, YZ19, ZLWH17, ZZY25, ZGGX14a, ZGGX14b]. time-transformed [NM20, NPM22]. Time-varying [dNGL16, MLDJ16, WG21, ZLWH17]. times [AHKS08, GHF23, IV05, NLP17, PORCGP00, Vil95]. tolerance [BBC10, VH07]. tool [AAMDR20]. total [BMRR14]. tour [AG16, BS16a, BS16b, BL16b, GW16, HM16, Wag16]. Tractable [KPJ22]. tradeoff [WC95]. trait [Rab14]. transform [BMRR14]. transformation [AHM18, AN19, CV16, KNV21, RJP11, sS12, WZYW18, Yan99, Yan00]. transformations [Lop10, Men94b]. transformed [Dag01, NM20, NPM22]. transition [PORCGP00, Rum03]. trees [BFFS09]. trend [FPI96, MPU03, MSK23, MK25]. TREX [BGLM19]. trials [ANZ22]. triangular [dBJM09]. trigonometric [Spe19]. Trimmed [FM01, Cíz13]. truncated [CJV05, LdUÁPF25, LdUÁdCIP11, LdUÁdCIP12, RSM06, sS12, She13]. truncation [ABA^{+}02, EM15, GS15a, GS15b]. tuning [GP14]. Tweedie [MK14]. Two [GS13, GLÁM21, JFCZ14, LLLZ24, PP19, RP09, WG07, BBC10, BOQ17, BH16, Bel14, BFR00, BCCAVG22, Bor01, dCCIS21, Cey14, CPW21, FS98, Fer04, GY96, GB16b, HL14, HN18, Jah03, LdUÁPF25, Lef03, LC06, MSK23, MK25, Mor14, Pér94, PST14a, PST14b, Ras95, SU14, VH07, WTZL17, WY23, XTZ20, ZY06]. two-component [Jah03]. two-dimensional [Lef03]. two-factor [VH07]. two-piece [BCCAVG22]. two-point [CPW21]. Two-sample [PP19, BBC10, BOQ17, BH16, FS98, Fer04, GB16b, LdUÁPF25, WTZL17]. two-sided [LC06]. Two-stage [GS13, Bel14, HL14, Mor14, PST14a, PST14b, SU14]. Two-step [LLLZ24]. Two-way [GLÁM21, MSK23, MK25, Ras95]. type [AHMJ92, AV00, AZ04, BBBV02, CQ05, CdU07, DGG14, FPRG17, GG11, IPR11, IKvdH23, Kal22, LdUÁPF25, MS11, MS15, Rad04, Rao01, SWZ25, BPD16, BP17, CI10, BBC10, WY09]. type-II [BBC10, WY09].

ultra [JS22]. ultra-high-dimensional [JS22]. ultrastructural [SS97]. unbiased [CG02, DGG14, SB92]. uncertain [BKM22, LIAV02]. uncertainty [LLJ21, LMJ25]. unconstrained [CD10]. uncorrelated [Raï10]. underdispersion [BW17]. Understanding [Cha95, DP23]. unequal [CVO02, Sah07, VD96]. unequally [AOV99]. unified [BCD23, ZG07]. Uniform [AM16, Lou05, CRV21a, CRV21b, GW99, MS11, MS15, Mon11]. uniformity [dMGP23]. unify [GCY21]. unimodal [AVR00]. Unit [PPT24b, BMRR21, HMS18, PPT24a]. unit-level [HMS18]. Unit-Weibull [PPT24b, PPT24a]. univeriate [GRT01]. Universal [DBC^{+}97]. unknown [AAMDR20, Rum03, YLL19]. unmatched [SKR18]. unordered [JV98]. unspecified [MWY21]. upcrossings [SMFP13]. updated [Det13, Gam13, GMC13a, GMC13b, Mei13, Spe13a, Van13a, dUÁ13]. updating [WMS25]. upper [RMLDG03]. urn [MPU13]. use [BDMPF22, Bou94, IP94]. used [CAM15, MCA11, MCA24, TS05]. useful [BdM22]. Using [CQ05, dlH92, AGMT23, Ava22, BPD16, BP17, BCS15a, BCS15b, BKM22, CG15, FvdW08, FG20, GR07, Gel15, GDVPP06, GMC93, GM95, Guo08, HMZ09, Kon13, LWML15, Lin15, NECA21, NR10, Ozt19, PJPW24, Rod94, RSW08a, RSW08b, RM15, SH08, SJD21, SG04, SCJS07, SAF21a, SAF21b, TS15, Tro08, UMG09, Yek08].

V [Rom94]. V-C [Rom94]. validation [GK19, dBIM20, dRF92]. validity [AQGSM05, CM12]. value [CQ05, DRB99, DGGJ01, GP03, GP14, GdW12, MF05, NA23, PR05, WG07]. valued [CZ18, HSC10, HLM22, LNZ23, SYL25]. values [GLS15, LBSM13, LBSM15, VK20, Wan19, WLL25, dlHRB01]. variability [AF24, GG21, Van12]. Variable [GM24, LTWY23, BPD16, BP17, CSL22, CBV24, GM95, MG08a, MG08b, MS17, WZYW18, ZLYH16]. Variable-dependent [LTWY23]. variables [BS97, BEMP20, CRV12, CRV21a, CRV21b, CMLB05, CFRG10, DP23, Dub99, FL24, GP08, GLSU15, GNDR09, HPO04, HCS17, IPR11, Nig06, NAV21, PN22, RAVL15, SP15, Vél01, WXH^{+}14, WSCH15, WW12, WWHY18, dBJM09]. Variance [ZPH18, ANZ22, BPY18, CY15, CL10, EE92, GR18, GMdP05, GM03, Kok94, MK14, NBGP22, NSS13, PJPW24, Rad98, RAP12, VD96]. variance-stabilized [ANZ22]. variances [Pau11]. variants [Yua24]. variates [Lop10]. variation [AH17]. variations [ASS07]. VARX [Duc05]. varying [AGV17, AGV14, Bra20, GSBS04, GWHY14, HJV18, LW24, LGW19, MMR04, MLDJ16, TZ15, WG21, YWLZ15, YLL19, ZLWH17, dNGL16]. varying-coefficient [TZ15, YWLZ15, YLL19]. vast [Kon13]. vector [HN18, MU23a, MU23b, Raï10, WMS25]. vectors [AN19, BP08, HLM22, Jar15, LHB^{+}17]. version [WWHY18]. via [BS97, BE20, CS97a, CWZ21, FFO20, GS24, GM24, GMNH24, GW99, HV14, LSW24, LL23, LLR^{+}24, LGW19, PJH^{+}24a, PJH^{+}24b, RJP11, STPC12, SZ24, SKR18, VPP22, VS09, WL22, WLL25, ZZY25, ZS24]. view [FL08, SCC24, Van95]. viewed [LD96]. viewpoint [WZR94]. viral [MCL16]. volatility [LLG14].

wage [AGMT23]. Wang [NA07]. Wasserstein [dW02]. Watson [GS07, MGP00, MMR04]. Wavelet [FPRMA04, Gan13]. Wavelet-based [FPRMA04]. wavelets [Bar09]. way [Bia95, GP14, GLÁM21, GALT23, LX16, MSK23, MK25, Ras95]. weak [MU23a, MU23b]. weakly [Ber11, Bra11, DPR11a, DPR11b, LM11, Pap11, Sch98, Vel11]. Weibull [CQ05, DGGG08, FBKV14, GLSU15, GG11, PPT24a, PPT24b]. Weibull-like [GLSU15]. Weibull-type [GG11]. Weight [FGBB^{+}22]. Weighted [AG19, Csö02, GMdP11, LD96, TZ15, WWHY18, BBS18, BS97, CRV12, CRV21a, CRV21b, CMLB05, EH20a, EH20b, GG11, HCS17, JG20, Mei20, NA23, Qin25, Rah09, Ric20, RC94, STPC12, SDZ20, SR20, WSCH15, dW02]. Weinberg [CGPV08]. Weiss [Alv01]. Weiss-Hill [Alv01]. well [KPB^{+}00]. well-concentrated [KPB^{+}00]. Where [Zha22]. which [GP08, LBW01]. white [De 08, GJL96, PAT04]. Whitney [DP18]. whose [ASLFP13, FS12]. wide [CWZ21]. widely [WXH^{+}14]. width [DL04]. WIKS [dCCIS21]. wind [Her14, Muñ14, Pin14, TL14, ZGGX14a, ZGGX14b]. winds [ARM08, Bic08, Bri08, Daw08, Fue08, GSG^{+}08a, GSG^{+}08b, Jol08, WJ08]. without [FL24]. Wold [FMR24]. work [Ano07, Ano08, Ano09, Ano10, Ano11, GT18, Mac18a, Mac18b, MB18, SF18]. Wu [Hor24].

XII [AHMJ92].

year [DT19b]. years [BR19, DT19a, Kot19, Mor19, WZ19, del19a]. Yu [BKM22].

zero [KWWZ23]. zero-and-one [KWWZ23]. zonal [DGGJ01].

References

Alvarez-Andrade:2018:SAF

[AABL18] Sergio Alvarez-Andrade, Salim Bouzebda, and Aimé Lachal. Strong approximations for the p-fold integrated empirical process with applications to statistical tests. TEST, 27 (4):826–849, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0572-0.

Ameijeiras-Alonso:2019:MTC

[AACRC19] Jose Ameijeiras-Alonso, Rosa M. Crujeiras, and Alberto Rodríguez-Casal. Mode testing, critical bandwidth and excess mass. TEST, 28(3):900–919, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0611-5.

Abu-Awwad:2020:FSM

[AAMDR20] A. Abu-Awwad, V. Maume-Deschamps, and P. Ribereau. Fitting spatial max-mixture processes with unknown extremal dependence class: an exploratory analysis tool. TEST, 29(2):479–522, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00663-5.

Arnold:2002:SMM

[ABA^{+}02] Barry C. Arnold, Robert J. Beaver, A. Azzalini, N. Balakrishnan, A. Bhaumik, D. K. Dey, C. M. Cuadras, J. M. Sarabia, Barry C. Arnold, and Robert J. Beaver. Skewed multivariate models related to hidden truncation and/or selective reporting. TEST, 11(1):7–54, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595728.

Aoki:2001:NIM

[ABS01] Reiko Aoki, Hereno Bolfarine, and Julio M. Singer. Null intercept measurement error regression models. TEST, 10 (2):441–457, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595707.

Atkinson:2017:RBR

[ACR17] Anthony C. Atkinson, Aldo Corbellini, and Marco Riani. Robust Bayesian regression with the forward search: theory and data analysis. TEST, 26(4):869–886, December 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0542-6;   http://link.
springer.com/content/pdf/10.1007/s11749-017-0542-
6.pdf.

Angulo:1998:D

[ACW^{+}98] José M. Angulo, N. Cressie, C. K. Wikle, P. García Soidán, M. Febrero Bande, C. A. Glasbey, John T. Kent, Ana F. Militino, and Michael L. Stein. Discussion. TEST, 7(2):283– 285, December 1998. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02565112.

Arias-Castro:2017:DFT

[ACW17] Ery Arias-Castro and Meng Wang. Distribution-free tests for sparse heterogeneous mixtures. TEST, 26(1):71–94, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0499-x.

Arcones:2006:EDP

[ACZ06] Miguel A. Arcones, Hengjian Cui, and Yijun Zuo. Empirical depth processes. TEST, 15(1):151–177, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595423.

Abraham:1999:AAI

[AD99] Christophe Abraham and Jean-Pierre Daurès. Analytic approximation of the interval of Bayes actions derived from a class of loss functions. TEST, 8(1):129–145, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595866.

Adrover:2023:ARC

[AD23] Jorge G. Adrover and Stella M. Donato. Aspects of robust canonical correlation analysis, principal components and association. TEST, 32(2):623–650, June 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00846-1.

Arranz:2006:BCT

[AE06] Miguel A. Arranz and Alvaro Escribano. Bootstrapping cointegration tests under structural co-breaks: A robust extended ECM test. TEST, 15(1):179–208, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595424.

Alvarez-Esteban:2016:CMS

[ÁEdBCAM16] P. C. Álvarez-Esteban, E. del Barrio, J. A. Cuesta-Albertos, and C. Matrán. A contamination model for the stochastic order. TEST, 25(4):751–774, December 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0494-2.

Assuncao:2007:ITP

[AF07] R. M. Assunção and P. A. Ferrari. Independence of thinned processes characterizes the Poisson process: An elementary proof and a statistical application. TEST, 16(2):333–345, August 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0012-z.

Asadi:2024:VMR

[AF24] Majid Asadi and Maxim Finkelstein. On variability of the mean remaining lifetime at random age. TEST, 33(3):717– 730, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00917-3.

Ascorbebeitia:2022:CTC

[AFO22a] Jone Ascorbebeitia, Eva Ferreira, and Susan Orbe. Correction to: Testing conditional multivariate rank correlations: the effect of institutional quality on factors influencing competitiveness. TEST, 31(4):1167, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00823-0. See [AFO22b].

Ascorbebeitia:2022:TCM

[AFO22b] Jone Ascorbebeitia, Eva Ferreira, and Susan Orbe. Testing conditional multivariate rank correlations: the effect of institutional quality on factors influencing competitiveness. TEST, 31(4):931–949, December 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00806-1. See correction [AFO22a].

Arnold:1998:DMN

[AG98] Barry C. Arnold and D. V. Gokhale. Distributions most nearly compatible with given families of conditional distributions. TEST, 7(2):377–390, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02565119.

Azais:2015:SPI

[AG15] Romain Azaïs and Alexandre Genadot. Semi-parametric inference for the absorption features of a growth-fragmentation model. TEST, 24(2):341–360, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-014-0410-
6.

Arlot:2016:CRF

[AG16] Sylvain Arlot and Robin Genuer. Comments on: “A random forest guided tour”. TEST, 25(2):228–238, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0484-4. See [BS16a] and rejoinder

[BS16b].

Agostinelli:2019:WLE

[AG19] Claudio Agostinelli and Luca Greco. Weighted likelihood estimation of multivariate location and scatter. TEST, 28 (3):756–784, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0596-0.

Anastasiade-Guinand:2023:GWD

[AGMT23] Mihaela-Catalina Anastasiade-Guinand, Alina Matei, and Yves Tillé. Gender wage difference estimation at quantile levels using sample survey data. TEST, 32(4):1392–1433, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00885-8.

Aguilera:2016:CPE

[Agu16] Ana M. Aguilera. Comments on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):23–26, March 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0475-x;   http://link.springer.com/
content/pdf/10.1007/s11749-015-0475-x.pdf. See [YWZ16a]

and rejoinder [YWZ16b].

Andriyana:2014:PSQ

[AGV14] Y. Andriyana, I. Gijbels, and A. Verhasselt. P-splines quantile regression estimation in varying coefficient models. TEST, 23(1):153–194, March 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-013-0346-
2.

Ahkim:2017:STV

[AGV17] M. Ahkim, I. Gijbels, and A. Verhasselt. Shape testing in varying coefficient models. TEST, 26(2):429–450, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0518-y.

Al-Hussaini:2010:IBC

[AH10] Essam K. Al-Hussaini. Inference based on censored samples from exponentiated populations. TEST, 19(3):487–513, November 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0183-5.

Aerts:2017:RAT

[AH17] Stephanie Aerts and Gentiane Haesbroeck. Robust asymptotic tests for the equality of multivariate coefficients of variation. TEST, 26(1):163–187, March 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0504-4.

Al-Hussaini:2003:BIP

[AHA03] Essam K. Al-Hussaini and Abd El-Baset A. Ahmad. On Bayesian interval prediction of future records. TEST, 12(1): 79–99, June 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595812.

Aue:2008:MSM

[AHKS08] Alexander Aue, Lajos Horváth, Piotr Kokoszka, and Josef Steinebach. Monitoring shifts in mean: asymptotic normality of stopping times. TEST, 17(3):515–530, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0041-7.

Allison:2018:TAS

[AHM18] J. S. Allison, M. Husková, and S. G. Meintanis. Testing the adequacy of semiparametric transformation models. TEST, 27(1):70–94, March 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0544-4.

Al-Hussaini:1992:EUB

[AHMJ92] E. K. Al-Hussaini, M. A. M. A. Mousa, and Z. F. Jaheen. Estimation under the Burr type XII failure model based on censored data: a comparative study. TEST, 1(1):47–60, December 1992. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562661.

Albano:2004:BIR

[AJS04] Gian Luigi Albano and Frédéric Jouneau-Sion. Bayesian inference in repeated English auctions. TEST, 13(1):193– 211, June 2004. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02603006.

Atienza:2022:HSB

[ALB22a] David Atienza, Pedro Larrañaga, and Concha Bielza. Hybrid semiparametric Bayesian networks. TEST, 31(2):299– 327, June 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00812-3. See comments

[Scu22, Sal22, Spe22, Mor22] and rejoinder [ALB22b].

Atienza:2022:RHS

[ALB22b] David Atienza, Pedro Larrañaga, and Concha Bielza. Rejoinder on: Hybrid semiparametric Bayesian networks. TEST, 31(2):344–347, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00821-2. See

[ALB22a, ALB22b].

Alves:2001:WHE

[Alv01] M. Isabel Fraga Alves. Weiss-Hill estimator. TEST, 10(1): 203–224, June 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595832.

Agostinelli:2015:RRE

[ALYZ15a] Claudio Agostinelli, Andy Leung, Victor J. Yohai, and Ruben H. Zamar. Rejoinder on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):484–488, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0457-z;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0457-

z.pdf. See [ALYZ15b, CÖ15, Far15, Mar15a, RV15, Van15, Wel15].

Agostinelli:2015:REM

[ALYZ15b] Claudio Agostinelli, Andy Leung, Victor J. Yohai, and Ruben H. Zamar. Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination. TEST, 24(3):441–461, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-015-0450-6. See comments [CÖ15, Far15, Mar15a, RV15, Van15, Wel15] and rejoinder [ALYZ15a].

Afendras:2016:UIO

[AM16] Georgios Afendras and Marianthi Markatou. Uniform integrability of the OLS estimators, and the convergence of their moments. TEST, 25(4):775–784, December 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0498-y.

Aguilera-Morillo:2013:PSA

[AMAEV13] M. Carmen Aguilera-Morillo, Ana M. Aguilera, Manuel Escabias, and Mariano J. Valderrama. Penalized spline approaches for functional logit regression. TEST, 22(2):251– 277, June 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0307-1.

Anzarmou:2023:SOL

[AMO23] Youssef Anzarmou, Abdallah Mkhadri, and Karim Oualkacha. Sparse overlapped linear discriminant analysis. TEST, 32(1):388–417, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00839-6.

Arevalillo:2019:SOB

[AN19] Jorge M. Arevalillo and Hilario Navarro. A stochastic ordering based on the canonical transformation of skewnormal vectors. TEST, 28(2):475–498, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0583-5.

Andel:1997:LSN

[And97] J. Andel. On least-squares and naïve extrapolations in a non-linear AR(1) process. TEST, 6(1):91–100, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02564427.

Aneiros:2012:CSR

[Ane12] Germán Aneiros. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):439–441, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0299-x. See [Tjø12b] and

rejoinder [Tjø12a].

Anonymous:2007:ARW

[Ano07] Anonymous. Acknowledgement of refereeing work. TEST, 16(3):613–614, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0079-1.

Anonymous:2008:ARW

[Ano08] Anonymous. Acknowledgment of refereeing work. TEST, 17 (3):628–629, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0132-8.

Anonymous:2009:ARW

[Ano09] Anonymous. Acknowledgement of refereeing work. TEST, 18(3):605–606, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0170-x.

Anonymous:2010:ARW

[Ano10] Anonymous. Acknowledgement of refereeing work. TEST, 19(3):609–610, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0221-3.

Anonymous:2011:ARW

[Ano11] Anonymous. Acknowledgement of refereeing work. TEST, 20(3):678–679, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0276-9.

Anonymous:2025:JNa

[Ano25a] Anonymous. Journal navigation. TEST, 34(1):??, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic).

Anonymous:2025:JNb

[Ano25b] Anonymous. Journal navigation. TEST, 34(2):502–507, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic).

Anonymous:2025:JNc

[Ano25c] Anonymous. Journal navigation. TEST, 34(3):786–814, September 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic).

Antoniadis:2010:CPM

[Ant10] Anestis Antoniadis. Comments on: “l_{1}-penalization for mixture regression models”. TEST, 19(2):257–258, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0198-y. See [SBvdG10a] and rejoin-

der [SBvdG10b].

Antognini:2022:NIA

[ANZ22] Alessandro Baldi Antognini, Marco Novelli, and Maroussa Zagoraiou. A new inferential approach for response-adaptive clinical trials: the variance-stabilized bootstrap. TEST, 31(1):235–254, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00777-9.

Aguilera:1999:FUS

[AOV99] Ana M. Aguilera, Francisco A. Ocaña, and Mariano J. Valderrama. Forecasting with unequally spaced data by a functional principal component approach. TEST, 8(1):233– 253, June 1999. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595871.

Aneiros:2001:APP

[AQ01] Germán Aneiros and Alejandro Quintela. Asymptotic properties in partial linear models under dependence. TEST, 10 (2):333–355, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595701.

Andres:2005:VCC

[AQGSM05] A. Martín Andrés, M. J. Sánchez Quevedo, J. M. Tapia García, and A. Silva-Mato. On the validity condition of the chi-squared test in 2×2 tables. TEST, 14(1):99–128, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595399.

Arcones:2002:MDM

[Arc02] Miguel A. Arcones. Moderate deviations for m-estimators. TEST, 11(2):465–500, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595717.

Arcones:2005:COE

[Arc05] Miguel A. Arcones. Convergence of the optimal M-estimator over a parametric family of M-estimators. TEST, 14(1):281– 315, June 2005. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595407.

Arellano:2007:CPD

[Are07] Manuel Arellano. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):23–27, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0047-9. See [Hsi07a] and rejoinder

[Hsi07b].

Arevalillo:2014:HOA

[Are14] Jorge M. Arevalillo. Higher-order approximations to the quantile of the distribution for a class of statistics in the first-order autoregression. TEST, 23(2):291–310, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0348-0.

Angulo:2008:CAP

[ARM08] José M. Angulo and María D. Ruiz-Medina. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):236–237, August 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0115-9.

Arnold:2007:CPC

[Arn07] Barry C. Arnold. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):268–270, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0065-7. See [Bal07a] and rejoinder

[Bal07b].

Arnold:2022:PDP

[ARS22] Barry C. Arnold, Tomasz Rychlik, and Magdalena Szymkowiak. Preservation of distributional properties of component lifetimes by system lifetimes. TEST, 31(4):901–930, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00808-z.

Aneiros:2018:BSF

[ARVV18] Germán Aneiros, Paula Raña, Philippe Vieu, and Juan Vilar. Bootstrap in semi-functional partial linear regression under dependence. TEST, 27(3):659–679, September 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-017-0566-y.

Amo-Salas:2013:ODS

[ASLFP13] Mariano Amo-Salas, Jesús López-Fidalgo, and Emilio Porcu. Optimal designs for some stochastic processes whose covariance is a function of the mean. TEST, 22(1):159–181, March 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0311-5.

Al-Saleh:2007:NIP

[ASS07] M. F. Al-Saleh and H. M. Samawi. A note on inclusion probability in ranked set sampling and some of its variations. TEST, 16(1):198–209, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0009-7.

Aston:2014:CES

[Ast14] John A. D. Aston. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2): 256–257, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0369-3;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0369-

3.pdf. See [HR14a] and rejoinder [HR14b].

Akkaya:2005:REH

[AT05] Aysen D. Akkaya and Moti L. Tiku. Robust estimation and hypothesis testing under short-tailedness and inliers. TEST, 14(1):129–150, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595400.

Akkaya:2008:STD

[AT08] Aysen D. Akkaya and Moti L. Tiku. Short-tailed distributions and inliers. TEST, 17(2):282–296, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0032-8.

Autin:2008:MRR

[Aut08] Florent Autin. Maxisets for μ-thresholding rules. TEST, 17(2):332–349, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0035-5.

Anido:2000:IEP

[AV00] Carmen Anido and Teófilo Valdés. An iterative estimating procedure for probit-type nonresponse models in surveys with call backs. TEST, 9(1):233–253, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595860.

Aneiros:2016:CPE

[AV16] Germán Aneiros and Philippe Vieu. Comments on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):27–32, March 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0471-1;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0471-

1.pdf. See [YWZ16a] and rejoinder [YWZ16b].

Avagyan:2022:PME

[Ava22] Vahe Avagyan. Precision matrix estimation using penalized Generalized Sylvester matrix equation. TEST, 31(4):950– 967, December 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00807-0.

Arellano-Valle:1994:PIM

[AVBI94] R. B. Arellano-Valle, H. Bolfarine, and P. L. Iglesias. A predictivistic interpretation of the multivariate t distribution. TEST, 3(2):221–236, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562703.

Anido:2000:MIE

[AVR00] Carmen Anido, Teófilo Valdés, and Carlos Rivero. Modal iterative estimation in linear models with unimodal errors and non-grouped and grouped data collected from different sources. TEST, 9(2):393–416, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595742.

Aleksandrov:2020:PED

[AW20] Boris Aleksandrov and Christian H. Weiß. Parameter estimation and diagnostic tests for INMA(1) processes. TEST, 29(1):196–232, March 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00653-7.

Aulogiaris:2004:MEC

[AZ04] G. Aulogiaris and K. Zografos. A maximum entropy characterization of symmetric Kotz type and Burr multivariate distributions. TEST, 13(1):65–83, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02603001.

Browne:2024:OSM

[BA24] Ryan P. Browne and Jeffrey L. Andrews. The orthogonal skew model: computationally efficient multivariate skewnormal and skew-t distributions with applications to modelbased clustering. TEST, 33(3):752–785, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00920-2.

Balakrishnan:2007:PCM

[Bal07a] N. Balakrishnan. Progressive censoring methodology: an appraisal. TEST, 16(2):211–259, August 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-007-0061-y. See comments [Nag07, Dem07, CS07, Arn07, KC07, Kun07, Gui07, Jos07, BL07, NC07] and rejoinder [Bal07b].

Balakrishnan:2007:RPC

[Bal07b] N. Balakrishnan. Rejoinder on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):290–296, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0072-8. See [Bal07a, Nag07, Dem07,

CS07, Arn07, KC07, Kun07, Gui07, Jos07, BL07, NC07].

Baltagi:2007:CPD

[Bal07c] Badi H. Baltagi. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):28–30, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0048-8. See [Hsi07a] and rejoinder

[Hsi07b].

Banerjee:2018:CPM

[Ban18] Sudipto Banerjee. Comments on: “Process modeling for slope and aspect with application to elevation data maps”. TEST, 27(4):773–775, December 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0621-3;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0621-3.pdf. See [WBG18a]

and rejoinder [WBG18b].

Barakat:1997:CLD

[Bar97] H. M. Barakat. On the continuation of the limit distributions of the extreme and central terms of a sample. TEST, 6(2):351–368, December 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02564703.

Barbedor:2009:ICA

[Bar09] Pascal Barbedor. Independent component analysis by wavelets. TEST, 18(1):136–155, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:/

/link.springer.com/article/10.1007/s11749-007-0073-
7.

Beamonte:2003:BSA

[BB03] Eduardo Beamonte and José D. Bermúdez. A Bayesian semiparametric analysis for additive hazard models with censored observations. TEST, 12(2):347–363, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595719.

Beirlant:2002:DTE

[BBBV02] Jan Beirlant, Alain Berlinet, Gérard Biau, and Igor Vajda. Divergence-type errors of smooth Barron-type density estimators. TEST, 11(1):191–217, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595736.

Balakrishnan:2010:ETS

[BBC10] N. Balakrishnan, E. Beutner, and E. Cramer. Exact twosample nonparametric confidence, prediction, and tolerance intervals based on ordinary and progressively type-II right censored data. TEST, 19(1):68–91, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0133-7.

Baragona:2022:DDP

[BBC22a] Roberto Baragona, Francesco Battaglia, and Domenico Cucina. Data-driven portmanteau tests for time series. TEST, 31(3):675–698, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
021-00794-8.

Bianco:2022:PRE

[BBC22b] Ana M. Bianco, Graciela Boente, and Gonzalo Chebi. Penalized robust estimators in sparse logistic regression. TEST, 31(3):563–594, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00792-w.

Bianco:2011:ABR

[BBGMPG11] Ana Bianco, Graciela Boente, Wenceslao González-Manteiga, and Ana Pérez-González. Asymptotic behavior of robust estimators in partially linear models with missing responses: the effect of estimating the missing probability on the simplified marginal estimators. TEST, 20(3):524–548, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0216-0.

Bianco:2019:PME

[BBGMPG19] Ana M. Bianco, Graciela Boente, Wenceslao González-Manteiga, and Ana Pérez-González. Plug-in marginal estimation under a general regression model with missing responses and covariates. TEST, 28(1):106–146, March 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0591-5.

Bunday:1997:NAN

[BBK97] B. D. Bunday, S. M. Husnain Bokhari, and K. H. Khan. A new algorithm for the normal distribution function. TEST, 6(2):369–377, December 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02564704.

Berrendero:2023:FLR

[BBLC23] José R. Berrendero, Beatriz Bueno-Larraz, and Antonio Cuevas. On functional logistic regression: some conceptual issues. TEST, 32(1):321–349, March 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00836-9.

Barron:2018:NWL

[BBS18] Andrew R. Barron, Mirta Bensić, and Kristian Sabo. A note on weighted least square distribution fitting and full standardization of the empirical distribution function. TEST, 27 (4):946–967, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0578-2.

Bertail:2007:SOP

[BC07] Patrice Bertail and Stéphan Clémençon. Second-order properties of regeneration-based bootstrap for atomic Markov chains. TEST, 16(1):109–122, May 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0004-z.

Borrajo:2021:NEB

[BC21] Laura Borrajo and Ricardo Cao. Nonparametric estimation for big-but-biased data. TEST, 30(4):861–883, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00749-5.

Bevilacqua:2022:CRF

[BCCAVG22] Moreno Bevilacqua, Christian Caamaño-Carrillo, Reinaldo B. Arellano-Valle, and Camilo Gómez. A class of random fields with two-piece marginal distributions for modeling pointreferenced data with spatial outliers. TEST, 31(3):644–674, September 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00797-5.

Boj:2016:GLD

[BCD^{+}16] Eva Boj, Adrià Caballé, Pedro Delicado, Anna Esteve, and Josep Fortiana. Global and local distance-based generalized linear models. TEST, 25(1):170–195, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0447-1.

Barabesi:2023:SMB

[BCD23] Lucio Barabesi, Andrea Cerioli, and Marco Di Marzio. Statistical models and the Benford hypothesis: a unified framework. TEST, 32(4):1479–1507, December 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00881-y.

Bourazas:2024:BSS

[BCD24] Konstantinos Bourazas, Guido Consonni, and Laura Deldossi. Bayesian sample size determination for detecting heterogeneity in multi-site replication studies. TEST, 33(3):697– 716, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00916-4.

Biau:2008:SCM

[BCDG08] Gérard Biau, Benoît Cadre, Luc Devroye, and László Györfi. Strongly consistent model selection for densities. TEST, 17 (3):531–545, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0042-6.

Bertail:2025:TIE

[BCF25] Patrice Bertail, Stephan Clémençon, and Carlos Fernández. Tail index estimation for discrete heavy-tailed distributions with application to statistical inference for regular Markov chains. TEST, 34(3):660–690, September 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00975-9.

Brunel:2009:NDE

[BCG09] E. Brunel, F. Comte, and A. Guilloux. Nonparametric density estimation in presence of bias and censoring. TEST, 18(1):166–194, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0075-5.

Brunel:2016:NDS

[BCGC16] Elodie Brunel, Fabienne Comte, and Valentine Genon-Catalot. Nonparametric density and survival function estimation in the multiplicative censoring model. TEST, 25 (3):570–590, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0479-1.

Bayer:2015:BBM

[BCN15] Fábio M. Bayer and Francisco Cribari-Neto. Bootstrapbased model selection criteria for beta regressions. TEST, 24(4):776–795, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0434-6.

Bradley:2015:CSS

[BCS15a] Jonathan R. Bradley, Noel Cressie, and Tao Shi. Comparing and selecting spatial predictors using local criteria. TEST, 24(1):1–28, March 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0415-1. See

comments [Gel15, CG15, Lin15, RM15, TS15] and rejoinder [BCS15b].

Bradley:2015:RCS

[BCS15b] Jonathan R. Bradley, Noel Cressie, and Tao Shi. Rejoinder on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1):54–60, March 2015. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-014-0414-2;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0414-

2.pdf. See [BCS15a, Gel15, CG15, Lin15, RM15, TS15].

Barahona:2020:SCG

[BCS20] S. Barahona, P. Centella, and A. Simó. Supervised classification of geometrical objects by integrating currents and functional data analysis. TEST, 29(3):637–660, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-019-00669-z.

Barbeito:2023:BSS

[BCS23] Inés Barbeito, Ricardo Cao, and Stefan Sperlich. Bandwidth selection for statistical matching and prediction. TEST, 32(1):418–446, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00838-7.

Biedermann:2000:TLR

[BD00] Stefanie Biedermann and Holger Dette. Testing linearity of regression models with dependent errors by kernel based methods. TEST, 9(2):417–438, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595743.

Biswas:2014:AAM

[BdCPG14] Atanu Biswas, Maria del Carmen Pardo, and Apratim Guha. Auto-association measures for stationary time series of categorical data. TEST, 23(3):487–514, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0364-8.

Bourguignon:2022:SUR

[BdM22] Marcelo Bourguignon and Rodrigo M. R. de Medeiros. A simple and useful regression model for fitting count data. TEST, 31(3):790–827, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00801-6.

Baz:2022:SRG

[BDMPF22] Juan Baz, Irene Díaz, Susana Montes, and Raúl Pérez-Fernández. Some results on the Gaussian Markov Random Field construction problem based on the use of invariant subgraphs. TEST, 31(3):856–874, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00804-3.

Betsch:2020:TND

[BE20] Steffen Betsch and Bruno Ebner. Testing normality via a distributional fixed point property in the Stein characterization. TEST, 29(1):105–138, March 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00630-0.

Bell:2014:CST

[Bel14] William R. Bell. Comments on: “Single- and two-stage cross-sectional and time series benchmarking procedures for small area estimation”. TEST, 23(4):667–669, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0382-6;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0382-

6.pdf. See [PST14b] and rejoinder [PST14a].

Burgard:2020:FHM

[BEMP20] Jan Pablo Burgard, María Dolores Esteban, Domingo Morales, and Agustín Pérez. A Fay–Herriot model when auxiliary variables are measured with error. TEST, 29(1): 166–195, March 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00649-3.

Bernardo:2005:ICR

[Ber05] José M. Bernardo. Intrinsic credible regions: an objective Bayesian approach to interval estimation. TEST, 14(2):317– 384, December 2005. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595408.

Bertail:2011:CSW

[Ber11] Patrice Bertail. Comments on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20 (3):487–490, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0272-0. See

[DPR11b] and rejoinder [DPR11a].

Berkes:2014:CES

[Ber14] István Berkes. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):258– 260, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0370-x;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0370-

x.pdf. See [HR14a] and rejoinder [HR14b].

Balding:2009:LTS

[BFFS09] David Balding, Pablo A. Ferrari, Ricardo Fraiman, and Mariela Sued. Limit theorems for sequences of random trees. TEST, 18(2):302–315, August 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0092-z.

Bartolucci:2014:LMM

[BFP14a] F. Bartolucci, A. Farcomeni, and F. Pennoni. Latent Markov models: a review of a general framework for the analysis of longitudinal data with covariates. TEST, 23(3):433–465, September 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0381-7. See comments

[Bia14, BM14, Paa14, VS14] and rejoinder [BFP14b].

Bartolucci:2014:RLM

[BFP14b] F. Bartolucci, A. Farcomeni, and F. Pennoni. Rejoinder on: “Latent Markov models: a review of a general framework for the analysis of longitudinal data with covariates”. TEST, 23(3):484–486, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0393-3;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0393-3.pdf. See [BFP14a, Bia14, BM14, Paa14, VS14].

Berti:2000:ENP

[BFR00] Patrizia Berti, Lorenzo Fattorini, and Pietro Rigo. Eliminating nuisance parameters: two characterizations. TEST, 9(1): 133–148, June 2000. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595855.

Bien:2019:PEB

[BGLM19] Jacob Bien, Irina Gaynanova, Johannes Lederer, and Christian L. Müller. Prediction error bounds for linear regression with the TREX. TEST, 28(2):451–474, June 2019.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0584-4.

Battagliese:2023:CMP

[BGLV23] Diego Battagliese, Clara Grazian, Brunero Liseo, and Cristiano Villa. Copula modelling with penalized complexity priors: the bivariate case. TEST, 32(2):542–565, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00843-w.

Bagkavos:2024:NCS

[BGN24] Dimitrios Bagkavos, Montserrat Guillen, and Jens P. Nielsen. Nonparametric conditional survival function estimation and plug-in bandwidth selection with multiple covariates. TEST, 33(4):1225–1257, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00945-7.

Baringhaus:2016:RTS

[BH16] Ludwig Baringhaus and Norbert Henze. Revisiting the two-sample runs test. TEST, 25(3):432–448, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0463-1.

Bhattacharya:2020:ICO

[Bha20] Ritwik Bhattacharya. Implementation of compound optimal design strategy in censored life-testing experiment. TEST, 29(4):1029–1050, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-019-00699-7.

Benhenni:2017:RMC

[BHGR17] Karim Benhenni, Sonia Hedli-Griche, and Mustapha Rachdi. Regression models with correlated errors based on functional random design. TEST, 26(1):1–21, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0495-1.

Boniece:2024:CPD

[BHJ24] B. Cooper Boniece, Lajos Horváth, and Peter M. Jacobs. Change point detection in high dimensional data with U-statistics. TEST, 33(2):400–452, June 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00900-y.

Bian:1995:RBE

[Bia95] G. Bian. Robust Bayesian estimators in a one-way ANOVA model. TEST, 4(1):115–135, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02563106.

Bian:1997:BIL

[Bia97] Gorui Bian. Bayesian inference in location-scale distributions with independent bivariate priors. TEST, 6(1):137– 157, June 1997. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02564431.

Bianconcini:2014:CLM

[Bia14] Silvia Bianconcini. Comments on: “Latent Markov models: a review of a general framework for the analysis of longitudinal data with covariates”. TEST, 23(3):466–468, September 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0390-6;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0390-

6.pdf. See [BFP14a] and rejoinder [BFP14b].

Bickel:2007:CNI

[Bic07] Peter J. Bickel. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):445– 447, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0081-7. See [FJ07a] and re-

joinder [FJ07b].

Bickel:2008:CAP

[Bic08] Peter J. Bickel. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):238– 239, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0116-8.

Bereswill:2013:ENM

[BJ13] Mareike Bereswill and Jan Johannes. On the effect of noisy measurements of the regressor in functional linear models. TEST, 22(3):488–513, September 2013. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0325-7.

Bermudez:2022:CBB

[BK22] Lluís Bermúdez and Dimitris Karlis. Copula-based bivariate finite mixture regression models with an application for insurance claim count data. TEST, 31(4):1082–1099, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00814-1.

Boulesteix:2008:CAB

[BKK08] Anne-Laure Boulesteix, Athanassios Kondylis, and Nicole Krämer. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):31–35, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0103-0.

Burgard:2022:MER

[BKM22] Jan Pablo Burgard, Joscha Krause, and Domingo Morales. A measurement error Rao–Yu model for regional prevalence estimation over time using uncertain data obtained from dependent survey estimates. TEST, 31(1):204–234, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-021-00776-w.

Balasooriya:2007:CPC

[BL07] Uditha Balasooriya and C.-K. Low. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16 (2):284–286, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0070-x. See

[Bal07a] and rejoinder [Bal07b].

Biau:2016:CPE

[BL16a] Gérard Biau and Clément Levrard. Comments on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):41–43, March 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0472-0;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0472-

0.pdf. See [YWZ16a] and rejoinder [YWZ16b].

Buhlmann:2016:CRF

[BL16b] Peter Bühlmann and Florencia Leonardi. Comments on: “A random forest guided tour”. TEST, 25(2):239–246, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0483-5. See [BS16a] and rejoinder

[BS16b].

Bar-Lev:2003:CCP

[BLBB03] Shaul K. Bar-Lev, Elizabeta Bobovich, and Benzion Boukai. A common conjugate prior structure for several randomized response models. TEST, 12(1):101–113, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595813.

Boubeta:2016:EBP

[BLM16] Miguel Boubeta, María José Lombardía, and Domingo Morales. Empirical best prediction under area-level Poisson mixed models. TEST, 25(3):548–569, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0469-8.

Bickel:2006:RS

[BLT^{+}06] Peter J. Bickel, Bo Li, Alexandre B. Tsybakov, Sara A. van de Geer, Bin Yu, Teófilo Valdés, Carlos Rivero, Jianqing Fan, and Aad van der Vaart. Regularization in statistics. TEST, 15(2):271–344, September 2006. CODEN ????

ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02607055.

Bockenholt:2014:CLM

[BM14] Ulf Böckenholt and Blakeley B. McShane. Comments on: “Latent Markov models: a review of the general framework for the analysis of longitudinal data with covariates”. TEST, 23(3):469–472, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0388-0;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0388-0.pdf. See [BFP14a] and rejoinder [BFP14b].

Boente:2017:MIE

[BM17] Graciela Boente and Alejandra Martínez. Marginal integration M-estimators for additive models. TEST, 26(2):231– 260, June 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0508-0.

Borgert:2024:CSB

[BM24] J. E. Borgert and J. S. Marron. Comments on: Shapebased functional data analysis. TEST, 33(1):66–70, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00914-6. See [WHS24b, WHS24a].

Berger:1994:ORB

[BMP^{+}94] James O. Berger, Elías Moreno, Luis Raul Pericchi, M. Jesús Bayarri, José M. Bernardo, Juan A. Cano, Julián De la Horra, Jacinto Martín, David Ríos-Insúa, Bruno Betrò, A. Dasgupta, Paul Gustafson, Larry Wasserman, Joseph B. Kadane, Cid Srinivasan, Michael Lavine, Anthony O’Hagan, Wolfgang Polasek, Christian P. Robert, Constantinos Goutis, Fabrizio Ruggeri, Gabriella Salinetti, and Siva Sivaganesan. An overview of robust Bayesian analysis. TEST, 3(1):5– 124, June 1994. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562676.

Belzunce:2014:CSC

[BMRR14] Félix Belzunce, Carolina Martínez-Riquelme, and José M. Ruiz. A characterization and sufficient conditions for the total time on test transform order. TEST, 23(1):72–85, March 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0342-6.

Belzunce:2021:CPB

[BMRR21] Félix Belzunce, Carolina Martínez-Riquelme, and José M. Ruiz. Comparisons of policies based on relevation and replacement by a new one unit in reliability. TEST, 30(1): 211–227, March 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00710-6.

Belzunce:2015:RSM

[BMRSL15] Félix Belzunce, Julio Mulero, José María Ruíz, and Alfonso Suárez-Llorens. On relative skewness for multivariate distributions. TEST, 24(4):813–834, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0436-4.

Belzunce:2008:GOE

[BNOR08] Félix Belzunce, Asok K. Nanda, Eva-María Ortega, and José M. Ruiz. Generalized orderings of excess lifetimes of renewal processes. TEST, 17(2):297–310, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0033-7.

Bergesio:2021:RPE

[BNY21] Andrea Bergesio, María Eugenia Szretter Noste, and Víctor J. Yohai. A robust proposal of estimation for the sufficient dimension reduction problem. TEST, 30(3):758–783, September 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00745-9.

Barcenas:2017:QFE

[BOQ17] R. Bárcenas, J. Ortega, and A. J. Quiroz. Quadratic forms of the empirical processes for the two-sample problem for functional data. TEST, 26(3):503–526, September 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-017-0522-x.

Borroni:2001:SNA

[Bor01] Claudio G. Borroni. Some notes about nonparametric tests for the equality of two populations. TEST, 10(1):147–159, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595829.

Bouza:1994:UAI

[Bou94] Carlos Bouza. The use of auxiliary information for solving non-response problems. TEST, 3(2):113–122, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562697.

Bradshaw:2008:DTC

[BP08] David Bradshaw and Marianna Pensky. Decision theory classification of high-dimensional vectors based on small samples. TEST, 17(1):83–100, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0024-8.

Bhattacharya:2017:COT

[BP17] Ritwik Bhattacharya and Biswabrata Pradhan. Computation of optimum Type–II progressively hybrid censoring schemes using variable neighborhood search algorithm. TEST, 26(4):802–821, December 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0534-6.

Burke:2021:LBA

[BP21] Kevin Burke and Valentin Patilea. A likelihood-based approach for cure regression models. TEST, 30(3):693–712, September 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00738-8.

Bhattacharya:2016:OLT

[BPD16] Ritwik Bhattacharya, Biswabrata Pradhan, and Anup Dewanji. On optimum life-testing plans under Type–II progressive censoring scheme using variable neighborhood search algorithm. TEST, 25(2):309–330, June 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0449-z.

Bai:2018:CLT

[BPY18] Zhidong Bai, Guangming Pan, and Yanqing Yin. A central limit theorem for sums of functions of residuals in a highdimensional regression model with an application to variance homoscedasticity test. TEST, 27(4):896–920, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-017-0575-x.

Berg:2020:RBN

[BPZ20] Arthur Berg, Dimitris Politis, and Hui Zeng. Reduced bias nonparametric lifetime density and hazard estimation. TEST, 29(3):704–727, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
019-00677-z.

Balakrishnan:2004:POI

[BQ04] N. Balakrishnan and A. J. Quiroz. A procedure for outlier identification in data sets from continuous distributions. TEST, 13(1):247–262, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02603008.

Bertin:2009:MSN

[BR09] Karine Bertin and Vincent Rivoirard. Maxiset in supnorm for kernel estimators. TEST, 18(3):475–496, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0109-7.

Boente:2012:REG

[BR12] Graciela Boente and Daniela Rodriguez. Robust estimates in generalized partially linear single-index models. TEST, 21(2):386–411, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0249-z.

Burkschat:2018:SIQ

[BR18] Marco Burkschat and Tomasz Rychlik. Sharp inequalities for quantiles of system lifetime distributions from failuredependent proportional hazard model. TEST, 27(3):618–638, September 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0564-0.

Beaumont:2019:CDS

[BR19] Jean-Francois Beaumont and J. N. K. Rao. Comments on: Deville and Särndal’s Calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1071–1076, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.springer.

com/article/10.1007/s11749-019-00687-x. See [DT19a,

DT19b].

Bagyan:2024:CAE

[BR24] Armine Bagyan and Donald Richards. Complete asymptotic expansions and the high-dimensional Bingham distributions. TEST, 33(2):540–563, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00910-w.

Bravo:2011:CSW

[Bra11] Francesco Bravo. Comment on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20 (3):483–486, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0271-1. See

[DPR11b] and rejoinder [DPR11a].

Bravo:2020:REI

[Bra20] Francesco Bravo. Robust estimation and inference for general varying coefficient models with missing observations. TEST, 29(4):966–988, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-019-00692-0.

Briggs:2008:CAP

[Bri08] William M. Briggs. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):240– 242, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0117-7.

Boente:2020:REG

[BRV20] Graciela Boente, Daniela Rodriguez, and Pablo Vena. Robust estimators in a generalized partly linear regression model under monotony constraints. TEST, 29(1):50–89, March 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00629-7.

Bhattacharya:1994:BPB

[BS94] Samir K. Bhattacharya and Nand K. Singh. Bayesian prediction for business mortality analysis. TEST, 3(2):207–220, December 1994. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562702.

Basu:1997:REE

[BS97] A. Basu and S. Sarkar. Robust estimation in the errors variables model via weighted likelihood estimating equations. TEST, 6(1):187–203, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02564433.

Biau:2016:RFG

[BS16a] Gérard Biau and Erwan Scornet. A random forest guided tour. TEST, 25(2):197–227, June 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-016-0481-7. See comments [AG16, BL16b, GW16, HM16, Wag16] and rejoinder [BS16b].

Biau:2016:RRF

[BS16b] Gérard Biau and Erwan Scornet. Rejoinder on: “A random forest guided tour”. TEST, 25(2):264–268, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-016-0488-0;   http://link.
springer.com/content/pdf/10.1007/s11749-016-0488-

0.pdf. See [BS16a, AG16, BL16b, GW16, HM16, Wag16].

Bianco:2019:RIN

[BS19] Ana M. Bianco and Paula M. Spano. Robust inference for nonlinear regression models. TEST, 28(2):369–398, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-017-0570-2.

Bhattacharya:2020:MCB

[BSFB20] Ritwik Bhattacharya, Baidya Nath Saha, Graceila González Farías, and Narayanaswamy Balakrishnan. Multi-criteria-based optimal life-testing plans under hybrid censoring scheme. TEST, 29(2):430–453, June 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00660-8.

Barreto-Souza:2023:NLI

[BSNSS23] Wagner Barreto-Souza, Sokol Ndreca, Rodrigo B. Silva, and Roger W. C. Silva. Non-linear INAR(1) processes under an alternative geometric thinning operator. TEST, 32(2):695– 725, June 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00849-y.

Bernardo:2011:OBR

[BT11] José Miguel Bernardo and Vera Lucia Tomazella. Objective Bayesian reference analysis for the Poisson process model in presence of recurrent events data. TEST, 20(1):204–221, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0195-1.

Buhlmann:2019:CDS

[Büh19] Peter Bühlmann. Comments on: Data science, big data and statistics. TEST, 28(2):330–333, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00646-6. See [GP19a, GP19b].

Bultel:1997:CFR

[Bül97] Dirk Bültel. Correction to “Foundations for a Robust Theory of Decision Making: the Simple Case” by David Ríos– Insua. TEST, 6(2):433–435, December 1997. CODEN ????

ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02564708. See

[RI92].

Beyhum:2023:RCR

[BV23] Jad Beyhum and Ingrid Van Keilegom. Robust censored regression with ℓ_{1}-norm regularization. TEST, 32(1):146– 162, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00829-8.

Bourguignon:2017:IPM

[BW17] Marcelo Bourguignon and Christian H. Weiß. An INAR(1) process for modeling count time series with equidispersion, underdispersion and overdispersion. TEST, 26(4):847–868, December 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0536-4.

Bivand:2018:CIG

[BW18] Roger S. Bivand and David W. S. Wong. Comparing implementations of global and local indicators of spatial association. TEST, 27(3):716–748, September 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0599-x.

Bradic:2017:CHD

[BZ17] Jelena Bradic and Yinchu Zhu. Comments on: “Highdimensional simultaneous inference with the bootstrap”. TEST, 26(4):720–728, December 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-017-0556-

0. See [DBZ17a] and rejoinder [DBZ17b].

Cristobal:2001:ONC

[CA01] José A. Cristóbal and José T. Alcalá. An overview of nonparametric contributions to the problem of functional estimation from biased data. TEST, 10(2):309–332, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595700.

Cuesta-Albertos:2013:CMF

[CA13] Juan A. Cuesta-Albertos. Comments on: “Model-free modelfitting and predictive distributions”. TEST, 22(2):222– 223, June 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0319-5;   http://link.
springer.com/content/pdf/10.1007/s11749-013-0319-

5.pdf. See [Pol13a] and rejoinder [Pol13b].

Cabana:2009:CGF

[Cab09] Alejandra Cabaña. Comments on: “Goodness-of-fit-tests in mixed models”. TEST, 18(2):240–243, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-009-0149-7. See [CH09a] and rejoinder [CH09a].

Cuesta-Albertos:2010:SMA

[CAFB10] J. A. Cuesta-Albertos and M. Febrero-Bande. A simple multiway ANOVA for functional data. TEST, 19(3):537–557, November 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0185-3.

Cuesta-Albertos:2017:CFS

[CAFBdlF17] J. A. Cuesta-Albertos, M. Febrero-Bande, and M. Oviedo de la Fuente. The DD^{G}-classifier in the functional setting. TEST, 26(1):119–142, March 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0502-6.

Casero-Alonso:2015:ODS

[CALF15] Víctor Casero-Alonso and Jesús López-Fidalgo. Optimal designs subject to cost constraints in simultaneous equations models. TEST, 24(4):701–713, December 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0430-x.

Coelho:2015:ENE

[CAM15] Carlos A. Coelho, Barry C. Arnold, and Filipe J. Marques. The exact and near-exact distributions of the main likelihood ratio test statistics used in the complex multivariate normal setting. TEST, 24(2):386–416, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0418-y.

Cao:1999:OBM

[Cao99] Ricardo Cao. An overview of bootstrap methods for estimating and predicting in time series. TEST, 8(1):95–116, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595864.

Cao:2007:CNI

[Cao07] Ricardo Cao. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):468–470, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0089-z. See [FJ07a] and re-

joinder [FJ07b].

Cao:2019:CDS

[Cao19] Ricardo Cao. Comments on: Data science, big data and statistics. TEST, 28(3):664–670, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00666-2. See [GP19a, GP19b].

Cao:2023:CNE

[Cao23] Ricardo Cao. Comments on: Nonparametric estimation in mixture cure models with covariates. TEST, 32(2):499–505, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00856-z. See [LCPJ23a, LCPJ23b].

Cardot:2010:CDR

[Car10] Hervé Cardot. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):30–33, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0182-6.

Castelo:2012:CSR

[Cas12] Robert Castelo. Comments on: “Sequences of regressions and their independences”. TEST, 21(2):253–254, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0284-4. See [WS12b] and rejoinder

[WS12a].

Conde-Amboage:2019:PBS

[CASS19] Mercedes Conde-Amboage and César Sánchez-Sellero. A plug-in bandwidth selector for nonparametric quantile regression. TEST, 28(2):423–450, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0582-6.

Cordeiro:2007:TOB

[CB07] Gauss M. Cordeiro and Lúcia P. Barroso. A third-order bias corrected estimate in generalized linear models. TEST, 16 (1):76–89, May 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0002-1.

Cameron:2022:THF

[CB22] James Cameron and Pramita Bagchi. A test for heteroscedasticity in functional linear models. TEST, 31(2):519–542, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-021-00786-8.

Cox:1995:RBT

[CBB^{+}95] D. R. Cox, M. J. Bayarri, M. J. Bayarri, C. M. Cuadras, José M. Bernadro, F. J. Girón, E. Moreno, N. Keiding, D. V. Lindley, L. R. Pericchi, L. Piccinato, N. Reid, and N. Wermuth. The relation between theory and application in statistics. TEST, 4(2):207–261, December 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562627.

Corpas-Burgos:2019:CHH

[CBBRMB19] F. Corpas-Burgos, P. Botella-Rocamora, and M. A. Martinez-Beneito. On the convenience of heteroscedasticity in highly multivariate disease mapping. TEST, 28(4):1229– 1250, December 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00628-8.

Crommen:2024:IVA

[CBV24] Gilles Crommen, Jad Beyhum, and Ingrid Van Keilegom. An instrumental variable approach under dependent censoring. TEST, 33(2):473–495, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00903-9.

Cabras:2021:GFM

[CCR21] Stefano Cabras, María Eugenia Castellanos, and Oliver Ratmann. Goodness of fit for models with intractable likelihood. TEST, 30(3):713–736, September 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00747-7.

Chacon:2010:MPB

[CD10] J. E. Chacón and T. Duong. Multivariate plug-in bandwidth selection with unconstrained pilot bandwidth matrices. TEST, 19(2):375–398, August 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0168-4.

Consonni:2016:OBM

[CD16] Guido Consonni and Laura Deldossi. Objective Bayesian model discrimination in follow-up experimental designs. TEST, 25(3):397–412, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-015-0461-
3.

Comas:2011:SOA

[CDM11] C. Comas, P. Delicado, and J. Mateu. A second order approach to analyse spatial point patterns with functional marks. TEST, 20(3):503–523, November 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0215-1.

Cao:2007:PTP

[CdU07] Ricardo Cao and Ignacio López de Ullibarri. Producttype and presmoothed hazard rate estimators with censored data. TEST, 16(2):355–382, August 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0014-x.

Chen:2024:JEL

[CDZ24] Da Chen, Linlin Dai, and Yichuan Zhao. Jackknife empirical likelihood for the correlation coefficient with additive distortion measurement errors. TEST, 33(4):1129–1159, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-024-00941-x.

Ceyhan:2014:CRD

[Cey14] Elvan Ceyhan. Comparison of relative density of two random geometric digraph families in testing spatial clustering. TEST, 23(1):100–134, March 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0344-4.

Cha:2020:PRA

[CF20] Ji Hwan Cha and Maxim Finkelstein. Is perfect repair always perfect? TEST, 29(1):90–104, March 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00645-7.

Cholaquidis:2023:LSD

[CFM23] A. Cholaquidis, R. Fraiman, and L. Moreno. Level sets of depth measures in abstract spaces. TEST, 32(3):942–957, September 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00858-x.

Cholaquidis:2024:SAM

[CFMPL24] Alejandro Cholaquidis, Ricardo Fraiman, Leonardo Moreno, and Beatriz Pateiro-López. Statistical analysis of measures of non-convexity. TEST, 33(1):180–203, March 2024.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00889-4.

Casella:1996:SIM

[CFP^{+}96] George Casella, Juan Ferrándiz, Daniel Peña, David Rios Insua, José M. Bernardo, P. A. García-López, A. González, J. Berger, A. P. Dawid, Thomas J. Diciccio, Martin T. Wells, Paul Gustafson, Larry Wasserman, Edward I. George, Jun S. Liu, and et al. Statistical inference and Monte Carlo algorithms. TEST, 5(2):249–344, December 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562621.

Croux:2010:FRE

[CFRG10] Christophe Croux, Mohammed Fekri, and Anne Ruiz-Gazen. Fast and robust estimation of the multivariate errors in variables model. TEST, 19(2):286–303, August 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0155-9.

Cholaquidis:2020:SSL

[CFS20] A. Cholaquidis, R. Fraiman, and M. Sued. On semisupervised learning. TEST, 29(4):914–937, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-019-00690-2.

Caeiro:2002:CAU

[CG02] Frederico Caeiro and M. Ivette Gomes. A class of asymptotically unbiased semi-parametric estimators of the tail index. TEST, 11(2):345–364, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595711.

Caeiro:2009:SPS

[CG09] Frederico Caeiro and M. Ivette Gomes. Semi-parametric second-order reduced-bias high quantile estimation. TEST, 18(2):392–413, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0108-8.

Castruccio:2015:CCS

[CG15] Stefano Castruccio and Marc G. Genton. Comments on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1):31–34, March 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0385-3;   http://link.springer.com/
content/pdf/10.1007/s11749-014-0385-3.pdf. See [BCS15a]

and rejoinder [BCS15b].

Colombi:2019:LBT

[CG19] Roberto Colombi and Sabrina Giordano. Likelihood-based tests for a class of misspecified finite mixture models for ordinal categorical data. TEST, 28(4):1175–1202, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00626-w.

Cardot:2017:FEM

[CGB17] Hervé Cardot and Antoine Godichon-Baggioni. Fast estimation of the median covariation matrix with application to online robust principal components analysis. TEST, 26(3):461– 480, September 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0519-x.

Comte:2011:MKF

[CGCK11] Fabienne Comte, Valentine Genon-Catalot, and Mathieu Kessler. Multiplicative Kalman filtering. TEST, 20(2):389– 411, August 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0208-0.

Consonni:2008:CPB

[CGPV08] Guido Consonni, Eduardo Gutiérrez-Peña, and Piero Veronese. Compatible priors for Bayesian model comparison with an application to the Hardy–Weinberg equilibrium model. TEST, 17(3):585–605, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0057-7.

Colubi:2010:CDR

[CGR10] Ana Colubi and Gil González-Rodríguez. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):34–36, May 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0179-1.

Cai:2023:RSI

[CGX23a] T. Tony Cai, Zijian Guo, and Yin Xia. Rejoinder on: Statistical inference and large-scale multiple testing for highdimensional regression models. TEST, 32(4):1187–1194, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00911-9. See [CGX23b].

Cai:2023:SIL

[CGX23b] T. Tony Cai, Zijian Guo, and Yin Xia. Statistical inference and large-scale multiple testing for high-dimensional regression models. TEST, 32(4):1135–1171, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00870-1. See comments [TF23, CJ23,

Rit23, dUÁ23] and rejoinder [CGX23a].

Claeskens:2009:GFT

[CH09a] Gerda Claeskens and Jeffrey D. Hart. Goodness-of-fit tests in mixed models. TEST, 18(2):213–239, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-009-0148-8. See comments [Cab09, Mol09, JN09, MK09, Tha09] and rejoinder [CH09b].

Claeskens:2009:RGF

[CH09b] Gerda Claeskens and Jeffrey D. Hart. Rejoinder on: “Goodness-of-fit tests in mixed models”. TEST, 18(2): 265–270, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0154-x. See

[CH09a, Cab09, Mol09, JN09, MK09, Tha09] and rejoinder [CH09a].

Chan:1995:UET

[Cha95] Wai-Sum Chan. Understanding the effect of time series outliers on sample autocorrelations. TEST, 4(1):179–186, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02563108.

Chatterjee:2017:CHD

[Cha17] Arindam Chatterjee. Comments on: “High-dimensional simultaneous inference with the bootstrap”. TEST, 26(4):729– 730, December 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0557-z;   http://link.
springer.com/content/pdf/10.1007/s11749-017-0557-

z.pdf. See [DBZ17a] and rejoinder [DBZ17b].

Chesneau:2007:MAG

[Che07] Christophe Chesneau. A maxiset approach of a Gaussian noise model. TEST, 16(3):523–546, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0018-6.

Chown:2020:NLS

[CHV20] Justin Chown, Cédric Heuchenne, and Ingrid Van Keilegom. The nonparametric location-scale mixture cure model. TEST, 29(4):1008–1028, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-019-00698-8.

Cramer:2010:APT

[CI10] Erhard Cramer and George Iliopoulos. Adaptive progressive Type–II censoring. TEST, 19(2):342–358, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0167-5.

Cano:2018:IPB

[CIS18] J. A. Cano, M. Iniesta, and D. Salmerón. Integral priors for Bayesian model selection: how they operate from simple to complex cases. TEST, 27(4):968–987, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0579-1.

Cizek:2013:RLT

[Cíz13] Pavel Cízek. Reweighted least trimmed squares: an alternative to one-step estimators. TEST, 22(3):514–533, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0335-5.

Claeskens:2023:CSI

[CJ23] Gerda Claeskens and Maarten Jansen. Comments on: Statistical inference and large-scale multiple testing for highdimensional regression models. TEST, 32(4):1177–1179, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00896-5. See [CGX23b].

Cai:2024:OEE

[CJMW24] Li Cai, Lei Jin, Jiuzhou Miao, and Suojin Wang. Oracleefficient M-estimation for single-index models with a smooth simultaneous confidence band. TEST, 33(4):1041–1061, December 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00935-9.

Cao:2005:RHR

[CJV05] Ricardo Cao, Paul Janssen, and Noël Veraverbeke. Relative hazard rate estimation for right censored and left truncated data. TEST, 14(1):257–280, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595406.

Cano:2004:IPN

[CKM04] Juan Antonio Cano, Mathieu Kessler, and Elías Moreno. On intrinsic priors for nonnested models. TEST, 13(2):445–463, December 2004. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595781.

Casalis:1994:CJS

[CL94] M. Casalis and G. Letac. Characterization of the Jorgensen set in generalized linear models. TEST, 3(1):145–162, June 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562678.

Cheung:2010:BVE

[CL10] K. Y. Cheung and Stephen M. S. Lee. Bootstrap variance estimation for Nadaraya quantile estimator. TEST, 19(1): 131–145, May 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0137-y;   http://link.
springer.com/content/pdf/10.1007/s11749-009-0137-
y.pdf.

Chagny:2015:OAE

[CL15] Gaëlle Chagny and Claire Lacour. Optimal adaptive estimation of the relative density. TEST, 24(3):605–631, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0426-6.

Cerqueti:2023:STB

[CL23] Roy Cerqueti and Claudio Lupi. Severe testing of Benford’s law. TEST, 32(2):677–694, June 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00848-z.

Clemen:2002:ICS

[Cle02] Robert T. Clemen. Incentive contrats and strictly proper scoring rules. TEST, 11(1):167–189, June 2002. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595735.

Cai:2020:OEE

[CLH^{+}20] Li Cai, Lisha Li, Simin Huang, Liang Ma, and Lijian Yang. Oracally efficient estimation for dense functional data with holiday effects. TEST, 29(1):282–306, March 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00655-5.

Carroll:2007:CNI

[CM07] Raymond J. Carroll and Arnab Maity. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):456–458, December 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-007-0085-3. See [FJ07a] and rejoinder [FJ07b].

Chattopadhyay:2010:PCU

[CM10] Gopaldeb Chattopadhyay and Indranil Mukhopadhyay. Progressive censoring under inverse sampling for nonparametric multi-sample location problem. TEST, 19(2):325–341, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0157-7.

Chang:2012:AFV

[CM12] In Hong Chang and Rahul Mukerjee. On the approximate frequentist validity of the posterior quantiles of a parametric function: results based on empirical and related likelihoods. TEST, 21(1):156–169, March 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0240-8.

Cuparic:2022:NCB

[CM22] Marija Cuparić and Bojana Milosević. New characterizationbased exponentiality tests for randomly censored data. TEST, 31(2):461–487, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00787-7.

Castillo-Mateo:2024:BJQ

[CMGA^{+}24] Jorge Castillo-Mateo, Alan E. Gelfand, Jesús Asín, Ana C. Cebrián, and Jesús Abaurrea. Bayesian joint quantile autoregression. TEST, 33(1):335–357, March 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00895-6.

Castano-Martinez:2005:DSW

[CMLB05] Antonia Castaño-Martínez and Fernando López-Blázquez. Distribution of a sum of weighted noncentral chi-square variables. TEST, 14(2):397–415, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595410.

Claeskens:2008:OST

[CNJ08] Gerda Claeskens, Rosemary Nguti, and Paul Janssen. Onesided tests in shared frailty models. TEST, 17(1):69–82, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0023-9.

Croux:2015:CRE

[CÖ15] Christophe Croux and Viktoria Öllerer. Comments on: “‘Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24 (3):462–466, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0455-1;
http://link.springer.com/content/pdf/10.1007/s11749-

015-0455-1.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Carota:1998:DPE

[CP98] Cinzia Carota and Giovanni Parmigiani. A Dirichlet process elaboration diagnostic for binomial goodness of fit. TEST, 7(1):133–145, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02565106.

Cadre:2012:HFC

[CP12] Benoît Cadre and Quentin Paris. On Hölder fields clustering. TEST, 21(2):301–316, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0244-4.

Cleanthous:2021:RAG

[CPW21] Galatia Cleanthous, Emilio Porcu, and Philip White. Regularity and approximation of Gaussian random fields evolving temporally over compact two-point homogeneous spaces. TEST, 30(4):836–860, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
021-00755-1.

Chan:2012:IET

[CPZ12] Ngai Hang Chan, Liang Peng, and Rongmao Zhang. Interval estimation of the tail index of a GARCH(1,1) model. TEST, 21(3):546–565, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0264-0.

Cabana:2005:UEM

[CQ05] Alejandra Cabaña and Adolfo J. Quiroz. Using the empirical moment generating function in testing for the Weibull and the type i extreme value distributions. TEST, 14(2):417– 431, December 2005. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595411.

Carota:1994:RBA

[CR94] Cinzia Carota and Fabrizio Ruggeri. Robust Bayesian analysis given priors on partition sets. TEST, 3(2):73–86, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562694.

Cipra:1997:KFO

[CR97] T. Cipra and R. Romera. Kalman filter with outliers and missing observations. TEST, 6(2):379–395, December 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02564705.

Coelho:2017:THB

[CR17] Carlos A. Coelho and Anuradha Roy. Testing the hypothesis of a block compound symmetric covariance matrix for elliptically contoured distributions. TEST, 26(2):308–330, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0512-4. See correction [CR22].

Coelho:2022:CTH

[CR22] Carlos A. Coelho and Anuradha Roy. Correction to: Testing the hypothesis of a block compound symmetric covariance matrix for elliptically contoured distributions. TEST, 31(3):877–878, September 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00796-6. See [CR17].

Cousido-Rocha:2019:TEL

[CRdUÁH19] Marta Cousido-Rocha, Jacobo de Uña-Álvarez, and Jeffrey D. Hart. Testing equality of a large number of densities under mixing conditions. TEST, 28(4):1203–1228, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-00625-3.

Crujeiras:2010:CGS

[Cru10] Rosa M. Crujeiras. Comments on: “A general sciencebased framework for nonlinear spatio-temporal dynamical models”. TEST, 19(3):456–458, November 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-010-0211-5. See [WH10b] and rejoinder [WH10b].

Cabrera:2012:STC

[CRV12] Manuel Ordóñez Cabrera, Andrew Rosalsky, and Andrei Volodin. Some theorems on conditional mean convergence and conditional almost sure convergence for randomly weighted sums of dependent random variables. TEST, 21(2): 369–385, June 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0248-0.

Cabrera:2021:CCE

[CRV21a] Manuel Ordóñez Cabrera, Andrew Rosalsky, and Andrei Volodin. Correction to: On the concept of B-statistical uniform integrability of weighted sums of random variables and the law of large numbers with mean convergence in the statistical sense. TEST, 30(4):1076–1077, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00793-9. See [CRV21b].

Cabrera:2021:CEB

[CRV21b] Manuel Ordóñez Cabrera, Andrew Rosalsky, and Andrei Volodin. On the concept of B-statistical uniform integrability of weighted sums of random variables and the law of large numbers with mean convergence in the statistical sense. TEST, 30(1):83–102, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00706-2. See correction [CRV21a].

Chen:1997:PSM

[CS97a] Ming-Hui Chen and Qi-Man Shao. Performance study of marginal posterior density estimation via Kullback–Leibler divergence. TEST, 6(2):321–350, December 1997. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02564702.

Choy:1997:HMS

[CS97b] S. T. B. Choy and A. F. M. Smith. Hierarchical models with scale mixtures of normal distributions. TEST, 6(1): 205–221, June 1997. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02564434.

Castillo:2007:CPC

[CS07] Enrique Castillo and José María Sarabia. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):265–267, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0064-8. See

[Bal07a] and rejoinder [Bal07b].

Canale:2016:BNL

[CS16a] Antonio Canale and Bruno Scarpa. Bayesian nonparametric location-scale-shape mixtures. TEST, 25(1):113–130, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0446-2.

Chen:2016:SPE

[CS16b] Ming Chen and Qiongxia Song. Semi-parametric estimation and forecasting for exogenous log-GARCH models. TEST, 25(1):93–112, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0442-6.

Chacon:2023:BHT

[CS23] José E. Chacón and Javier Fernández Serrano. Bump hunting through density curvature features. TEST, 32(4):1251–1275, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00872-z.

Cai:2022:ABL

[CSL22] Kaida Cai, Hua Shen, and Xuewen Lu. Adaptive bi-level variable selection for multivariate failure time model with a diverging number of covariates. TEST, 31(4):968–993, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00809-y.

Crujeiras:2021:CRA

[CSN21] Rosa M. Crujeiras and Paula Saavedra-Nieves. Comments on: Recent advances in directional statistics. TEST, 30(1): 64–67, March 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00761-3. See [PGP21a,

PGP21b].

Csorgo:2002:WCT

[Csö02] Sándor Csörgö. Weighted correlation tests for scale families. TEST, 11(1):219–248, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595737.

Cano:2008:IES

[CSR08] J. A. Cano, D. Salmerón, and C. P. Robert. Integral equation solutions as prior distributions for Bayesian model selection. TEST, 17(3):493–504, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0040-8.

Comte:2018:HEC

[CSS18] Fabienne Comte, Adeline Samson, and Julien J. Stirnemann. Hazard estimation with censoring and measurement error: application to length of pregnancy. TEST, 27(2):338–359, June 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-017-0548-0.

Cao:2012:NSD

[CTC12] Min Cao, Allen H. Tai, and Ling-Yau Chan. New statistical distributions for group counting in Bernoulli and Poisson processes. TEST, 21(1):29–53, March 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0232-0.

Chen:2009:RRE

[CV09a] Song Xi Chen and Ingrid Van Keilegom. Rejoinder on: “A review on empirical likelihood methods for regression”. TEST, 18(3):468–474, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0166-6. See

[CV09b, Spe09, PZ09, Vel09, SS09, McK09, LL09].

Chen:2009:REL

[CV09b] Song Xi Chen and Ingrid Van Keilegom. A review on empirical likelihood methods for regression. TEST, 18(3): 415–447, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0159-5. See

[Spe09, PZ09, Vel09, SS09, McK09, LL09] and [CV09a].

Colling:2016:GFT

[CV16] Benjamin Colling and Ingrid Van Keilegom. Goodness-of-fit tests in semiparametric transformation models. TEST, 25(2): 291–308, June 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0448-0.

Calvo:2002:BMM

[CVO02] Miquel Calvo, Angel Villarroya, and Josep M. Oller. A biplot method for multivariate normal populations with unequal covariance matrices. TEST, 11(1):143–165, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595734.

Chen:2023:IEP

[CW23] Zijuan Chen and Suojin Wang. Inferences for extended partially linear single-index models. TEST, 32(2):602–622, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00845-8.

Cook:1993:ERS

[CWB^{+}93] R. Dennis Cook, Nate Wetzel, José D. Bermúdez, J. De La Horra, Frank Critchley, Michael Lavine, Ker-Chau Li, R. E. McCulloch, Sally C. Morton, X. Shen, Sanford Weisberg, and S. Zacks. Exploring regression structure with graphics. TEST, 2(1–2):33–100, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562669.

Cui:2021:CAM

[CWZ21] Tingting Cui, Pengfei Wang, and Wensheng Zhu. Covariateadjusted multiple testing in genome-wide association studies via factorial hidden Markov models. TEST, 30(3):737–757, September 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00746-8.

Cai:2015:SSC

[CY15] Li Cai and Lijian Yang. A smooth simultaneous confidence band for conditional variance function. TEST, 24(3):632– 655, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0427-5.

Cui:2018:NBI

[CZ18] Yan Cui and Fukang Zhu. A new bivariate integer-valued GARCH model allowing for negative cross-correlation. TEST, 27(2):428–452, June 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0552-4.

Dagne:2001:BTM

[Dag01] Getachew A. Dagne. Bayesian transformed models for small area estimation. TEST, 10(2):375–391, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595703.

delMarRueda:2003:DEQ

[dAM03] M. del Mar Rueda, Antonio Arcos, and M. Dolores Martínez. Difference estimators of quantiles in finite populations. TEST, 12(2):481–496, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595726.

Dawid:2008:CAP

[Daw08] A. Philip Dawid. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):243– 244, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0118-6.

delBarrio:2010:CPM

[dB10] Eustasio del Barrio. Comments on: “l_{1}-penalization for mixture regression models”. TEST, 19(2):276–279, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0202-6.

Devroye:1997:USF

[DBC^{+}97] Duc Devroye, J. Beirlant, R. Cao, R. Fraiman, P. Hall, M. C. Jones, Gábor Lugosi, E. Mammen, J. S. Marron, C. Sánchez-Sellero, J. de Uña, F. Udina, and L. Devroye. Universal smoothing factor selection in density estimation: theory and practice. TEST, 6(2):223–320, December 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02564701.

delBarrio:2000:CEQ

[dBCAM^{+}00] Eustasio del Barrio, Juan A. Cuesta-Albertos, Carlos Matrán, Sándor Csörgö, Carles M. Cuadras, Tertius de Wet, Evarist Giné, Richard Lockhart, Axel Munk, and Winfried Stute. Contributions of empirical and quantile processes to the asymptotic theory of goodness-of-fit tests. TEST, 9(1): 1–96, June 2000. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595852.

delBarrio:2020:AVM

[dBIM20] E. del Barrio, H. Inouzhe, and C. Matrán. On approximate validation of models: a Kolmogorov–Smirnov-based approach. TEST, 29(4):938–965, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-019-00691-1.

delBarrio:2009:LIB

[dBJM09] Eustasio del Barrio, Arnold Janssen, and Carlos Matrán. On the low intensity bootstrap for triangular arrays of independent identically distributed random variables. TEST, 18(2):283–301, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0077-3.

Dezeure:2017:HDS

[DBZ17a] Ruben Dezeure, Peter Bühlmann, and Cun-Hui Zhang. High-dimensional simultaneous inference with the bootstrap. TEST, 26(4):685–719, December 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-017-0554-2. See comments [BZ17, Cha17, LN17, LS17, LY17] and rejoinder [DBZ17b].

Dezeure:2017:RHD

[DBZ17b] Ruben Dezeure, Peter Bühlmann, and Cun-Hui Zhang. Rejoinder on: “High-dimensional simultaneous inference with the bootstrap”. TEST, 26(4):751–758, December 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-017-0560-4. See [DBZ17a, BZ17, Cha17, LN17, LS17, LY17].

Dean:2011:CNI

[DC11] C. B. Dean and E. Juarez Colunga. Comments on: “Nonparametric inference based on panel count data”. TEST, 20 (1):43–45, May 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0229-8. See [ZBS11a] and

rejoinder [ZBS11b].

Ceregatti:2021:WGB

[dCCIS21] Rafael de Carvalho Ceregatti, Rafael Izbicki, and Luis Ernesto Bueno Salasar. WIKS: a general Bayesian nonparametric index for quantifying differences between two populations. TEST, 30(1):274–291, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00718-y.

Dawid:1995:CCE

[DDM^{+}95] A. P. Dawid, M. H. DeGroot, J. Mortera, R. Cooke, S. French, C. Genest, M. J. Schervish, D. V. Lindley, K. J. McConway, and R. L. Winkler. Coherent combination of experts’ opinions. TEST, 4(2):263–313, December 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562628.

Dauxois:2006:BCB

[DDP06] Jean-Yves Dauxois, Pierre Druilhet, and Denys Pommeret. A Bayesian choice between Poisson, binomial and negative binomial models. TEST, 15(2):423–432, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02607060.

DeSantis:2007:ABF

[De 07] Fulvio De Santis. Alternative Bayes factors: Sample size determination and discriminatory power assessment. TEST, 16(3):504–522, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0017-7.

DeGooijer:2008:PSL

[De 08] Jan G. De Gooijer. Partial sums of lagged cross-products of AR residuals and a test for white noise. TEST, 17(3):567– 584, November 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0058-6.

DEpifanio:1996:NRP

[D’E96] G. D’Epifanio. Notes on a recursive procedure for point estimation. TEST, 5(1):203–225, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562689.

Dehling:2014:CES

[Deh14] Herold Dehling. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):261– 264, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0376-4;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0376-
4.pdf.

delMarRueda:2019:CDS

[del19a] Maria del Mar Rueda. Comments on: Deville and Särndal’s Calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1077–1081, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00683-1. See [DT19a, DT19b].

Delicado:2019:CDS

[Del19b] Pedro Delicado. Comments on: Data science, big data and statistics. TEST, 28(2):334–337, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00639-5. See [GP19a, GP19b].

Delicado:2024:CSB

[Del24] Pedro Delicado. Comments on: Shape-based functional data analysis. TEST, 33(1):62–65, March 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00904-8. See [WHS24b, WHS24a].

Dembinska:2007:CPC

[Dem07] Anna Dembińska. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):262–264, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0063-9. See [Bal07a] and rejoinder

[Bal07b].

Dette:2013:CUR

[Det13] Holger Dette. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22 (3):437–441, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0333-7;
http://link.springer.com/content/pdf/10.1007/s11749-
013-0333-7.pdf.

Dehghan:2019:AID

[DF19] Sakineh Dehghan and Mohammad Reza Faridrohani. Affine invariant depth-based tests for the multivariate one-sample location problem. TEST, 28(3):671–693, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0593-3.

Drton:2012:CSR

[DFK12] Mathias Drton, Chris Fox, and Andreas Käufl. Comments on: “Sequences of regressions and their independencies”. TEST, 21(2):255–261, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0285-3. See

[WS12b] and rejoinder [WS12a].

Ditzhaus:2021:QQB

[DFP21] Marc Ditzhaus, Roland Fried, and Markus Pauly. QANOVA: quantile-based permutation methods for general factorial designs. TEST, 30(4):960–979, December 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00758-y.

DiMarzio:2018:CLL

[DFPT18] Marco Di Marzio, Stefania Fensore, Agnese Panzera, and Charles C. Taylor. Circular local likelihood. TEST, 27 (4):921–945, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0576-9;
http://link.springer.com/content/pdf/10.1007/s11749-
017-0576-9.pdf.

Datta:1995:NPM

[DG95] G. S. Datta and J. K. Ghosh. Noninformative priors for maximal invariant parameter in group models. TEST, 4(1): 95–114, June 1995. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02563105.

Ditzhaus:2022:TMH

[DG22] Marc Ditzhaus and Daniel Gaigall. Testing marginal homogeneity in Hilbert spaces with applications to stock market returns. TEST, 31(3):749–770, September 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00802-5.

Dierckx:2014:LRA

[DGG14] Goedele Dierckx, Yuri Goegebeur, and Armelle Guillou. Local robust and asymptotically unbiased estimation of conditional Pareto-type tails. TEST, 23(2):330–355, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0350-6.

Diebolt:2008:BRE

[DGGG08] Jean Diebolt, Laurent Gardes, Stéphane Girard, and Armelle Guillou. Bias-reduced estimators of the Weibull tailcoefficient. TEST, 17(2):311–331, August 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0034-6.

Diaz-Garcia:2001:EVZ

[DGGJ01] José A. Díaz-García and Ramón Gutiérrez-Jaimez. The expected value of zonal polynomials. TEST, 10(1):133–145, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595828.

Diaz-Garcia:2005:FSR

[DGGJ05] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Functions of singular random matrices with applications. TEST, 14(2):475–487, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595414.

Daouia:2011:KEE

[DGGL11] Abdelaati Daouia, Laurent Gardes, Stéphane Girard, and Alexandre Lekina. Kernel estimators of extreme level curves. TEST, 20(2):311–333, August 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0196-0.

Datta:2011:BBA

[DGSM11] G. S. Datta, M. Ghosh, R. Steorts, and J. Maples. Bayesian benchmarking with applications to small area estimation. TEST, 20(3):574–588, November 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0218-y.

Dey:1998:SIA

[DGSV98] Dipak K. Dey, Alan E. Gelfand, Tim B. Swartz, and Pantelis K. Vlachos. A simulation-intensive approach for checking hierarchical models. TEST, 7(2):325–346, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02565116.

Debicki:2015:EOS

[DHJL15] Krzysztof Dȩbicki, Enkelejd Hashorva, Lanpeng Ji, and Chengxiu Ling. Extremes of order statistics of stationary processes. TEST, 24(2):229–248, June 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0404-4.

DiCecco:2012:CET

[Di 12] Davide Di Cecco. Conditional exact tests for Markovianity and reversibility in multiple categorical sequences. TEST, 21(1):170–187, March 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0241-7.

Dembinska:2021:MLE

[DJ21] Anna Dembińska and Krzysztof Jasiński. Maximum likelihood estimators based on discrete component lifetimes of a k-out-of-n system. TEST, 30(2):407–428, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00724-0.

Diop:2023:GPC

[DK23] Mamadou Lamine Diop and William Kengne. A general procedure for change-point detection in multivariate time series. TEST, 32(1):1–33, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00824-z.

Datta:2011:EMS

[DKMR11] Gauri Sankar Datta, Tatsuya Kubokawa, Isabel Molina, and J. N. K. Rao. Estimation of mean squared error of modelbased small area estimators. TEST, 20(2):367–388, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0206-2.

Devroye:2004:WSM

[DL04] Luc Devroye and Gábor Lugosi. Bin width selection in multivariate histograms by the combinatorial method. TEST, 13(1):129–145, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02603004.

delaHorra:1992:UPM

[dlH92] Julián de la Horra. Using the prior mean of a nuisance parameter. TEST, 1(1):31–38, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562659.

delaHorra:2001:PPP

[dlHRB01] Julián de la Horra and M. Teresa Rodriguez-Bernal. Posterior predictive p-values: what they are and what they are not.

TEST, 10(1):75–86, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595824.

deLima:2021:EBP

[dLNA21] W. V. Félix de Lima, A. D. C. Nascimento, and G. J. A. Amaral. Entropy-based pivotal statistics for multi-sample problems in planar shape. TEST, 30(1):153–178, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00707-1.

Du:2024:EHT

[DLZ24] Bin Du, Xiumin Liu, and Junlong Zhao. Extended Hotelling T^{2} test in distributed frameworks. TEST, 33(4):1160–1179, December 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00939-5.

Fernandez-de-Marcos:2023:NOT

[dMGP23] Alberto Fernández de Marcos and Eduardo García-Portugués. On new omnibus tests of uniformity on the hypersphere. TEST, 32(4):1508–1529, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
023-00882-x.

DeBastiani:2015:IDE

[DMUOG15] Fernanda De Bastiani, Audrey Helen Mariz de Aquino Cysneiros, Miguel Angel Uribe-Opazo, and Manuel Galea. Influence diagnostics in elliptical spatial linear models. TEST, 24(2):322–340, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0409-z.

doNascimento:2016:TVE

[dNGL16] Fernando Ferraz do Nascimento, Dani Gamerman, and Hedibert Freitas Lopes. Time-varying extreme pattern with dynamic models. TEST, 25(1):131–149, March 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0444-4.

Dauxois:2007:ICA

[DNR07] Jacques Dauxois, Guy Martial Nkiet, and Yves Romain. On invariance in canonical analysis with applications to covariate discriminant analysis. TEST, 16(2):314–332, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0011-0.

Dolado:2012:CSR

[Dol12] Juan J. Dolado. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):442–446, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0303-5. See [Tjø12b] and

rejoinder [Tjø12a].

Deldossi:2019:ODD

[DOT19] Laura Deldossi, Silvia Angela Osmetti, and Chiara Tommasi. Optimal design to discriminate between rival copula models for a bivariate binary response. TEST, 28(1):147– 165, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0595-1.

Doukhan:2012:CSR

[Dou12] Paul Doukhan. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):447–450, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0298-y. See [Tjø12b] and

rejoinder [Tjø12a].

Dobler:2018:BPB

[DP18]  Dennis Dobler and Markus Pauly.  Bootstrap- and

permutation-based inference for the Mann–Whitney effect for right-censored and tied data. TEST, 27(3):639–658, September 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0565-z.

Delicado:2023:UCP

[DP23] Pedro Delicado and Daniel Peña. Understanding complex predictive models with ghost variables. TEST, 32(1):107– 145, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00826-x.

Doukhan:2011:RSW

[DPR11a] Paul Doukhan, Silika Prohl, and Christian Y. Robert. Rejoinder on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20(3):499–502, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0275-x. See [DPR11a, Vel11, Bra11,

Ber11, LM11, Pap11].

Doukhan:2011:SWD

[DPR11b] Paul Doukhan, Silika Prohl, and Christian Y. Robert. Subsampling weakly dependent time series and application to extremes. TEST, 20(3):447–479, November 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-011-0269-8. See comments [Vel11, Bra11, Ber11, LM11, Pap11] and rejoinder [DPR11a].

DeLaHorra:1999:PPV

[DRB99] Julián De La Horra and María Teresa Rodríguez-Bernal. The posterior predictive p-value for the problem of goodness of fit. TEST, 8(1):117–128, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595865.

delRio:1992:LCV

[dRF92] A. Quintela del Río and J. M. Vilar Fernández. A local crossvalidation algorithm for dependent data. TEST, 1(1):123– 153, December 1992. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562667.

Dryden:2024:CSB

[Dry24] Ian L. Dryden. Comments on: Shape-based functional data analysis. TEST, 33(1):59–61, March 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00907-5. See [WHS24b, WHS24a].

Dube:1995:PEI

[DS95] M. Dube and V. Singh. Predictive efficiency of improved estimators in restricted regression models. TEST, 4(2):323– 331, December 1995. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562630.

daSilva:2018:RCM

[dS18] João Lita da Silva. On the rates of convergence for moments convergence in regression models. TEST, 27(2):477– 495, June 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0561-3.

Devaud:2019:DSC

[DT19a] Denis Devaud and Yves Tillé. Deville and Särndal’s calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1033–1065, December 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-019-00681-3. See comments [Kot19, Mor19, BR19, del19a, WZ19] and rejoinder [DT19b].

Devaud:2019:RDS

[DT19b] Denis Devaud and Yves Tillé. Rejoinder on: Deville and Särndal’s calibration: revisiting a 25-year-old successful optimization problem. TEST, 28(4):1087–1091, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00685-z. See [DT19a, Kot19, Mor19,

BR19, del19a, WZ19].

deUna-Alvarez:2002:PLE

[dUÁ02] Jacobo de Uña-Álvarez. Product-limit estimation for lengthbiased censored data. TEST, 11(1):109–125, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595732.

deUna-Alvarez:2011:CNI

[dUÁ11] Jacobo de Uña-Álvarez. Comments on: “Nonparametric inference based on panel count data”. TEST, 20(1):62–64, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0225-z. See [ZBS11a] and rejoinder

[ZBS11b].

deUna-Alvarez:2013:CUR

[dUÁ13] Jacobo de Uña-Álvarez. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22 (3):414–418, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0329-3;
http://link.springer.com/content/pdf/10.1007/s11749-
013-0329-3.pdf.

deUna-Alvarez:2023:CSI

[dUÁ23] Jacobo de Uña-Álvarez. Comments on: Statistical inference and large-scale multiple testing for high-dimensional regression models. TEST, 32(4):1184–1186, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00908-4. See [CGX23b].

deUna-Alvarez:2013:GCG

[dUÁV13] Jacobo de Uña-Álvarez and Noël Veraverbeke. Generalized copula-graphic estimator. TEST, 22(2):343–360, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0314-2.

Dube:1999:MRE

[Dub99] Madhulika Dube. Mixed regression estimator under inclusion of some superfluous variables. TEST, 8(2):411–417, December 1999. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595878.

Ducharme:2001:GFT

[Duc01] Gilles R. Ducharme. Goodness-of-fit tests for the inverse Gaussian and related distributions. TEST, 10(2):271–290, December 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595697.

Duchesne:2005:ADR

[Duc05] Pierre Duchesne. On the asymptotic distribution of residual autocovariances in VARX models with applications. TEST, 14(2):449–473, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595413.

Delhelle:2025:CBD

[DV25] Morine Delhelle and Ingrid Van Keilegom. Copula based dependent censoring in cure models. TEST, 34(2):325–360, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-024-00961-7.

deWet:2002:GFT

[dW02] T. de Wet. Goodness-of-fit tests for location and scale families based on a weighted L_{2}–Wasserstein distance measure. TEST, 11(1):89–107, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595731.

Daniels:2009:CMD

[DW09] Michael J. Daniels and Chenguang Wang. Comments on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):51–58, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0141-2. See

[IM09a] and rejoinder [IM09b].

Bermudez:2003:BAP

[dZBT03] P. de Zea Bermudez and M. A. Amaral Turkman. Bayesian approach to parameter estimation of the generalized Pareto distribution. TEST, 12(1):259–277, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595822.

Escobar-Bach:2017:BCR

[EBGGY17] Mikael Escobar-Bach, Yuri Goegebeur, Armelle Guillou, and Alexandre You. Bias-corrected and robust estimation of the bivariate stable tail dependence function. TEST, 26(2):284– 307, June 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0511-5.

Espinheira:2020:RIA

[EdOS20] Patrícia L. Espinheira and Alisson de Oliveira Silva. Residual and influence analysis to a general class of simplex regression. TEST, 29(2):523–552, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00665-3.

Espejo:1992:ESP

[EE92] Mariano Ruiz Espejo and Mercedes Ruiz Espejo. Equilibrated strategy for population variance estimation. TEST, 1(1):79–91, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562664.

Efromovich:2007:CNI

[Efr07] Sam Efromovich. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):465– 467, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0088-0. See [FJ07a] and re-

joinder [FJ07b].

Einmahl:2012:TBS

[EG12] John H. J. Einmahl and Maria Gantner. Testing for bivariate spherical symmetry. TEST, 21(1):54–73, March 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0235-5.

Elfadaly:2013:EDC

[EG13] Fadlalla G. Elfadaly and Paul H. Garthwaite. Eliciting Dirichlet and Connor–Mosimann prior distributions for multinomial models. TEST, 22(4):628–646, November 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0336-4.

Embrechts:2011:CIM

[EH11] Paul Embrechts and Marius Hofert. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):263–270, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0252-4. See

[GNZ11a] and rejoinder [GNZ11b].

Ebner:2020:RTM

[EH20a] Bruno Ebner and Norbert Henze. Rejoinder on: Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):911–913, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00744-w. See [EH20b] and comments

[JG20, Mei20, Ric20, SR20].

Ebner:2020:TMN

[EH20b] Bruno Ebner and Norbert Henze. Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):845–892, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-020-00740-0. See comments [JG20, Mei20, Ric20, SR20] and rejoinder [EH20a].

Eilers:2020:CIC

[Eil20] Paul Eilers. Comments on: Inference and computation with Generalized Additive Models and their extensions. TEST, 29(2):340–342, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-020-00715-1. See

[Woo20a] and rejoinder [Woo20b].

Esteban:2023:SAE

[ELLV^{+}23] María Dolores Esteban, María José Lombardía, Esther López-Vizcaíno, Domingo Morales, and Agustín Pérez. Small area estimation of average compositions under multivariate nested error regression models. TEST, 32(2):651–676, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00847-0.

Espejo:2015:CMC

[ELORM15] Rosa M. Espejo, Nikolai N. Leonenko, Andriy Olenko, and María D. Ruiz-Medina. On a class of minimum contrast estimators for Gegenbauer random fields. TEST, 24(4):657– 680, December 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0428-4.

Esteban:2020:SAE

[ELP20] María Dolores Esteban, María José Lombardía, and Agustín Pérez. Small area estimation of proportions under area-level compositional mixed models. TEST, 29(3):793–818, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-019-00688-w. See correction [ELP21].

Esteban:2021:CSA

[ELP21] María Dolores Esteban, María José Lombardía, and Agustín Pérez. Correction to: Small area estimation of proportions under area-level compositional mixed models. TEST, 30(2): 527–528, June 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00767-x. See [ELP20].

Emura:2015:AES

[EM15] T. Emura and K. Murotani. An algorithm for estimating survival under a copula-based dependent truncation model. TEST, 24(4):734–751, December 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0432-8.

Eckardt:2021:SOL

[EM21] Matthias Eckardt and Jorge Mateu. Second-order and local characteristics of network intensity functions. TEST, 30(2): 318–340, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.springer.

com/article/10.1007/s11749-020-00720-4. See correc-

tion [EM22].

Eckardt:2022:CSO

[EM22] Matthias Eckardt and Jorge Mateu. Correction to: Secondorder and local characteristics of network intensity functions. TEST, 31(3):875–876, September 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00791-x. See [EM21].

Ekstrom:2020:CAE

[EMJ20] M. Ekström, S. M. Mirakhmedov, and S. Rao Jammalamadaka. A class of asymptotically efficient estimators based on sample spacings. TEST, 29(3):617–636, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-019-00637-7.

Elias:2025:SPP

[EN25] Antonio Elías and Stanislav Nagy. Statistical properties of partially observed integrated functional depths. TEST, 34(1):125–150, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00954-6.

Egozcue:2019:CDS

[EPG19a] Juan José Egozcue and Vera Pawlowsky-Glahn. Compositional data: the sample space and its structure. TEST, 28 (3):599–638, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00670-6. See

comments [FH19, Gre19, MF19] and rejoinder [EPG19b].

Egozcue:2019:RCD

[EPG19b] Juan José Egozcue and Vera Pawlowsky-Glahn. Rejoinder on: Compositional data: the sample space and its structure. TEST, 28(3):658–663, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00674-2. See [EPG19a, FH19, Gre19,

MF19].

Espejo:2006:PSM

[EPS06] M. Ruiz Espejo, M. Delgado Pineda, and H. P. Singh. Postgrouped sampling method of estimation. TEST, 15(1):209– 226, June 2006. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595425.

Eno:2000:BRP

[EY00] Daniel R. Eno and Keying Ye. Bayesian reference prior analysis for polynomial calibration models. TEST, 9(1):191–208, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595858.

Fan:2001:NBE

[Fan01] Tsai-Hung Fan. Noninformative Bayesian estimation for the optimum in a single factor quadratic response model. TEST, 10(2):225–240, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595694.

Farcomeni:2015:CRE

[Far15] Alessio Farcomeni. Comments on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):467–470, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0451-5;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0451-

5.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Febrero-Bande:2013:CMF

[FB13] Manuel Febrero-Bande. Comments on: “Model-free modelfitting and predictive distributions”. TEST, 22(2):224–226, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0321-y. See [Pol13a] and rejoinder

[Pol13b].

Febrero-Bande:2016:CPE

[FB16] Manuel Febrero-Bande. Comments on: “Probabilityenhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):35–40, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-015-0476-9;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0476-

9.pdf. See [YWZ16a] and rejoinder [YWZ16b].

Frostig:2022:TEM

[FB22] Tzviel Frostig and Yoav Benjamini. Testing the equality of multivariate means when p > n by combining the Hotelling and Simes tests. TEST, 31(2):390–415, June 2022.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00781-z.

Febrero-Bande:2013:GAM

[FBGM13] Manuel Febrero-Bande and Wenceslao González-Manteiga. Generalized additive models for functional data. TEST, 22(2):278–292, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0308-0.

Franco:2014:GMW

[FBKV14] Manuel Franco, Narayanaswamy Balakrishnan, Debasis Kundu, and Juana-María Vivo. Generalized mixtures of Weibull components. TEST, 23(3):515–535, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0362-x.

Finkelstein:2021:DBI

[FC21] Maxim Finkelstein and Ji Hwan Cha. On degradationbased imperfect repair and induced generalized renewal processes. TEST, 30(4):1026–1045, December 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00765-z.

Finkelstein:2022:RDA

[FC22] Maxim Finkelstein and Ji Hwan Cha. Reducing degradation and age of items in imperfect repair modeling. TEST, 31 (4):1058–1081, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00813-2.

Ferreira:1999:LDP

[Fer99] Helena Ferreira. Limit distributions for point processes of exceedances of random levels. TEST, 8(1):191–200, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595869.

Fernandez:2004:OTS

[Fer04] Arturo J. Fernández. One- and two-sample prediction based on doubly censored exponential data and prior information.

TEST, 13(2):403–416, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595779.

Fernandez:2002:LPR

[FF02] Juan M. Vilar Fernández and Mario Francisco Fernández. Local polynomial regression smoothers with AR-error structure. TEST, 11(2):439–464, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595716.

Ferreira:2012:EDS

[FF12] Marta Ferreira and Helena Ferreira. On extremal dependence: some contributions. TEST, 21(3):566–583, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0261-3.

Ferreira:2013:EMA

[FF13] Marta Ferreira and Helena Ferreira. Extremes of multivariate ARMAX processes. TEST, 22(4):606–627, November 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0326-6.

Ferreira:2021:TDS

[FF21] Helena Ferreira and Marta Ferreira. Tail dependence and smoothness of time series. TEST, 30(1):198–210, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00709-z.

Fraiman:2017:NSD

[FFF17] Daniel Fraiman, Nicolas Fraiman, and Ricardo Fraiman. Nonparametric statistics of dynamic networks with distinguishable nodes. TEST, 26(3):546–573, September 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0524-8.

Forzani:2013:DES

[FFL13] Liliana Forzani, Ricardo Fraiman, and Pamela Llop. Density estimation for spatial-temporal models. TEST, 22(2):321– 342, June 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0313-3.

Fontaine:2020:MDC

[FFO20] Charles Fontaine, Ron D. Frostig, and Hernando Ombao. Modeling dependence via copula of functionals of Fourier coefficients. TEST, 29(4):1125–1144, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00703-5.

Ferrante:2016:SES

[FFR16] Marco Ferrante, Elisabetta Ferraris, and Carles Rovira. On a stochastic epidemic SEIHR model and its diffusion approximation. TEST, 25(3):482–502, September 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0465-z.

Ferreira:2015:PLB

[FFZdC15] Guillermo Ferreira, Jorge I. Figueroa-Zúñiga, and Mário de Castro. Partially linear beta regression model with autoregressive errors. TEST, 24(4):752–775, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0433-7.

Frery:2020:CSM

[FG20] Alejandro C. Frery and Juliana Gambini. Comparing samples from the G^{0} distribution using a geodesic distance. TEST, 29(2):359–378, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00658-2.

Ferri-Garcia:2022:WSN

[FGBB^{+}22] Ramón Ferri-García, Jean-François Beaumont, Keven Bosa, Joanne Charlebois, and Kenneth Chu. Weight smoothing for nonprobability surveys. TEST, 31(3):619–643, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-021-00795-7.

Freijeiro-Gonzalez:2023:CNS

[FGFBGM23a] Laura Freijeiro-González, Manuel Febrero-Bande, and Wenceslao González-Manteiga. Correction: Novel specification tests for synchronous additive concurrent model formulation based on martingale difference divergence. TEST, 32(3):1131, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00863-0. See

[FGFBGM23b].

Freijeiro-Gonzalez:2023:NST

[FGFBGM23b] Laura Freijeiro-González, Manuel Febrero-Bande, and Wenceslao González-Manteiga. Novel specification tests for synchronous additive concurrent model formulation based on martingale difference divergence. TEST, 32(3):908–941, September 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00857-y. See correction

[FGFBGM23a].

Freise:2021:ODP

[FGS21] Fritjof Freise, Ulrike Graßhoff, and Rainer Schwabe. Doptimal designs for Poisson regression with synergetic interaction effect. TEST, 30(4):1004–1025, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00752-w.

Ferraty:2013:FPP

[FGSV13] F. Ferraty, A. Goia, E. Salinelli, and P. Vieu. Functional projection pursuit regression. TEST, 22(2):293–320, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0306-2.

Ferraty:2002:FNM

[FGV02] Frédéric Ferraty, Aldo Goia, and Philippe Vieu. Functional nonparametric model for time series: a fractal approach for dimension reduction. TEST, 11(2):317–344, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595710.

Filzmoser:2019:CCD

[FH19] Peter Filzmoser and Karel Hron. Comments on: Compositional data: the sample space and its structure. TEST, 28 (3):639–643, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00671-5. See

[EPG19a, EPG19b].

Freitas:2012:LLM

[FHT12] A. Freitas, J. Hüsler, and M. G. Temido. Limit laws for maxima of a stationary random sequence with random sample size. TEST, 21(1):116–131, March 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0238-2.

Fenoy:2003:SSP

[FI03] M. Mar Fenoy and Pilar Ibarrola. Sufficiency in sequentially planned decision procedures. TEST, 12(2):365–384, December 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595720.

Fan:2007:NIG

[FJ07a] Jianqing Fan and Jiancheng Jiang. Nonparametric inference with generalized likelihood ratio tests. TEST, 16(3): 409–444, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0080-8. See

comments [Bic07, Hal07, Mül07, LW07, CM07, Hor07, Mam07, Efr07, Cao07] and rejoinder [FJ07b].

Fan:2007:RNI

[FJ07b] Jianqing Fan and Jiancheng Jiang. Rejoinder on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):471–478, December 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-007-0090-6. See [FJ07a, Bic07, Hal07, Mül07, LW07, CM07, Hor07, Mam07, Efr07, Cao07].

Filipiak:2024:TCS

[FJL24] Katarzyna Filipiak, Mateusz John, and Yuli Liang. Testing covariance structures belonging to a quadratic subspace under a doubly multivariate model. TEST, 33(3):847–876, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00922-0.

Fermin:2008:SVI

[FL08] Ana K. Fermín and C. Ludeña. A statistical view of iterative methods for linear inverse problems. TEST, 17(2):381– 400, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0038-2.

Fan:2010:CPM

[FL10] Jianqing Fan and Jinchi Lv. Comments on: “l_{1}-penalization for mixture regression models”. TEST, 19(2):264–269, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0200-8.

Florens:2024:PLI

[FL24] Jean-Pierre Florens and Elia Lapenta. Partly linear instrumental variables regressions without smoothing on the instruments. TEST, 33(3):897–920, September 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00931-z.

Flores:2015:SNA

[Flo15] Salvador Flores. Sharp non-asymptotic performance bounds for ℓ_{1} and Huber robust regression estimators. TEST, 24 (4):796–812, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0435-5.

Feng:2025:TLS

[FLS25] Yongzhen Feng, Jie Li, and Xiaojun Song. Testing linearity in semi-functional partially linear regression models. TEST, 34(3):742–785, September 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-025-00979-5.

Fraiman:2001:TMF

[FM01] Ricardo Fraiman and Graciela Muniz. Trimmed means for functional data. TEST, 10(2):419–440, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595706.

Fraiman:1999:ME

[FMGE^{+}99] Ricardo Fraiman, Jean Meloche, Luis A. García-Escudero, Alfonso Gordaliza, Xuming He, Ricardo Maronna, Víctor J. Yohai, Simon J. Sheather, Joseph W. McKean, Christopher G. Small, Andrew Wood, R. Fraiman, and Jean Meloche. Multivariate l-estimation. TEST, 8(2):255–317, December 1999. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595872.

Ferreira:2021:BRM

[FMM21] Guillermo Ferreira, Jorge Mateu, and Joel Muñoz. Bootstrapping regression models with locally stationary disturbances. TEST, 30(2):341–363, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
020-00721-3.

Ferreira:2018:STA

[FMP18] Guillermo Ferreira, Jorge Mateu, and Emilio Porcu. Spatiotemporal analysis with short- and long-memory dependence: a state-space approach. TEST, 27(1):221–245, March 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0541-7.

Fraiman:2024:ACW

[FMR24] Ricardo Fraiman, Leonardo Moreno, and Thomas Ransford. Application of the Cramér–Wold theorem to testing for invariance under group actions. TEST, 33(2):379–399, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00899-2.

Fokianos:2012:CSR

[Fok12] Konstantinos Fokianos. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):451–454, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0297-z. See [Tjø12b] and

rejoinder [Tjø12a].

Farcomeni:2020:RMB

[FP20] Alessio Farcomeni and Antonio Punzo. Robust model-based clustering with mild and gross outliers. TEST, 29(4):989– 1007, December 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-019-00693-z.

Fraiman:1996:NCB

[FPI96] R. Fraiman and G. Pérez-Iribarren. Nonparametric conservative bands for the trend of Gaussian AR(p) models. TEST, 5(1):125–144, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562685.

Fernandez-Ponce:2017:NPO

[FPRG17] J. M. Fernández-Ponce and M. R. Rodríguez-Griñolo. New properties of the orthant convex-type stochastic orders. TEST, 26(3):618–637, September 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0527-5.

Fernandez-Pascual:2004:WBF

[FPRMA04] Rosaura Fernández-Pascual, María D. Ruiz-Medina, and Jose M. Angulo. Wavelet-based functional reconstruction and extrapolation of fractional random fields. TEST, 13 (2):417–444, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595780.

Forzani:2024:ARN

[FRS24] Liliana Forzani, Daniela Rodriguez, and Mariela Sued. Asymptotic results for nonparametric regression estimators after sufficient dimension reduction estimation. TEST, 33 (4):987–1013, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00932-y.

Farcomeni:2021:DRL

[FRV21] Alessio Farcomeni, Monia Ranalli, and Sara Viviani. Dimension reduction for longitudinal multivariate data by optimizing class separation of projected latent Markov models. TEST, 30(2):462–480, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-020-00727-x.

Fernandez:1998:RPN

[FS98] Carmen Fernández and Mark F. J. Steel. Reference priors for non-normal two-sample problems. TEST, 7(1):179–205, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02565109.

Fujisawa:2012:OSA

[FS12] Hironori Fujisawa and Takayuki Sakaguchi. Optimal significance analysis of microarray data in a class of tests whose null statistic can be constructed. TEST, 21(2):280–300, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0243-5.

Ferraccioli:2023:NTS

[FSF23] Federico Ferraccioli, Laura M. Sangalli, and Livio Finos. Nonparametric tests for semiparametric regression models. TEST, 32(3):1106–1130, September 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00868-9.

Farcomeni:2010:RBM

[FT10] Alessio Farcomeni and Luca Tardella. Reference Bayesian methods for recapture models with heterogeneity. TEST, 19(1):187–208, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0147-9.

Frias:2022:SCP

[FTSM22] María P. Frías, Antoni Torres-Signes, and Jorge Mateu. Spatial Cox processes in an infinite-dimensional framework. TEST, 31(1):175–203, March 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00773-z.

Fuentes:2008:CAP

[Fue08] Montserrat Fuentes. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):245– 248, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0119-5.

Ferreira:2008:CCF

[FvdW08] José A. Ferreira and Mark A. van de Wiel. Comments on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17(3):443–445, November 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0127-5. See [RSW08a] and

rejoinder [RSW08b].

Francq:2018:GFT

[FWZ18] Christian Francq, Olivier Wintenberger, and Jean-Michel Zakoïan. Goodness-of-fit tests for Log–GARCH and EGARCH models. TEST, 27(1):27–51, March 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0506-2.

Feng:2015:RCR

[FZWZ15] Long Feng, Changliang Zou, Zhaojun Wang, and Lixing Zhu. Robust comparison of regression curves. TEST, 24(1):185– 204, March 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0394-2.

Fan:2010:CDR

[FZZ10] Jianqing Fan, Jin-Ting Zhang, and Wenyang Zhang. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):37–42, May 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0180-8.

Galeano:2012:CSR

[Gal12] Pedro Galeano. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):455–458, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0300-8. See [Tjø12b] and

rejoinder [Tjø12a].

Goto:2023:HTO

[GALT23] Yuichi Goto, Koichi Arakaki, Yan Liu, and Masanobu Taniguchi. Homogeneity tests for one-way models with dependent errors under correlated groups. TEST, 32(1):163– 183, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00828-9.

Gamero:2013:CUR

[Gam13] M. D. Jiménez Gamero. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22(3):412–413, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0328-4;
http://link.springer.com/content/pdf/10.1007/s11749-

013-0328-4.pdf. See [GMC13b] and rejoinder [GMC13a].

Gamero:2014:ECF

[Gam14] M. Dolores Jiménez Gamero. On the empirical characteristic function process of the residuals in GARCH models and applications. TEST, 23(2):409–432, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0359-5.

Gannaz:2013:WPL

[Gan13] Irène Gannaz. Wavelet penalized likelihood estimation in generalized functional models. TEST, 22(1):122–158, March 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0310-6.

Gao:2012:CSR

[Gao12] Jiti Gao. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):459–463, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0301-7. See [Tjø12b] and rejoinder

[Tjø12a].

Ghosh:2016:REG

[GB16a] Abhik Ghosh and Ayanendranath Basu. Robust estimation in generalized linear models: the density power divergence approach. TEST, 25(2):269–290, June 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0445-3.

Ghosh:2016:DFH

[GB16b] Anil K. Ghosh and Munmun Biswas. Distribution-free high-dimensional two-sample tests based on discriminating hyperplanes. TEST, 25(3):525–547, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0467-x.

Ghosh:1995:PMP

[GCS95] M. Ghosh, B. P. Carlin, and M. S. Srivastava. Probability matching priors for linear calibration. TEST, 4(2):333–357, December 1995. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562631.

Ghorbani:2021:FMP

[GCY21] Mohammad Ghorbani, Ottmar Cronie, and Jun Yu. Functional marked point processes: a natural structure to unify spatio-temporal frameworks and to analyse dependent functional data. TEST, 30(3):529–568, September 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00730-2.

Gomez-Deniz:2010:AGG

[GD10] E. Gómez-Déniz. Another generalization of the geometric distribution. TEST, 19(2):399–415, August 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0169-3.

Ge:2003:RBM

[GDS03] Youngchao Ge, Sandrine Dudoit, and Terence P. Speed. Resampling-based multiple testing for microarray data analysis. TEST, 12(1):1–77, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595811.

Gomez-Deniz:2006:NCB

[GDVPP06] E. Gómez-Déniz, F. J. Vázquez-Polo, and J. M. Pérez. A note on computing bonus-malus insurance premiums using a hierarchical Bayesian framework. TEST, 15(2):345–359, September 2006. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02607056.

Goegebeur:2012:ETO

[GdW12] Yuri Goegebeur and Tertius de Wet. Estimation of the thirdorder parameter in extreme value statistics. TEST, 21(2): 330–354, June 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0246-2.

Geffray:2009:SAD

[Gef09] Ségolen Geffray. Strong approximations for dependent competing risks with independent censoring. TEST, 18(1):76–95, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0113-y.

Gelfand:2015:CCS

[Gel15] Alan E. Gelfand. Comments on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1): 29–30, March 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0412-4;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0412-

4.pdf. See [BCS15a] and rejoinder [BCS15b].

Gomes:2006:BRR

[GF06] M. Ivette Gomes and Fernanda Figueiredo. Bias reduction in risk modelling: Semi-parametric quantile estimation. TEST, 15(2):375–396, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02607058.

Greco:2016:PAS

[GF16] Luca Greco and Alessio Farcomeni. A plug-in approach to sparse and robust principal component analysis. TEST, 25 (3):449–481, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0464-0.

Gupta:2004:GSN

[GG04] Ramesh C. Gupta and Rameshwar D. Gupta. Generalized skew normal model. TEST, 13(2):501–524, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595784.

Gupta:2008:ASD

[GG08] Rameshwar D. Gupta and Ramesh C. Gupta. Analyzing skewed data by power normal model. TEST, 17(1):197– 210, May 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0030-x.

Goegebeur:2011:WME

[GG11] Yuri Goegebeur and Armelle Guillou. A weighted mean excess function approach to the estimation of Weibulltype tails. TEST, 20(1):138–162, May 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0190-6.

Gardes:2021:EVD

[GG21] Laurent Gardes and Stéphane Girard. On the estimation of the variability in the distribution tail. TEST, 30(4):884– 907, December 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00754-2.

Goegebeur:2024:CTM

[GGQ24] Yuri Goegebeur, Armelle Guillou, and Jing Qin. Conditional tail moment and reinsurance premium estimation under random right censoring. TEST, 33(1):230–250, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00890-x.

Girard:2012:EEH

[GGS12] Stéphane Girard, Armelle Guillou, and Gilles Stupfler. Estimating an endpoint with high-order moments. TEST, 21 (4):697–729, December 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0277-8.

Gupta:2008:MLH

[GHF08] Arjun K. Gupta, Solomon W. Harrar, and Yasunori Fujikoshi. MANOVA for large hypothesis degrees of freedom under non-normality. TEST, 17(1):120–137, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0026-6.

Goyal:2022:GDS

[GHF22] Dheeraj Goyal, Nil Kamal Hazra, and Maxim Finkelstein. On the general δ-shock model. TEST, 31(4):994–1029, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00810-5.

Goyal:2023:GCS

[GHF23] Dheeraj Goyal, Nil Kamal Hazra, and Maxim Finkelstein. A general class of shock models with dependent interarrival times. TEST, 32(3):1079–1105, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00867-w.

Ghosal:1997:RPM

[Gho97] S. Ghosal. Reference priors in multiparameter nonregular cases. TEST, 6(1):159–186, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02564432.

Ghosh:2020:CAL

[Gho20] Abhik Ghosh. Comments on: On active learning methods for manifold data. TEST, 29(1):34–37, March 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00695-x. See [LDR20a].

Ghosh:2022:RPI

[Gho22] Abhik Ghosh. Robust parametric inference for finite Markov chains. TEST, 31(1):118–147, March 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00771-1.

Goncalves:1996:NTA

[GJL96] E. Gonçalves, P. Jacob, and N. Mendes Lopes. A new test for ARMA models with errors following a general white noise process. TEST, 5(1):187–202, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562688.

Geerdens:2020:GFT

[GJV20] Candida Geerdens, Paul Janssen, and Ingrid Van Keilegom. Goodness-of-fit test for a parametric survival function with cure fraction. TEST, 29(3):768–792, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-019-00680-4.

Geng:2019:MDM

[GK19] Pei Geng and Hira L. Koul. Minimum distance model checking in Berkson measurement error models with validation data. TEST, 28(3):879–899, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0610-6.

Gamboa:2007:EPM

[GL07] Fabrice Gamboa and Jean-Michel Loubes. Estimation of parameters of a multifractal process. TEST, 16(2):383–407, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0015-9.

Gonzalez:2021:TWL

[GLÁM21] Jonatan A. González, Bernardo M. Lagos-Álvarez, and Jorge Mateu. Two-way layout factorial experiments of spatial point pattern responses in mineral flotation. TEST, 30(4):1046– 1075, December 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00768-w.

Gil:2001:RCA

[GLGLM01] M. Ángeles Gil, M. Teresa López-García, M. Asunción Lubiano, and Manuel Montenegro. Regression and correlation analyses of a linear relation between random intervals. TEST, 10(1):183–201, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595831.

Gouet:2012:ROA

[GLS12] Raúl Gouet, F. Javier López, and Gerardo Sanz. On δ-record observations: asymptotic rates for the counting process and elements of maximum likelihood estimation. TEST, 21(1): 188–214, March 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0242-6.

Gouet:2015:PPN

[GLS15] Raúl Gouet, F. Javier López, and Gerardo Sanz. On the point process of near-record values. TEST, 24(2):302–321, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0408-0.

Gasull:2015:MGR

[GLSU15] Armengol Gasull, José A. López-Salcedo, and Frederic Utzet. Maxima of Gamma random variables and other Weibulllike distributions and the Lambert W function. TEST, 24 (4):714–733, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0431-9.

Gyorfi:2025:DFT

[GLW25] László Györfi, Tamás Linder, and Harro Walk. Distributionfree tests for lossless feature selection in classification and regression. TEST, 34(1):262–287, March 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00958-2.

Guttman:1994:PBR

[GM94] Irwin Guttman and Ulrich Menzefricke. Prediction based on response surface data obtained with random blocking. TEST, 3(2):87–99, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562695.

Guerrero:1995:RAB

[GM95] V. M. Guerrero and J. Martínez. A recursive ARIMAbased procedure for disaggregating a time series variable using concurrent data. TEST, 4(2):359–376, December 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562632.

Giron:1998:NAG

[GM98a] F. Javier Girón and Elías Moreno. A note on an assertion by E. Gutiérrez–Peña and A. F. M. Smith (Test, 1997, p. 87). TEST, 7(2):427–429, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02565122.

Giudici:1998:NES

[GM98b] Paolo Giudici and Maura Mezzetti. Nonparametric estimation of survival functions by means of partial exchangeability structures. TEST, 7(1):111–132, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02565105.

Guttman:2003:PDF

[GM03] Irwin Guttman and Ulrich Menzefricke. Posterior distributions for functions of variance components. TEST, 12(1): 115–123, June 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595814.

Ghosal:2024:VSF

[GM24] Rahul Ghosal and Arnab Maity. Variable selection in function-on-scalar single-index model via the alternating direction method of multipliers. TEST, 33(1):106–126, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00884-9.

Gonzalez-Manteiga:1993:THG

[GMC93] W. González-Manteiga and R. Cao. Testing the hypothesis of a general linear model using nonparametric regression estimation. TEST, 2(1–2):161–188, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562674.

Gonzalez-Manteiga:2008:CAB

[GMC08] Wenceslao González-Manteiga and Rosa M. Crujeiras. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):40–42, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0105-y.

Gonzalez-Manteiga:2013:RUR

[GMC13a] Wenceslao González-Manteiga and Rosa M. Crujeiras. Rejoinder on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22(3):442–447, September 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0334-6;   http://link.
springer.com/content/pdf/10.1007/s11749-013-0334-

6.pdf. See [GMC13b, Gam13, dUÁ13, Spe13a, Van13a, Mei13, Det13].

Gonzalez-Manteiga:2013:URG

[GMC13b] Wenceslao González-Manteiga and Rosa M. Crujeiras. An updated review of goodness-of-fit tests for regression models. TEST, 22(3):361–411, September 2013. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-013-0327-5. See comments [Gam13, dUÁ13, Spe13a, Van13a, Mei13, Det13] and rejoinder [GMC13a].

Gonzalez:2004:NEO

[GMdP04] Miguel González, Rodrigo Martínez, and Inés del Puerto. Nonparametric estimation of the offspring distribution and the mean for a controlled branching process. TEST, 13 (2):465–479, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595782.

Gonzalez:2005:EVC

[GMdP05] Miguel González, Rodrigo Martínez, and Iné del Puerto. Estimation of the variance for a controlled branching process. TEST, 14(1):199–213, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595403.

Gonzalez:2011:WCL

[GMdP11] M. González, M. Mota, and I. del Puerto. Weighted conditional least square estimators for bisexual branching processes with immigration. TEST, 20(3):607–629, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0220-4.

Graf:2019:GMM

[GMM19] Monique Graf, J. Miguel Marín, and Isabel Molina. A generalized mixed model for skewed distributions applied to small area estimation. TEST, 28(2):565–597, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0594-2.

Gonzalez-Manteiga:2010:CDR

[GMMC10] Wenceslao González-Manteiga and Adela Martínez-Calvo. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):43–45, May 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0177-3.

Godtliebsen:2024:CDI

[GMNH24] Fred Godtliebsen, Eirik Myrvoll-Nilsen, and Lasse Holmström. Comments on: Data integration via analysis of subspaces (DIVAS). TEST, 33(3):683–685, September 2024.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00937-7. See [PJH^{+}24a].

Giummole:2019:OBI

[GMRV19] F. Giummolè, V. Mameli, E. Ruli, and L. Ventura. Objective Bayesian inference with proper scoring rules. TEST, 28 (3):728–755, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0597-z.

Guerrero:1999:TCD

[GN99] Víctor M. Guerrero and Fabio H. Nieto. Temporal and contemporaneous disaggregation of multiple economic time series. TEST, 8(2):459–489, December 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595880.

Ghannam:2024:CPD

[GN24] Mai Ghannam and Sévérien Nkurunziza. Change-point detection in a tensor regression model. TEST, 33(2):609–630, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00915-5.

Guillou:2009:RLB

[GNDR09] A. Guillou, P. Naveau, J. Diebolt, and P. Ribereau. Return level bounds for discrete and continuous random variables. TEST, 18(3):584–604, November 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0125-7.

Genest:2011:IMA

[GNZ11a] Christian Genest, Johanna Neslehová, and Johanna Ziegel. Inference in multivariate Archimedean copula models. TEST, 20(2):223–256, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-011-0250-

6. See comments [Val11, Val11, EH11, JD11, WE11, Seg11, Lam11, Tsu11] and rejoinder [GNZ11b].

Genest:2011:RIM

[GNZ11b] Christian Genest, Johanna Neslehová, and Johanna Ziegel. Rejoinder on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):290–292, August 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-011-0258-y. See [GNZ11a, Val11, Val11, EH11, JD11, WE11, Seg11, Lam11, Tsu11].

Goicoa:2019:CMR

[Goi19] T. Goicoa. Comments on: Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1): 40–42, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00633-x. See [KKLU19a,

KKLU19b].

Goldsmith:2025:CEF

[Gol25] Jeff Goldsmith. Comments on: Exploratory functional data analysis. TEST, 34(2):483–487, June 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00964-y. See [QDG25a], comments

[Hyn25, Lil25, LP25, PS25], and rejoinder [QDG25b].

Garcia-Perez:1994:BRE

[GP94] Alfonso García-Pérez. Bayesian robustness of the empirical distribution. TEST, 3(1):183–194, June 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562680.

Garcia-Perez:2003:MAC

[GP03] Alfonso García-Pérez. Von Mises approximation of the critical value of a test. TEST, 12(2):385–411, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595721.

Garcia-Perez:2008:ATW

[GP08] Alfonso García-Pérez. Approximations for F -tests which are ratios of sums of squares of independent variables with a model close to the normal. TEST, 17(2):350–369, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0036-4.

Garcia-Perez:2014:VLW

[GP14] A. García-Pérez. The p value line: a way to choose the tuning constant in tests based on the Huber M-estimator. TEST, 23(3):536–555, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0361-y.

Galeano:2019:DSB

[GP19a] Pedro Galeano and Daniel Peña. Data science, big data and statistics. TEST, 28(2):289–329, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-019-00651-9. See comments [Büh19, Del19b, GS19, Mar19, NS19, RACC19, SH19, Tsa19, VZ19, Cao19] and rejoinder [GP19b].

Galeano:2019:RDS

[GP19b] Pedro Galeano and Daniel Peña. Rejoinder on: Data science, big data and statistics. TEST, 28(2):363–368, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00652-8. See [GP19a, Büh19, Del19b,

GS19, Mar19, NS19, RACC19, SH19, Tsa19, VZ19].

Gijbels:2010:NEM

[GPC10] I. Gijbels, I. Prosdocimi, and G. Claeskens. Nonparametric estimation of mean and dispersion functions in extended generalized linear models. TEST, 19(3):580–608, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0187-1.

Gil:2000:SRE

[GPR00] Juan A. Gil, Daniel Peña, and Julio Rodríguez. Statistical research in Europe: 1985-1997. TEST, 9(1):255–281, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595861.

Gutierrez-Pena:1997:EBC

[GPSB^{+}97] E. Gutiérrez-Peña, A. F. M. Smith, José M. Bernardo, Guido Consonni, Piero Veronese, E. I. George, F. J. Girón, M. L. Martínez, G. Letac, and Carl N. Morris. Exponential and Bayesian conjugate families: Review and extensions. TEST, 6(1):1–90, June 1997. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02564426.

Genest:2004:TIR

[GR04] Christian Genest and Bruno Rémillard. Test of independence and randomness based on the empirical copula process. TEST, 13(2):335–369, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595777.

Gayraud:2007:CRN

[GR07] Ghislaine Gayraud and Judith Rousseau. Consistency results on nonparametric Bayesian estimation of level sets using spatial priors. TEST, 16(1):90–108, May 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0003-0.

Ghoudi:2018:SIT

[GR18] Kilani Ghoudi and Bruno Rémillard. Serial independence tests for innovations of conditional mean and variance models. TEST, 27(1):3–26, March 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0521-3.

Greco:2011:MHD

[Gre11] Luca Greco. Minimum Hellinger distance based inference for scalar skew-normal and skew-t distributions. TEST, 20(1): 120–137, May 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0191-5.

Greenacre:2019:CCD

[Gre19] Michael Greenacre. Comments on: Compositional data: the sample space and its structure. TEST, 28(3):644–652, September 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00673-3. See [EPG19a,

EPG19b].

Gutierrez:2001:ISP

[GRT01] Ramón Gutiérrez, Patrica Román, and Francisco Torres. Inference on some parametric functions in the univeriate lognormal diffusion process with exogenous factors. TEST, 10 (2):357–373, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595702.

Garcia-Soidan:1996:AES

[GS96] P. García-Soidán. Asymptotic expansions for statistics computed from spatial data. TEST, 5(1):61–76, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562682.

Ghosh:2006:CDP

[GS06] Subir Ghosh and Yun Shen. Comparison of designs in presence of a possible correlation in observations. TEST, 15 (2):485–504, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02607063.

Garcia-Soidan:2007:ANN

[GS07] Pilar García-Soidán. Asymptotic normality of the Nadaraya– Watson semivariogram estimators. TEST, 16(3):479–503, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0016-8.

Ghosh:2013:TSB

[GS13] Malay Ghosh and Rebecca C. Steorts. Two-stage benchmarking as applied to small area estimation. TEST, 22 (4):670–687, November 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0338-2.

Gardes:2015:EEE

[GS15a] Laurent Gardes and Gilles Stupfler. Erratum to: “Estimating extreme quantiles under random truncation”. TEST, 24 (2):228, June 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0437-3;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0437-

3.pdf. See [GS15b].

Gardes:2015:EEQ

[GS15b] Laurent Gardes and Gilles Stupfler. Estimating extreme quantiles under random truncation. TEST, 24(2):207–227, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0403-5. See erratum [GS15a].

Genton:2019:CDS

[GS19] Marc G. Genton and Ying Sun. Comments on: Data science, big data and statistics. TEST, 28(2):338–341, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00642-w. See [GP19a, GP19b].

Greven:2020:CIC

[GS20] Sonja Greven and Fabian Scheipl. Comments on: Inference and computation with Generalized Additive Models and their extensions. TEST, 29(2):343–350, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-020-00714-2;   http://link.springer.com/
content/pdf/10.1007/s11749-020-00714-2.pdf.  See

[Woo20a] and rejoinder [Woo20b].

Gaynanova:2024:CDI

[GS24]  Irina Gaynanova and Renat Sergazinov.  Comments

on: Data integration via analysis of subspaces (DIVAS). TEST, 33(3):675–682, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

024-00936-8. See [PJH^{+}24a].

Gelfand:2004:NMP

[GSBS04] Alan E. Gelfand, Alexandra M. Schmidt, Sudipto Banerjee, and C. F. Sirmans. Nonstationary multivariate process modeling through spatially varying coregionalization. TEST, 13 (2):263–312, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595775.

Ganguly:1992:BEG

[GSCB92] Arup Ganguly, Nand Kishore Singh, Haren Choudhuri, and Samir K. Bhattacharya. Bayesian estimation of the Gini index for the PID. TEST, 1(1):93–104, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562665.

Gneiting:2008:APF

[GSG^{+}08a] Tilmann Gneiting, Larissa I. Stanberry, Eric P. Grimit, Leonhard Held, and Nicholas A. Johnson. Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds. TEST, 17(2):211–235, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0114-x. See

comments [ARM08, Bic08, Bri08, Daw08, Fue08, Jol08, WJ08] and rejoinder [GSG^{+}08b].

Gneiting:2008:RAP

[GSG^{+}08b] Tilmann Gneiting, Larissa I. Stanberry, Eric P. Grimit, Leonhard Held, and Nicholas A. Johnson. Rejoinder on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):256–264, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0122-x. See

[GSG^{+}08a, ARM08, Bic08, Bri08, Daw08, Fue08, Jol08, WJ08].

Goia:2011:ESA

[GSS11] Aldo Goia, Ernesto Salinelli, and Pascal Sarda. Exploring the statistical applicability of the Poincaré inequality: a test of normality. TEST, 20(2):334–352, August 2011.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0205-3.

Gayraud:2011:GFT

[GT11] Ghislaine Gayraud and Karine Tribouley. A goodness-of-fit test for copula densities. TEST, 20(3):549–573, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0217-z.

Greco:2018:CSR

[GT18] Fedele Greco and Carlo Trivisano. Comments on: “Some recent work on multivariate Gaussian Markov random fields”. TEST, 27(3):549–553, September 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0607-1;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0607-1.pdf. See [Mac18b]

and rejoinder [Mac18a].

Guilbaud:2007:CPC

[Gui07] Olivier Guilbaud. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):279–281, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0068-4. See [Bal07a] and rejoinder

[Bal07b].

Guo:2008:CCF

[Guo08] Wenge Guo. Comments on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17(3):446–449, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-008-0128-4. See [RSW08a] and rejoinder [RSW08b].

Gomez-Villegas:2008:CPC

[GVGP08] Miguel A. Gómez-Villegas and Beatriz González-Pérez. ϵ-Contaminated priors in contingency tables. TEST, 17(1): 163–178, May 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0028-4.

Ghorbani:2025:KBT

[GVM25] Mohammad Ghorbani, Nafiseh Vafaei, and Mari Myllymäki. A kernel-based test for the first-order separability of spatiotemporal point processes. TEST, 34(3):530–579, September 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-025-00972-y.

Gomez-Villegas:1998:RBF

[GVS98] Miguel A. Gómez-Villegas and Luis Sanz. Reconciling Bayesian and frequentist evidence in the point null testing problem. TEST, 7(1):207–216, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02565110.

Gupta:1999:DUM

[GW99] Arjun K. Gupta and Jacek Wesolowski. Discrete uniform mixtures via posterior means. TEST, 8(2):399–409, December 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595877.

Geurts:2016:CRF

[GW16] Pierre Geurts and Louis Wehenkel. Comments on: “A random forest guided tour”. TEST, 25(2):247–253, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0487-1. See [BS16a] and rejoinder

[BS16b].

Galeano:2017:DMC

[GW17] Pedro Galeano and Dominik Wied. Dating multiple change points in the correlation matrix. TEST, 26(2):331–352, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0513-3.

Gu:2014:SCC

[GWHY14] Lijie Gu, Li Wang, Wolfgang K. Härdle, and Lijian Yang. A simultaneous confidence corridor for varying coefficient regression with sparse functional data. TEST, 23(4):806–843, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0392-4.

Ghosh:1996:NPT

[GY96] M. Ghosh and M-Ch. Yang. Noninformative priors for the two sample normal problem. TEST, 5(1):145–157, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562686.

Gahrooei:2020:CAL

[GYP20] Mostafa Reisi Gahrooei, Hao Yan, and Kamran Paynabar. Comments on: On Active Learning Methods for Manifold Data. TEST, 29(1):38–41, March 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00696-w. See [LDR20a].

Guo:2019:VLN

[GZCZ19] Jia Guo, Bu Zhou, Jianwei Chen, and Jin-Ting Zhang. An L^{2}-norm-based test for equality of several covariance functions: a further study. TEST, 28(4):1092–1112, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0617-z.

Hall:2007:CNI

[Hal07] Peter Hall. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):448–449, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0082-6. See [FJ07a] and re-

joinder [FJ07b].

Hayter:2014:ICN

[Hay14] Anthony Hayter. Identifying common normal distributions. TEST, 23(1):135–152, March 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0345-3.

Hadjicostas:1999:IPP

[HB99] Petros Hadjicostas and Scott M. Berry. Improper and proper posteriors with improper priors in a Poisson-gamma hierarchical model. TEST, 8(1):147–166, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595867.

Hermanns:2018:IPC

[HC18] M. Hermanns and E. Cramer. Inference with progressively censored k-out-of-n system lifetime data. TEST, 27(4):787– 810, December 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0569-8.

Huo:2021:TSF

[HC21] Lijuan Huo and Jin Seo Cho. Testing for the sandwichform covariance matrix of the quasi-maximum likelihood estimator. TEST, 30(2):293–317, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
020-00719-x.

Hu:2017:SLW

[HCS17] Di Hu, Pingyan Chen, and Soo Hak Sung. Strong laws for weighted sums of ψ-mixing random variables and applications in errors-in-variables regression models. TEST, 26 (3):600–617, September 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0526-6.

He:2011:CNI

[He11] Xin He. Comments on: “Nonparametric inference based on panel count data”. TEST, 20(1):46–47, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-010-0227-x. See [ZBS11a] and rejoinder [ZBS11b].

Heckman:2010:CDR

[Hec10] Nancy E. Heckman. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):46–49, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0181-7.

Heinen:2012:CSR

[Hei12] Andréas Heinen. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):464–466, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0304-4. See [Tjø12b] and

rejoinder [Tjø12a].

Hering:2014:CST

[Her14] Amanda S. Hering. Comments on: “Space–time wind speed forecasting for improved power system dispatch”. TEST, 23(1):34–44, March 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0355-9;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0355-9.pdf. See [ZGGX14b] and rejoinder [ZGGX14a].

Hazra:2018:SCF

[HF18] Nil Kamal Hazra and Maxim Finkelstein. On stochastic comparisons of finite mixtures for some semiparametric families of distributions. TEST, 27(4):988–1006, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0581-7.

Huang:2019:NEL

[HF19] Na Huang and Piotr Fryzlewicz. NOVELIST estimator of large correlation and covariance matrices and their inverses. TEST, 28(3):694–727, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0592-4;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0592-4.pdf.

Hazra:2018:SOP

[HFC18] Nil Kamal Hazra, Maxim Finkelstein, and Ji Hwan Cha. Stochastic ordering for populations of manufactured items. TEST, 27(1):173–196, March 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0539-1.

Hudson:2008:BAR

[HG08] Brent G. Hudson and Richard H. Gerlach. A Bayesian approach to relaxing parameter restrictions in multivariate GARCH models. TEST, 17(3):606–627, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0056-8.

Hemerik:2018:ETR

[HG18] Jesse Hemerik and Jelle Goeman. Exact testing with random permutations. TEST, 27(4):811–825, December 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0571-1;   http://link.
springer.com/content/pdf/10.1007/s11749-017-0571-
1.pdf.

Hubert:2013:RMS

[HGV13] Mia Hubert, Irène Gijbels, and Dina Vanpaemel. Reducing the mean squared error of quantile-based estimators by smoothing. TEST, 22(3):448–465, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-012-0293-3.

Hlavka:2012:MCE

[HHKM12] Zdenek Hlávka, Marie Husková, Claudia Kirch, and Simos G. Meintanis. Monitoring changes in the error distribution of autoregressive models based on Fourier methods. TEST, 21 (4):605–634, December 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0265-z.

Hlavka:2021:TSI

[HHM21] Zdenek Hlávka, Marie Husková, and Simos G. Meintanis. Testing serial independence with functional data. TEST, 30 (3):603–629, September 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-020-00732-0.

Hudecova:2024:STC

[HHM24] Sárka Hudecová, Marie Husková, and Simos G. Meintanis. Specifications tests for count time series models with covariates. TEST, 33(4):1014–1040, December 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00933-x.

Hidalgo:1999:NTM

[Hid99] Javier Hidalgo. Nonparametric tests for model selection with time series data. TEST, 8(2):365–398, December 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595876.

Higdon:2010:CGS

[Hig10] Dave Higdon. Comments on: “A general science-based framework for dynamical spatio-temporal models”. TEST, 19(3):462–465, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0213-3. See

[WH10b] and rejoinder [WH10b].

Henze:2019:NCT

[HJG19] Norbert Henze and María Dolores Jiménez-Gamero. A new class of tests for multinormality with i.i.d. and GARCH data based on the empirical moment generating function. TEST, 28(2):499–521, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0589-z.

Hendrickx:2018:PSE

[HJV18] K. Hendrickx, P. Janssen, and A. Verhasselt. Penalized spline estimation in varying coefficient models with censored data. TEST, 27(4):871–895, December 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0574-y.

Horvath:2009:EAE

[HL09] Lajos Horváth and Remigijus Leipus. Effect of aggregation on estimators in AR(1) sequence. TEST, 18(3):546–567, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0123-9.

Ha:2014:CST

[HL14] Neung Soo Ha and Partha Lahiri. Comments on: “Singleand two-stage cross-sectional and time series benchmarking procedures for small area estimation”. TEST, 23(4):670– 673, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0400-8;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0400-

8.pdf. See [PST14b] and rejoinder [PST14a].

Hossain:2021:OSE

[HL21] Shakhawat Hossain and Le An Lac. Optimal shrinkage estimations in partially linear single-index models for binary longitudinal data. TEST, 30(4):811–835, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00753-3.

Henze:2022:TCS

[HLM22] Norbert Henze, Pierre Lafaye De Micheaux, and Simos G. Meintanis. Tests for circular symmetry of complex-valued random vectors. TEST, 31(2):488–518, June 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00788-6.

Huskova:2010:TED

[HM10] Marie Husková and Simos G. Meintanis. Tests for the error distribution in nonparametric possibly heteroscedastic regression models. TEST, 19(1):92–112, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0135-5.

Hooker:2016:CRF

[HM16] Giles Hooker and Lucas Mentch. Comments on: “A random forest guided tour”. TEST, 25(2):254–260, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0485-3. See [BS16a] and rejoinder

[BS16b].

Hobza:2018:SAE

[HMS18] Tomás Hobza, Domingo Morales, and Laureano Santamaría. Small area estimation of poverty proportions under unit-level temporal binomial-logit mixed models. TEST, 27(2):270– 294, June 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0545-3.

Hobza:2005:CFI

[HMV05] Tomás Hobza, Isabel Molina, and Igor Vajda. On convergence of Fisher informations in continuous models with quantized observations. TEST, 14(1):151–179, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595401.

Hou:2023:SDL

[HMW23] Zhaohan Hou, Wei Ma, and Lei Wang. Sparse and debiased lasso estimation and inference for high-dimensional composite quantile regression with distributed data. TEST, 32(4): 1230–1250, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00875-w.

Hui:2009:PML

[HMZ09] Terrence P. Hui, Reza Modarres, and Gang Zheng. Pseudo maximum likelihood estimates using ranked set sampling with applications to estimating correlation. TEST, 18 (2):365–380, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0096-8.

Hyodo:2018:STM

[HN18] Masashi Hyodo and Takahiro Nishiyama. A simultaneous testing of the mean vector and the covariance matrix among two populations for high-dimensional data. TEST, 27(3): 680–699, September 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0567-x.

Hogan:2009:CMD

[Hog09a] Joseph W. Hogan. Comments on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):59– 64, May 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0142-1. See [IM09a], rejoin-

der [IM09b], and erratum [Hog09b].

Hogan:2009:ECS

[Hog09b] Joseph W. Hogan. Erratum to: “Considerations for sensitivity analysis with likelihood-based models”. TEST, 18(3): 607, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0172-8;   http://link.
springer.com/content/pdf/10.1007/s11749-009-0172-

8.pdf. See [Hog09a].

Hooker:2010:CDR

[Hoo10] Giles Hooker. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):50–53, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0178-2.

Horowitz:2007:CNI

[Hor07] Joel L. Horowitz. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):459– 461, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0086-2. See [FJ07a] and re-

joinder [FJ07b].

Horvath:2024:CSB

[Hor24] Lajos Horváth. Comments on: Shape-based functional data analysis by Wu, Huang and Srivastava. TEST, 33(1):71– 72, March 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00918-2. See [WHS24b,

WHS24a].

Huskova:2014:CES

[HP14] Marie Husková and Zuzana Prásková. Comments on: “Extensions of some classical methods in change point analysis”.

TEST, 23(2):265–269, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0373-7.

Heuchenne:2012:TOS

[HPF12] Cédric Heuchenne and Juan Carlos Pardo-Fernández. Testing for one-sided alternatives in nonparametric censored regression. TEST, 21(3):498–518, September 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0260-4.

Hotta:2004:EOF

[HPO04] Luiz K. Hotta, Pedro L. Valls Pereira, and Rissa Ota. Effect of outliers on forecasting temporally aggregated flow variables. TEST, 13(2):371–402, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595778.

Ham:2023:HTA

[HQ23] Dae Woong Ham and Jiaze Qiu. Hypothesis testing in adaptively sampled data: ART to maximize power beyond iid sampling. TEST, 32(3):998–1037, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00861-2.

Horvath:2014:ESC

[HR14a] Lajos Horváth and Gregory Rice. Extensions of some classical methods in change point analysis. TEST, 23(2):219–255, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0368-4. See comments [Ast14, Ber14,

Deh14, HP14, Kir14, Kok14, MP14, Tra14] and rejoinder [HR14b].

Horvath:2014:RES

[HR14b] Lajos Horváth and Gregory Rice. Rejoinder on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):287–290, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0375-5;

http://link.springer.com/content/pdf/10.1007/s11749-

014-0375-5.pdf. See [HR14a, Ast14, Ber14, Deh14, HP14, Kir14, Kok14, MP14, Tra14].

Hlubinka:2015:GEQ

[HS15] Daniel Hlubinka and Miroslav Siman. On generalized elliptical quantiles in the nonlinear quantile regression setup. TEST, 24(2):249–264, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0405-3.

Hou:2022:ISS

[HS22] Jiawei Hou and Yunquan Song. Interquantile shrinkage in spatial additive autoregressive models. TEST, 31(4):1030– 1057, December 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00811-4.

Hall:2010:EIV

[HSC10] Andreia Hall, Manuel Scotto, and João Cruz. Extremes of integer-valued moving average sequences. TEST, 19(2):359– 374, August 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0158-6.

Hsiao:2007:PDA

[Hsi07a] Cheng Hsiao. Panel data analysis — advantages and challenges. TEST, 16(1):1–22, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-007-0046-

x. See comments [Are07, Bal07c, Sic07, WM07, Mai07, Ner07, PS07, Shi07] and rejoinder [Hsi07b].

Hsiao:2007:RPD

[Hsi07b] Cheng Hsiao. Rejoinder on: “Panel data analysis — advantages and challenges”. TEST, 16(1):56–57, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-007-0055-9. See [Hsi07a, Are07, Bal07c, Sic07, WM07, Mai07, Ner07, PS07, Shi07].

Hwang:2005:LPD

[HSK05] Hyungtae Hwang, Beongsoo So, and Yongdai Kim. On limiting posterior distributions. TEST, 14(2):567–580, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595418.

Huckemann:2021:CRA

[Huc21] Stephan F. Huckemann. Comments on: Recent advances in directional statistics. TEST, 30(1):71–75, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00764-0. See [PGP21a, PGP21b].

Hooker:2014:BMR

[HV14] Giles Hooker and Anand N. Vidyashankar. Bayesian model robustness via disparities. TEST, 23(3):556–584, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0360-z.

Han:2023:CNE

[HW23] Bo Han and Xiaoguang Wang. Comments on: Nonparametric estimation in mixture cure models with covariates. TEST, 32(2):496–498, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

023-00853-2. See [LCPJ23a, LCPJ23b].

He:2013:GFT

[HX13] Daojiang He and Xingzhong Xu. A goodness-of-fit testing approach for normality based on the posterior predictive distribution. TEST, 22(1):1–18, March 2013. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-012-0282-6.

Hyndman:2025:CEF

[Hyn25] Rob J. Hyndman. Comments on: Exploratory functional data analysis. TEST, 34(2):459–482, June 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-025-00963-z. See [QDG25a], comments

[Gol25, Lil25, LP25, PS25], and rejoinder [QDG25b].

Huang:2022:IDE

[HZY22] Kun Huang, Sijie Zheng, and Lijian Yang. Inference for dependent error functional data with application to eventrelated potentials. TEST, 31(4):1100–1120, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00820-3.

Irwin:2002:STN

[ICJ02] Mark E. Irwin, Noel Cressie, and Gardar Johannesson. Spatial-temporal nonlinear filtering based on hierarchical statistical models. TEST, 11(2):249–302, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595708.

Iyengar:2002:SMC

[ID02] Malini Iyengar and Dipak K. Dey. A semiparametric model for compositional data analysis in presence of covariates on the simplex. TEST, 11(2):303–315, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595709.

Ion:2023:SIB

[IKvdH23] Roxana A. Ion, Chris A. J. Klaassen, and Edwin R. van den Heuvel. Sharp inequalities of Bienaymé–Chebyshev and Gauß type for possibly asymmetric intervals around the mean. TEST, 32(2):566–601, June 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00844-9.

Ibrahim:2009:MDM

[IM09a] Joseph G. Ibrahim and Geert Molenberghs. Missing data methods in longitudinal studies: a review. TEST, 18(1):1–43, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0138-x. See comments [Uga09, Lit09,

DW09, Hog09a, KC09] and rejoinder [IM09b].

Ibrahim:2009:RMD

[IM09b] Joseph G. Ibrahim and Geert Molenberghs. Rejoinder on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):68–75, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0144-z. See

[IM09a, Uga09, Lit09, DW09, Hog09a, KC09].

Iannario:2016:RIC

[IMP16] Maria Iannario, Anna Clara Monti, and Domenico Piccolo. Robustness issues for cub models. TEST, 25(4):731–750, December 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0493-3.

Irony:1994:MUD

[IP94] Telba Z. Irony and Carlos A. B. Pereira. Motivation for the use of discrete distributions in quality assurance. TEST, 3 (2):181–193, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562700.

Ibacache-Pulgar:2013:SAM

[IPPC13] Germán Ibacache-Pulgar, Gilberto A. Paula, and Francisco José A. Cysneiros. Semiparametric additive models under symmetric distributions. TEST, 22(1):103–121, March 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0309-z.

Ibarrola:2011:FTT

[IPR11] Ricardo Vélez Ibarrola and Tomás Prieto-Rumeau. De Finetti-type theorems for nonexchangeable 0–1 random variables. TEST, 20(2):293–310, August 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0192-4.

Iglesias:1998:CMS

[IPT98] Pilar L. Iglesias, Carlos A. B. Pereira, and Nelson I. Tanaka. Characterizations of multivariate spherical distributions in l_{∞}-norm. TEST, 7(2):307–324, December 1998.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02565115.

Ishwaran:2011:CNI

[Ish11] Hemant Ishwaran. Comments on: “Nonparametric inference based on panel count data”. TEST, 20(1):48–53, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0226-y. See [CH09a] and rejoinder

[CH09a].

Ibarrola:2005:MAB

[IV05] Pilar Ibarrola and Ricardo Vélez. Multi-armed bandit processes with optimal selection of the operating times. TEST, 14(1):239–255, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595405.

Jaheen:2003:BPU

[Jah03] Zeinhum F. Jaheen. Bayesian prediction under a mixture of two-component Gompertz lifetime model. TEST, 12(2):413– 426, December 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595722.

Jaruskova:2015:DNS

[Jar15] Daniela Jarusková. Detecting non-simultaneous changes in means of vectors. TEST, 24(4):681–700, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0429-3.

Janssen:2011:CIM

[JD11] Paul Janssen and Luc Duchateau. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):271–275, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0253-3. See

[GNZ11a] and rejoinder [GNZ11b].

Jin:2014:TNM

[JFCZ14] Hua Jin, Xiaobo Feng, Mingming Chen, and Chenling Zhang. Two new methods for non-inferiority testing of the ratio in matched-pair setting. TEST, 23(4):691–707, December 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0374-6.

Jimenez-Gamero:2020:CTM

[JG20] M. Dolores Jiménez-Gamero. Comments on: Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):893–897, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00742-y. See [EH20b, EH20a].

Jimenez-Gamero:2022:TEL

[JGCRJJ22] M. D. Jiménez-Gamero, M. Cousido-Rocha, and F. Jiménez-Jiménez. Testing the equality of a large number of populations. TEST, 31(1):1–21, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
021-00769-9.

Jimenez-Gamero:2024:TPL

[JGdUÁ24] M. D. Jiménez-Gamero and J. de Uña-Álvarez. Testing Poissonity of a large number of populations. TEST, 33(1):81– 105, March 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00883-w.

Jimenez-Gamero:2020:GFT

[JGLM20] M. Dolores Jiménez-Gamero, Sangyeol Lee, and Simos G. Meintanis. Goodness-of-fit tests for parametric specifications of conditionally heteroscedastic models. TEST, 29(3):682– 703, September 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-019-00676-0.

Jimenez-Gamero:2018:ECF

[JGMRMG18] M. D. Jiménez-Gamero, J. L. Moreno-Rebollo, and J. A. Mayor-Gallego. On the estimation of the characteristic function in finite populations with applications. TEST, 27(1): 95–121, March 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0514-2.

Jimenez-Gamero:2005:IDR

[JGMRMPMR05] M. Dolores Jiménez-Gamero, Juan Luis Moreno-Rebollo, Juan M. Muñoz-Pichardo, and Ana M. Muñoz-Reyes. Influence diagnostics in regression with complex designs through conditional bias. TEST, 14(2):515–542, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595416.

Jiang:2006:MMP

[JL06] Jiming Jiang and P. Lahiri. Mixed model prediction and small area estimation. TEST, 15(1):1–96, June 2006. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595419.

Jongbloed:2021:SES

[JMS21] Geurt Jongbloed, Kimberly S. McGarrity, and Jilt Sietsma. Smooth estimation of size distributions in an oriented cylinder model. TEST, 30(2):505–526, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00729-9.

Jiang:2009:CGF

[JN09] Jiming Jiang and Thuan Nguyen. Comments on: “Goodnessof-fit-tests in mixed models”. TEST, 18(2):248–255, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0151-0. See [CH09a] and rejoinder

[CH09a].

Jolliffe:2008:CAP

[Jol08] Ian T. Jolliffe. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):249– 250, August 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0120-z.

Jones:1996:CRL

[Jon96] M. C. Jones. On close relations of local likelihood density estimation. TEST, 5(2):345–356, December 1996. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562622.

Jones:2004:FDA

[Jon04] M. C. Jones. Families of distributions arising from distributions of order statistics. TEST, 13(1):1–43, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02602999.

Joshi:2007:CPC

[Jos07] Prakash C. Joshi. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):282–283, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0069-3. See [Bal07a] and rejoinder

[Bal07b].

Jimenz:2001:REM

[JS01] Raúl Jiménz and Yongzhao Shao. On robustness and efficiency of minimum divergence estimators. TEST, 10(2):241– 248, December 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595695.

Jiang:2022:SIC

[JS22] Rong Jiang and Mengxian Sun. Single-index composite quantile regression for ultra-high-dimensional data. TEST, 31(2): 443–460, June 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00785-9.

Janssen:2016:BEC

[JSV16] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Bernstein estimation for a copula derivative with application to conditional distribution and regression functionals. TEST, 25(2):351–374, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0459-x.

Jiang:2022:GFT

[JT22] Jiming Jiang and Mahmoud Torabi. Goodness-of-fit test with a robustness feature. TEST, 31(1):76–100, March 2022.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00772-0.

Jones:1998:CSP

[JV98] M. C. Jones and S. K. Vines. Choosing the smoothing parameter for unordered multinomial data. TEST, 7(2):413– 426, December 1998. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02565121.

Jankova:2017:HCR

[JvdG17] Jana Janková and Sara van de Geer. Honest confidence regions and optimality in high-dimensional precision matrix estimation. TEST, 26(1):143–162, March 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0503-5.

Kalogridis:2022:AMT

[Kal22] Ioannis Kalogridis. Asymptotics for m-type smoothing splines with non-smooth objective functions. TEST, 31(2): 373–389, June 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00782-y.

Karagrigoriou:2000:AEO

[Kar00] Alex Karagrigoriou. Asymptotically efficient order selection in nonstationary AR processes. TEST, 9(2):371–391, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595741.

Kuchibhotla:2017:AMD

[KB17] Arun Kumar Kuchibhotla and Ayanendranath Basu. On the asymptotics of minimum disparity estimation. TEST, 26 (3):481–502, September 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0520-4.

Kamps:2007:CPC

[KC07] Udo Kamps and Erhard Cramer. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16 (2):271–275, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0066-6. See

[Bal07a] and rejoinder [Bal07b].

Keles:2008:CAB

[KC08] Sündüz Keles and Hyonho Chun. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):36–39, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0104-z.

Kenward:2009:CMD

[KC09] Michael G. Kenward and James R. Carpenter. Comments on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):65–67, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0143-0. See

[IM09a] and rejoinder [IM09b].

Kalogridis:2023:REE

[KCV23] Ioannis Kalogridis, Gerda Claeskens, and Stefan Van Aelst. Robust and efficient estimation of nonparametric generalized linear models. TEST, 32(3):1055–1078, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00866-x.

Kedem:2012:CSR

[Ked12] Benjamin Kedem. Comments on: “Some recent theory for autoregressive count time series”. TEST, 21(3):467–468, September 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0302-6. See [Tjø12b] and

rejoinder [Tjø12a].

Kadane:1993:SBR

[KGP^{+}93] Joseph B. Kadane, Javier Girón, Daniel Peña, Peter Fishburn, Simon French, D. V. Lindley, Giovanni Parmigiani, and Robert L. Winkler. Several Bayesians: a review. TEST, 2(1–2):1–32, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562668.

Kirch:2014:CES

[Kir14] Claudia Kirch. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):270– 275, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0377-3.

Kneib:2019:MRL

[KKLU19a] Thomas Kneib, Nadja Klein, Stefan Lang, and Nikolaus Umlauf. Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1):1–39, March 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-019-00631-z. See comments [Goi19, Rei19, Sch19, SRHD19] and rejoinder [KKLU19b].

Kneib:2019:RMR

[KKLU19b] Thomas Kneib, Nadja Klein, Stefan Lang, and Nikolaus Umlauf. Rejoinder on: Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1):55– 59, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00636-8. See [KKLU19a,

Goi19, Rei19, Sch19, SRHD19].

Koudou:2014:ECS

[KL14] Angelo Efoevi Koudou and Christophe Ley. Efficiency combined with simplicity: new testing procedures for Generalized Inverse Gaussian models. TEST, 23(4):708–724, December 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0378-2.

Klar:2005:TEA

[Kla05] Bernhard Klar. Tests for exponentiality against the M and LM–Classes of life distributions. TEST, 14(2):543–565, December 2005. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595417.

Kong:2017:SSR

[KLYZ17] Xin-Bing Kong, Zhi Liu, Yuan Yao, and Wang Zhou. Sure screening by ranking the canonical correlations. TEST, 26(1):46–70, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0497-z.

Kiwitt:2013:NTI

[KN13] Sebastian Kiwitt and Natalie Neumeyer. A note on testing independence by a copula-based order selection approach. TEST, 22(1):62–82, March 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0294-2.

Kneib:2020:CIC

[Kne20] Thomas Kneib. Comments on: Inference and computation with Generalized Additive Models and their extensions. TEST, 29(2):351–353, June 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-020-00713-3;   http://link.springer.com/
content/pdf/10.1007/s11749-020-00713-3.pdf.  See

[Woo20a] and rejoinder [Woo20b].

Kloodt:2021:STS

[KNV21] Nick Kloodt, Natalie Neumeyer, and Ingrid Van Keilegom. Specification testing in semi-parametric transformation models. TEST, 30(4):980–1003, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00756-0.

Kokonendji:1994:EFV

[Kok94] Célestin C. Kokonendji. Exponential families with variance functions in \sqrt{ΔP} (\sqrt{Δ}): Seshadri’s class: Seshadri’s class. TEST, 3(2):123–172, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02562698.

Kokoszka:2014:CES

[Kok14] Piotr Kokoszka. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):276– 278, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0371-9;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0371-
9.pdf.

Kong:2013:DAR

[Kon13] Xin-Bing Kong. A direct approach to risk approximation for vast portfolios under gross-exposure constraint using high-frequency data. TEST, 22(4):647–669, November 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0337-3.

Kott:2019:CDS

[Kot19] Phillip S. Kott. Comments on: Deville and Särndal’s Calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1066–1067, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00686-y. See [DT19a, DT19b].

Koudou:1998:LBP

[Kou98] Angelo Efoévi Koudou. Lancaster bivariate probability distributions with Poisson, negative binomial and gamma margins. TEST, 7(1):95–110, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02565104.

Kerkyacharian:2000:TAM

[KPB^{+}00] Gérard Kerkyacharian, Dominique Picard, Lucien Birgé, Peter Hall, Oleg Lepski, Enno Mammen, Alexandre Tsybakov, G. Kerkyacharian, and D. Picard. Thresholding algorithms, maxisets and well-concentrated bases. TEST, 9(2):283–344, December 2000. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595738.

Kato:2022:TCD

[KPJ22] Shogo Kato, Arthur Pewsey, and M. C. Jones. Tractable circula densities from Fourier series. TEST, 31(3):595–618, September 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00790-y.

Kratschmer:2006:IRF

[Krä06] Volker Krätschmer. Integrals of random fuzzy sets. TEST, 15(2):433–469, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02607061.

Krauczi:2009:SQC

[Kra09] Éva Krauczi. A study of the quantile correlation test for normality. TEST, 18(1):156–165, May 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0074-6.

Kundu:2007:CPC

[Kun07] Debasis Kundu. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):276–278, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0067-5. See [Bal07a] and rejoinder

[Bal07b].

Kalogridis:2025:RDE

[KV25] Ioannis Kalogridis and Stefan Van Aelst. Resistant dispersion estimation for nonparametric regression. TEST, 34(3):612– 659, September 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-025-00974-w.

Kang:2023:AZO

[KWWZ23] Yao Kang, Shuhui Wang, Dehui Wang, and Fukang Zhu. Analysis of zero-and-one inflated bounded count time series with applications to climate and crime data. TEST, 32(1): 34–73, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00825-y.

Liu:2005:AOC

[LA05] Ivy Liu and Alan Agresti. The analysis of ordered categorical data: an overview and a survey of recent developments. TEST, 14(1):1–73, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595397.

Lazaro:2020:ABI

[LAGR20] E. Lázaro, C. Armero, and V. Gómez-Rubio. Approximate Bayesian inference for mixture cure models. TEST, 29(3):750–767, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-019-00679-x.

Lambert:2011:CIM

[Lam11] Philippe Lambert. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):284–286, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0256-0. See [GNZ11a] and rejoinder

[GNZ11b].

Lambert:2023:CNE

[Lam23] Philippe Lambert. Comments on: Nonparametric estimation in mixture cure models with covariates. TEST, 32(2): 506–509, June 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00860-3. See [LCPJ23a,

LCPJ23b].

Lemonte:2018:NLB

[LB18] Artur J. Lemonte and Jorge L. Bazán. New links for binary regression: an application to coca cultivation in Peru. TEST, 27(3):597–617, September 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0563-1.

Lopez-Blazquez:2013:DTR

[LBSM13] F. López-Blázquez and B. Salamanca-Miño. Distribution theory of δ-record values. Case δ ≤= 0. TEST, 22(4):715– 738, November 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0340-8.

Lopez-Blazquez:2015:DTR

[LBSM15] Fernando López-Blázquez and Begoña Salamanca-Miño. Distribution theory of δ-record values: case δ ≥ 0. TEST, 24 (3):558–582, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0424-0.

Lopez-Blazquez:2001:DDW

[LBW01] Fernando López-Blázquez and Jack Wesolowski. Discrete distributions for which the regression of the first record on the second is linear. TEST, 10(1):121–131, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595827.

Lui:2003:PSS

[LC03] Kung-Jong Lui and William G. Cumberland. Power and sample size calculation for 2×2 tables under multinomial sampling with random loss. TEST, 12(1):141–152, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595816.

Luceno:2006:RIF

[LC06] Alberto Luceno and Antonio S. Cofiño. The random intrinsic fast initial response of two-sided CUSUM charts. TEST, 15 (2):505–524, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02607064.

Louzada:2012:MTS

[LC12] Francisco Louzada and Juliana Cobre. A multiple time scale survival model with a cure fraction. TEST, 21(2):355–368, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0247-1.

Lima:2019:RSI

[LCB19] Italo R. Lima, Guanqun Cao, and Nedret Billor. Robust simultaneous inference for the mean function of functional data. TEST, 28(3):785–803, September 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0598-y.

Lopez-Cheda:2017:NLE

[LCJC17] Ana López-Cheda, M. Amalia Jácome, and Ricardo Cao. Nonparametric latency estimation for mixture cure models.

TEST, 26(2):353–376, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0515-1.

Lopez-Cheda:2023:NEM

[LCPJ23a] Ana López-Cheda, Yingwei Peng, and María Amalia Jácome. Nonparametric estimation in mixture cure models with covariates. TEST, 32(2):467–495, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

022-00840-z. See comments [HW23, Cao23, Lam23, SSGM23] and rejoinder [LCPJ23b].

Lopez-Cheda:2023:RNE

[LCPJ23b] Ana López-Cheda, Yingwei Peng, and María Amalia Jácome. Rejoinder on: Nonparametric estimation in mixture cure models with covariates. TEST, 32(2):513–520, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00871-0. See [HW23, Cao23, Lam23,

SSGM23].

Lindley:1993:OAS

[LD93] Dennis V. Lindley and John J. Deely. Optimal allocation in stratified sampling with partial information. TEST, 2 (1–2):147–160, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562673.

Larose:1996:WDV

[LD96] D. T. Larose and D. K. Dey. Weighted distributions viewed in the context of model selection: A Bayesian perspective. TEST, 5(1):227–246, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562690.

Lopez-Diaz:2018:SOA

[LDLDMF18] María Concepción López-Díaz, Miguel López-Díaz, and Sergio Martínez-Fernández. Stochastic orders to approach investments in condor financial derivatives. TEST, 27(1):122– 146, March 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0537-3.

Li:2020:ALM

[LDR20a] Hang Li, Enrique Del Castillo, and George Runger. On active learning methods for manifold data. TEST, 29(1):1– 33, March 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00694-y. See comments

[Gho20, GYP20] and rejoinder [LDR20b].

Li:2020:RAL

[LDR20b] Hang Li, Enrique Del Castillo, and George Runger. Rejoinder on: “On active learning methods for manifold data”. TEST, 29(1):42–49, March 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-019-00697-9. See [GYP20, Gho20, LDR20a].

Liang:2011:LPE

[LdUÁdCIP11] Han-Ying Liang, Jacobo de Uña-Álvarez, and María del Carmen Iglesias-Pérez. Local polynomial estimation of a conditional mean function with dependent truncated data. TEST, 20(3):653–677, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0234-6.

Liang:2012:APC

[LdUÁdCIP12] Han-Ying Liang, Jacobo de Uña-Álvarez, and María del Carmen Iglesias-Pérez. Asymptotic properties of conditional distribution estimator with truncated, censored and dependent data. TEST, 21(4):790–810, December 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-012-0281-7.

Lago:2025:KST

[LdUÁPF25] Adrián Lago, Jacobo de Uña-Álvarez, and Juan Carlos Pardo-Fernández. A Kolmogorov–Smirnov-type test for the two-sample problem with left-truncated data. TEST, 34(1): 69–90, March 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00948-4.

Lefebvre:2003:FPP

[Lef03] Mario Lefebvre. First-passage problems for degenerate twodimensional diffusion processes. TEST, 12(1):125–139, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595815.

Lin:2025:THD

[LGL25] Yinan Lin, Zijian Guo, and Zhenhua Lin. Testing high-dimensional mediation effect with arbitrary exposuremediator coefficients. TEST, 34(3):509–529, September 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-025-00971-z.

Lv:2019:SEL

[LGW19] Jing Lv, Chaohui Guo, and Jibo Wu. Smoothed empirical likelihood inference via the modified Cholesky decomposition for quantile varying coefficient models with longitudinal data. TEST, 28(3):999–1032, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0616-0.

Li:2017:TLC

[LHB^{+}17] Huiqin Li, Jiang Hu, Zhidong Bai, Yanqing Yin, and Kexin Zou. Test on the linear combinations of mean vectors in high-dimensional data. TEST, 26(1):188–208, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0505-3.

Li:2008:CABa

[Li08a] Bing Li. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):19–21, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0099-5.

Li:2008:CABb

[Li08b] Lexin Li. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):22–24, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0100-3.

Li:2017:INM

[Li17] Zaixing Li. Inference of nonlinear mixed models for clustered data under moment conditions. TEST, 26(4):759–781, December 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0532-8.

Loschi:2002:CUE

[LIAV02] Rosangela Helena Loschi, Pilar Loreto Iglesias, and Rinaldo Boris Arellano-Valle. Conditioning on uncertain event: Extensions to Bayesian inference. TEST, 11(2):365–383, December 2002. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595712.

Liebscher:2012:MCP

[Lie12] Eckhard Liebscher. Model checks for parametric regression models. TEST, 21(1):132–155, March 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0239-1.

Lillo:2025:CEF

[Lil25] Rosa E. Lillo. Comments on: Exploratory functional data analysis. TEST, 34(2):488–489, June 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00966-w. See [QDG25a], comments

[Hyn25, Gol25, LP25, PS25], and rejoinder [QDG25b].

Lindgren:2015:CCS

[Lin15] Finn Lindgren. Comments on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1):35– 44, March 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0417-z. See [BCS15a] and

rejoinder [BCS15b].

Little:2009:CMD

[Lit09] Roderick J. Little. Comments on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):47–50, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0140-3. See [IM09a] and rejoinder

[IM09b].

Linton:2010:ICS

[LJC10] Oliver B. Linton and David T. Jacho-Chávez. On internally corrected and symmetrized kernel estimators for nonparametric regression. TEST, 19(1):166–186, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0145-y.

Lu:2019:PFP

[LJW^{+}19] Hezhi Lu, Hua Jin, Zhining Wang, Chao Chen, and Ying Lu. Prior-free probabilistic interval estimation for binomial proportion. TEST, 28(2):522–542, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0588-0.

Lazecka:2023:ACR

[LKM23] Małgorzata Lazȩcka, Bartosz Kołodziejek, and Jan Mielniczuk. Analysis of conditional randomisation and permutation schemes with application to conditional independence testing. TEST, 32(4):1459–1478, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00878-7.

Li:2009:CRE

[LL09] Gang Li and Xuyang Lu. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18(3):463– 467, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0165-7. See [CV09b] and

[CV09a].

Lin:2023:GDD

[LL23] Lu Lin and Feng Li. Global debiased DC estimations for biased estimators via pro forma regression. TEST, 32(2): 726–758, June 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00850-5.

Lin:2014:BRT

[LLG14] Liang-Ching Lin, Sangyeol Lee, and Meihui Guo. The Bickel–Rosenblatt test for continuous time stochastic volatility models. TEST, 23(1):195–218, March 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0347-1.

Li:2024:MCG

[LLHL24] Xinmin Li, Haozhe Liang, Wolfgang Härdle, and Hua Liang. Model checking for generalized partially linear models. TEST, 33(2):361–378, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00897-4.

Liu:2021:SMU

[LLJ21] Xiaohui Liu, Yuanyuan Li, and Jiming Jiang. Simple measures of uncertainty for model selection. TEST, 30(3):673– 692, September 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00737-9.

Liu:2024:TSS

[LLLZ24] Yang Liu, Yukun Liu, Pengfei Li, and Riquan Zhang. Two-step semiparametric empirical likelihood inference from capture-recapture data with missing covariates. TEST, 33 (3):786–808, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00921-1.

Liu:2024:NSD

[LLR^{+}24] Tianqing Liu, Danning Li, Fengjiao Ren, Jianguo Sun, and Xiaohui Yuan. A new sufficient dimension reduction method via rank divergence. TEST, 33(3):921–950, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00929-7.

Lee:2018:ANP

[LLT18] Youngmi Lee, Sangyeol Lee, and Dag Tjøstheim. Asymptotic normality and parameter change test for bivariate Poisson INGARCH models. TEST, 27(1):52–69, March 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0510-6.

Li:2024:MPB

[LLTS24] Xun-Jian Li, Shuang Li, Guo-Liang Tian, and Jianhua Shi. Modeling paired binary data by a new bivariate Bernoulli model with flexible beta kernel correlation. TEST, 33(4): 1180–1224, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00947-5.

Lombardia:2021:SMD

[LLVR21] María José Lombardía, Esther López-Vizcaíno, and Cristina Rueda. Selection model for domains across time: application to labour force survey by economic activities. TEST, 30(1): 228–254, March 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00712-4.

Li:2014:ELL

[LLZZ14] Zhouping Li, Yuanyuan Lin, Guoliang Zhou, and Wang Zhou. Empirical likelihood for least absolute relative error regression. TEST, 23(1):86–99, March 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-013-0343-5.

Lahiri:2011:CSW

[LM11] S. N. Lahiri and S. Mukhopadhyay. Comments on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20(3):491–496, November 2011. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-011-0273-z. See [DPR11b] and rejoinder [DPR11a].

Lasinio:2018:CPM

[LM18] Giovanna Jona Lasinio and Gianluca Mastrantonio. Comments on: “Process modeling for slope and aspect with application to elevation data maps”. TEST, 27(4):776–777, December 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0622-2;   http://link.
springer.com/content/pdf/10.1007/s11749-018-0622-

2.pdf. See [WBG18a] and rejoinder [WBG18b].

Liu:2025:OBO

[LMJ25] Yuzi Liu, Haiqiang Ma, and Jiming Jiang. A one-bring-one route for assessing the uncertainty of small area estimation in nested-error regression models. TEST, 34(2):361–382, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-025-00965-x.

Locantore:1999:RPC

[LMS^{+}99] N. Locantore, J. S. Marron, D. G. Simpson, N. Tripoli, J. T. Zhang, K. L. Cohen, Graciela Boente, Ricardo Fraiman, Babette Brumback, Christophe Croux, Jianqing Fan, Alois Kneip, John I. Marden, Daniel Peña, Javier Prieto, and et al. Robust principal component analysis for functional data. TEST, 8(1):1–73, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595862.

Loffler:2017:CHD

[LN17] Matthias Löffler and Richard Nickl. Comments on: “High-dimensional simultaneous inference with the bootstrap”. TEST, 26(4):731–733, December 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0558-y;   http://link.springer.com/
content/pdf/10.1007/s11749-017-0558-y.pdf. See [DBZ17a]

and rejoinder [DBZ17b].

Liu:2023:SIS

[LNZ23] Xiumin Liu, Lu Niu, and Junlong Zhao. Statistical inference on the significance of rows and columns for matrix-valued data in an additive model. TEST, 32(3):785–828, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00852-3.

Loperfido:2008:MML

[Lop08] Nicola Loperfido. Modeling maxima of longitudinal contralateral observations. TEST, 17(2):370–380, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0037-3.

Loperfido:2010:CTS

[Lop10] Nicola Loperfido. Canonical transformations of skewnormal variates. TEST, 19(1):146–165, May 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0146-x.

Loperfido:2024:TEP

[Lop24] Nicola Loperfido. Tensor eigenvectors for projection pursuit. TEST, 33(2):453–472, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00902-w.

Louani:2005:UDL

[Lou05] Djamal Louani. Uniform L_{1}-distance large deviations in nonparametric density estimation. TEST, 14(1):75–98, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595398.

Li:2018:DRA

[LP18] Weiyu Li and Valentin Patilea. A dimension reduction approach for conditional Kaplan–Meier estimators. TEST, 27(2):295–315, June 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0546-2.

Lopez-Pintado:2025:CEF

[LP25] Sara Lopez-Pintado. Comments on: Exploratory functional data analysis. TEST, 34(2):490–493, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00968-8. See [QDG25a], comments

[Hyn25, Gol25, Lil25, PS25], and rejoinder [QDG25b].

Launay:2015:CIH

[LPL15] Tristan Launay, Anne Philippe, and Sophie Lamarche. Construction of an informative hierarchical prior for a small sample with the help of historical data and application to electricity load forecasting. TEST, 24(2):361–385, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0416-0.

Li:2011:ELC

[LPQ11] Deyuan Li, Liang Peng, and Yongcheng Qi. Empirical likelihood confidence intervals for the endpoint of a distribution function. TEST, 20(2):353–366, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0204-4.

Llatas:1997:FPB

[LQR97] I. Llatas, A. J. Quiroz, and J. M. Renóm. A fast permutationbased algorithm for block clustering. TEST, 6(2):397–418, December 1997. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02564706.

Lupparelli:2012:CSR

[LR12] Monia Lupparelli and Alberto Roverato. Comments on: “Sequences of regressions and their independences”. TEST, 21(2):262–264, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0286-2. See

[WS12b] and rejoinder [WS12a].

Lee:2009:MDP

[LS09] Sangyeol Lee and Junmo Song. Minimum density power divergence estimator for GARCH models. TEST, 18(2):316– 341, August 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0093-y.

Lockhart:2017:CHD

[LS17] Richard A. Lockhart and Richard J. Samworth. Comments on: “High-dimensional simultaneous inference with the bootstrap”. TEST, 26(4):734–739, December 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0555-1;   http://link.springer.com/
content/pdf/10.1007/s11749-017-0555-1.pdf. See [DBZ17a]

and rejoinder [DBZ17b].

Li:2024:OSQ

[LSW24] Xing Li, Yujing Shao, and Lei Wang. Optimal subsampling for L_{p}-quantile regression via decorrelated score. TEST, 33 (4):1084–1104, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00940-y.

Li:2023:VDP

[LTWY23] Lu Li, Kai Tan, Xuerong Meggie Wen, and Zhou Yu. Variable-dependent partial dimension reduction. TEST, 32(2):521–541, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00841-y.

Lugosi:2010:CPM

[Lug10] Gábor Lugosi. Comment on: “l_{1}-penalization for mixture regression models”. TEST, 19(2):259–263, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0199-x.

Lafferty:2007:CNI

[LW07] John Lafferty and Larry Wasserman. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):453–455, December 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-007-0084-4. See [FJ07a] and rejoinder [FJ07b].

Ledwina:2012:NTS

[LW12] Teresa Ledwina and Grzegorz Wylupek. Nonparametric tests for stochastic ordering. TEST, 21(4):730–756, December 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0278-7.

Ly:2022:BFP

[LW22] Alexander Ly and Eric-Jan Wagenmakers. Bayes factors for peri-null hypotheses. TEST, 31(4):1121–1142, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00819-w.

Li:2024:HOS

[LW24] Tizheng Li and Yuping Wang. Higher-order spatial autoregressive varying coefficient model: estimation and specification test. TEST, 33(4):1258–1299, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00944-8.

Lin:2015:RFA

[LWML15] Tsung-I Lin, Pal H. Wu, Geoffrey J. McLachlan, and Sharon X. Lee. A robust factor analysis model using the restricted skew-t distribution. TEST, 24(3):510–531, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0422-2.

Lai:2023:BDT

[LWZZ23] Jiayu Lai, Xiaoyi Wang, Kaige Zhao, and Shurong Zheng. Block-diagonal test for high-dimensional covariance matrices. TEST, 32(1):447–466, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
022-00842-x.

Liu:2016:CDI

[LX16] Xuhua Liu and Xingzhong Xu. Confidence distribution inferences in one-way random effects model. TEST, 25(1):59– 74, March 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0440-8.

Liu:2017:CHD

[LY17] Hanzhong Liu and Bin Yu. Comments on: “Highdimensional simultaneous inference with the bootstrap”. TEST, 26(4):740–750, December 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-017-0559-

x. See [DBZ17a] and rejoinder [DBZ17b].

Liu:2021:ELI

[LZ21] Rong Liu and Yichuan Zhao. Empirical likelihood inference for generalized additive partially linear models. TEST, 30 (3):569–585, September 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-020-00731-1.

Liu:2023:STP

[LZYZ23] Junmin Liu, Deli Zhu, Luoyao Yu, and Xuehu Zhu. Specification testing of partially linear single-index models: a groupwise dimension reduction-based adaptive-to-model approach. TEST, 32(1):232–262, March 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00833-y.

MacNab:2018:RSR

[Mac18a] Ying C. MacNab. Rejoinder on: “Some recent work on multivariate Gaussian Markov random fields”. TEST, 27(3):554– 569, September 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0608-0;   http://link.
springer.com/content/pdf/10.1007/s11749-018-0608-

0.pdf. See [Mac18b, MB18, SF18, GT18, Mac18a].

MacNab:2018:SRW

[Mac18b] Ying C. MacNab. Some recent work on multivariate Gaussian Markov random fields. TEST, 27(3):497–541, September 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0605-3. See comments [MB18, SF18,

GT18, Mac18a] and rejoinder [Mac18a].

MacNab:2023:CMG

[Mac23] Ying C. MacNab. On coregionalized multivariate Gaussian Markov random fields: construction, parameterization, and Bayesian estimation and inference. TEST, 32(1):263– 293, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00832-z.

Mairesse:2007:CPD

[Mai07] Jacques Mairesse. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):37–41, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0051-0. See [Hsi07a] and rejoinder

[Hsi07b].

Malouche:1998:NEF

[Mal98] Dhafer Malouche. Natural exponential families associated to Pick functions. TEST, 7(2):391–412, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02565120.

Mammen:2007:CNI

[Mam07] Enno Mammen. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16(3):462– 464, December 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0087-1. See [FJ07a] and re-

joinder [FJ07b].

Maronna:2015:CRE

[Mar15a] Ricardo A. Maronna. Comments on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):471–472, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0453-3;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0453-

3.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Maruotti:2015:HNI

[Mar15b] Antonello Maruotti. Handling non-ignorable dropouts in longitudinal data: a conditional model based on a latent Markov heterogeneity structure. TEST, 24(1):84–109, March 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0397-z.

Marron:2019:CDS

[Mar19] J. S. Marron. Comments on: Data science, big data and statistics. TEST, 28(2):342–344, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00644-8. See [GP19a, GP19b].

Mardia:2021:CRA

[Mar21] Kanti V. Mardia. Comments on: Recent advances in directional statistics. TEST, 30(1):59–63, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00760-4. See [PGP21a, PGP21b].

Marandon:2024:CLP

[Mar24] Ariane Marandon. Conformal link prediction for false discovery rate control. TEST, 33(4):1062–1083, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00934-w.

Mateu:2010:CGS

[Mat10] J. Mateu. Comments on: “A general science-based framework for dynamical spatio-temporal models”. TEST, 19 (3):452–455, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0210-6. See

[WH10b] and rejoinder [WH10b].

Mayor:2002:OCS

[May02] José A. Mayor. Optimal cluster selection probabilities to estimate the finite population distribution function under PPS cluster sampling. TEST, 11(1):73–88, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595730.

Martinez-Beneito:2018:CSR

[MB18] Miguel A. Martinez-Beneito. Comments on: “Some recent work on multivariate Gaussian Markov random fields”. TEST, 27(3):542–544, September 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0606-2;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0606-2.pdf. See [Mac18b]

and rejoinder [Mac18a].

Mulder:2021:PII

[MBB21] Joris Mulder, James O. Berger, and M. J. Bayarri. On the prevalence of information inconsistency in normal linear models. TEST, 30(1):103–132, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00704-4.

Mukerjee:2009:CIB

[MC09] Rahul Mukerjee and Ling-Yau Chan. Confidence intervals based on empirical statistics: existence of a probability matching prior and connection with frequentist Bartlett adjustability. TEST, 18(2):271–282, August 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0076-4.

Marques:2011:GNE

[MCA11] Filipe J. Marques, Carlos A. Coelho, and Barry C. Arnold. A general near-exact distribution theory for the most common likelihood ratio test statistics used in multivariate analysis. TEST, 20(1):180–203, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0193-3. See

[MCA24].

Marques:2024:CGN

[MCA24] Filipe J. Marques, Carlos A. Coelho, and Barry C. Arnold. Correction to: A general near-exact distribution theory for the most common likelihood ratio test statistics used in Multivariate Analysis. TEST, 33(4):1300–1301, December 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00927-9. See [MCA11].

McKeague:2009:CRE

[McK09] Ian W. McKeague. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18(3):461–462, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0164-8. See [CV09b] and

[CV09a].

Matos:2016:CME

[MCL16] Larissa A. Matos, Luis M. Castro, and Víctor H. Lachos. Censored mixed-effects models for irregularly observed repeated measures with applications to HIV viral loads. TEST, 25(4):627–653, December 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0486-2.

Molina:2022:EPI

[MCN22] Isabel Molina, Paul Corral, and Minh Nguyen. Estimation of poverty and inequality in small areas: review and discussion. TEST, 31(4):1143–1166, December 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00822-1.

Magalhaes:2003:PNA

[MD03] Fernando Magalhães and Iran R. Dunsmore. Predicting the number of accidents at a road junction. TEST, 12(1):153– 172, June 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595817.

Meintanis:2013:CUR

[Mei13] Simos G. Meintanis. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22 (3):432–436, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0332-8;
http://link.springer.com/content/pdf/10.1007/s11749-
013-0332-8.pdf.

Meintanis:2020:CTM

[Mei20] Simos G. Meintanis. Comments on: Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):898–902, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00743-x. See [EH20b, EH20a].

Mendel:1994:OPB

[Men94a] Max B. Mendel. Operational parameters in Bayesian models. TEST, 3(2):195–206, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562701.

Mendoza:1994:ANU

[Men94b] Manuel Mendoza. Asymptotic normality under transformations. A result with Bayesian applications. TEST, 3(2):173– 180, December 1994. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562699.

Menzefricke:1999:BPG

[Men99] Ulrich Menzefricke. Bayesian prediction in growth-curve models with correlated errors. TEST, 8(1):75–93, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595863.

Mendo:2012:EPI

[Men12] Luis Mendo. Estimation of a probability in inverse binomial sampling under normalized linear-linear and inverselinear loss. TEST, 21(4):656–675, December 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0267-x.

Madruga:2001:BPS

[MEW01] M. Regina Madruga, Luis G. Esteves, and Sergio Wechsler. On the Bayesianity of pereira-stern tests. TEST, 10(2):291– 299, December 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595698.

Martins:2005:MEI

[MF05] A. P. Martins and H. Ferreira. The multivariate extremal index and the dependence structure of a multivariate extreme value distribution. TEST, 14(2):433–448, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595412.

Martins:2014:EPM

[MF14] A. P. Martins and H. Ferreira. Extremal properties of M4 processes. TEST, 23(2):388–408, June 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0358-6.

Martin-Fernandez:2019:CCD

[MF19] J. A. Martín-Fernández. Comments on: Compositional data: the sample space and its structure. TEST, 28(3):653–657, September 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00672-4. See [EPG19a,

EPG19b].

Martinez-Florez:2015:DCP

[MFBG15] Guillermo Martínez-Flórez, Heleno Bolfarine, and Héctor W. Gómez. Doubly censored power-normal regression models with inflation. TEST, 24(2):265–286, June 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0406-2.

Moreno:2008:CBOa

[MG08a] Elías Moreno and F. Javier Girón. Comparison of Bayesian objective procedures for variable selection in linear regression. TEST, 17(3):472–490, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0039-1.

Moreno:2008:CBOb

[MG08b] Elías Moreno and F. Javier Girón. Comparison of Bayesian objective procedures for variable selection in linear regression. TEST, 17(3):491–492, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0095-9;   http://link.springer.com/
content/pdf/10.1007/s11749-008-0095-9.pdf.

Molina:2021:ADI

[MG21] Isabel Molina and Malay Ghosh. Accounting for dependent informative sampling in model-based finite population inference. TEST, 30(1):179–197, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00708-0.

Martins:2004:AHE

[MGN04] M. João Martins, M. Ivette Gopmes, and M. Manuela Neves. Averages of Hill estimators. TEST, 13(1):113–128, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02603003.

Mendoza:2000:BCA

[MGP00] Manuel Mendoza and Eduardo Gutiérrez-Peña. Bayesian conjugate analysis of the Galton–Watson process. TEST, 9(1):149–171, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595856.

Mardia:1998:KKF

[MGRA98] Kanti V. Mardia, Colin Goodall, Edwin J. Redfern, and Francisco J. Alonso. The kriged Kalman filter. TEST, 7 (2):217–282, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02565111.

Meintanis:2018:E

[MHJG18] S. G. Meintanis, M. Husková, and M. D. Jiménez-Gamero. Editorial. TEST, 27(1):1–2, March 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0577-3;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0577-3.pdf.

Manteiga:2020:GFT

[MHSB20] Wenceslao González Manteiga, Cédric Heuchenne, César Sánchez Sellero, and Alessandro Beretta. Goodness-of-fit tests for censored regression based on artificial data points. TEST, 29(2):599–615, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00662-6.

Martin:2003:JSB

[MIR03] Jacinto Martín, David Ríos Insua, and Fabrizio Ruggeri. Joint sensitivity in Bayesian decision theory. TEST, 12(1): 173–194, June 2003. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595818.

Mitchell:1992:EPD

[Mit92] Ann F. S. Mitchell. Estimative and predictive distances. TEST, 1(1):105–121, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02562666.

Molina:2008:BBP

[MJR08] Manuel Molina, Christine Jacob, and Alfonso Ramos. Bisexual branching processes with offspring and mating depending on the number of couples in the population. TEST, 17(2):265–281, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0031-9.

Munk:2009:CGF

[MK09] Axel Munk and Tatyana Krivobokova. Comments on: “Goodness-of-fit-tests in mixed models”. TEST, 18(2):256– 259, August 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0152-z. See [CH09a] and

rejoinder [CH09a].

Mainassara:2014:NST

[MK14] Yacouba Boubacar Maïnassara and Célestin C. Kokonendji. On normal stable Tweedie models and power-generalized variance functions of only one component. TEST, 23(3):585– 606, September 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0363-9.

Mondal:2025:TTT

[MK25] Anjana Mondal and Somesh Kumar. Testing for trend in twoway heteroscedastic ANCOVA models. TEST, 34(3):691– 713, September 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-025-00977-7.

Munda:2016:TDH

[MLDJ16] Marco Munda, Catherine Legrand, Luc Duchateau, and Paul Janssen. Testing for decreasing heterogeneity in a new time-varying frailty model. TEST, 25(4):591–606, December 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0468-9.

Mastrantonio:2016:STC

[MLG16] Gianluca Mastrantonio, Giovanna Jona Lasinio, and Alan E. Gelfand. Spatio-temporal circular models with non-separable covariance structure. TEST, 25(2):331–350, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0458-y.

Matos:2019:HTL

[MLLC19] Larissa A. Matos, Víctor H. Lachos, Tsung-I Lin, and Luis M. Castro. Heavy-tailed longitudinal regression models for censored data: a robust parametric approach. TEST, 28 (3):844–878, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0603-5.

Mentz:2002:RET

[MM02] Raúl P. Mentz and Carlos I. Martínez. Robust estimation in time series. TEST, 11(2):385–404, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595713.

Moreno:2022:BFE

[MM22] Elías Moreno and Carmen Martínez. Bayesian and frequentist evidence in one-sided hypothesis testing. TEST, 31(1): 278–297, March 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00778-8.

Mahmoud:2000:AEB

[MMI00] Mohamed Mahmoud, Nahed A. Mokhlis, and Sahar A. N. Ibrahim. Assessing the error in bootstrap estimates with dependent data. TEST, 9(2):471–486, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595746.

Molina:2004:LBS

[MMR04] Manuel Molina, Manuel Mota, and Alfonso Ramos. Limiting behaviour for superadditive bisexual Galton–Watson processes in varying environments. TEST, 13(2):481–499, December 2004. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595783.

Molina:2008:BEC

[MMR08] Manuel Molina, Manuel Mota, and Alfonso Ramos. Bayesian estimation in the class of bisexual branching processes with population-size dependent mating. TEST, 17(1):179–196, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0029-3.

Monteiro:2021:MIT

[MMS21] Andreia Monteiro, Raquel Menezes, and Maria Eduarda Silva. Modelling informative time points: an evolutionary process approach. TEST, 30(2):364–382, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00722-2.

Moreno:2021:OBM

[MMVP21] Elías Moreno, Carmen Martínez, and Francisco-José Vázquez-Polo. Objective Bayesian model choice for non-nested families: the case of the Poisson and the negative binomial. TEST, 30(1):255–273, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
020-00717-z.

Montanero:2002:ECA

[MNOP02] J. Montanero, A. Nogales, J. A. Oyola, and P. Pérez. Exhaustivity, completeness and almost invariance. TEST, 11 (2):405–411, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595714.

Meintanis:2024:SPM

[MNP24] Simos G. Meintanis, John P. Nolan, and Charl Pretorius. Specification procedures for multivariate stable-paretian laws for independent and for conditionally heteroskedastic data. TEST, 33(2):517–539, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00909-3.

Miguel:1999:BFI

[MO99] Jesús A. Miguel and Pilar Olave. Bootstrapping forecast intervals in ARCH models. TEST, 8(2):345–364, December 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595875.

Molina:2009:CGF

[Mol09] Isabel Molina. Comments on: “Goodness-of-fit-tests in mixed models”. TEST, 18(2):244–247, August 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-009-0150-1. See [CH09a] and rejoinder [CH09a].

Monnier:2011:NRH

[Mon11] Jean-Baptiste Monnier. Nonparametric regression on the hyper-sphere with uniform design. TEST, 20(2):412–446, August 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0233-7.

Moreno:2005:OBM

[Mor05] Elías Moreno. Objective Bayesian methods for one-sided testing. TEST, 14(1):181–198, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595402.

Morales:2014:CST

[Mor14] Domingo Morales. Comments on: “Single- and two-stage cross-sectional and time series benchmarking procedures for small area estimation”. TEST, 23(4):674–679, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0384-4;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0384-

4.pdf. See [PST14b] and rejoinder [PST14a].

Morales:2019:CDS

[Mor19] Domingo Morales. Comments on: Deville and Särndal’s Calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1068–1070, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00684-0. See [DT19a, DT19b].

Moral:2022:CHS

[Mor22] Serafín Moral. Comments on: Hybrid semiparametric Bayesian networks. TEST, 31(2):340–343, June 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00816-z. See [ALB22a, ALB22b].

Moreno:1993:PAB

[MP93] Elías Moreno and Luís Raúl Pericchi. Prior assessments for bands of probability measures: Empirical Bayes analysis. TEST, 2(1–2):101–110, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562670.

Morales:2000:NST

[MP00] Domingo Morales and Leandro Pardo. New smooth test statistics of goodness-of-fit for categorized composite null hypotheses. TEST, 9(1):173–190, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595857.

Morales:2001:SAP

[MP01] Domingo Morales and Leandro Pardo. Some approximations to power functions of φ-divergence tests in parametric models. TEST, 10(2):249–269, December 2001. CODEN ????

ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595696.

Martin:2014:CES

[MP14] Nirian Martín and Leandro Pardo. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):279–282, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0372-8;
http://link.springer.com/content/pdf/10.1007/s11749-
014-0372-8.pdf.

Moler:2000:MQS

[MPS00] José A. Moler, Fernando Plo, and Miguel San Miguel. Minimal quasi-stationary distributions under null R-recurrence. TEST, 9(2):455–470, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595745.

Militino:2003:RTP

[MPU03] Ana F. Militino, M. Blanca Palacios, and M. Dolores Ugarte. Robust trend parameters in a multivariate spatial linear model. TEST, 12(2):445–457, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595724.

Moler:2013:GPU

[MPU13] José Moler, Fernando Plo, and Henar Urmeneta. A generalized Pólya urn and limit laws for the number of outputs in a family of random circuits. TEST, 22(1):46–61, March 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0292-4.

Manzotti:2001:SHQ

[MQ01] Alessandro Manzotti and Adolfo J. Quiroz. Spherical harmonics in quadratic forms for testing multivariate normality. TEST, 10(1):87–104, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595825.

Moreno-Roldan:2007:IDM

[MRMPEG07] D. Moreno-Roldán, J. M. Muñoz-Pichardo, and A. Enguix-González. Influence diagnostics in multiple discriminant analysis. TEST, 16(1):172–187, May 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0007-9.

Mallor:2003:RSS

[MS03] Fermín Mallor and Javier Santos. Reliability of systems subject to shocks with a stochastic dependence for the damages. TEST, 12(2):427–444, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595723.

Majerski:2010:AMP

[MS10] Piotr Majerski and Zbigniew Szkutnik. Approximations to most powerful invariant tests for multinormality against some irregular alternatives. TEST, 19(1):113–130, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0136-4.

Mason:2011:GRU

[MS11] David M. Mason and Jan W. H. Swanepoel. A general result on the uniform in bandwidth consistency of kerneltype function estimators. TEST, 20(1):72–94, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0188-0. See erratum [MS15].

Mason:2015:EGR

[MS15] David M. Mason and Jan W. H. Swanepoel. Erratum to: “A general result on the uniform in bandwidth consistency of kernel-type function estimators”. TEST, 24(1): 205–206, March 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0420-4;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0420-

4.pdf. See [MS11].

Mukhopadhyay:2017:MPV

[MS17] Minerva Mukhopadhyay and Tapas Samanta. A mixture of g-priors for variable selection when the number of regressors grows with the sample size. TEST, 26(2):377–404, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0516-0.

Mondal:2023:TTT

[MSK23] Anjana Mondal, Paavo Sattler, and Somesh Kumar. Testing for trend in two-way crossed effects model under heteroscedasticity. TEST, 32(4):1434–1458, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00879-6.

Matsui:2008:GFT

[MT08] Muneya Matsui and Akimichi Takemura. Goodness-of-fit tests for symmetric stable distributions —- empirical characteristic function approach. TEST, 17(3):546–566, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0045-y.

Mainassara:2023:CEW

[MU23a] Yacouba Boubacar Maïnassara and Eugen Ursu. Correction: Estimating weak periodic vector autoregressive time series. TEST, 32(3):1132–1133, September 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00865-y. See [MU23b].

Mainassara:2023:EWP

[MU23b] Yacouba Boubacar Maïnassara and Eugen Ursu. Estimating weak periodic vector autoregressive time series. TEST, 32(3):958–997, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

023-00859-w. See correction [MU23a].

Muller:2007:CNI

[Mül07] Hans-Georg Müller. Comments on: “Nonparametric inference with generalized likelihood ratio tests”. TEST, 16 (3):450–452, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0083-5. See

[FJ07a] and rejoinder [FJ07b].

Munoz:2014:CST

[Muñ14] M. Pilar Muñoz. Comments on: “Space–time wind speed forecasting for improved power system dispatch”. TEST, 23(1):30–31, March 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0353-y;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0353-y.pdf. See [ZGGX14b] and rejoinder [ZGGX14a].

Murakami:2016:MGF

[Mur16] Hidetoshi Murakami. A moment generating function of a combination of linear rank tests and its asymptotic efficiency. TEST, 25(4):674–691, December 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0490-6.

Muller:2016:CLM

[MvdG16] Patric Müller and Sara van de Geer. Censored linear model in high dimensions. TEST, 25(1):75–92, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-015-0441-7.

Meilan-Vila:2021:NMR

[MVFFP21] Andrea Meilán-Vila, Mario Francisco-Fernández, and Agnese Panzera. Nonparametric multiple regression estimation for circular response. TEST, 30(3):650–672, September 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00736-w.

Meilan-Vila:2020:GFT

[MVOC20] Andrea Meilán-Vila, Jean D. Opsomer, and Rosa M. Crujeiras. A goodness-of-fit test for regression models with spatially correlated errors. TEST, 29(3):728–749, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-019-00678-y.

Marazzi:2019:RCM

[MVYA19] Alfio Marazzi, Marina Valdora, Victor Yohai, and Michael Amiguet. A robust conditional maximum likelihood estimator for generalized linear models with a dispersion parameter. TEST, 28(1):223–241, March 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0624-0.

Morvai:2004:IES

[MW04] Gusztáv Morvai and Benjamin Weiss. Intermittent estimation of stationary time series. TEST, 13(2):525–542, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595785.

Maleki:2019:FCP

[MWAV19] Mohsen Maleki, Darren Wraith, and Reinaldo B. Arellano-Valle. A flexible class of parametric distributions for Bayesian linear mixed models. TEST, 28(2):543–564, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0590-6.

McLachlan:2008:CAB

[MWN08] Geoffrey J. McLachlan, K. Wang, and S. K. Ng. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):43–46, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0106-x.

Ma:2021:SMR

[MWY21] Yanyuan Ma, Shaoli Wang, and Weixin Yao. Semiparametric mixture regression with unspecified error distributions. TEST, 30(2):429–444, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-020-00725-z.

Muller:2010:DRS

[MY10a] Hans-Georg Müller and Wenjing Yang. Dynamic relations for sparsely sampled Gaussian processes. TEST, 19(1):1– 29, May 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0176-4. See comments

[Car10, CGR10, CGR10, GMMC10, Hec10, Hoo10, Sen10, Wan10] and rejoinder [MY10b].

Muller:2010:RDR

[MY10b] Hans-Georg Müller and Wenjing Yang. Rejoinder on: “Dynamic relations for sparsely sampled Gaussian processes. TEST, 19(1):60–67, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0173-7. See

[MY10a, Car10, CGR10, CGR10, GMMC10, Hec10, Hoo10, Sen10, Wan10].

Neves:2007:SPA

[NA07] Cláudia Neves and M. Isabel Fraga Alves. Semi-parametric approach to the Hasofer–Wang and Greenwood statistics in extremes. TEST, 16(2):297–313, August 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0010-1.

Navarro:2023:CBS

[NA23] Jorge Navarro and Jorge M. Arevalillo. On connections between skewed, weighted and distorted distributions: applications to model extreme value distributions. TEST, 32(4): 1307–1335, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00874-x.

Nagaraja:2007:CPC

[Nag07] Haikady N. Nagaraja. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):260–261, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0062-x. See [Bal07a] and rejoinder

[Bal07b].

Nunez-Anton:1999:LDN

[NARPV99] Vicente Núñez-Antón, Juan M. Rodríguez-Póo, and Philippe Vieu. Longitudinal data with nonstationary errors: a nonparametric three-stage approach. TEST, 8(1):201–231, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595870.

Navarro:2016:SCG

[Nav16] Jorge Navarro. Stochastic comparisons of generalized mixtures and coherent systems. TEST, 25(1):150–169, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0443-5.

Novo:2021:SSR

[NAV21] Silvia Novo, Germán Aneiros, and Philippe Vieu. Sparse semiparametric regression when predictors are mixture of functional and high-dimensional variables. TEST, 30(2):481– 504, June 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00728-w.

Novo:2025:SFP

[NAV25] Silvia Novo, Germán Aneiros, and Philippe Vieu. Semifunctional partial linear regression with measurement error: an approach based on k NN estimation. TEST, 34(1):235– 261, March 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00957-3.

Nieto-Barajas:2022:GDS

[NBGP22] Luis Nieto-Barajas and Eduardo Gutiérrez-Peña. General dependence structures for some models based on exponential families with quadratic variance functions. TEST, 31(3):699– 716, September 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00798-4.

Ng:2007:CPC

[NC07] Hon Keung Tony Ng and Ping-Shing Chan. Comments on: “Progressive censoring methodology: an appraisal”. TEST, 16(2):287–289, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-007-0071-9. See

[Bal07a] and rejoinder [Bal07b].

Navarro-Esteban:2021:HDO

[NECA21] P. Navarro-Esteban and J. A. Cuesta-Albertos. Highdimensional outlier detection using random projections. TEST, 30(4):908–934, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-
020-00750-y.

Nerlove:2007:CPD

[Ner07] Marc Nerlove. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):42–46, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0052-z. See [Hsi07a] and rejoinder

[Hsi07b].

Nel:1993:BAM

[NG93] D. G. Nel and P. C. N. Groenewald. A Bayesian approach to the multivariate Behrens–Fisher problem under the assumption of proportional covariance matrices. TEST, 2(1–2):111– 124, December 1993. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562671.

Nigm:2006:BER

[Nig06] E. M. Nigm. Bootstrapping extremes of random variables under power normalization. TEST, 15(1):257–269, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595427.

Nik:2024:MAC

[Nik24] Simon Nik. Marginal analysis of count time series in the presence of missing observations. TEST, 33(4):1105–1128, December 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00938-6.

Nadarajah:2006:GPD

[NK06] Saralees Nadarajah and Samuel Kotz. A generalized Planck distribution. TEST, 15(2):361–374, September 2006. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02607057.

Navarro:2017:CRI

[NLP17] Jorge Navarro, Maria Longobardi, and Franco Pellerey. Comparison results for inactivity times of k-out-of-n and general coherent systems with dependent components. TEST, 26 (4):822–846, December 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0535-5.

Navarro:2020:CCS

[NM20] Jorge Navarro and Julio Mulero. Comparisons of coherent systems under the time-transformed exponential model. TEST, 29(1):255–281, March 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00656-4.

Navarro:2022:SDR

[NPM22] Jorge Navarro, Franco Pellerey, and Julio Mulero. On sums of dependent random lifetimes under the time-transformed exponential model. TEST, 31(4):879–900, December 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00805-2.

Navarro:2010:CCS

[NR10] Jorge Navarro and Rafael Rubio. Comparisons of coherent systems using stochastic precedence. TEST, 19(3):469–486, November 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0207-1.

Navarro:2006:RPS

[NRS06] Jorge Navarro, José M. Ruiz, and Carlos J. Sandoval. Reliability properties of systems with exchangeable components and exponential conditional distributions. TEST, 15(2):471– 484, September 2006. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02607062.

Nachtsheim:2019:CDS

[NS19] Abigael C. Nachtsheim and John Stufken. Comments on: Data science, big data and statistics. TEST, 28(2):345–348, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00643-9. See [GP19a, GP19b].

Nobre:2013:TVC

[NSS13] Juvêncio S. Nobre, Julio M. Singer, and Pranab K. Sen. Utests for variance components in linear mixed models. TEST, 22(4):580–605, November 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0316-8.

OHagan:1997:PIF

[O’H97] A. O’Hagan. Properties of intrinsic and fractional Bayes factors. TEST, 6(1):101–118, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02564428.

Ohtani:1998:MCR

[Oht98] Kazuhiro Ohtani. An MSE comparison of the restricted Stein-rule and minimum mean squared error estimators in regression. TEST, 7(2):361–376, December 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02565118.

Ovalle-Munoz:2024:CLS

[OMRM24a] Diana P. Ovalle-Muñoz and M. Dolores Ruiz-Medina. Correction to: LRD spectral analysis of multifractional functional time series on manifolds. TEST, 33(2):631, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-024-00924-y. See [OMRM24b].

Ovalle-Munoz:2024:LSA

[OMRM24b] Diana P. Ovalle-Muñoz and M. Dolores Ruiz-Medina. LRD spectral analysis of multifractional functional time series on manifolds. TEST, 33(2):564–588, June 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-023-00913-7. See correction [OMRM24a].

Olsen:2021:FDR

[OPV21] Niels Lundtorp Olsen, Alessia Pini, and Simone Vantini. False discovery rate for functional data. TEST, 30(3):784– 809, September 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00751-x.

Ortiz:1998:ODR

[OR98] Isabel Ortiz and Carmelo Rodríguez. Optimal designs with respect to Elfving’s partial minimax criterion for heteroscedastic polynomial regression. TEST, 7(2):347–360, December 1998. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02565117.

Osiewalski:1999:BAN

[Osi99] Jacek Osiewalski. Bayesian analysis of nonlinear regression with equicorrelated elliptical errors. TEST, 8(2):339–344, December 1999. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595874.

Otsu:2009:RQR

[Ots09] Taisuke Otsu. RESET for quantile regression. TEST, 18 (2):381–391, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0097-7.

Ozkut:2023:ROR

[Ozk23] Murat Ozkut. Reliability and optimal replacement policy for a generalized mixed shock model. TEST, 32(3):1038–1054, September 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00864-z.

Ozturk:2019:SIU

[Ozt19] Omer Ozturk. Statistical inference using rank-based poststratified samples in a finite population. TEST, 28(4):1113– 1143, December 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0618-y.

Paas:2014:CLM

[Paa14] Leonard J. Paas. Comments on: “Latent Markov models: a review of a general framework for the analysis of longitudinal data with covariates”. TEST, 23(3):473–477, September 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0387-1;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0387-

1.pdf. See [BFP14a] and rejoinder [BFP14b].

Pawel:2024:PPR

[PAHW24] Samuel Pawel, Frederik Aust, Leonhard Held, and Eric-Jan Wagenmakers. Power priors for replication studies. TEST, 33(1):127–154, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00888-5.

Paparoditis:2011:CSW

[Pap11] Efstathios Paparoditis. Comments on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20(3):497–498, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0274-y. See

[DPR11b] and rejoinder [DPR11a].

Papathomas:2018:CBL

[Pap18] Michail Papathomas. On the correspondence from Bayesian log-linear modelling to logistic regression modelling with g-priors. TEST, 27(1):197–220, March 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0540-8;   http://link.springer.com/
content/pdf/10.1007/s11749-017-0540-8.pdf.

Pardo:2011:CNI

[Par11] M. C. Pardo. Comments on: “Nonparametric inference based on panel count data”. TEST, 20(1):54–57, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0224-0. See [ZBS11a] and rejoinder

[ZBS11b].

Pereira:2004:BPT

[PAT04] Isabel M. S. Pereira and M. Antonia Amaral-Turkman. Bayesian prediction in threshold autoregressive models with exponential white noise. TEST, 13(1):45–64, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02603000.

Pauly:2011:DAQ

[Pau11] Markus Pauly. Discussion about the quality of F -ratio resampling tests for comparing variances. TEST, 20(1):163– 179, May 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0194-2.

Paroli:2020:OBC

[PC20] Roberta Paroli and Guido Consonni. Objective Bayesian comparison of order-constrained models in contingency tables. TEST, 29(1):139–165, March 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-019-00650-
w.

Perrot-Dockes:2023:SIF

[PDBNR23] Marie Perrot-Dockès, Gilles Blanchard, Pierre Neuvial, and Etienne Roquain. Selective inference for false discovery proportion in a hidden Markov model. TEST, 32(4):1365–1391, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00886-7.

Pearl:2003:SCI

[Pea03] Judea Pearl. Statistics and causal inference: a review. TEST, 12(2):281–345, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595718.

Perez:1994:APB

[Pér94] María-Eglée Pérez. An automatic and proper Bayesian estimation analysis of 2×2 contingency tables with one and two fixed margins. TEST, 3(2):101–112, December 1994.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562696.

Pewsey:2018:PBE

[Pew18] Arthur Pewsey. Parametric bootstrap edf-based goodness-of-fit testing for sinh-arcsinh distributions. TEST, 27(1):147– 172, March 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0538-2.

Pewsey:2012:LBI

[PGB12] Arthur Pewsey, Héctor W. Gómez, and Heleno Bolfarine. Likelihood-based inference for power distributions. TEST, 21(4):775–789, December 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0280-0.

Pewsey:2021:RAD

[PGP21a] Arthur Pewsey and Eduardo García-Portugués. Recent advances in directional statistics. TEST, 30(1):1–58, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-021-00759-x. See comments [Mar21, CSN21, Sce21, Huc21] and rejoinder [PGP21b].

Pewsey:2021:RRA

[PGP21b] Arthur Pewsey and Eduardo García-Portugués. Rejoinder on: Recent advances in directional statistics. TEST, 30(1): 76–82, March 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00762-2. See [PGP21a,

Mar21, CSN21, Sce21, Huc21].

Pinson:2014:CST

[Pin14] Pierre Pinson. Comments on: “Space–time wind speed forecasting for improved power system dispatch”. TEST, 23(1): 26–29, March 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0352-z;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0352-

z.pdf. See [ZGGX14b] and rejoinder [ZGGX14a].

Prothero:2024:DIA

[PJH^{+}24a] Jack Prothero, Meilei Jiang, Jan Hannig, Quoc Tran-Dinh, Andrew Ackerman, and J. S. Marron. Data integration via analysis of subspaces (DIVAS). TEST, 33(3):633–674, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00923-z. See comments

[GS24, GMNH24, SZ24, ZS24, ?] and rejoinder [PJH^{+}24b].

Prothero:2024:RDI

[PJH^{+}24b] Jack Prothero, Meilei Jiang, Jan Hannig, Quoc Tran-Dinh, Andrew Ackerman, and J. S. Marron. Rejoinder on: Data integration via analysis of subspaces (DI-VAS). TEST, 33(3):693–696, September 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00950-w. See [PJH^{+}24a].

Paczek:2024:ESI

[PJPW24] Kewin Pa̧czek, Damian Jelito, Marcin Pitera, and Agnieszka Wyłomańska. Estimation of stability index for symmetric α-stable distribution using quantile conditional variance ratios. TEST, 33(1):297–334, March 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00894-7.

Pradhan:2009:PCG

[PK09] Biswabrata Pradhan and Debasis Kundu. On progressively censored generalized exponential distribution. TEST, 18 (3):497–515, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0110-1.

Peterlin:2023:CSD

[PKB23] Jakob Peterlin, Natasa Kejzar, and Rok Blagus. Correct specification of design matrices in linear mixed effects models: tests with graphical representation. TEST, 32(1):184– 210, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00830-1.

Presno:2003:TSS

[PL03] María José Presno and Anna Jusés López. Testing for stationarity in series with a shift in the mean. A Fredholm approach. TEST, 12(1):195–213, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595819.

Pedroso:2023:MMM

[PLQ23] Ricardo C. Pedroso, Rosangela H. Loschi, and Fernando Andrés Quintana. Multipartition model for multiple change point identification. TEST, 32(2):759–783, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00851-4.

Pateiro-Lopez:2013:RSP

[PLRC13] Beatriz Pateiro-López and Alberto Rodríguez-Casal. Recovering the shape of a point cloud in the plane. TEST, 22(1): 19–45, March 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-012-0283-5.

Pei:2025:MCN

[PLZ25] Jian Pei, Yang Lu, and Fukang Zhu. Mixed causalnoncausal count process. TEST, 34(2):??, June 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00960-8.

Papastamoulis:2025:BIC

[PM25] Panagiotis Papastamoulis and Fotios S. Milienos. Bayesian inference and cure rate modeling for event history data. TEST, 34(1):1–27, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00942-w.

Poleto:2011:SAI

[PMPS11] Frederico Z. Poleto, Geert Molenberghs, Carlos Daniel Paulino, and Julio M. Singer. Sensitivity analysis for incomplete continuous data. TEST, 20(3):589–606, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0219-x.

Pellerey:2022:SMD

[PN22] Franco Pellerey and Jorge Navarro. Stochastic monotonicity of dependent variables given their sum. TEST, 31(2):543– 561, June 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00789-5.

Politis:2013:MFM

[Pol13a] Dimitris N. Politis. Model-free model-fitting and predictive distributions. TEST, 22(2):183–221, June 2013. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-013-0317-7. See comments [CA13, FB13, Spe13b, Van13b, Vel13] and rejoinder [Pol13b].

Politis:2013:RMF

[Pol13b] Dimitris N. Politis. Rejoinder on: “Model-free model-fitting and predictive distributions”. TEST, 22(2):240–250, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0323-9. See [Pol13a, CA13, FB13,

Spe13b, Van13b, Vel13].

Pommeret:1996:OPN

[Pom96] D. Pommeret. Orthogonal polynomials and natural exponential families. TEST, 5(1):77–111, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562683.

Perez-Ocon:2000:MML

[PORCGP00] Rafael Pérez-Ocón, J. Eloy Ruiz-Castro, and M. Luz Gámiz-Pérez. Markov models with lognormal transition rates in the analysis of survival times. TEST, 9(2):353–370, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595740.

Paparoditis:2000:LSI

[PP00] Efstathios Paparoditis and Dimitris N. Politis. Large-sample inference in the general AR(1) model. TEST, 9(2):487–509, December 2000. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595747.

Plunkett:2019:TST

[PP19] Amanda Plunkett and Junyong Park. Two-sample test for sparse high-dimensional multinomial distributions. TEST, 28(3):804–826, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0600-8.

Pennoni:2023:CHM

[PPB23] Fulvia Pennoni, Leonard J. Paas, and Francesco Bartolucci. A causal hidden Markov model for assessing effects of multiple direct mail campaigns. TEST, 32(4):1336–1364, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00877-8.

Perone-Pacifico:1996:BRC

[PPST96] M. Perone-Pacifico, G. Salinetti, and L. Tardella. Bayesian robustness on constrained density band classes. TEST, 5 (2):395–409, December 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562625.

Pumi:2024:PCU

[PPT24a] Guilherme Pumi, Taiane Schaedler Prass, and Cleiton Guollo Taufemback. Publisher correction: Unit-Weibull autoregressive moving average models. TEST, 33(1):358–359, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-023-00905-7. See [PPT24b].

Pumi:2024:UWA

[PPT24b] Guilherme Pumi, Taiane Schaedler Prass, and Cleiton Guollo Taufemback. Unit-Weibull autoregressive moving average models. TEST, 33(1):204–229, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

023-00893-8. See publisher correction [PPT24a].

Peng:2010:SJE

[PQ10] Liang Peng and Yongcheng Qi. Smoothed jackknife empirical likelihood method for tail copulas. TEST, 19(3):514–536, November 2010. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0184-4.

Peng:2012:JEL

[PQV12] Liang Peng, Yongcheng Qi, and Ingrid Van Keilegom. Jackknife empirical likelihood method for copulas. TEST, 21(1): 74–92, March 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0236-4.

Philippe:1998:NCP

[PR98] Anne Philippe and Christian P. Robert. A note on the confidence properties of reference priors for the calibration model. TEST, 7(1):147–160, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02565107.

Prieto-Rumeau:2005:CLT

[PR05] Tomás Prieto-Rumeau. Central limit theorem for the estimator of the value of an optimal stopping problem. TEST, 14(1):215–237, June 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595404.

Pumi:2020:KRM

[PRB20] Guilherme Pumi, Cristine Rauber, and Fábio M. Bayer. Kumaraswamy regression model with Aranda–Ordaz link function. TEST, 29(4):1051–1071, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00700-8.

Pietrzak:2016:LTE

[PRSW16] Maciej Pietrzak, Grzegorz A. Rempala, Michal Seweryn, and Jacek Wesolowski. Limit theorems for empirical Rényi entropy and divergence with applications to molecular diversity analysis. TEST, 25(4):654–673, December 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0489-z.

Prus:2020:ODM

[Pru20] Maryna Prus. Optimal designs in multiple group random coefficient regression models. TEST, 29(1):233–254, March 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00654-6.

Park:2007:CPD

[PS07] Byeong U. Park and Seuck Heun Song. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1): 47–51, May 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0053-y. See [Hsi07a] and re-

joinder [Hsi07b].

Paganoni:2025:CEF

[PS25] Anna Maria Paganoni and Laura M. Sangalli. Comments on: Exploratory functional data analysis. TEST, 34(2):494–497, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-025-00969-7. See [QDG25a], comments

[Hyn25, Gol25, Lil25, LP25], and rejoinder [QDG25b].

Pal:2025:SAS

[PSK25] Ayan Pal, Debashis Samanta, and Debasis Kundu. A semiparametric approach for simple step-stress model. TEST, 34(1):91–124, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00951-9.

Pfeffermann:2014:RST

[PST14a] Danny Pfeffermann, Anna Sikov, and Richard Tiller. Rejoinder on: “Single- and two-stage cross-sectional and time series benchmarking procedures for small area estimation”. TEST, 23(4):686–690, December 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0399-x;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0399-x.pdf. See [PST14b, Bel14, HL14, Mor14, SU14, PST14a].

Pfeffermann:2014:STS

[PST14b] Danny Pfeffermann, Anna Sikov, and Richard Tiller. Singleand two-stage cross-sectional and time series benchmarking procedures for small area estimation. TEST, 23(4):631–666, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0398-y. See comments

[Bel14, HL14, Mor14, SU14] and rejoinder [PST14a].

Pesta:2020:NPF

[PW20] Michal Pesta and Martin Wendler. Nuisance-parameter-free changepoint detection in non-stationary series. TEST, 29(2): 379–408, June 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00659-1.

Peng:2009:CRE

[PZ09] Liang Peng and Rongmao Zhang. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18 (3):452–454, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-009-0161-y. See

[CV09b] and [CV09a].

Qu:2025:EFD

[QDG25a] Zhuo Qu, Wenlin Dai, and Marc G. Genton. Exploratory functional data analysis. TEST, 34(2):431–458, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-024-00952-8. See comments [Hyn25, Gol25, Lil25, LP25, PS25] and rejoinder [QDG25b].

Qu:2025:RDE

[QDG25b] Zhuo Qu, Wenlin Dai, and Marc G. Genton. Rejoinder to the discussion on “Exploratory Functional Data Analysis”. TEST, 34(2):498–501, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL

https://link.springer.com/article/10.1007/s11749-

025-00976-8. See [QDG25a] and comments [Hyn25, Gol25, Lil25, LP25, PS25].

Qing:2025:MME

[Qin25] Huan Qing. Mixed membership estimation for categorical data with weighted responses. TEST, 34(3):580–611, September 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-025-00973-x.

Quinlan:2021:CRM

[QQP21] José J. Quinlan, Fernando A. Quintana, and Garritt L. Page. On a class of repulsive mixture models. TEST, 30(2):445– 461, June 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00726-y.

Rabena:1998:DRD

[Rab98] M. T. Rabena. Deriving reference decisions. TEST, 7(1): 161–177, June 1998. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02565108.

Rabier:2014:QTL

[Rab14] Charles-Elie Rabier. On quantitative trait locus mapping with an interference phenomenon. TEST, 23(2):311–329, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0349-z.

Riani:2019:CDS

[RACC19] Marco Riani, Anthony C. Atkinson, Andrea Cerioli, and Aldo Corbellini. Comments on: Data science, big data and statistics. TEST, 28(2):349–352, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00647-5. See [GP19a, GP19b].

Radulovic:1998:SBV

[Rad98] Dragan Radulović. On the subsample bootstrap variance estimation. TEST, 7(2):295–306, December 1998. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02565114.

Radulovic:2004:RTB

[Rad04] Dragan Radulović. Renewal type bootstrap for Markov chains. TEST, 13(1):147–192, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02603005.

Radulovic:2009:ALD

[Rad09] Dragan Radulović. Another look at the disjoint blocks bootstrap. TEST, 18(1):195–212, May 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0091-5.

Rahimov:2009:ADW

[Rah09] I. Rahimov. Asymptotic distributions for weighted estimators of the offspring mean in a branching process. TEST, 18 (3):568–583, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0124-8.

Raissi:2010:ABT

[Raï10] Hamdi Raïssi. Autocorrelation-based tests for vector error correction models with uncorrelated but nonindependent errors. TEST, 19(2):304–324, August 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-009-0156-8.

Rajaratnam:2012:CSR

[Raj12] Bala Rajaratnam. Comment on: “Sequences of regressions and their independences”. TEST, 21(2):268–273, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0288-0. See [WS12b] and rejoinder

[WS12a].

Rao:2001:CRT

[Rao01] B. L. S. Prakasa Rao. Cramer-Rao type integral inequalities for general loss functions. TEST, 10(1):105–120, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595826.

Russo:2012:AVC

[RAP12] Cibele M. Russo, Reiko Aoki, and Gilberto A. Paula. Assessment of variance components in nonlinear mixed-effects elliptical models. TEST, 21(3):519–545, September 2012.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0262-2.

Rashid:1995:RAT

[Ras95] M. Mushfiqur Rashid. Robust analysis of two-way models with repeated measures on both factors. TEST, 4(1): 39–62, June 1995. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02563102.

Rocha:2015:MLM

[RAVL15] Gustavo H. M. A. Rocha, Reinaldo B. Arellano-Valle, and Rosangela H. Loschi. Maximum likelihood methods in a robust censored errors-in-variables model. TEST, 24(4):857– 877, December 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0439-1.

Robert:1994:DWL

[RC94] Christian P. Robert and George Casella. Distance weighted losses for testing and confidence set evaluation. TEST, 3(1): 163–182, June 1994. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562679.

Robert:1996:NBT

[RC96] Christian P. Robert and Nathalie Caron. Noninformative Bayesian testing and neutral Bayes factors. TEST, 5(2):411– 437, December 1996. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562626.

Rocha:2009:BAM

[RCN09] Andréa V. Rocha and Francisco Cribari-Neto. Beta autoregressive moving average models. TEST, 18(3):529– 545, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0112-z. See

erratum [RCN17].

Rocha:2017:EBA

[RCN17] Andréa V. Rocha and Francisco Cribari-Neto. Erratum to: “Beta autoregressive moving average models”. TEST, 26(2):

451–459, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.springer.

com/article/10.1007/s11749-017-0528-4;   https://
link.springer.com/content/pdf/10.1007/s11749-017-

0528-4.pdf. See [RCN09].

Rodriguez-Casal:2022:CSD

[RCSN22a] Alberto Rodríguez-Casal and Paula Saavedra-Nieves. Correction to: Spatial distribution of invasive species: an extent of occurrence approach. TEST, 31(2):442, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00784-w.

Rodriguez-Casal:2022:SDI

[RCSN22b] Alberto Rodríguez-Casal and Paula Saavedra-Nieves. Spatial distribution of invasive species: an extent of occurrence approach. TEST, 31(2):416–441, June 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00783-x.

Rodriguez:2003:DMD

[RCSO03] J. Rodríguez, A. Conde, A. J. Sáez, and M. J. Olmo. On discrete multivariate distributions symmetric in frequencies. TEST, 12(2):459–480, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595725.

Riani:2014:CFE

[RCT14] Marco Riani, Andrea Cerioli, and Francesca Torti. On consistency factors and efficiency of robust s-estimators. TEST, 23(2):356–387, June 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0357-7.

Reimherr:2019:CMR

[Rei19] Matthew Reimherr. Comments on: Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1):43–45, March 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00635-9. See [KKLU19a, KKLU19b].

Russo:2024:CFM

[RF24] Alfonso Russo and Alessio Farcomeni. A copula formulation for multivariate latent Markov models. TEST, 33(3):731– 751, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00919-9.

Reyes:2017:BSK

[RFFC17] Miguel Reyes, Mario Francisco-Fernández, and Ricardo Cao. Bandwidth selection in kernel density estimation for intervalgrouped data. TEST, 26(3):527–545, September 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0523-9.

Ruwet:2013:BBT

[RGEGMI13] C. Ruwet, L. A. García-Escudero, A. Gordaliza, and A. Mayo-Iscar. On the breakdown behavior of the TCLUST clustering procedure. TEST, 22(3):466–487, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-012-0312-4.

Rosa:2017:OAD

[RH17] Samuel Rosa and Radoslav Harman. Optimal approximate designs for comparison with control in dose-escalation studies. TEST, 26(3):638–660, September 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0529-3.

Rios-Insua:1992:FRT

[RI92] David Rios-Insua. Foundations for a robust theory of decision making: the simple case. TEST, 1(1):69–78, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/

10.1007/BF02562663. See correction [Bül97].

Richards:2020:CTM

[Ric20] Donald Richards. Comments on: Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):903–906, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00739-7. See [EH20b, EH20a].

Ritz:2013:PLR

[Rit13] Christian Ritz. Penalized likelihood ratio tests for repeated measurement models. TEST, 22(3):534–547, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0324-8.

Ritov:2023:CSI

[Rit23] Ya’acov Ritov. Comments on: Statistical inference and large-scale multiple testing for high-dimensional regression models. TEST, 32(4):1180–1183, December 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00898-3. See [CGX23b].

Rivoirard:2004:TPP

[Riv04] Vincent Rivoirard. Thresholding procedure with priors based on Pareto distributions. TEST, 13(1):213–246, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02603007.

Rosco:2011:SDT

[RJP11] J. F. Rosco, M. C. Jones, and Arthur Pewsey. Skew t distributions via the sinh–arcsinh transformation. TEST, 20 (3):630–652, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0222-2.

Ruiz-Medina:2015:CCS

[RM15] M. D. Ruiz-Medina. Comments on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1): 45–46, March 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0411-5;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0411-

5.pdf. See [BCS15a] and rejoinder [BCS15b].

Rueda:2010:SAE

[RMG10] Cristina Rueda, José A. Menéndez, and Federico Gómez. Small area estimators based on restricted mixed models. TEST, 19(3):558–579, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-010-0186-
2.

Rodriguez-Muniz:2003:DRU

[RMLDG03] Luis J. Rodríguez-Muñiz, Miguel López-Díaz, and M. Ángeles Gil. Differentiating random upper semicontinuous functions under the integral sign. TEST, 12(1):241–258, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595821.

Ruiz-Medina:2019:DMR

[RMME19] M. D. Ruiz-Medina, D. Miranda, and R. M. Espejo. Dynamical multiple regression in function spaces, under kernel regressors, with ARH(1) errors. TEST, 28(3):943–968, September 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0614-2.

Rodrigues:1994:BEN

[Rod94] Josemar Rodrigues. Bayesian estimation of a normal mean parameter using the Linex loss function and robustness considerations. TEST, 3(2):237–246, December 1994. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562704.

Romo:1994:CLT

[Rom94] Juan Romo. The central limit theorem for empirical processes on V-C classes: A majorizing measure approach. TEST, 3(2):47–72, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02562693.

Reed:2009:TNF

[RP09] William J. Reed and Arthur Pewsey. Two nested families of skew-symmetric circular distributions. TEST, 18(3):516– 528, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0111-0.

Rodriguez-Poo:2001:NFA

[RPL01] Juan M. Rodríguez-Poo and Oliver Linton. Nonparametric factor analysis of residual time series. TEST, 10(1):161– 182, June 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595830.

Robert:1993:ENM

[RS93] Christian P. Robert and Caroline Soubiran. Estimation of a normal mixture model through Gibbs sampling and prior feedback. TEST, 2(1–2):125–146, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562672.

Rocha:2011:IDG

[RS11] Andréa V. Rocha and Alexandre B. Simas. Influence diagnostics in a general class of beta regression models. TEST, 20 (1):95–119, May 2011. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-010-0189-z.

Rueda:1997:GPM

[RSF97] C. Rueda, B. Salvador, and M. A. Fernández. A good property of the maximum likelihood estimator in a restricted normal model. TEST, 6(1):127–135, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02564430.

Rumi:2006:EMT

[RSM06] Rafael Rumí, Antonio Salmerón, and Serafín Moral. Estimating mixtures of truncated exponentials in hybrid Bayesian networks. TEST, 15(2):397–421, September 2006. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02607059.

Ristic:2019:GBI

[RSMJ19] Miroslav M. Ristić, Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. A GQL-based inference in non-stationary BINMA(1) time series. TEST, 28(3):969–998, September 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0615-1.

Ruggeri:2025:MBS

[RSSSL25] F. Ruggeri, M. Sánchez-Sánchez, and A. Suárez-Llorens. Measuring Bayesian sensitivity in the compound Poisson process. TEST, 34(3):??, September 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00970-0.

Romano:2008:CFD

[RSW08a] Joseph P. Romano, Azeem M. Shaikh, and Michael Wolf. Control of the false discovery rate under dependence using the bootstrap and subsampling. TEST, 17(3):417– 442, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0126-6. See

comments [FvdW08, Guo08, SH08, Tro08, Yek08] and rejoinder [RSW08b].

Romano:2008:RCF

[RSW08b] Joseph P. Romano, Azeem M. Shaikh, and Michael Wolf. Rejoinder on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17 (3):461–471, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0134-6. See

[FvdW08, Guo08, SH08, Tro08, Yek08] and rejoinder [RSW08b].

Rueda:1992:BAP

[Rue92] Raul Rueda. A Bayesian alternative to parametric hypothesis testing. TEST, 1(1):61–67, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562662.

Rumeau:2003:SIF

[Rum03] Tomás Prieto Rumeau. Statistical inference for a finite optimal stopping problem with unknown transition probabilities.

TEST, 12(1):215–239, June 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595820.

Rousseeuw:2015:CRE

[RV15] Peter J. Rousseeuw and Wannes Van den Bossche. Comments on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):473–477, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-015-0454-2;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0454-

2.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Ryabko:2012:TCH

[Rya12] Daniil Ryabko. Testing composite hypotheses about discrete ergodic processes. TEST, 21(2):317–329, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0245-3.

Rychlik:2019:SBD

[Ryc19] Tomasz Rychlik. Sharp bounds on distribution functions and expectations of mixtures of ordered families of distributions. TEST, 28(1):166–195, March 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0604-4;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0604-4.pdf.

Saraceno:2021:RME

[SA21] Giovanni Saraceno and Claudio Agostinelli. Robust multivariate estimation based on statistical depth filters. TEST, 30(4):935–959, December 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00757-z.

Salehi:2012:MRL

[SAE12] E. T. Salehi, M. Asadi, and S. Eryilmaz. On the mean residual lifetime of consecutive k-out-of-n systems. TEST, 21(1): 93–115, March 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0237-3.

Suarez:2021:CPD

[SAF21a] Rebeca Peláez Suárez, Ricardo Cao Abad, and Juan M. Vilar Fernández. Correction to: Probability of default estimation in credit risk using a nonparametric approach. TEST, 30 (2):406, June 2021. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00748-6. See [SAF21b].

Suarez:2021:PDE

[SAF21b] Rebeca Peláez Suárez, Ricardo Cao Abad, and Juan M. Vilar Fernández. Probability of default estimation in credit risk using a nonparametric approach. TEST, 30(2):383–405, June 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00723-1. See correction [SAF21a].

Saha:2007:ORR

[Sah07] Amitava Saha. Optional randomized response in stratified unequal probability sampling — a simulation based numerical study with Kuk’s method. TEST, 16(2):346–354, August 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-006-0013-y.

Salmeron:2022:CHS

[Sal22] Antonio Salmerón. Comments on: Hybrid semiparametric Bayesian networks. TEST, 31(2):331–334, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00818-x. See [ALB22a, ALB22b].

Sanso:1997:SAL

[San97] B. Sansó. Simple approximations for location and ANOVA models with non-conjugate priors. TEST, 6(1):119–126, June 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02564429.

Sanso:2010:CGS

[San10] Bruno Sansó. Comments on: “A general science-based framework for dynamical spatio-temporal models”. TEST, 19 (3):459–461, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0212-4. See

[WH10b] and rejoinder [WH10b].

Singh:1992:AUR

[SB92] Housila P. Singh and R. S. Biradar. Almost unbiased ratio-cum-product estimators for the finite population mean. TEST, 1(1):19–29, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02562658.

Singpurwalla:1998:SCP

[SBC^{+}98] Nozer D. Singpurwalla, G. Box, D. R. Cox, D. K. Dey, A. Fries, J. K. Ghosh, M. A. Gómez-Villegas, T. Z. Irony, W. Kliemann, S. Kotz, D. V. Lindley, M. F. McGrath, D. Peña, and N. D. Singpurwalla. The stochastic control of process capability indices. TEST, 7(1):1–74, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02565102.

Stadler:2010:PMR

[SBvdG10a] Nicolas Städler, Peter Bühlmann, and Sara van de Geer. l_{1}-penalization for mixture regression models. TEST, 19 (2):209–256, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0197-z. See

comments [Ant10, Lug10, FL10, SZ10, dB10, dB10] and rejoinder [SBvdG10b].

Stadler:2010:RPM

[SBvdG10b] Nicolas Städler, Peter Bühlmann, and Sara van de Geer. Rejoinder: “l_{1}-penalization for mixture regression models”. TEST, 19(2):280–285, August 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-010-0203-5. See [SBvdG10a, Ant10, Lug10, FL10, SZ10, dB10, dB10].

Siviero:2024:SLV

[SCC24] Emilia Siviero, Emilie Chautru, and Stephan Clémençon. A statistical learning view of simple kriging. TEST, 33(1):271– 296, March 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00891-w.

Scealy:2021:CRA

[Sce21] Janice L. Scealy. Comments on: Recent advances in directional statistics. TEST, 30(1):68–70, March 2021. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00763-1. See [PGP21a, PGP21b].

Schmidli:1996:BAR

[Sch96] H. Schmidli. Bayesian analysis of reduced rank regression. TEST, 5(1):159–186, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562687.

Scholl:1998:SOP

[Sch98] Holger R. Scholl. Shannon optimal priors on independent identically distributed statistical experiments converge weakly to Jeffreys’ prior. TEST, 7(1):75–94, June 1998. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02565103.

Schafer:2008:CAB

[Sch08] Juliane Schäfer. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1): 28–30, May 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0102-1.

Schliep:2018:CPM

[Sch18] Erin M. Schliep. Comments on: “Process modeling for slope and aspect with application to elevation data maps”. TEST, 27(4):778–782, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-018-0620-4. See

[WBG18a] and rejoinder [WBG18b].

Schnell:2019:CMR

[Sch19] Patrick Schnell. Comments on: Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1): 46–51, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00632-y. See [KKLU19a,

KKLU19b].

Singh:2007:FEM

[SCJS07] H. P. Singh, P. Chandra, Anwar H. Joarder, and Sarjinder Singh. Family of estimators of mean, ratio and product of a finite population using random nonresponse. TEST, 16 (3):565–597, December 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0020-z.

Scutari:2022:CHS

[Scu22] Marco Scutari. Comments on: Hybrid semiparametric Bayesian networks. TEST, 31(2):328–330, June 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00815-0. See [ALB22a, ALB22b].

Sabillon:2023:DDR

[SCZ23] Gustavo Alexis Sabillón, Luiz Gabriel Fernandes Cotrim, and Daiane Aparecida Zuanetti. A data-driven reversible jump for estimating a finite mixture of regression models. TEST, 32(1):350–369, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00835-w.

Senarathne:2020:BSD

[SDM20] S. G. J. Senarathne, C. C. Drovandi, and J. M. McGree. Bayesian sequential design for copula models. TEST, 29(2): 454–478, June 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00661-7.

Sang:2020:DBW

[SDZ20] Yongli Sang, Xin Dang, and Yichuan Zhao. Depth-based weighted jackknife empirical likelihood for non-smooth Ustructure equations. TEST, 29(2):573–598, June 2020.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00667-1.

Segers:2011:CIM

[Seg11] Johan Segers. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):281–283, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0255-1. See [GNZ11a] and rejoinder

[GNZ11b].

Senturk:2010:CDR

[Sen10] Damla Sentürk. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):54–55, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0174-6.

Sain:2018:CSR

[SF18] Stephan R. Sain and Reinhard Furrer. Comments on: “Some recent work on multivariate Gaussian Markov random fields”. TEST, 27(3):545–548, September 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0609-z;   http://link.springer.com/
content/pdf/10.1007/s11749-018-0609-z.pdf. See [Mac18b]

and rejoinder [Mac18a].

Singh:2004:SMC

[SG04] Parminder Singh and A. N. Gill. Some multiple comparisons using sample quasi ranges on censored data. TEST, 13 (2):313–334, December 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595776.

Sguera:2014:SDB

[SGL14] Carlo Sguera, Pedro Galeano, and Rosa Lillo. Spatial depthbased classification for functional data. TEST, 23(4):725– 750, December 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0379-1.

Said:2007:STI

[SGR07] Mériem Saïd, Nadia Ghazzali, and Louis-Paul Rivest. Score tests for independence in parametric competing risks models. TEST, 16(3):547–564, December 2007. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0019-5.

Sarkar:2008:CCF

[SH08] Sanat K. Sarkar and Ruth Heller. Comments on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17(3):450–455, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-008-0129-3. See [RSW08a] and rejoinder

[RSW08b].

Shi:2019:CDS

[SH19] Jian Qing Shi and Shane Halloran. Comments on: Data science, big data and statistics. TEST, 28(2):353–356, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00640-y. See [GP19a, GP19b].

Shalabh:2001:LSE

[Sha01] Shalabh. Least squares estimators in measurement error models under the balanced loss function. TEST, 10(2):301– 308, December 2001. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595699.

Shen:2013:CRB

[She13] Pao-Sheng Shen. A class of rank-based tests for doublytruncated data. TEST, 22(1):83–102, March 2013. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-012-0295-1.

Shin:2007:CPD

[Shi07] Yongcheol Shin. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):52–55, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0054-x. See [Hsi07a] and rejoinder

[Hsi07b].

Sickles:2007:CPD

[Sic07] Robin C. Sickles. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1):31–32, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0049-7. See [Hsi07a] and rejoinder

[Hsi07b].

Simone:2022:FMD

[Sim22] Rosaria Simone. On finite mixtures of Discretized Beta model for ordered responses. TEST, 31(3):828–855, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00800-7.

Schifano:2021:FEB

[SJD21] Elizabeth D. Schifano, Himchan Jeong, and Dipak K. Dey. Fully and empirical Bayes approaches to estimating copulabased models for bivariate mixed outcomes using Hamiltonian Monte Carlo. TEST, 30(1):133–152, March 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00705-3.

Sweeting:2003:SNF

[SK03] Trevor J. Sweeting and Samer A. Kharroubi. Some new formulae for posterior expectations and Bartlett corrections. TEST, 12(2):497–521, December 2003. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595727.

Sugasawa:2018:SAE

[SKR18] Shonosuke Sugasawa, Tatsuya Kubokawa, and J. N. K. Rao. Small area estimation via unmatched sampling and linking models. TEST, 27(2):407–427, June 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0551-5.

Strzalkowska-Kominiak:2013:ECC

[SKS13] E. Strzalkowska-Kominiak and W. Stute. Empirical copulas for consecutive survival data. TEST, 22(4):688–714, November 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0339-1.

Stramer:2002:ITA

[SL02] Osnat Stramer and Yu-Jau Lin. On inference for threshold autoregressive models. TEST, 11(1):55–71, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595729.

Santos:2017:MLE

[SL17] Cristiano C. Santos and Rosangela H. Loschi. Maximum likelihood estimation and parameter interpretation in elliptical mixed logistic regression. TEST, 26(1):209–230, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0507-1.

Sperlich:1999:IBM

[SLH99] Stefan Sperlich, Oliver B. Linton, and Wolfgang Härdle. Integration and backfitting methods in additive models-finite sample properties and comparison. TEST, 8(2):419–458, December 1999. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595879.

Surjanovic:2024:GHL

[SLL24] Nikola Surjanovic, Richard A. Lockhart, and Thomas M. Loughin. A generalized Hosmer–Lemeshow goodness-of-fit test for a family of generalized linear models. TEST, 33(2): 589–608, June 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00912-8.

Sguera:2021:NDS

[SLP21] Carlo Sguera and Sara López-Pintado. A notion of depth for sparse functional data. TEST, 30(3):630–649, September 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00734-y.

Singh:2023:SPT

[SM23] Rahul Singh and Neeraj Misra. Some parametric tests based on sample spacings. TEST, 32(1):211–231, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00831-0.

Sebastiao:2013:EUI

[SMFP13] J. R. Sebastião, A. P. Martins, H. Ferreira, and L. Pereira. Estimating the upcrossings index. TEST, 22(4):549–579, November 2013. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0315-9.

Shamma:2023:TMN

[SMS23] Nisreen Shamma, Mehrnaz Mohammadpour, and Masoumeh Shirozhan. A threshold modeling for nonlinear time series of counts: application to COVID-19 data. TEST, 32(4):1195– 1229, December 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-023-00869-8.

Sugasawa:2022:BSM

[SMT22] Shonosuke Sugasawa, Kosuke Morikawa, and Keisuke Takahata. Bayesian semiparametric modeling of response mechanism for nonignorable missing data. TEST, 31(1):101– 117, March 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00774-y.

Sang:2016:CPE

[SNC16] Peijun Sang, Yunlong Nie, and Jiguo Cao. Comments on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):33– 34, March 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0473-z;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0473-

z.pdf. See [YWZ16a] and rejoinder [YWZ16b].

Sanso:1992:NIC

[SP92] B. Sanso and L. R. Pericchi. Near ignorance classes of logconcave priors for the location model. TEST, 1(1):39–46, December 1992. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02562660.

Stanghellini:2015:ICE

[SP15] Elena Stanghellini and Eduwin Pakpahan. Identification of causal effects in linear models: beyond instrumental variables. TEST, 24(3):489–509, September 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0421-3.

Sattler:2024:THA

[SP24] Paavo Sattler and Markus Pauly. Testing hypotheses about correlation matrices in general MANOVA designs. TEST, 33(2):496–516, June 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00906-6.

Sperlich:2009:CRE

[Spe09] Stefan Sperlich. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18(3):448–451, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0160-z. See [CV09b] and [CV09a].

Sperlich:2013:CUR

[Spe13a] Stefan Sperlich. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22 (3):419–427, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0330-x.

Sperlich:2013:CMF

[Spe13b] Stefan Sperlich. Comments on: “Model-free model-fitting and predictive distributions”. TEST, 22(2):227–233, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0318-6. See [Pol13a] and rejoinder

[Pol13b].

Spezia:2019:MCM

[Spe19] Luigi Spezia. Modelling covariance matrices by the trigonometric separation strategy with application to hidden Markov models. TEST, 28(2):399–422, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0580-8.

Sperlich:2022:CHS

[Spe22] Stefan Sperlich. Comments on: Hybrid semiparametric Bayesian networks. TEST, 31(2):335–339, June 2022. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00817-y. See [ALB22a, ALB22b].

Szekely:2020:CTM

[SR20] Gábor J. Székely and Maria L. Rizzo. Comments on: Tests for multivariate normality — a critical review with emphasis on weighted L^{2}-statistics. TEST, 29(4):907–910, December 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-020-00741-z. See [EH20b, EH20a].

Stehlik:2008:OAB

[SRDMLF08] Milan Stehlík, Juan M. Rodríguez-Díaz, Werner G. Müller, and Jesús López-Fidalgo. Optimal allocation of bioassays in the case of parametrized covariance functions: an application to Lung’s retention of radioactive particles. TEST, 17(1): 56–68, May 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0022-x.

Suarez-Rancel:2000:LDD

[SRGS00] M. Mercedes Suárez-Rancel and Miguel A. González-Sierra. Local and deletion diagnostic. TEST, 9(2):345–352, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595739.

Stasinopoulos:2019:CMR

[SRHD19] M. D. Stasinopoulos, R. A. Rigby, G. Z. Heller, and F. De Bastiani. Comments on: Modular regression — a Lego system for building structured additive distributional regression models with tensor product interactions. TEST, 28(1):52– 54, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00634-w. See [KKLU19a,

KKLU19b].

Stephens:1992:SRT

[SS92] D. A. Stephens and A. F. M. Smith. Sampling-resampling techniques for the computation of posterior densities in normal means problems. TEST, 1(1):1–18, December 1992. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562657.

Srivastava:1997:AEP

[SS97] A. K. Srivastava and Shalabh. Asymptotic efficiency properties of least squares in an ultrastructural model. TEST, 6(2):419–431, December 1997. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02564707.

Sanchez-Sellero:2009:CRE

[SS09] César Sánchez-Sellero. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18(3):458– 460, November 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-009-0163-9. See [CV09b] and

[CV09a].

Shen:2012:ALT

[sS12] Pao sheng Shen. Analysis of left-truncated right-censored or doubly censored data with linear transformation models. TEST, 21(3):584–603, September 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0263-1.

Stocker:2024:CSB

[SSG24] Almond Stöcker, Lisa Steyer, and Sonja Greven. Comments on: Shape-based functional data analysis. TEST, 33(1):48– 58, March 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/

article/10.1007/s11749-023-00901-x. See [WHS24b,

WHS24a].

Sanchez-Sellero:2023:CNE

[SSGM23] César Sánchez-Sellero and Wenceslao González-Manteiga. Comments on: Nonparametric estimation in mixture cure models with covariates. TEST, 32(2):510–512, June 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00862-1. See [LCPJ23a, LCPJ23b].

Stanghellini:2012:CSR

[Sta12] Elena Stanghellini. Comments on: “Sequence of regressions and their independences”. TEST, 21(2):265–267, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-012-0287-1. See [WS12b] and rejoinder

[WS12a].

Staudte:2014:IQM

[Sta14] Robert G. Staudte. Inference for quantile measures of skewness. TEST, 23(4):751–768, December 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0391-5.

Salvati:2012:SAE

[STPC12] N. Salvati, N. Tzavidis, M. Pratesi, and R. Chambers. Small area estimation via m-quantile geographically weighted regression. TEST, 21(1):1–28, March 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-010-0231-1.

Strimmer:2008:CAB

[Str08] Korbinian Strimmer. Comments on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17 (1):25–27, May 2008. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-008-0101-2.

Stute:2001:RAT

[Stu01] Winfried Stute. Residual analysis for ARCH(p)-time series. TEST, 10(2):393–403, December 2001. CODEN ????

ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595704.

Steorts:2014:CST

[SU14] Rebecca C. Steorts and M. Dolores Ugarte. Comments on: “Single and two-stage cross-sectional and time series benchmarking procedures for small area estimation”. TEST, 23 (4):680–685, December 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0386-2;
http://link.springer.com/content/pdf/10.1007/s11749-

014-0386-2.pdf. See [PST14b] and rejoinder [PST14a].

Sued:2020:RDP

[SVY20] Mariela Sued, Marina Valdora, and Víctor Yohai. Robust doubly protected estimators for quantiles with missing data. TEST, 29(3):819–843, September 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-019-00689-9.

Schweer:2016:TPA

[SW16] Sebastian Schweer and Christian H. Weiß. Testing for Poisson arrivals in INAR(1) processes. TEST, 25(3):503–524, September 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0466-y.

Scotto:2018:MIM

[SWMG18] Manuel G. Scotto, Christian H. Weiß, Tobias A. Möller, and Sónia Gouveia. The max-INAR(1) model for count processes. TEST, 27(4):850–870, December 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0573-z.

Song:2025:ISA

[SWZ25] Shuyang Song, Jiaqi Wu, and Weiping Zhang. Integrative subgroup analysis for high-dimensional mixed-type multiresponse data. TEST, 34(1):151–197, March 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00953-7.

Sun:2025:CMD

[SYL25] Mingyu Sun, Kai Yang, and Ang Li. Conditional minimum density power divergence estimator for self-exciting integervalued threshold autoregressive models. TEST, 34(1):198– 234, March 2025. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00956-4.

Sun:2010:CPM

[SZ10] Tingni Sun and Cun-Hui Zhang. Comments on: “l_{1}- penalization for mixture regression models”. TEST, 19 (2):270–275, August 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0201-7.

Shu:2024:CDI

[SZ24] Hai Shu and Hongtu Zhu. Comments on: Data integration via analysis of subspaces (DIVAS). TEST, 33(3):686–688, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00943-9. See [PJH^{+}24a].

Su:2024:LIA

[SZL24] Zhonghao Su, Fukang Zhu, and Shuangzhe Liu. Local influence analysis in the softplus INGARCH model. TEST, 33 (3):951–985, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00930-0.

Tavares:2006:IRT

[TA06] Heliton R. Tavares and Dalton F. Andrade. Item response theory for longitudinal data: Item and population ability parameters estimation. TEST, 15(1):97–123, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595420.

Tyekucheva:2008:ABA

[TC08a] Svitlana Tyekucheva and Francesca Chiaromonte. Augmenting the bootstrap to analyze high dimensional genomic data. TEST, 17(1):1–18, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-008-0098-6. See

comments [Li08a, Li08b, Str08, Sch08, BKK08, KC08, GMC08, MWN08] and rejoinder [TC08b].

Tyekucheva:2008:RAB

[TC08b] Svitlana Tyekucheva and Francesca Chiaromonte. Rejoinder on: “Augmenting the bootstrap to analyze high dimensional genomic data”. TEST, 17(1):47–55, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-008-0107-9. See [TC08a, Li08a, Li08b, Str08, Sch08, BKK08, KC08, GMC08, MWN08].

Temido:2000:MRE

[Tem00] M. Graça Temido. Mixture results for extremal behaviour of strongly dependent nonstationary Gaussian sequences. TEST, 9(2):439–453, December 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595744.

Tenreiro:2022:AKD

[Ten22] Carlos Tenreiro. On automatic kernel density estimate-based tests for goodness-of-fit. TEST, 31(3):717–748, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-021-00799-3.

Teran:2008:EAH

[Ter08] Pedro Terán. On the equivalence of Aumann and Herer expectations of random sets. TEST, 17(3):505–514, November 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-007-0043-0.

Tian:2023:CSI

[TF23] Ye Tian and Yang Feng. Comments on: Statistical inference and large-scale multiple testing for high-dimensional regression models. TEST, 32(4):1172–1176, December 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00880-z. See [CGX23b].

Thas:2009:CGF

[Tha09] Olivier Thas. Comments on: “Goodness-of-fit-tests in mixed models”. TEST, 18(2):260–264, August 2009. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-009-0153-y. See [CH09a] and rejoinder [CH09a].

Thanh:2023:NCS

[Thà23] Lê Văn Thành. On a new concept of stochastic domination and the laws of large numbers. TEST, 32(1):74–106, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-022-00827-w.

Tjostheim:2012:RSR

[Tjø12a] Dag Tjøstheim. Rejoinder on: “Some recent theory for autoregressive count time series”. TEST, 21(3):469– 476, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0305-3. See

[Tjø12b, Ane12, Dol12, Dou12, Fok12, Gal12, Gao12, Hei12, Ked12].

Tjostheim:2012:SRT

[Tjø12b] Dag Tjøstheim. Some recent theory for autoregressive count time series. TEST, 21(3):413–438, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-012-0296-0. See comments [Ane12, Dol12, Dou12, Fok12, Gal12, Gao12, Hei12, Ked12] and rejoinder [Tjø12a].

Torkashvand:2016:CBE

[TJT16] Elaheh Torkashvand, Mohammad Jafari Jozani, and Mahmoud Torabi. Constrained Bayes estimation in small area models with functional measurement error. TEST, 25(4): 710–730, December 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0492-4.

Thorarinsdottir:2014:CST

[TL14] Thordis L. Thorarinsdottir and Anders Løland. Comments on: “Space–time wind speed forecasting for improved power system dispatch”. TEST, 23(1):32–33, March 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0354-x;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0354-

x.pdf. See [ZGGX14b] and rejoinder [ZGGX14a].

Tapia:2019:IDM

[TLdPDG19] Alejandra Tapia, Victor Leiva, Maria del Pilar Diaz, and Viviana Giampaoli. Influence diagnostics in mixed effects logistic regression models. TEST, 28(3):920–942, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0613-3.

Tracy:1995:PRR

[TO95] D. S. Tracy and S. S. Osahan. A partial randomized response strategy. TEST, 4(2):315–321, December 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562629.

Tong:2011:CNI

[Ton11] Xingwei Tong. Comments on: “Nonparametric inference based on panel count data”. TEST, 20(1):58–61, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-010-0228-9. See [ZBS11a] and rejoinder

[ZBS11b].

Tsuyuguchi:2020:ACB

[TPB20] Aline B. Tsuyuguchi, Gilberto A. Paula, and Michelli Barros. Analysis of correlated Birnbaum–Saunders data based on estimating equations. TEST, 29(3):661–681, September 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-019-00675-1.

Trapani:2014:CES

[Tra14] Lorenzo Trapani. Comments on: “Extensions of some classical methods in change point analysis”. TEST, 23(2):283– 286, June 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0367-5;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0367-
5.pdf.

Troendle:2008:CCF

[Tro08] James F. Troendle. Comments on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17(3):456–457, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-008-0130-x. See [RSW08a] and rejoinder [RSW08b].

Toutenburg:2005:ERC

[TS05] H. Toutenburg and Shalabh. Estimation of regression coefficients subject to exact linear restrictions when some observations are missing and quadratic error balanced loss function is used. TEST, 14(2):385–396, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595409.

Tingley:2015:CCS

[TS15] Martin P. Tingley and Benjamin A. Shaby. Comments on: “Comparing and selecting spatial predictors using local criteria”. TEST, 24(1):47–53, March 2015. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-014-0413-3;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0413-

3.pdf. See [BCS15a] and rejoinder [BCS15b].

Tsao:2006:NLP

[Tsa06] C. Andy Tsao. A note on Lindley’s paradox. TEST, 15(1): 125–139, June 2006. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02595421.

Tsay:2019:CDS

[Tsa19] Ruey S. Tsay. Comments on: Data science, big data and statistics. TEST, 28(2):357–359, June 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00641-x. See [GP19a, GP19b].

Tseng:2002:OCS

[Tse02] Yu-Ling Tseng. Optimal confidence sets for testing average bioequivalence. TEST, 11(1):127–141, June 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595733.

Tsukahara:2011:CIM

[Tsu11] Hideatsu Tsukahara. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):287–289, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0257-z. See [GNZ11a] and rejoinder

[GNZ11b].

Tedesco:2023:CQR

[TV23] Lorenzo Tedesco and Ingrid Van Keilegom. Comparison of quantile regression curves with censored data. TEST, 32 (3):829–864, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00854-1.

Trenkler:1996:BEM

[TW96] G. Trenkler and L. Wei. The Bayes estimator in a misspecified linear regression model. TEST, 5(1):113–123, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02562684.

Tan:2014:LLM

[TW14] Zhongquan Tan and Changchun Wu. Limit laws for the maxima of stationary chi-processes under random index. TEST, 23(4):769–786, December 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0366-6.

Tang:2012:ISB

[TWHZ12] Yanlin Tang, Huixia Judy Wang, Xuming He, and Zhongyi Zhu. An informative subset-based estimator for censored quantile regression. TEST, 21(4):635–655, December 2012.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0266-y.

Tang:2015:WLL

[TZ15] Linjun Tang and Zhangong Zhou. Weighted local linear CQR for varying-coefficient models with missing covariates. TEST, 24(3):583–604, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0425-z.

Ugarte:2009:CMD

[Uga09] M. D. Ugarte. Comments on: “Missing data methods in longitudinal studies: a review”. TEST, 18(1):44–46, May 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0139-9. See [IM09a] and rejoinder

[IM09b].

Ugarte:2009:BES

[UMG09] M. D. Ugarte, A. F. Militino, and T. Goicoa. Benchmarked estimates in small areas using linear mixed models with restrictions. TEST, 18(2):342–364, August 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0094-x.

Valdora:2025:RPE

[VA25] Marina Valdora and Claudio Agostinelli. Robust penalized estimators for high-dimensional generalized linear models. TEST, 34(3):714–741, September 2025. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-025-00978-6.

Valdez:2011:CIM

[Val11] Emiliano A. Valdez. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):257–262, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-011-0251-5. See [GNZ11a] and rejoinder

[GNZ11b].

VanDerLinde:1995:SBP

[Van95] A. Van Der Linde. Splines from a Bayesian point of view. TEST, 4(1):63–81, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02563103.

Vantini:2012:DPA

[Van12] Simone Vantini. On the definition of phase and amplitude variability in functional data analysis. TEST, 21(4):676–696, December 2012. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-011-0268-9.

VanKeilegom:2013:CUR

[Van13a] Ingrid Van Keilegom. Comments on: “An updated review of Goodness-of-Fit tests for regression models”. TEST, 22 (3):428–431, September 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-013-0331-9;
http://link.springer.com/content/pdf/10.1007/s11749-
013-0331-9.pdf.

VanKeilegom:2013:CMF

[Van13b] Ingrid Van Keilegom. Comments on: “Model-free modelfitting and predictive distributions”. TEST, 22(2):234–236, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-013-0322-x. See [Pol13a] and rejoinder

[Pol13b].

VanAelst:2015:CRE

[Van15] Stefan Van Aelst. Comments on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):478–481, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0456-0;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0456-

0.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Veraverbeke:2000:ECD

[VCS00] Noel Veraverbeke and Carmen Cadarso-Suárez. Estimation of the conditional distribution in a conditional Koziol–Green model. TEST, 9(1):97–122, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595853.

VanderMerwe:1996:BAS

[VD96] A. J. Van der Merwe and J. L. Du Plessis. A Bayesian approach to selection and ranking procedures: the unequal variance case. TEST, 5(2):357–377, December 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562623.

vanderLinde:2004:ABR

[vdL04] Angelika van der Linde. On the association between a random parameter and an observable. TEST, 13(1):85–111, June 2004. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02603002.

Velez:2001:NAS

[Vél01] Ricardo Vélez. A new approach to series of independent random variables. TEST, 10(2):405–418, December 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595705.

Velasco:2009:CRE

[Vel09] Carlos Velasco. Comments on: “A review on empirical likelihood methods for regression”. TEST, 18(3):455–457, November 2009. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0162-x. See [CV09b] and [CV09a].

Velasco:2011:CSW

[Vel11] Carlos Velasco. Comments on: “Subsampling weakly dependent time series and application to extremes”. TEST, 20 (3):480–482, November 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0270-2. See

[DPR11b] and rejoinder [DPR11a].

Velasco:2013:CMF

[Vel13] Carlos Velasco. Comments on: “Model-free model-fitting and predictive distributions”. TEST, 22(2):237–239, June 2013. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-013-0320-z;   http://link.
springer.com/content/pdf/10.1007/s11749-013-0320-

z.pdf. See [Pol13a] and rejoinder [Pol13b].

vanEeden:1994:CGB

[vEZ94a] Constance van Eeden and James V. Zidek. Corrections to “Group-Bayes estimation of the exponential mean: a preposterior analysis”. TEST, 3(2):247, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562705;   http://link.springer.com/content/
pdf/10.1007/BF02562705.pdf. See [vEZ94b].

vanEeden:1994:GBE

[vEZ94b] Constance van Eeden and James V. Zidek. Group-Bayes estimation of the exponential mean: a preposterior analysis. TEST, 3(1):125–143, June 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562677. See correc-

tions [vEZ94a].

Vilar-Fernandez:2000:RLP

[VFVF00] Juan M. Vilar-Fernández and José A. Vilar-Fernández. Recursive local polynomial regression under dependence conditions. TEST, 9(1):209–232, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595859.

Vilar-Fernandez:2007:BTN

[VFVFGM07] J. M. Vilar-Fernández, J. A. Vilar-Fernández, and W. González-Manteiga. Bootstrap tests for nonparametric comparison of regression curves with dependent errors. TEST, 16(1):123– 144, May 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-006-0005-y.

VanderMerwe:2007:BTI

[VH07] Abrie J. Van der Merwe and Johan Hugo. Bayesian tolerance intervals for the balanced two-factor nested random effects model. TEST, 16(3):598–612, December 2007.

CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0021-y.

Vilar:1995:KER

[Vil95] J. A. Vilar. Kernel estimation of the regression function with random sampling times. TEST, 4(1):137–178, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02563107.

Villa:2017:BET

[Vil17] Cristiano Villa. Bayesian estimation of the threshold of a generalised Pareto distribution for heavy-tailed observations. TEST, 26(1):95–118, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0501-7;
http://link.springer.com/content/pdf/10.1007/s11749-
016-0501-7.pdf.

Volovskiy:2020:MOL

[VK20] Grigoriy Volovskiy and Udo Kamps. Maximum observed likelihood prediction of future record values. TEST, 29(4):1072– 1097, December 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-020-00701-7.

Villegas:1999:CAC

[VM99] Cesáreo Villegas and Camilo J. Martínez. On the concepts of admissibility and coherence. TEST, 8(2):319–338, December 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595873.

VanKeilegom:2008:GFT

[VMS08] Ingrid Van Keilegom, Wenceslao González Manteiga, and César Sánchez Sellero. Goodness-of-fit tests in parametric regression based on the estimation of the error distribution. TEST, 17(2):401–415, August 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-007-0044-z.

Vovk:1993:FPC

[Vov93] V. G. Vovk. Forecasting point and continuous processes: Prequential analysis. TEST, 2(1–2):189–217, December 1993. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562675.

Vanegas:2015:SAJ

[VP15] Luis Hernando Vanegas and Gilberto A. Paula. A semiparametric approach for joint modeling of median and skewness. TEST, 24(1):110–135, March 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0401-7.

Vanegas:2017:LSR

[VP17] Luis Hernando Vanegas and Gilberto A. Paula. Logsymmetric regression models under the presence of noninformative left- or right-censored observations. TEST, 26(2):405–428, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-016-0517-z.

Velez:2022:IRL

[VPP22] D. Vélez, M. E. Pérez, and L. R. Pericchi. Increasing the replicability for linear models via adaptive significance levels. TEST, 31(3):771–789, September 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-022-00803-4.

Velez:2015:RAP

[VPR15] Ricardo Vélez and Tomás Prieto-Rumeau. Random assignment processes: strong law of large numbers and De Finetti theorem. TEST, 24(1):136–165, March 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0396-0.

Vidal-Sanz:2009:ASD

[VS09] Jose M. Vidal-Sanz. Automatic spectral density estimation for random fields on a lattice via bootstrap. TEST, 18(1):

96–114, May 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0059-5.

Visser:2014:CLM

[VS14] Ingmar Visser and Maarten Speekenbrink. Comments on: “Latent Markov models: a review of a general framework for the analysis of longitudinal data with covariates”. TEST, 23 (3):478–483, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-014-0389-z. See

[BFP14a] and rejoinder [BFP14b].

Villegas:2002:PM

[VSM02] Cesareo Villegas, Tim Swartz, and Carmillo Martínez. On the probability of a model. TEST, 11(2):413–438, December 2002. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/BF02595715.

VanAelst:2019:CDS

[VZ19] Stefan Van Aelst and Ruben H. Zamar. Comments on: Data science, big data and statistics. TEST, 28(2):360–362, June 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00648-4. See [GP19a, GP19b].

Wager:2016:CRF

[Wag16] Stefan Wager. Comments on: “A random forest guided tour”. TEST, 25(2):261–263, June 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/s11749-016-0482-
6;   http://link.springer.com/content/pdf/10.1007/

s11749-016-0482-6.pdf. See [BS16a] and rejoinder [BS16b].

Wang:2010:CDR

[Wan10] Naisyin Wang. Comments on: “Dynamic relations for sparsely sampled Gaussian processes”. TEST, 19(1):56–59, May 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-009-0175-5.

Wang:2019:MMN

[Wan19] Wan-Lun Wang. Mixture of multivariate t nonlinear mixed models for multiple longitudinal data with heterogeneity and missing values. TEST, 28(1):196–222, March 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0612-4.

Wang:2022:MPE

[Wan22] Zhu Wang. MM for penalized estimation. TEST, 31(1):54– 75, March 2022. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-021-00770-2.

Wang:2018:PMS

[WBG18a] Fangpo Wang, Anirban Bhattacharya, and Alan E. Gelfand. Process modeling for slope and aspect with application to elevation data maps. TEST, 27(4):749–772, December 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-018-0619-x. See comments [Ban18, LM18, Sch18] and rejoinder [WBG18b].

Wang:2018:RPM

[WBG18b] Fangpo Wang, Anirban Bhattacharya, and Alan E. Gelfand. Rejoinder on: “Process modeling for slope and aspect with application to elevation data maps”. TEST, 27(4):783–786, December 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0623-1;   http://link.
springer.com/content/pdf/10.1007/s11749-018-0623-

1.pdf. See [WBG18b, Ban18, LM18, Sch18].

Wasserman:1995:IT

[WC95] L. Wasserman and B. Clarke. Information tradeoff. TEST, 4 (1):19–38, June 1995. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02563101.

Wong:1998:NRI

[WC98] M. Y. Wong and D. R. Cox. A note on the robust interpretation of regression coefficients. TEST, 7(2):287–294, December 1998. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/BF02565113.

Wang:2011:CIM

[WE11] Weijing Wang and Takeshi Emura. Comments on: “Inference in multivariate Archimedean copula models”. TEST, 20(2):276–280, August 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0254-2. See

[GNZ11a] and rejoinder [GNZ11b].

Welsch:2015:CRE

[Wel15] Roy E. Welsch. Comments on: “Robust estimation of multivariate location and scatter in the presence of cellwise and casewise contamination”. TEST, 24(3):482–483, September 2015. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0452-4;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0452-

4.pdf. See [ALYZ15b] and rejoinder [ALYZ15a].

Wang:2007:CPD

[WG07] Chen-Pin Wang and Malay Ghosh. Change-point diagnostics in competing risks models: Two posterior predictive p-value approaches. TEST, 16(1):145–171, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0006-x.

White:2021:MFD

[WG21] Philip A. White and Alan E. Gelfand. Multivariate functional data modeling with time-varying clustering. TEST, 30(3):586–602, September 2021. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-020-00733-z.

Walker:2007:BPI

[WGP07] Stephen G. Walker and Eduardo Gutiérrez-Peña. Bayesian parametric inference in a nonparametric framework. TEST, 16(1):188–197, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0008-8.

Wikle:2010:GSB

[WH10a] Christopher K. Wikle and Mevin B. Hooten. A general science-based framework for dynamical spatio-temporal models. TEST, 19(3):417–451, November 2010. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-010-0209-z. See comments [Mat10, Cru10, San10, Hig10] and rejoinder [WH10b].

Wikle:2010:RGS

[WH10b] Christopher K. Wikle and Mevin B. Hooten. Rejoinder on: “A general science-based framework for dynamical spatiotemporal models”. TEST, 19(3):466–468, November 2010. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-010-0214-2. See [WH10b, Mat10, Cru10, San10, Hig10].

Wei:2014:CTC

[WH14] Wei Wei and Leonhard Held. Calibration tests for count data. TEST, 23(4):787–805, December 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0380-8.

Wu:2024:RSB

[WHS24a] Yuexuan Wu, Chao Huang, and Anuj Srivastava. Rejoinder on: Shape-based functional data analysis. TEST, 33(1): 73–80, March 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00925-x. See [WHS24b,

WHS24a].

Wu:2024:SBF

[WHS24b] Yuexuan Wu, Chao Huang, and Anuj Srivastava. Shapebased functional data analysis. TEST, 33(1):1–47, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11749-023-00876-9. See comments [SSG24, Dry24, Del24, BM24, Hor24] and rejoinder [WHS24a].

Winkler:2008:CAP

[WJ08] Robert L. Winkler and Victor Richmond R. Jose. Comments on: “Assessing probabilistic forecasts of multivariate quantities, with an application to ensemble predictions of surface winds”. TEST, 17(2):251–255, August 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-008-0121-y.

Wang:2017:RED

[WL17] Kangning Wang and Lu Lin. Robust and efficient direction identification for groupwise additive multiple-index models and its applications. TEST, 26(1):22–45, March 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0496-0.

Wang:2020:ALM

[WL20] Wan-Lun Wang and Tsung-I Lin. Automated learning of mixtures of factor analysis models with missing information. TEST, 29(4):1098–1124, December 2020. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-020-00702-6.

Wang:2022:RCM

[WL22] Wan-Lun Wang and Tsung-I Lin. Robust clustering of multiply censored data via mixtures of t factor analyzers. TEST, 31(1):22–53, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00766-y.

Wang:2015:CPP

[WLL15] Zhanfeng Wang, Wenxin Liu, and Yuanyuan Lin. A changepoint problem in relative error-based regression. TEST, 24 (4):835–856, December 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0438-2.

Wang:2025:FCG

[WLL25] Wan-Lun Wang, Victor Hugo Lachos, and Tsung-I Lin. Flexible clustering via Gaussian parsimonious mixture models with censored and missing values. TEST, 34(2):383–430, June 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL https://link.springer.com/article/
10.1007/s11749-025-00967-9.

Wu:2025:CQE

[WLVF25] Chengxin Wu, Nengxiang Ling, Philippe Vieu, and Guoliang Fan. Composite quantile estimation in partially functional linear regression model with randomly censored responses. TEST, 34(1):28–47, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00946-6.

Wansbeek:2007:CPD

[WM07] Tom Wansbeek and Erik Meijer. Comments on: “Panel data analysis — advantages and challenges”. TEST, 16(1): 33–36, May 2007. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0050-1;   http://link.
springer.com/content/pdf/10.1007/s11749-007-0050-

1.pdf. See [Hsi07a] and rejoinder [Hsi07b].

Winkler:1996:SRE

[WMC^{+}96] R. L. Winkler, Javier Muñoz, José L. Cervera, José M. Bernardo, Gail Blattenberger, Joseph B. Kadane, Dennis V. Lindley, Allan H. Murphy, Robert M. Oliver, and David Ríos-Insua. Scoring rules and the evaluation of probabilities. TEST, 5(1):1–60, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562681.

Wang:2025:CSO

[WMS25] Kangning Wang, Xiaoqing Meng, and Xiaofei Sun. Convolution smoothing and online updating estimation for support vector machine. TEST, 34(1):288–323, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00959-1.

Wang:2020:LEE

[WMY20] Qianqian Wang, Yanyuan Ma, and Guangren Yang. Locally efficient estimation in generalized partially linear model with measurement error in nonlinear function. TEST, 29(2):553– 572, June 2020. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-019-00668-0.

Wood:2020:ICG

[Woo20a] Simon N. Wood. Inference and computation with generalized additive models and their extensions. TEST, 29(2): 307–339, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.springer.

com/article/10.1007/s11749-020-00711-5;   http://
link.springer.com/content/pdf/10.1007/s11749-020-

00711-5.pdf. See comments [Eil20, GS20, Kne20] and rejoinder [Woo20b].

Wood:2020:RIC

[Woo20b] Simon N. Wood. Rejoinder on: Inference and computation with generalized additive models and their extensions. TEST, 29(2):354–358, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-020-00716-0. See

[Woo20a, Eil20, GS20, Kne20].

Wermuth:2012:RSR

[WS12a] Nanny Wermuth and Kayvan Sadeghi. Rejoinder on: “Sequences of regressions and their independences”. TEST, 21(2):274–279, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-012-0291-5. See

[WS12b, Cas12, DFK12, LR12, Sta12, Raj12].

Wermuth:2012:SRT

[WS12b] Nanny Wermuth and Kayvan Sadeghi. Sequences of regressions and their independences. TEST, 21(2):215–252, June 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-012-0290-6. See comments [Cas12, DFK12, LR12, Sta12, Raj12] and rejoinder [WS12a].

Wang:2024:PPP

[WS24] Zhijian Wang and Yunquan Song. Privacy-preserving parametric inference for spatial autoregressive model. TEST, 33 (3):877–896, September 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-024-00928-8.

Wang:2015:CCW

[WSCH15] Xuejun Wang, Aiting Shen, Zhiyong Chen, and Shuhe Hu. Complete convergence for weighted sums of NSD random variables and its application in the EV regression model. TEST, 24(1):166–184, March 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0402-6.

Wei:1995:MSE

[WT95] L. Wei and G. Trenkler. Mean square error matrix superiority of Empirical Bayes Estimators under misspecification. TEST, 4(1):187–205, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02563109. See correc-

tions [WT96].

Wei:1996:CMS

[WT96] L. Wei and G. Trenkler. Corrections to “Mean square error matrix superiority of empirical Bayes estimators under misspecification. TEST, 5(1):247–248, June 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02562691;
http://link.springer.com/content/pdf/10.1007/BF02562691.

pdf. See [WT95].

Wang:2017:NTS

[WTZL17] Haiyan Wang, Bo Tong, Huaiyu Zhang, and Xukun Li. New two-sample tests for skewed populations and their connection to theoretical power of bootstrap-t test. TEST, 26(3):661– 683, September 2017. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0530-x.

Wellalage:2023:HOA

[WVT23] Basitha K. Hewa Wellalage, Igor Volobouev, and A. Alexandre Trindade. High-order asymptotic approximations for improved inference under exceptionally low false positive error rates. TEST, 32(4):1276–1306, December 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00873-y.

Wiper:2006:BAB

[WW06] Michael P. Wiper and Simon P. Wilson. A Bayesian analysis of beta testing. TEST, 15(1):227–255, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595426.

Wei:2012:SIR

[WW12] Chuanhua Wei and Qihua Wang. Statistical inference on restricted partially linear additive errors-in-variables models. TEST, 21(4):757–774, December 2012. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-011-0279-6.

Wu:2018:WVS

[WWHY18] Yi Wu, Xuejun Wang, Shuhe Hu, and Lianqiang Yang. Weighted version of strong law of large numbers for a class of random variables and its applications. TEST, 27(2):379– 406, June 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0550-6.

Wang:2022:MST

[WWT22] Jun Wang, Dianpeng Wang, and Yubin Tian. Multidimensional specification test based on non-stationary time series. TEST, 31(2):348–372, June 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00780-0.

Wang:2016:SCB

[WWY16] Jiangyan Wang, Suojin Wang, and Lijian Yang. Simultaneous confidence bands for the distribution function of a finite population and of its superpopulation. TEST, 25(4):692– 709, December 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-016-0491-5.

Wang:2019:ALK

[WWY^{+}19] Xuejun Wang, Yi Wu, Wei Yu, Wenzhi Yang, and Shuhe Hu. Asymptotics for the linear kernel quantile estimator. TEST, 28(4):1144–1174, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-019-00627-9.

Wang:2014:CCW

[WXH^{+}14] Xuejun Wang, Chen Xu, Tien-Chung Hu, Andrei Volodin, and Shuhe Hu. On complete convergence for widely orthantdependent random variables and its applications in nonparametric regression models. TEST, 23(3):607–629, September 2014. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0365-7.

Wang:2009:OPS

[WY09] Bing Xing Wang and Keming Yu. Optimum plan for stepstress model with progressive type-II censoring. TEST, 18(1): 115–135, May 2009. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-007-0060-z.

Wang:2023:NTL

[WY23] WenWu Wang and Ping Yu. Nonequivalence of two leastabsolute-deviation estimators for mediation effects. TEST, 32(1):370–387, March 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-022-00837-8.

Wang:2024:BAT

[WYH24] Min Wang, Keying Ye, and Zifei Han. Bayesian analysis of testing general hypotheses in linear models with spherically symmetric errors. TEST, 33(1):251–270, March 2024. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00892-9.

Wu:2012:ETR

[WZ12] Jianhong Wu and Lixing Zhu. Estimation of and testing for random effects in dynamic panel data models. TEST, 21 (3):477–497, September 2012. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-011-0259-x.

Wu:2019:CDS

[WZ19] Changbao Wu and Shixiao Zhang. Comments on: Deville and Särndal’s Calibration: revisiting a 25-years-old successful optimization problem. TEST, 28(4):1082–1086, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00682-2. See [DT19a, DT19b].

Wynn:1994:TSS

[WZR94] Henry P. Wynn, Anatoly A. Zhigljavsky, and Juan Romo. The theory of search from a statistical viewpoint. TEST, 3(2):1–45, December 1994. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02562692.

Wu:2018:EVS

[WZYW18] Mingzhe Wu, Ming Zheng, Wen Yu, and Ruofan Wu. Estimation and variable selection for semiparametric transformation models under a more efficient cohort sampling design. TEST, 27(3):570–596, September 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0562-2.

Xia:2017:TSR

[Xia17] Yin Xia. Testing and support recovery of multiple highdimensional covariance matrices with false discovery rate control. TEST, 26(4):782–801, December 2017. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0533-7.

Xun:2020:EQD

[XTZ20] Li Xun, Li Tao, and Yong Zhou. Estimators of quantile difference between two samples with length-biased and right-censored data. TEST, 29(2):409–429, June 2020. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-019-00657-3.

Xia:2022:SLS

[XYY22] Siwei Xia, Yuehan Yang, and Hu Yang. Sparse Laplacian shrinkage with the graphical lasso estimator for regression problems. TEST, 31(1):255–277, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-021-00779-7.

Xie:2018:MPD

[XZ18] Chuanlong Xie and Lixing Zhu. A minimum projecteddistance test for parametric single-index Berkson models. TEST, 27(3):700–715, September 2018. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0568-9.

Xu:2016:EEM

[XZHW16] Peirong Xu, Jun Zhang, Xingfang Huang, and Tao Wang. Efficient estimation for marginal generalized partially linear single-index models with longitudinal data. TEST, 25 (3):413–431, September 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-015-0462-2.

Yang:1995:IRP

[Yan95] Y. Yang. Invariance of the reference prior under reparametrization. TEST, 4(1):83–94, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02563104.

Yang:1999:ETE

[Yan99] Zhelin Yang. Estimating a transformation and its effect on Box–Cox T -ratio. TEST, 8(1):167–190, June 1999. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595868.

Yang:2000:NSR

[Yan00] Zhenlin Yang. A new statistic for regression transformation. TEST, 9(1):123–131, June 2000. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/BF02595854.

Ye:1995:SRP

[Ye95] K. Ye. Selection of the reference priors for a balanced random effects model. TEST, 4(1):1–17, June 1995. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02563100.

Yekutieli:2008:CCF

[Yek08] Daniel Yekutieli. Comments on: “Control of the false discovery rate under dependence using the bootstrap and subsampling”. TEST, 17(3):458–460, November 2008. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-008-0131-9. See [RSW08a] and rejoinder [RSW08b].

Yekutieli:2015:BTC

[Yek15] Daniel Yekutieli. Bayesian tests for composite alternative hypotheses in cross-tabulated data. TEST, 24(2):287–301, June 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0407-1.

Yu:2019:LMG

[YH19] Shun Yu and Xianzheng Huang. Link misspecification in generalized linear mixed models with a random intercept for binary responses. TEST, 28(3):827–843, September 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-018-0602-6.

Yue:2019:IEQ

[YLL19] Lili Yue, Gaorong Li, and Heng Lian. Identification and estimation in quantile varying-coefficient models with unknown link function. TEST, 28(4):1251–1275, December 2019. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-019-00638-6.

Yamada:2015:KTM

[YRR15] Tomoya Yamada, Megan M. Romer, and Donald St. P. Richards. Kurtosis tests for multivariate normality with monotone incomplete data. TEST, 24(3):532–557, September 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0423-1.

Youndje:1996:OSH

[YSV96] É. Youndjé, P. Sarda, and P. Vieu. Optimal smooth hazard estimates. TEST, 5(2):379–394, December 1996. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02562624.

Yuan:2005:SCM

[Yua05] Ming Yuan. Semiparametric censorship model with covariates. TEST, 14(2):489–514, December 2005. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http:

//link.springer.com/article/10.1007/BF02595415.

Yuan:2024:RIV

[Yua24] Mingao Yuan. On the Randić index and its variants of network data. TEST, 33(1):155–179, March 2024. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00887-6.

Youndje:2008:OBS

[YW08] Élie Youndjé and Martin T. Wells. Optimal bandwidth selection for multivariate kernel deconvolution density estimation. TEST, 17(1):138–162, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0027-5.

You:2015:VCA

[YWLZ15] Jinhong You, Alan T. K. Wan, Shu Liu, and Yong Zhou. A varying-coefficient approach to estimating multi-level clustered data models. TEST, 24(2):417–440, June 2015. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-014-0419-x.

Yao:2016:PEE

[YWZ16a] Fang Yao, Yichao Wu, and Jialin Zou. Probabilityenhanced effective dimension reduction for classifying sparse functional data. TEST, 25(1):1–22, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-015-0470-2. See comments [Agu16,

AV16, SNC16, FB16, BL16a, ZL16, Zha16] and rejoinder [YWZ16b].

Yao:2016:RPE

[YWZ16b] Fang Yao, Yichao Wu, and Jialin Zou. Rejoinder on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):52–58, March 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0478-7;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0478-

7.pdf. See [YWZ16a, Agu16, AV16, SNC16, FB16, BL16a, ZL16, Zha16].

Yu:2019:JEL

[YZ19] Xue Yu and Yichuan Zhao. Jackknife empirical likelihood inference for the accelerated failure time model. TEST, 28(1): 269–288, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0601-7.

Yuan:2019:MFM

[YZHG19] Chaofeng Yuan, Wensheng Zhu, Xuming He, and Jianhua Guo. A mixture factor model with applications to microarray data. TEST, 28(1):60–76, March 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0585-3.

Yang:2017:LDR

[YZWH17] Wenzhi Yang, Zhangrui Zhao, Xinghui Wang, and Shuhe Hu. The large deviation results for the nonlinear regression model with dependent errors. TEST, 26(2):261–283, June 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-016-0509-z.

Zong:2023:CCB

[ZB23] Qingying Zong and Jonathan R. Bradley. Criterion constrained Bayesian hierarchical models. TEST, 32(1):294– 320, March 2023. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-022-00834-x.

Zhao:2011:NIB

[ZBS11a] Xingqiu Zhao, N. Balakrishnan, and Jianguo Sun. Nonparametric inference based on panel count data. TEST, 20(1):1–42, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0223-1. See

comments [DC11, He11, Ish11, Par11, Ton11, dUÁ11] and rejoinder [ZBS11b].

Zhao:2011:RNI

[ZBS11b] Xingqiu Zhao, N. Balakrishnan, and Jianguo Sun. Rejoinder on: “Nonparametric inference based on panel count data”. TEST, 20(1):65–71, May 2011. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-010-0230-2. See

[ZBS11b, DC11, He11, Ish11, Par11, Ton11, dUÁ11].

Zeller:2016:RMR

[ZCL16] Camila B. Zeller, Celso R. B. Cabral, and Víctor H. Lachos. Robust mixture regression modeling based on scale mixtures of skew-normal distributions. TEST, 25(2):375– 396, June 2016. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-015-0460-4.

Zhao:2018:CMD

[ZF18] Haibing Zhao and Wing Kam Fung. Controlling mixed directional false discovery rate in multidimensional decisions with applications to microarray studies. TEST, 27(2):316– 337, June 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0547-1.

Zhang:2015:ECS

[ZFX15] Jun Zhang, Zhenghui Feng, and Peirong Xu. Estimating the conditional single-index error distribution with a partial linear mean regression. TEST, 24(1):61–83, March 2015. CODEN ???? ISSN 1133-0686 (print), 1863-8260

(electronic). URL http://link.springer.com/article/
10.1007/s11749-014-0395-1.

Zewotir:2007:UAR

[ZG07] Temesgen Zewotir and Jacky S. Galpin. A unified approach on residuals, leverages and outliers in the linear mixed model. TEST, 16(1):58–75, May 2007. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-006-0001-2.

Zhu:2014:RST

[ZGGX14a] Xinxin Zhu, Marc G. Genton, Yingzhong Gu, and Le Xie. Rejoinder on: “Space–time wind speed forecasting for improved power system dispatch”. TEST, 23(1):45–50, March 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-014-0356-8;   http://link.
springer.com/content/pdf/10.1007/s11749-014-0356-

8.pdf. See [ZGGX14b, Pin14, Muñ14, TL14, Her14].

Zhu:2014:STW

[ZGGX14b] Xinxin Zhu, Marc G. Genton, Yingzhong Gu, and Le Xie. Space–time wind speed forecasting for improved power system dispatch. TEST, 23(1):1–25, March 2014. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-014-0351-0. See comments [Pin14, Muñ14, TL14, Her14] and rejoinder [ZGGX14a].

Zhang:2014:AMM

[Zha14] Chunming Zhang. Assessing mean and median filters in multiple testing for large-scale imaging data. TEST, 23(1):51– 71, March 2014. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-013-0341-7.

Zhang:2016:CPEb

[Zha16] Hao Helen Zhang. Comments on: “Probability-enhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):47–51, March 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11749-015-0477-8. See [YWZ16a] and rejoinder [YWZ16b].

Zhao:2022:WFN

[Zha22] Zhigen Zhao. Where to find needles in a haystack? TEST, 31(1):148–174, March 2022. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-021-00775-x.

Zhao:2008:IOR

[ZJ08] Yinshan Zhao and Harry Joe. Inferences for odds ratio with dependent pairs. TEST, 17(1):101–119, May 2008. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-006-0025-7.

Zezula:2018:TMR

[ZKR18] Ivan Zezula, Daniel Klein, and Anuradha Roy. Testing of multivariate repeated measures data with block exchangeable covariance structure. TEST, 27(2):360–378, June 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-017-0549-z.

Zhang:2016:CPEa

[ZL16] Chong Zhang and Yufeng Liu. Comments on: “Probabilityenhanced effective dimension reduction for classifying sparse functional data”. TEST, 25(1):44–46, March 2016. CODEN ???? ISSN 1133-0686 (print), 1863-

8260 (electronic).  URL http://link.springer.com/
article/10.1007/s11749-015-0474-y;   http://link.
springer.com/content/pdf/10.1007/s11749-015-0474-

y.pdf. See [YWZ16a] and rejoinder [YWZ16b].

Zhao:2017:JDB

[ZLWH17] Yan-Yong Zhao, Jin-Guan Lin, Hong-Xia Wang, and Xing-Fang Huang. Jump-detection-based estimation in timevarying coefficient models and empirical applications. TEST, 26(3):574–599, September 2017. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0525-7.

Zheng:2016:SIG

[ZLYH16] Shuzhuan Zheng, Rong Liu, Lijian Yang, and Wolfgang K. Härdle. Statistical inference for generalized additive models: simultaneous confidence corridors and variable selection. TEST, 25(4):607–626, December 2016. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-016-0480-8.

Zimmerman:2001:PMG

[ZNAG^{+}01] Dale L. Zimmerman, Vicente Núñez-Antón, Timothy G. Gregoire, Oliver Schabenberger, Jeffrey D. Hart, Michael G. Kenward, Geert Molenberghs, Geert Verbeke, Mohsen Pourahmadi, Philippe Vieu, Dela L. Zimmerman, Vicente Núñez-Antón, Timothy G. Gregoire, Oliver Schabenberger, Jeffrey D. Hart, and et al. Parametric modelling of growth curve data: an overview. TEST, 10(1):1–73, June 2001. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02595823.

Zhang:2019:ACS

[ZNSW19] Fode Zhang, Hon Keung Tony Ng, Yimin Shi, and Ruibing Wang. Amari–Chentsov structure on the statistical manifold of models for accelerated life tests. TEST, 28(1):77– 105, March 2019. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-018-0587-1.

Zhao:2018:VES

[ZPH18] Jingxin Zhao, Heng Peng, and Tao Huang. Variance estimation for semiparametric regression models by local averaging. TEST, 27(2):453–476, June 2018. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://link.

springer.com/article/10.1007/s11749-017-0553-3.

Zhou:2024:CDI

[ZS24] Ling Zhou and Peter X. K. Song. Comments on: Data integration via analysis of subspaces (DIVAS). TEST, 33(3):689– 692, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00955-5. See [PJH^{+}24a].

Zhang:2006:NDL

[ZY06] Jin Zhang and Keming Yu. The null distribution of the likelihood-ratio test for one or two outliers in a normal sample. TEST, 15(1):141–150, June 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02595422.

Zhang:2018:SSC

[ZY18] Yuanyuan Zhang and Lijian Yang. A smooth simultaneous confidence band for correlation curve. TEST, 27(2):247– 269, June 2018. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL http://link.springer.com/
article/10.1007/s11749-017-0543-5.

Zhang:2023:SCR

[ZZ23] Jiandong Zhang and Yiying Zhang. Stochastic comparisons of relevation allocation policies in coherent systems. TEST, 32(3):865–907, September 2023. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL https://link.

springer.com/article/10.1007/s11749-023-00855-0.

Zhang:2019:EHT

[ZZF19] Jun Zhang, Junpeng Zhu, and Zhenghui Feng. Estimation and hypothesis test for single-index multiplicative models. TEST, 28(1):242–268, March 2019. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11749-018-0586-2.

Zhao:2024:MCP

[ZZT24] Wenbiao Zhao, Lixing Zhu, and Falong Tan. Multiple change point detection for high-dimensional data. TEST, 33(3):809– 846, September 2024. CODEN ???? ISSN 1133-0686 (print),

1863-8260 (electronic). URL https://link.springer.com/
article/10.1007/s11749-024-00926-w.

Zhong:2025:IPA

[ZZY25] Chen Zhong, Yuanyuan Zhang, and Lijian Yang. Inference and prediction for ARCH time series via innovation distribution function. TEST, 34(1):48–68, March 2025. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-024-00949-3.